A Bibliography of Publications in ACM SIGPLAN Notices, 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 3.30

Title word cross-reference

#16G [797].

+ [1636]. 2 [1821]. = [328, 1636]. ^{∗} [1585]. ^{3} [1104]. ^{T M} [1826]. _{Ada} [140]. ∀ [1849]. λ [1886]. n [1107]. N ≤ 8 [997, 1144, 1010]. ω [1913]. ′ [1636].

* [918].

-calculus [1886]. -dimensional [1821]. -ively [1520].

/information [195].

0-201-17928-8 [1613]. 002R [1355].

1 [1624, 1283, 1077]. 1/83 [509]. 100 [1189]. 11 [95, 390, 139, 393]. 11/780 [139]. 1100 [58]. 16-bit [1827]. 164 [721]. 1838 [1611]. 1978 [39]. 1980 [1942, 245]. 1980’s [760]. 1981 [1945, 312, 311, 345, 354]. 1982 [1947]. 1983 [1951, 1952]. 1984 [1954, 1955]. 1985 [1141]. 1986 [1959, 1960, 1961, 18, 19]. 1987 [20, 25, 1238, 1356]. 1989 [39]. 198x [39, 1387].

2’ [832, 1003, 941, 1382, 918, 1190, 1499, 1005, 1192, 919, 920, 949, 1006, 1144, 1195, 950, 1375, 526, 843, 1157, 837, 1632, 800, 782, 913, 748, 1154, 1148, 791, 790, 1553, 452, 634, 459, 1182, 688, 842, 966]. 2000 [1832, 1843, 1846].

3 [1923, 1476]. 32 [375, 371]. 32-bit [1827]. 370 [1202].

432 [650, 387]. 4381 [1268].

6 [619]. 68 [513, 66, 107]. 6th [768].

74 [408]. 754 [11, 1938, 1141]. 754-1985 [1937, 1141, 1938]. 77 [340, 1226]. 780 [139].

80 [1065, 1639, 1803, 1272, 1063, 1666, 493, 1333, 1099, 1793, 1110, 1064, 1086, 1523, 1521]. ’81 [1949]. ’82 [1947]. ’83 [1951, 760, 1953]. ’84 [1957]. ’86 [1959, 1961, 1962, 424, 1068]. ’87 [1963, 1965]. ’88 [1966, 1967, 1355, 36]. 88-002R [1355]. 8800 [1280]. ’89 [1971, 1972, 1973, 1976, 1979]. 8X [22, 13, 39, 31, 29, 28, 24, 32, 21, 30, 35, 14, 1482].

= [787, 491].

abandon [496, 443]. ABCL [1624, 1077]. ABCL/1 [1624, 1077]. Abelson [1160]. ABF [1229]. abridgement [30]. absolute [789]. Abstract [1421, 217, 663, 713, 1032, 656, 1817, 1922, 1190, 1033, 678, 1034, 1485, 1035, 1180, 1920, 252, 1917, 1896, 1037, 1195, 1919, 1852, 1039, 73, 1040, 1894, 1041, 948, 1859, 1042, 1098, 1043, 1928, 1452, 1453, 1044, 1148, 69, 1046, 352, 1047, 1048, 1181, 957, 1049, 966, 1050, 675, 1036, 229]. Abstracting [642]. Abstraction [650, 176, 558, 223, 1537, 1451, 1729, 192, 643, 1499, 658, 554, 1916, 1868, 205, 209, 1340, 1399, 567, 551, 178, 222, 226, 1876, 227, 540, 857]. Abstractions [860, 520, 1545, 201, 568]. abundance [1414]. abuse [581]. acce [781]. acce-specifications [781]. Accelerator [1081]. acceptance [928]. accepted [957]. Acceptive [575, 671]. access [1716, 206, 1632, 1178, 1738, 278]. access-control [278]. accesses [1434]. according [927]. ACM [1942, 1945, 1954, 1953, 1944, 1975, 1958, 1977, 1956, 1964, 1968, 1952, 1949, 1969, 1943, 1957, 1978, 1946, 1970, 246].

ACM-SIGPLAN [1942]. ACM/SIGPLAN [1970]. Acore [1644]. acquisition [1750]. acronym [1690]. Action [559, 1563]. actions [683, 1203, 1204, 1422]. Active [1315, 1595]. activities [191, 1685]. actor [1019, 1638, 1642, 1643, 1644]. Actors [687, 1032, 1639, 1640, 1641, 1036, 1020]. Actra [1645]. ad [1911, 1911]. ad-hoc [1911]. Ada [1942, 1808, 523, 127, 918, 64, 236, 65, 235, 115, 254, 1312, 169, 246, 245, 247, 100, 149, 438, 135, 151, 343, 154, 742, 322, 323, 744, 121, 140, 1107, 144, 250, 659, 102, 324, 150, 1696, 440, 132, 131, 949, 148, 129, 124, 133, 950, 153, 910, 138, 356, 85, 143, 152, 141, 145, 243, 1187, 1122, 241, 341, 106, 130, 108, 134, 256, 387, 137, 1701, 1450, 139, 136, 244, 700, 147, 1185, 329, 320, 338, 142, 1002, 665, 319, 370]. Adding [1625, 1428, 902, 1557]. Addison [1613]. Addison-Wesley [1613]. Addition [270]. address [1266, 1340, 1399]. addressable [1827]. addressing [1273, 1788, 382]. ADE [1129]. adequacy [1890]. ADTs [1157]. advanced [454, 607, 700]. advancement [1686]. advances [1334]. advice [255]. after [631, 632]. again [1190, 652]. against [121, 1698, 479, 1181, 966]. agent [1773]. agents [1774]. ages [1773]. aggregate [972]. ahead [982]. AI [1651]. aid [612]. Aided [706]. aids [289, 469, 1888]. Aldus [1779]. Alex [954, 1112]. algebra [574, 336, 542, 859, 1778]. algebraic [1079, 1030, 1034, 645, 514, 98, 1366]. algebras [230]. ALGOL [66, 107, 513, 1423, 441]. Algol68 [168]. algorithm [498, 234, 1135, 236, 1512, 298, 1752, 414, 495, 328, 1010, 777, 55, 712, 1510, 1547, 1754]. Algorithmic [1104, 1584]. Algorithms [147, 233, 1549, 5, 981, 1248, 1170, 1497, 689]. alias [1910]. aliases [1873]. allocated [1902]. Allocation [794, 442, 398, 1737, 416, 1431, 739, 1233, 377, 1736, 994, 692, 995, 1439, 70, 1510]. allows [421]. alone [1523]. Alonzo [1760]. Alternate [785, 29, 334]. alternative [1, 378, 896, 1203, 1204, 516, 1096, 1097, 169, 801, 636]. Alternatives [838]. Alto [1961, 1964, 1962]. Ambiguities [452, 1477, 1182, 771]. Ambiguity [370, 474]. ambiguous [1705]. American [39, 1141]. Amit [894]. Amoeba [1561]. among [185, 744, 647]. AMPL [775]. analog [1470]. analogy [1166]. analyser [659]. analysers [1506]. Analysis [10, 850, 677, 1544, 1451, 1368, 322, 985, 1436, 1570, 1135, 1901, 1363, 942, 1603, 537, 854, 1437, 1438, 1910, 1900, 1359, 1576, 1163, 263, 1715, 1179, 1448, 979, 751, 1294, 1753, 735, 1861, 556, 1486, 899, 59, 392, 624, 533, 1895, 1859, 1886, 1226, 1902, 1909, 1456, 1449, 168, 391, 612, 1713, 1888, 937, 731]. analytic [1092]. analyzer [417, 51, 136, 1113]. Analyzing [1677, 729, 427]. Anatomy [1442, 951]. Andrew [1613]. Angeles [231, 1949, 1943, 1946]. Animated [1531]. ANNA [143, 1701]. annotated [302]. annotating [143]. Annual [1963, 1971, 1972]. anomalies [1738]. anomaly [1821]. Anonymous [1841]. ANS [771]. ANSI [577, 1937, 1938, 1141]. ANSI/IEEE [1937, 1141, 1938]. answer [267, 268]. anybody [769]. APL [483, 1549, 273, 927, 447, 959, 400, 168]. APLOMB [400]. Application [186, 1759, 349, 643, 1129, 590, 1678, 1779, 1071, 1090, 1903, 222, 1519, 1011, 1081]. Applications [1959, 1493, 11, 1520, 1976, 192, 132, 1137, 1652, 1282, 1070, 1748, 1909, 168, 1523, 1970]. applicative [1552, 1494, 470, 434, 1426, 1215]. Applying [1329]. Appraising [1865]. Approach [809, 708, 601, 1898, 1640, 942, 78, 1445, 817, 199, 1576, 607, 290, 5, 554, 1090, 1448, 591, 1223, 1643, 52, 209, 1157, 946, 215, 1328, 1338, 732, 1765, 1165, 560, 1540, 1134, 1767, 630, 1808]. Approaches [1050, 1263, 842]. Approximations [1358]. April [1954, 1974, 1956]. APSE [121, 1611]. arbitrary [982]. ARC [783]. Arcadia [1588]. Architectural [1974, 1260, 384, 378, 389, 726, 394, 1275]. Architecture [1833, 396, 156, 1078, 1281, 1071, 969, 1262, 1203, 1204, 374, 1277, 1753, 1070, 390, 1009, 388, 1588, 1646, 476, 1269]. architectures [1647, 1648, 1671, 901, 1261, 384, 1258]. Arcturus [700]. area [85, 556]. aren’t [1184]. argument [455]. Arguments [490, 491]. Arithmetic [7, 9, 11, 1938, 1940, 8, 1936, 1872, 485, 1939, 1935, 1937, 1141]. array [1482, 1481, 1577, 971, 1904, 636]. arrays [959, 1836, 45]. arrow [255]. article [113]. artificial [175]. Arun [635]. ascent [1470]. Aspects [1748, 431]. Aspenäs [476]. assembler [451, 1484, 238]. assembly [729, 663]. Assertions [56, 1620]. assessed [1820]. assessment [1688, 469, 868]. assessments [827]. assigning [498]. assignment [1908, 1882]. assignments [1877, 787]. assist [1259]. assistance [611]. assistant [1597, 1680]. assisted [608]. associate [1534]. Associating [208]. assumptions [326, 75]. astro [1256]. Asynchronous [1836, 1548, 1904, 1459]. Atlanta [1967]. atomic [555]. attainable [686]. attitudes [97]. Attribute [531, 848, 553, 733, 1901, 1712, 440, 424, 723, 977, 1723, 972, 623, 724, 734, 423, 1599, 74, 976, 1289, 1909, 1724, 1454, 1547]. attributed [630]. Attributes [847, 498, 530, 536, 853, 1132, 1288, 1539]. attribution [1600, 1289]. Augmentation [1098]. August [18]. Austin [1971, 1972]. authorization [1428]. AUTO [1615]. automata [1844]. Automated [17, 330, 1198]. Automatic [740, 1854, 1485, 1717, 727, 1569, 977, 728, 1440, 1886, 469, 1205, 1378, 128, 778, 1376, 1133, 1777, 412]. automatically [1455, 737, 931]. Automating [1689]. automation [563]. autonomy [101]. available [246]. AVANCE [1532]. average [1739]. avoid [484]. avoids [33]. AWB [1129]. AWB-ADE [1129]. AWK [649]. Axiomatic [211, 221]. Axioms [1850].

B [486, 815, 905]. back [1732, 1462, 1463]. background [302, 1284]. backing [1787]. backtracking [1524]. Backus [1171]. backward [761]. backward-directed [761]. bad [1769, 1184]. Banquet [1351, 1410]. base [189, 1126, 207, 228]. Based [1975, 1977, 1797, 1451, 1562, 1556, 429, 1033, 1080, 1695, 1305, 1766, 101, 1173, 739, 1593, 1058, 1522, 424, 1723, 1792, 696, 1073, 701, 253, 694, 1145, 1597, 1680, 824, 1753, 73, 426, 1084, 526, 843, 1874, 1563, 586, 545, 863, 1840, 1328, 1338, 935, 1785, 1684, 1253, 1460, 1199, 840, 1393, 587, 1540, 716, 278, 1591, 1518, 288, 792, 1515, 1309, 1634, 1851, 1693, 1686, 1635, 967, 571, 546, 864, 87, 225]. bases [188, 640, 208, 210, 218, 226, 1118, 217]. Basic [1026, 1042, 234, 835, 577, 521]. BASICs [1193]. basis [536, 853, 528, 845, 314, 480]. BASIX [571]. BBN [1492]. be [180, 1897, 1869, 1146, 242, 957]. BEG [1732]. beginning [907]. behavior [1562, 1495, 1072]. behaviors [1670]. Behaviour [182]. behavioural [197]. being [518]. Belady [653]. BELLMACTM [375]. BELLMACTM-32 [375]. Benchmark [1504, 1374, 1430, 304, 1381, 1503, 1472, 1528]. benefits [1619, 1700]. best [1617]. BETA [825]. better [1699, 1165, 1160]. between [185, 234, 874, 1292, 1434, 1067, 782, 1598, 1336, 1101, 956, 320, 338, 771]. Beyond [460, 880, 241]. Bibliography [638, 330, 456, 759, 74, 313, 162, 667, 472, 126, 948, 753, 302, 1371, 868]. big [92, 316, 15]. Binary [1937, 1938, 1141, 1940, 8, 499, 263, 1939, 830, 98, 1287, 1935]. Binding [1856, 1197, 1886]. Bindings [1388]. BINS [811]. birds [144]. Bisimulation [1897, 1934]. Bit [117, 1827, 359]. blackbox [1586]. BLAS [906]. BLAZE [1632]. Blit [616, 617]. Block [1027, 1040, 1203, 1204, 285]. block-and-actions [1203, 1204]. block-structured [285]. blocks [1239, 841]. Bn [891]. BNF [125]. BOF [1342, 1401]. Book [1808, 1613, 940, 1238]. books [280, 321, 353, 361, 272, 312, 311, 345, 354]. Boolean [1095]. Boston [1942, 1947, 1974, 1968]. bottom [1863]. bottom-up [1863]. bound [678]. boundary [1067]. boxcharts [1247]. bracketed [637]. Bracketing [264]. branches [263]. breaking [297]. Brevity [333]. brigade [1686]. BrouHaHa [1325]. bucket [1686]. bucket-brigade [1686]. BUFFERS [477]. bug [1701]. bugs [579]. build [1311, 1545]. builder [1697, 788]. Building [399, 1071, 1524, 435, 1478, 648, 1769, 898]. BURS [1903]. Butterfly [1492]. bypass [1743]. bytecodes [1062].

C [1429, 1501, 1457, 72, 442, 398, 885, 1749, 510, 678, 908, 1704, 493, 1333, 1227, 377, 1278, 817, 342, 1006, 998, 951, 1734, 1777, 824, 1613, 1511, 930, 1369, 1199, 1121, 1546, 1823, 352, 1014, 1047, 963, 964, 1370, 1519, 1196, 1755, 1566]. C* [1483]. C-based [1199]. C-language [442, 398]. C. [523]. CA [1951, 1961, 1953, 1950, 1966, 1949]. cache [1743, 377]. caches [1266]. caching [1933]. CAD [229]. CAI [828]. CAJOLE [305]. CAJOLE-an [305]. calculating [1160]. calculation [1008]. calculus [1877, 1871, 1886, 1913, 1918, 1899]. Calif. [1946, 1962]. California [1952, 1977, 1964]. call [419, 478, 242, 986]. CALLed [242]. CALLing [242]. calls [1747, 1441, 1698, 1177, 379, 1459]. Can [1033, 341, 13, 1869, 1703, 1146, 242]. Canada [1955, 1957]. candidate [754, 755]. can’t [1897]. Cantor [1638]. capabilities [1835]. capability [46, 129, 386]. capacity [747]. Capsules [566]. card [1693]. card-marking [1693]. Case [1808, 1447, 1747, 1276, 121, 907, 1088, 477, 1698, 479, 819, 1124, 1460, 1336, 1370, 62, 393, 1196]. cases [500]. categorized [126]. category [667]. CCAL [1237]. Cedar [1741, 549, 867, 548, 866]. cell [488, 310]. cellular [892]. Centaur [1589]. central [1283]. certain [870, 1422]. chaining [968]. CHAMIL [62]. change [480, 1050]. changeable [1091]. changes [37]. changing [255, 1103, 1025]. chaotic [1810]. character [166, 787]. characteristic [505, 797]. characterization [334]. Characterizing [1495]. Charrette [137]. chart [324, 114]. Chatin [498]. Cheap [885, 1267]. check [1691]. Checking [288, 746, 874, 1428, 1816, 1083, 745, 418, 1912, 1263, 1317]. children [1773]. chill [390]. chip [383]. choice [956]. choosing [403]. Cichelli [3, 4]. Cint [1227]. circuit [1490]. circular [977, 724, 1454]. circularity [1909]. CISCs [1276]. CLAM [1318]. CLANG [924]. clarified [771]. Clarify [1419]. clarity [347, 333]. Class [1476, 1303, 1308, 1893, 1792, 512, 1857, 80, 1224, 1225, 1891, 1055, 1730, 1458, 1454]. Classes [352, 191, 1801, 1095, 1067, 1805]. classical [1869]. Classification [1031, 1049, 1179]. clause [1387, 1819]. clauses [1106, 635]. cleanliness [1236]. clique [1736]. CLOS [1783, 1785]. closure [1931, 965]. closure-passing [1931]. CLP [1916]. Clu [1751]. coagulation [717]. COBOL [1186, 878, 408, 242, 1198, 901, 518, 751]. Code [717, 718, 1735, 411, 744, 1173, 719, 1445, 428, 1512, 1245, 729, 992, 1440, 1733, 330, 410, 1471, 1469, 993, 426, 412, 686, 390, 427, 1903, 1221, 1222, 430, 384, 737, 139, 1151, 664, 405, 1509]. code-generator [412]. codes [166]. Coding [304, 373]. coercion [1209]. cognitively [716]. cognitively-based [716]. cohabit [1876]. Coherency [1266]. collecting [1203, 1204, 1887]. collection [1432, 1433, 1235, 1234, 1765, 368, 1554, 1693, 1686]. Collections [621, 399]. collector [1176, 1764]. color [1521]. Coloring [1737, 498, 416, 739]. colouring [1752]. combination [1785]. combinator [1722, 1393]. combinator-based [1393]. Combinatory [1173]. combine [1699]. Combined [683, 1221, 1222]. Combining [1334, 1658, 112, 1397, 490, 491]. Command [641, 121, 573, 685, 87, 333]. commands [481, 496, 443]. Comment [1387, 690, 115, 1500, 71]. Comments [950, 1620, 1560, 1201, 1546, 127, 355, 246, 1614, 100, 1108, 113, 620, 243, 106, 475, 1386, 631, 632, 35, 1689, 635, 637, 666]. commercial [1384]. committee [1391]. committees [1558]. Common [474, 120, 1622, 1623, 91, 1611, 1053, 1355, 1396, 1753, 1015, 1045]. Common-LISP [1396]. CommonLoops [1054, 1313]. CommonObjects [1015, 1045]. Comms [823]. communicating [820, 1918]. Communication [1497, 396, 746, 1740, 811, 204, 556, 1538, 1548, 169, 100]. Communication-sensitive [1497]. Compact [114]. compacting [1234]. Comparative [466, 1179, 471, 937]. Compared [951, 524, 927, 932]. Comparing [576, 271, 276]. Comparison [1249, 1824, 1696, 1528, 773, 1246, 455, 1369, 1022, 1046]. COMPAS [1159]. compatibility [1791, 622, 167]. compatible [1159, 945]. Compilation [971, 1020, 607, 1720, 989, 722, 1930, 401, 745, 732, 436, 730, 1271]. compilations [1240]. Compile [1498, 970]. Compile-time [1498, 970]. compiled [874, 763, 1782, 829, 622]. Compiler [1947, 1955, 1961, 638, 383, 1006, 456, 423, 1904, 435, 725, 406, 1957, 852, 1962, 778, 1443, 1066, 1618, 409, 500, 1364, 429, 432, 140, 763, 1725, 1281, 659, 468, 895, 645, 348, 1723, 606, 720, 94, 425, 129, 133, 433, 1833, 1516, 879, 953, 980, 1442, 408, 990, 421, 994, 1874, 86, 1511, 95, 434, 611, 753, 1009, 1455, 748, 747, 137, 975, 903, 1701, 737, 1456, 422, 868, 1113, 769, 721, 1462, 1463, 535, 393]. compiler-generator [659]. Compilers [155, 1735, 984, 378, 952, 945, 1504, 1381, 1503, 255, 1497, 1818, 915]. Compiling [1482, 1457, 1272, 1860, 552, 1200, 1483, 1036, 1127, 707, 1734, 1346, 1405, 431]. complete [1929]. Completeness [303, 346, 1851, 1870]. complex [1383, 368, 654, 172, 248]. Complexity [505, 850, 1264, 818, 647, 65, 253, 263, 667, 812, 813, 897, 1179, 307, 362, 1008, 533, 1702, 461, 1744, 1372, 44, 627, 119, 269, 689, 71, 1746]. component [310]. Composite [1304, 1667]. composition [292, 1320, 1533]. compositional [1898, 1917]. compound [113]. compressing [729]. Computable [1869, 259, 260]. Computation [1873, 1542, 741, 1389, 1277, 1385, 662, 341, 1933, 1635]. Computational [1792, 1173, 647, 336, 1307]. computationally [1216]. computations [1879, 904, 1444]. Computer [289, 1006, 476, 706, 1482, 1648, 574, 23, 263, 1155, 1162, 774, 892, 1591, 518, 721, 288]. Computer-Aided [706]. computer-based [288]. computers [945, 1282, 1658]. Computing [1641, 1824, 1637, 1638, 1908, 1049]. concatenation [1703]. concept [152, 791]. Concepts [1307, 1529, 146, 463, 206, 1026, 1042, 1098, 228]. conceptual [207, 214, 216, 225, 226]. Concerning [24, 1387]. concerns [1005, 876, 800, 832]. concise [342]. concrete [65, 1812]. Concurrency [1663, 170, 1032, 885, 1625, 1629, 1662, 1354, 1413, 1867]. Concurrent [1975, 1642, 960, 1977, 1681, 1655, 1394, 1584, 1662, 1632, 1631, 1670, 1866, 1447, 466, 1017, 1637, 1432, 1433, 1922, 600, 1360, 1649, 1628, 1678, 146, 1293, 1586, 1850, 1659, 1583, 1707, 1074, 1076, 1016, 1039, 1075, 1660, 1675, 1237, 1668, 1643, 1487, 326, 1669, 1671, 1849, 1672, 1853, 1456, 1540, 1673, 1010, 1582, 471, 1048, 1646, 1676, 1541, 602, 1634, 1331, 1077, 1635, 1002, 1664, 875, 404, 1085]. Concurrently [1531]. conditional [113, 1828, 824]. conditions [669]. Conference [1959, 1967, 1973, 1974, 1949, 1979, 1966, 1976, 1943, 1946, 246]. Confessions [955]. configuration [711, 1609, 1506, 1526]. configurations [1607]. configurator [1679]. Configuring [1523]. Confinement [80]. conflicts [1434, 1813]. conformant [636]. conformant-array [636]. conjunctive [1926]. Connection [1482]. consideration [1431]. considerations [199, 892]. Considered [448, 79, 164, 165, 832, 834]. Consistency [184, 69, 1025]. Consistent [941, 1606]. constancy [488]. constant [984]. constants [166]. constrained [1251, 1664]. Constraint [1298, 1858, 1799, 1916, 1656]. Constraints [198, 197, 1522, 1800, 488, 1518]. construct [465, 516]. constructing [1821, 528, 845]. Construction [1947, 1955, 1961, 638, 659, 1962, 640, 1759, 1833, 1229, 423, 1070, 1799, 753, 1957, 1713, 1378]. constructions [1913]. Constructive [1230, 229, 5, 221]. Constructor [847, 530]. constructs [264, 1560, 559, 1337]. Consul [212]. Containment [512]. Context [503]. Continuation [1931, 1895]. Continuation-passing [1931]. Continuations [852, 829, 1224, 1225, 1891, 1731, 535]. Contour [122]. contrasts [199]. contribution [164, 693]. Contributions [1763, 17, 1012]. Control [331, 431, 1449, 44, 71, 657, 1585, 705, 198, 711, 1332, 554, 453, 1731, 1237, 1560, 522, 571, 1662, 1844, 1147, 1663, 654, 278, 517, 172, 1616, 882, 784, 613, 248]. control-flow [613]. Controllable [1798]. controlled [554, 242]. controller [1831, 1839, 1795]. Controlling [1539, 325, 1693]. convenience [1892]. conventional [1875]. convergence [1032]. conversational [47]. conversion [17, 814, 271]. Convert [943]. converting [895]. cooperating [1663]. cooperation [371]. Cooperative [1846]. coordinates [789]. copies [1067]. Coping [1147]. coprocessor [1268]. Copy [1932]. core [1644]. Cornell [282]. Coroutines [646, 804]. Correct [1895, 255]. Correction [509, 1389]. Corrections [1811]. correctness [1804, 454, 494, 1661, 651]. Correspondence [3, 26, 4, 168]. cost [802, 271]. Counterview [167]. counting [1740, 1187, 365]. coupled [733]. course [1162]. CP [1853, 619]. CP-6 [619]. Cray [1006, 407]. Creating [1828, 1741, 385, 903]. CRISP [1278]. criteria [1661]. Critical [1807, 1156, 824, 1784, 364]. critique [91, 1152, 1160]. cross [250, 124, 238]. crunching [964]. CSP [1192, 1120, 1144]. CST [1628, 1721]. CT [1970]. CUL [795]. CUL-de-SAC [795]. Current [433, 801]. customizable [301]. Customization [1725]. Cyber [831]. CYBIL [831]. Cycle [1591, 1685, 1124, 1184]. cyclomatic [269].

D [829]. daemons [1679]. DAG [1717]. Darwin [704]. Data [1537, 191, 192, 196, 658, 1359, 204, 1583, 979, 751, 567, 218, 178, 222, 116, 226, 1544, 188, 569, 189, 1716, 1126, 259, 260, 520, 678, 643, 193, 1695, 1436, 1901, 640, 417, 644, 1642, 198, 1290, 1180, 553, 512, 283, 252, 1896, 5, 402, 576, 781, 1297, 331, 305, 201, 812, 554, 1383, 1448, 1812, 1316, 1484, 207, 880, 208, 464, 759, 209, 274, 1340, 1399, 482, 332, 1662, 213, 215, 1000, 363, 1862, 1098, 1421, 217, 1291, 891, 276, 663, 1909, 27, 1658, 223, 1205, 69, 352, 504, 1876, 227, 228, 555, 842, 490, 491, 642, 1300]. data [731, 1750, 675, 449, 540, 857, 217, 229, 491, 490]. data-driven [1658]. data-flow [553]. data-oriented [1583, 504]. data-type [27]. Database [888, 224, 1334, 177, 1529, 1335, 192, 193, 195, 196, 1290, 1117, 200, 203, 1304, 1527, 209, 1596, 212, 214, 1116, 594, 559, 220, 558, 1103, 1663, 1336, 1781, 227, 1050, 1025]. database-intensive [192]. databases [1528, 1191, 1152, 203, 205, 176, 1657]. dataflow [1900, 1919, 1189, 1821]. DBMS [1102, 1303]. deadlock [33, 786]. deadlock-free [786]. deadlocks [929]. dealing [275]. debug [585, 614]. debuggable [385]. debugger [600, 614, 1577, 615, 616, 617, 1561, 682, 1579, 610, 593, 613, 619]. Debugging [1950, 1562, 1573, 582, 583, 599, 532, 849, 584, 1952, 1446, 1969, 1581, 1447, 1978, 1566, 1568, 1575, 1213, 1214, 601, 588, 589, 1585, 615, 1267, 1216, 590, 1445, 131, 1571, 606, 1578, 1576, 607, 608, 1586, 598, 609, 1583, 360, 389, 1584, 1563, 586, 1574, 596, 611, 594, 1701, 587, 1215, 612, 602]. Dec. [1964]. December [1942, 231, 345, 245]. Decentralised [571]. Decision [448, 66, 1697, 691, 1702, 834]. declaration [681]. declarative [1864]. decomposition [1854, 1718, 1490, 883]. Deeply [633, 657, 894, 453, 522, 1147]. Default [194]. defense [690, 518]. define [781]. defined [977, 166]. defines [512]. Defining [365, 675, 1065]. definite [1819]. Definition [1707, 676, 151, 93, 670, 454, 1809, 40, 455, 1884, 59, 106, 439, 1853]. definitions [576, 840]. déjà [1193]. Delegation [1306, 1797, 1798]. DELTA [619]. Demeter [1546, 1619]. Demonic [1742]. demonstration [492]. Denormalized [12]. Denotational [1363, 430, 429, 1804, 1868, 1224, 1225, 1884, 1859, 975]. Dependence [1715, 1906, 985, 1714, 1890, 1435, 714, 1907]. dependency [233]. dependent [457]. derivation [648, 1604, 1754]. Deriving [1855]. Descartes [560]. descent [1470, 357]. describability [114]. describe [1746]. Describing [323]. Description [695, 815, 1290, 342, 726, 1511, 215, 1270]. descriptors [154]. Design [1967, 1973, 1443, 299, 1624, 1228, 1099, 1278, 546, 864, 888, 1199, 1979, 1764, 770, 1544, 188, 189, 437, 1311, 588, 1322, 1335, 669, 323, 1577, 1695, 1856, 110, 640, 417, 537, 854, 1142, 1299, 565, 340, 1667, 29, 1832, 5, 286, 153, 1751, 1674, 672, 897, 380, 1339, 1398, 1763, 285, 1461, 1762, 220, 1535, 1787, 1450, 116, 1205, 1350, 1409, 127, 1818, 1309, 62, 1767, 1085, 967, 222]. design2 [1068]. designed [996]. Designers [1936]. Designing [999, 192, 562, 716]. designs [1836, 541, 858]. detailed [1930]. Detecting [1878, 1195, 1434]. Detection [929, 1569, 1701, 1738]. Determining [1384, 1739, 689]. deterministic [495]. developing [1568, 1244, 1591]. Development [1954, 600, 703, 1956, 1964, 1968, 1102, 1840, 708, 1334, 1481, 702, 1129, 537, 854, 1257, 807, 1769, 1652, 1597, 1310, 1155, 1605, 1594, 1124, 1136, 1598, 714, 697, 1790, 1606, 1131, 1615, 1139]. Developments [1708, 1475]. device [1423]. device-independent [1423]. Dhrystone [1472]. diagnosis [413, 415]. diagram [1089, 102]. diagrams [473, 322, 643]. Dialect [1238, 1065]. dialects [931, 1193]. Dialogues [528, 845]. DIANA [1288, 498]. DICE [707]. dichotomy [525]. dictionaries [204]. Diego [1966, 1977]. difference [146]. Differences [782, 320, 338, 771, 168, 1336]. different [657, 249, 1137, 922]. differentiated [1461]. differentiation [1777]. Digital [1936]. Dijkstra [631, 632]. dimension [1428]. Dimensional [677, 1821]. Dimensions [1309, 165]. Direct [388, 986, 1262]. directed [668, 438, 761, 1231, 283, 425, 285, 561, 788, 975, 388, 431, 693]. directions [1807, 1488, 1941]. Directory [1848]. disadvantage [764]. DISC [1707]. discourse [628]. Discovering [993]. discovery [1485]. discrete [1489]. discrete-event [1489]. Discussion [516, 31, 475]. disease [631]. dispatch [1517]. dispatcher [1782]. display [299, 560, 301]. Displaying [1285]. displays [877, 43]. disruptive [712]. Distributed [1235, 598, 599, 1969, 1978, 1069, 1865, 1647, 601, 1562, 1322, 1542, 1624, 1080, 1305, 1649, 1640, 101, 1058, 705, 1099, 1650, 1561, 707, 1651, 1740, 1652, 1653, 1654, 356, 1059, 1757, 1656, 550, 706, 1459, 1244, 1538, 1657, 820, 1199, 1765, 1540, 310, 557, 833, 1567, 1606, 1100, 1566, 1848, 883]. DistributedConcurrentSmalltalk [1633]. distribution [1765]. division [1269]. do [13, 397, 79, 164, 564, 1892, 506, 165]. DO-SELECT [506]. DOC [1445]. Document [1355, 710, 291, 292, 290, 300, 1229, 287]. documentation [1688, 40, 106]. documenting [141, 301]. documents [287]. DOD [91, 1611]. DOD-STD-1838 [1611]. Does [769, 1382, 1143]. domain [454]. domains [208]. dominator [1901]. Dost [1133]. doublet [665]. down [322, 1115]. downward [420]. dpANS [1387]. Draft [1622, 616, 1623, 306, 22, 37, 76, 601, 1030, 39, 617, 608, 24, 587, 1935]. dragon [1265]. dramatic [795]. draw [1892]. Driven [1641, 411, 410, 726, 332, 1658, 1795, 1767, 490, 491]. DSM [1780]. DUIF [504]. dumb [720]. Dumont [113]. during [1124, 1821, 316]. DYMOS [618]. Dynamic [1924, 190, 1506, 794, 1509, 46, 1368, 942, 719, 618, 1197, 421, 392, 1429, 692, 675]. dynamically [1725, 1766, 1902]. dynamically-typed [1725, 1766].

each [931]. eager [1032]. early [1124]. economical [276]. ECOOP [1356]. ECSP [746]. Edison [1169]. Editing [809, 1127, 1291, 1231, 283, 536, 853, 693]. editor [668, 291, 873, 1451, 299, 534, 851, 1292, 696, 1071, 959, 285, 1821, 788, 301, 284, 287, 281, 731, 967, 693, 540, 857]. Editorial [958]. Editors [860, 481, 1137, 701, 568, 694, 496, 443]. eductive [1212]. Edward [35]. Effect [849, 1437, 1438, 1264, 532, 457, 1881, 390]. Effective [154, 1882, 1467, 662]. Effectiveness [978, 1710]. effects [80, 1731, 1385]. efficiency [146, 910, 932, 96, 1221, 1222, 328]. Efficient [1219, 741, 1389, 1900, 1920, 969, 989, 568, 860, 972, 722, 1486, 1008, 1446, 954, 1136, 1287, 1559, 1211, 742, 1699, 1766, 1908, 1732, 1498, 1448, 1459, 1574, 1112, 384, 436, 1506, 1510]. effort [91, 1358]. efforts [1347, 1406]. Eiffel [1149]. elaboration [681]. Eliminating [806, 974, 871]. Elimination [794, 1932, 1621, 692]. Elint [1678]. ELL [659]. ELSE [657, 453, 522, 635, 652, 633, 834]. ELSEs [894]. EMACS [301]. embedded [1786, 132, 1253]. Emerald [1060]. emerging [21]. EMPIRIC [996]. Empirical [391, 235]. Empty [1871]. emulate [405]. enabled [1771]. enabled-sets [1771]. Encapsulation [1056, 658, 539, 856]. Encapsulators [1086]. encoding [1699]. end [438, 140, 133]. ends [1732]. engineered [125]. Engineering [1950, 1954, 428, 1956, 1964, 1968, 1952, 1155, 1122, 747, 1117, 1166, 1592, 1607, 651, 596, 1116, 1610, 770, 706]. English [671, 478, 448, 840, 834]. Enhanced [639, 485, 251]. Enhancement [1023, 1208]. enhancements [1038]. enhancing [1716]. enough [1273]. entity [1033]. entity-based [1033]. entry [66]. entry/multi [810]. enumerated [445]. Enumerations [241, 487]. Environment [849, 435, 708, 1591, 1520, 1334, 1478, 1392, 193, 195, 1129, 1603, 1593, 537, 854, 1228, 699, 1120, 1133, 1290, 1119, 543, 861, 707, 715, 701, 532, 1769, 40, 1530, 1310, 1891, 593, 544, 862, 526, 843, 550, 706, 1605, 1596, 545, 863, 1772, 296, 704, 1594, 1122, 1136, 1633, 1301, 714, 1789, 1460, 1657, 551, 527, 844, 1794, 1595, 1121, 1533, 317, 43, 700, 547, 865, 1588, 504, 508, 173, 287, 1125, 1131, 597, 1139, 319, 703]. Environments [1954, 1958, 1857, 1956, 1964, 1968, 762, 162, 131, 822, 1609, 536, 853, 973, 592, 1590, 596, 1598, 528, 845, 1116, 1610, 1140, 1790, 716, 1134, 1606, 1138]. EPIC [980]. EPILOG [491]. EPILOG=PROLOG [490]. EPN [749, 750, 1560]. equality [1878]. equation [942, 932]. equational [303, 346, 1172, 1366]. equations [42]. equivalence [237]. equivalent [70]. ergo [1590, 1599]. Errata [9]. Error [686, 1154, 778, 1105, 413, 590, 1726, 1710, 1842, 935, 1358, 357, 792]. error-recovery [1105]. error-sensitive [590]. Errors [778, 442, 398, 415, 288]. ESM [1311]. ESN [749, 750]. especially [168]. essence [1880]. estimate [1504, 1381]. estimators [1358]. Etude [286, 300]. Euclid [875, 404, 1694]. Evaluating [1661, 1138]. Evaluation [940, 1032, 1603, 113, 235, 972, 620, 722, 879, 724, 988, 1171, 1371, 1275, 1713, 1454, 1547]. evaluator [623]. evaluators [723, 977]. Even [255]. Event [1563, 1562, 1489, 1795]. Event-based [1563, 1562]. event-driven [1795]. events [1662]. evolution [482, 704, 793, 1331]. Evolving [145, 1125]. examination [364]. example [817, 1855, 1201, 1823, 1755]. examples [1241]. Exception [904, 1565, 1535, 792, 881]. exception-based [792]. Exceptions [1024, 1033, 578, 275]. excerpts [1943]. Execution [156, 1854, 1542, 1392, 1571, 419, 1262, 1206, 1207, 825, 1487, 595, 1572, 1221, 1222, 388, 1739, 1658, 956, 393]. executions [605]. exemplar [1084]. exercise [1852, 1798]. exhaustive [426, 1509]. existing [88]. exit [907, 810, 666, 1184, 795, 882]. exotic [427]. expansion [1457, 399, 1734, 921]. Experience [718, 1069, 1126, 429, 1721, 934, 1313, 408, 412, 439, 1564, 1331, 1493, 711, 707, 1073, 1488, 1769, 1492, 130, 1098, 1270, 1970]. Experiences [1741, 772, 774, 1479, 778, 1579, 736, 309, 1450, 422, 284, 1520]. experiment [1485, 382, 410, 879, 1844, 1601]. Experimental [1246, 1006, 429, 644, 58, 1268, 683, 593]. experimentation [1237]. experiments [1745, 1330, 1307, 119]. expert [898, 953, 1415, 1508]. explanation [541, 858]. explicit [1803]. Exploiting [1496, 191]. exploratory [1505]. Explorer [1783]. express [1665]. expression [968, 1903]. expressions [589, 113, 307, 1557, 1211, 1887]. expressive [1929]. Extended [1111, 1183, 1709, 1605, 1706, 713, 1922, 1775, 1485, 1396, 1696, 1920, 1241, 1555, 1387, 1928, 1386, 1506, 957]. extendibility [1149]. Extending [810, 1293, 1772, 1802, 1796]. extensible [301, 1131]. extension [754, 755, 501, 351, 1882, 89, 1208, 787, 1912, 902, 327, 1744, 1015, 278, 420, 1819, 227]. Extensions [1875, 570, 1029, 1652, 911, 1037, 938, 935]. external [69]. Extracting [1913].

Fable [563]. Fabrik [1530]. facilitate [1257, 996]. facilities [1793, 235, 254, 767, 790, 1209, 127, 319]. facility [1565, 348, 1007, 401, 1524, 1130, 352, 420]. failed [103]. failures [1870]. Fair [1032]. fairness [1865]. false [525]. family [1783, 701]. Fast [1910, 379, 1517, 1512, 1782, 1440, 1064, 1202, 983, 903, 1818, 1361]. faster [1146]. fault [587]. favour [167]. FC [1360, 1708]. FCFS [1489]. feasibility [1384]. Feature [1163, 937]. feature-analysis [937]. features [1482, 375, 1074, 1111, 1183, 1709, 626]. featuring [829]. Feb [1971]. February [20, 354]. Felix [1682]. FFP [1514]. Fibonacci [1200]. Fibonacci-like [1200]. Fighting [631]. File [123, 339, 514]. files [1695, 807, 88]. final [320, 338]. Finding [839, 442, 398, 977, 335]. finds [400]. fine [996]. finite [1712, 261, 1469]. Firefly [1279]. Firmware [394, 395]. first [1971, 1308, 1893, 1857, 1224, 1225, 1891, 1055, 1730, 1458, 1323]. first-class [1893, 1224, 1225, 1891, 1730, 1458]. fixed [977]. fixed-point-finding [977]. Fixing [82]. fixpoint [1899]. Flamingo [1069]. Flavors [1051]. fleshy [1870]. Flexible [888, 1651, 421, 136]. FLIC [1471]. Floating [10, 7, 9, 11, 1937, 1938, 1141, 1940, 8, 1939, 99, 1935, 1700]. Floating-Point [10, 1937, 1938, 1141, 1940, 8, 1939, 99, 1935, 1700]. flow [1544, 1436, 1901, 417, 741, 1389, 1359, 553, 1297, 331, 305, 1448, 979, 1316, 751, 464, 571, 482, 1895, 1909, 431, 1449, 44, 490, 731, 613, 71, 449, 491]. flow-insensitive [1389]. flow-sensitive [1436]. flowchart [48, 504]. flowcharts [122]. fly [1738]. focus [1604]. fold [1464, 1465]. folk [1690]. folklore [286]. foreground [1838]. Foreword [754, 756, 757]. form [1908, 248]. formal [873, 151, 670, 494, 65, 781, 1383, 152, 335, 725, 439, 581, 1011]. formalism [1809, 444]. Formalization [1616]. formatting [297, 105, 503, 128]. Forms [1007, 112]. Formulations [1619]. forth [1825, 1827, 1830, 1835, 1838, 1840, 1845, 934, 1971, 1826, 1829, 1262, 1844]. FORTH-like [934]. forthcoming [1386]. Fortran [39, 232, 761, 340, 306, 1162, 68, 662, 341, 1009, 359, 277, 793, 173, 117, 22, 37, 17, 1482, 13, 39, 485, 31, 29, 23, 28, 88, 24, 32, 36, 21, 30, 1226, 35, 14, 721, 15]. Fortrans [16]. forward [755, 414]. forwarding [1323]. found [1821]. foundation [1140]. Foundational [1637]. Foundations [1037, 1588, 1029]. Four [1971, 1365]. Fourier [1361]. Fourteenth [1963]. Fourth [1692]. FP [112, 1153, 1171].

FP-systems [1171]. FPS [721]. FPS-164 [721]. fragments [542, 859]. frame [1774, 1775]. frames [1232]. framework [1779, 1592, 1631, 215, 1070, 562, 1018, 1795, 1591, 1713, 713, 1519, 1536, 1300]. Francisco [1951, 1953]. Free [678, 849, 377, 532, 686, 1167, 786]. frequencies [392]. frequency [1429]. fresh [1722]. front [438, 133]. front-end [133]. frontends [903]. FSMs [495]. Full [1868, 318, 1187, 1559]. Fully [1917, 1382, 1919, 1894]. Function [458, 1777, 1419, 1734, 455, 653, 1933, 1617]. Functional [451, 476, 1106, 511, 1379, 822, 1932, 552, 933, 672, 1471, 1362, 988, 1760, 1210, 1393, 572, 497, 1361]. functionality [27]. functionally [1506]. Functions [2, 3, 4, 99, 6, 1824, 1549, 236, 1233, 989, 923]. Further [893, 666, 869]. fusion [1679]. future [16, 1648].

Gabriel [940]. GALILEI [796]. Ganzinger [115]. garbage [1235, 1176, 1234, 1765, 368, 1554, 1693, 1764, 1686]. garden [1044]. Garp [1675]. GemStone [1303]. general [894, 1649, 283, 623, 381, 1454]. generalization [894, 880, 1802]. Generalized [589, 773, 912, 1926, 290, 1021, 1041]. generate [719]. generated [1455, 737]. generating [500, 644, 836]. Generation [727, 973, 712, 1648, 1692, 1063, 1717, 1133, 1390, 977, 1292, 728, 992, 1440, 330, 410, 1728, 879, 717, 426, 1161, 390, 1903, 384, 975, 318, 1151, 1559, 1515, 1693, 1686, 1509]. generation-based [1515, 1693, 1686]. Generational [1740]. generative [425]. Generator [954, 6, 718, 1213, 1214, 411, 429, 719, 659, 428, 1245, 909, 1732, 425, 1145, 1203, 1204, 879, 412, 1874, 427, 1112, 1420, 139, 529, 846, 698]. generators [411, 1440, 1733, 1810, 430, 737]. Generic [998, 1000, 920, 960, 1157, 688, 842, 1139]. Genericity [1093]. Generics [142, 941, 678, 744]. Georgia [1967]. Germany [1963]. get [1143, 1700]. getting [484, 15]. GKS [519]. Glanville [718, 719, 1151]. GLISP [567]. Global [1879, 995, 70, 1647, 1468, 1448, 978, 1490]. globality [326]. globalizing [976]. globally [1445]. glossary [360]. GNOME [701]. GO [761]. goal [349]. Gödel [78]. goes [1143]. Good [1769]. Gopalakrishna [71]. ’GOTO [516, 80, 1143, 1621, 447]. goto-less [447]. Goto-patches [1621]. Gough [909]. governed [1798, 1594]. Gprof [419]. Graham [718, 719, 1151]. Graham-Glanville [718, 719, 1151]. grain [1496, 996]. grammar [440, 1723, 671, 41, 256, 1819, 1168, 329, 460, 523]. Grammars [1514, 313, 1712, 424, 553, 723, 977, 575, 733, 972, 623, 724, 734, 423, 74, 976, 1289, 1909, 1477, 1115, 1724, 1454, 1547, 630]. Grammatical [1451]. GRAMOL [1507]. Granularity [1634]. Graph [1600, 435, 1289, 498, 1906, 805, 1436, 416, 419, 1609, 1752, 1362, 1722, 714, 168, 369]. graph-analysis [168]. graph-colouring [1752]. Graphical [697, 656, 668, 1817, 1318, 1675, 541, 858, 561, 1153, 1018, 1580, 1582, 1518, 1131]. graphical/textual [1675]. graphics [293, 1282, 1748, 294, 295, 1300]. graphs [648, 1890, 1435, 1821, 1907]. GraphTrace [1531]. GRASE [668, 872]. great [92, 316]. Gries [824]. group [1971, 164]. Grove [1950, 1952]. Guide [7, 9]. guided [464]. Guidelines [350, 385, 266, 373]. Gypsy [1608].

H. [1186]. HAL [439, 257]. HAL/S

[439, 257]. HAL/S/V [257]. HALF [769]. Hall [1238]. Halstead [927, 1358]. Hamming [1395, 1424]. Handling [1705, 172, 1501, 1565, 123, 1087, 455, 767, 1154, 1535, 904, 881]. Hanoi [804, 887, 765, 922, 674, 830, 799, 1167, 777]. Hardware [608, 372, 371, 386, 1258, 396, 1267, 383, 1442, 1517, 1263, 1847, 1693]. Hardware/software [372, 371]. Harmful [1815, 448, 80, 242, 834]. harmonic [580]. Harmony [1683]. Harris [1832]. Hartmann [1105]. Harvard [1009]. Hash [2, 3, 4, 6, 923]. Hashing [458, 1824]. Haven [1970]. Having [1722]. heap [441, 1510]. heart [1669, 395]. Heights [1960]. Heraklit [1653]. here [952]. Heterogeneous [205, 1562, 1260, 1573, 1657]. heuristics [1737, 1497]. hiding [101]. hierarchical [291, 52]. hierarchies [569, 191, 1298, 194]. hierarchy [1179, 1340, 1399]. High [1950, 1575, 1749, 919, 606, 1952, 611, 1825, 601, 589, 615, 573, 608, 971, 114, 104, 339, 205, 1425, 1217, 1218, 1719, 556, 1874, 1838, 938, 60, 625, 559, 747, 238, 564, 170, 420, 712, 1713, 613, 1651]. high-capacity [747]. High-Level [1950, 1952, 1575, 611, 1825, 601, 615, 573, 1874, 625, 420, 1713, 613, 1651]. high-performance [971, 1719, 556]. High-Speed [1749, 1217, 1218]. Higher [1452, 1453, 1724, 794, 1233, 1877, 692, 1918]. Higher-order [1452, 1453, 1233, 1877]. highlevel [682]. highly [980, 1461]. histories [1742]. History [1943, 1946, 685, 467, 118]. hoc [1911]. Holophrasted [43]. HOPL [1943]. horizontal [1004, 384]. host [574, 1827]. Hotel [1971]. HP [1269]. human [125]. Hurricane [1066]. hurt [397]. hybrid [808, 1315, 1221, 1222]. Hypercube [1745, 1483]. hypermedia [1070].

I/O [1747, 259, 260, 82, 837, 1287, 267, 268, 914]. I/O-computable [259, 260]. iAPX [650, 387, 371]. iAPX-32 [371]. iAPX-432 [650, 387]. IBM [763, 1268, 945, 914]. IBM-PC [914]. IC [1585, 563]. Icon [808, 1220]. iconic [1602]. identical [762]. identification [65, 115, 108]. identifiers [1003, 63, 403, 270]. identity [1094].

Ideographic [695]. IDEOSY [695]. IDL [1290, 1292, 1293, 1287, 1291, 1285, 1284, 1286]. IEEE [1141, 11, 1937, 1141, 1935]. IEEE-754 [11]. IF [657, 894, 453, 522, 635, 652, 633, 834]. IF-THEN-ELSE [657, 453, 522, 635, 652, 633, 834]. IF-THEN-ELSEs [894]. IGOR [1571]. II [604, 1643, 1799, 868, 1209]. III [1974]. ILLIAC [309]. Illinois [1210]. Illustrated [1443]. Illustration [1081]. illustrative [1241, 1111, 1183]. images [1101]. imbedded [979]. imitate [1703]. immutable [1100]. impact [537, 854, 1833]. imperative [1455, 1397]. implement [1484, 1072]. Implementation [1967, 1973, 1243, 1364, 7, 9, 860, 888, 610, 604, 1217, 1218, 1288, 831, 892, 603, 1979, 873, 1, 1443, 135, 151, 1322, 411, 742, 1624, 1577, 1499, 1396, 1063, 1766, 101, 1192, 1783, 1188, 383, 65, 300, 235, 568, 1721, 66, 339, 1074, 457, 544, 862, 1494, 650, 60, 775, 1821, 776, 1199, 1490, 1534, 1205, 1540, 284, 127, 1349, 1408, 828, 555, 1085, 665]. implementation-hiding [101]. implementation-oriented [151]. implementations [1367, 252, 911, 75, 937, 386, 1366]. implemented [233, 680, 395]. Implementing [499, 520, 987, 920, 1752, 1132, 709, 1178, 1240, 1393, 1268, 842, 1002]. implementor [245]. Implications [1198]. imported [1003]. improve [382]. improved [1711]. improvement [1863, 453, 522, 635]. Improvements [1547, 993, 758, 779]. Improving [909, 242, 657, 1368, 694]. Impulse [1068]. Impulse-86 [1068]. in- [625]. include [1802]. Incompleteness [210, 1851]. increase [826]. increased [372]. Incremental [1392, 1901, 1728, 1206, 1207, 1234, 1933, 1765, 730, 1454, 731, 1451, 1135, 468, 895, 607, 707, 972, 732, 642, 1378, 1547, 1444]. incrementally [1234]. indentation [1004].

Indenting [86]. indention [128]. independence [331]. independent [908, 1157, 1423, 1261]. index [667]. Indexed [1647, 313]. Indirect [1757, 405]. Inductive [1896]. industrial [132, 812, 1310, 306, 1592, 1509]. industrial-strength [1592, 1509]. inefficient [1426]. inference [1729, 1805, 1885, 1209]. INFIX [328]. influence [897]. informal [162, 494, 305]. Information [39, 886, 1033, 193, 1305, 195, 741, 1389, 1123, 211, 1845]. Inheritance [1884, 1771, 194, 1788, 1342, 1401, 1804, 1782, 1023, 880, 912, 1093, 1802, 1823, 1056, 1306, 1341, 1400, 1755]. Initial [1579, 1387]. initiated [1110]. Inline [1734, 921, 743, 1457]. Inmos [772]. innovation [72]. innovative [1560]. inorder [334]. input [639, 351, 919, 1698, 104, 50, 38, 963]. input-output [351]. input/output [639, 919, 1698, 104]. insecurities [452, 1182]. insensitive [741, 1389]. insertion [1]. insertion-only [1]. INSIST [1326]. instances [1190, 744, 1067, 1285]. instantiation [1506]. instruction [1556, 991, 1264, 969, 392, 391, 393, 1509]. instructions [427]. Instruments [1082]. Integer [1269]. Integral [1121]. Integral-C [1121]. Integrated [592, 188, 437, 1129, 1119, 1576, 300, 1176, 1592, 1590, 1607, 1657, 1595, 1490, 557, 1131, 597, 1521]. Integrating [1568, 195, 992, 200, 1889, 1527]. Integration [189, 549, 867, 1602, 1106, 1513]. Intel [650]. intelligence [175]. intelligent [1078, 705, 1597, 1534, 1756]. Intensional [1529]. intensive [192, 1216]. Intent [1657]. inter [811, 1165]. inter-procedural [1165]. inter-process [811]. Interacting [1595]. interaction [1602, 561, 55]. Interactive [703, 1586, 50, 595, 585, 602, 188, 1480, 614, 193, 1585, 292, 1129, 47, 300, 296, 551, 1533, 560, 1215, 43, 1300, 613, 695, 1326]. interactively [1817, 660]. interchange [740]. Interconnection [766, 897]. interest [1971]. interface [1069, 1602, 1826, 1057, 286, 1132, 242, 1799, 1785, 1611, 1068, 1520, 1518, 1676]. Interfaces [1806, 874, 1318, 1133, 1332, 1320, 672, 1136, 562, 527, 844, 560]. intergenerational [1693]. intering [1889]. Intermedia [1070, 1336]. Intermediate [638, 753, 46, 1471, 1461, 1270]. Internal [872]. International [1974, 768]. interpersonal [1633]. interpretation [1219, 1230, 1216, 1852, 1223, 948, 1859, 1211, 1887, 1754]. interpretative [66]. interpreted [1237]. interpreter [1231, 835, 1227, 150, 1212, 1188, 1236, 932, 898, 1217, 1218, 1325, 1208, 1220, 1226, 1221, 1222, 1210, 1215, 1209, 932]. interpreter/native [1221, 1222]. Interpreters [1965, 1759, 945, 685]. Interpreting [1229]. Interpretive [1965, 1228, 822]. Interprocedural [985, 984, 974, 1437, 1438, 1435, 1436, 537, 741, 854, 1389, 1910, 1486]. interprocess [746]. interrupt [914]. interrupts [1825, 1826, 820]. interview [962]. intractable [979]. introduce [274]. introducing [1252]. Introduction [943, 240, 118, 1936, 1013, 305, 404, 1052, 1286, 1635]. introductory [701, 1162]. intrusive [1480]. intuitive [647]. invariant [349]. invent [1540]. investigating [1491]. invocation [251, 172]. Io [1822]. iostream [1823]. IPC [1687]. ISBN [1613]. ISDE [703]. ISEF [1592]. ISO [473, 1700, 758, 779, 771]. ISO-Pascal [1700, 758, 779]. issue [1013, 509]. Issues [1951, 1335, 1958, 220, 1637, 928, 1156, 1667, 546, 864, 610, 604, 1610, 760, 603, 1953]. ISTAR [1119]. items [399]. Iteration [1180, 1552]. iterations [248]. iterative [765, 799]. IV [662, 309]. ively [1520]. ivory [949].

JADE [1231]. January [1963, 1966, 1972]. JANUS [292]. Jasmine [1130]. job [172]. job-control [172]. John [18, 19, 20, 34]. Jones [278]. Joyce [1745, 1697]. July [124, 1970]. junction [516]. June [1945, 1947, 1951, 1955, 1960, 126, 1943, 1356, 1953, 1967, 1944, 1958, 1957, 1946, 312, 1962, 1965, 1979]. Junek [1961].

KAREL [788]. KEE [1295]. kernel [780, 1832, 1059, 1524, 1139]. Keynote [1340, 1399]. keypunch [1821]. Keyword [963, 838]. kill [16]. kit [776]. Knowing [1165]. Knowledge [586, 587, 1593, 1145, 1074, 206, 210, 1017, 1048, 174, 225]. Knowledge-based [586, 587, 1593, 1145, 225]. Koenig [1613].

labels [516]. laboratory [1761]. Lakhotias [635]. LALR [981, 1248, 444, 1170, 256]. lambda [1871]. Lamina [1678]. Language [1942, 1951, 1967, 1973, 39, 1958, 1667, 888, 1091, 1558, 1104, 626, 1238, 1979, 1924, 906, 656, 1817, 72, 1334, 179, 796, 1213, 1214, 600, 1451, 442, 398, 884, 917, 375, 500, 411, 1775, 93, 805, 121, 1625, 1695, 1360, 1856, 574, 1640, 1626, 1864, 492, 1725, 1766, 943, 644, 670, 1142, 493, 454, 1227, 1678, 1290, 1278, 565, 829, 146, 262, 1212, 511, 671, 606, 696, 573, 1733, 1759, 1073, 91, 486, 200, 474, 463, 305, 239, 1229, 1262, 1007, 1653, 694, 1583, 1868, 1707, 672, 1074, 336, 1891, 934, 1059, 1471, 999, 1061, 898, 1675, 1237]. language [293, 1425, 735, 1223, 143, 825, 556, 87, 930, 1394, 482, 1161, 1244, 650, 243, 1644, 515, 59, 450, 60, 1178, 1187, 1122, 816, 285, 1633, 1821, 106, 767, 1253, 1220, 563, 1153, 562, 1455, 1760, 315, 1391, 559, 788, 1199, 840, 178, 760, 663, 1337, 238, 1502, 1853, 737, 1780, 560, 1205, 1540, 244, 827, 497, 676, 278, 61, 557, 833, 352, 924, 1397, 1017, 1048, 1185, 420, 1770, 294, 295, 802, 1270, 279, 1541, 1309, 62, 1953, 246, 996, 1444, 967, 1251, 1664, 369, 831, 770, 127]. language-based [1451, 1695, 696, 694, 278, 967]. language-directed [285]. language-specific [1213, 1214]. Languages [1959, 1974, 638, 766, 476, 760, 1943, 466, 794, 1963, 1966, 1972, 1019, 189, 1807, 1865, 690, 1239, 1335, 1024, 1976, 1030, 1034, 1640, 1875, 112, 1788, 927, 1782, 1427, 1257, 1390, 553, 1877, 1792, 1552, 1197, 123, 1586, 1932, 1816, 648, 951, 552, 104, 339, 622, 897, 999, 1757, 1500, 1076, 1016, 1039, 1075, 1155, 325, 673, 1494, 470, 489, 819, 1027, 1040, 333, 938, 1538, 285, 1388, 625, 309, 1426, 1461, 220, 1727, 1393, 75, 1215, 564, 1056, 716, 471, 170, 355, 56, 664, 692, 1676, 1946, 1970, 1258, 795, 369, 1811]. languages/environments/methodologies [716]. Large [1492, 1257, 1576, 1090, 1091, 1605, 1028, 1043, 436, 172, 173]. Large-scale [1492, 1576]. Law [1546, 1328, 1338, 1798, 1594, 1619]. law-based [1328, 1338]. law-governed [1798, 1594]. layered [1028, 1043]. Lazy [267, 268, 1171, 722]. LDF [1189]. learning [1491]. leave [816]. legacy [1768]. length [1699, 942, 878]. less [1005, 447, 263, 1911]. lesson [72]. Let [16]. Letter [458, 1824]. Level [1950, 1952, 794, 1825, 1575, 233, 601, 46, 589, 1570, 615, 1445, 919, 606, 573, 641, 608, 41, 253, 989, 104, 339, 610, 978, 205, 1425, 1874, 938, 60, 215, 625, 611, 901, 1221, 1222, 461, 559, 238, 564, 170, 692, 420, 1713, 613, 584, 1651]. lexical [51, 1506, 1113]. lexically [432]. Lexing [634]. Libraries [535, 852, 1704]. library [754, 755, 1823, 172]. Life [1591, 1685, 1124, 653, 955]. life-time [653]. Lifetime [1902]. like [1868, 1200, 934, 787, 627, 957, 308, 664]. likely [957]. LILA [659]. Limitations [1151, 64]. limited [1639, 66, 593]. Linda [1493]. line [297, 946, 1892]. line-oriented [946]. Linear [435, 1437, 1438, 989, 623, 1778]. linear-time [623]. Linearizable [1659].

LINGUIST [424]. LINGUIST-86 [424]. Linguistic [1459]. link [995]. Linked [1476, 642]. linker [1468]. lion [1001]. Liskov [278]. LISP [1355, 940, 1238, 432, 1396, 58, 1188, 433, 455, 524, 653, 388, 892, 1263, 1370, 1053, 1054, 84, 1236, 980, 1753, 1719, 994, 1487, 87, 1015, 1045, 580, 597]. Lisp-based [87]. Lispedit [595]. LISPTALK [1394]. List [1476, 388, 368, 916]. list-directed [388]. listing [124]. lists [923]. literate [1794]. live [1033]. liveness [1743]. LL [909, 836, 414, 1317]. Load [1835, 1872]. local [556, 1894, 109]. local-area [556]. locality [662, 1718]. localization [587]. locals [1841]. Logic [1948, 1191, 1152, 1929, 1900, 1464, 1465, 237, 543, 861, 1917, 303, 346, 1172, 114, 1858, 1175, 1076, 1016, 1039, 1075, 1656, 544, 862, 1513, 1631, 1862, 1853, 1397, 1915, 1366]. Logical [1320, 202, 218]. Logo [789]. look [982, 445, 1722, 1274]. look-ahead [982]. lookahead [399, 909, 836, 981, 1248, 1170]. loop [1466, 740, 48, 1464, 1465, 882, 666, 53]. looping [869, 801]. Loops [1815, 763, 907, 806, 989, 870, 871, 1053]. Loral [1189]. Loslan [775]. lottery [826]. loves [955]. low [1570, 901]. low-level [1570]. LR [683, 399, 982, 313, 1726, 1705, 961, 414, 983, 1813, 1470, 318, 1559, 1378]. LR-parser [683]. LR-parsers [1705]. LSE [767]. LSI [95]. LSI-11 [95]. Lucid [120]. LUSTRE [1864].

MA [1942, 1947, 1613]. MACE [1651]. Mach [1565, 1059]. Machine [908, 1261, 154, 261, 377, 720, 1720, 153, 978, 1469, 1061, 395, 1189, 993, 1753, 434, 901, 1393, 1275, 405, 1270, 1482]. Machine-independent [908, 1261]. machine-level [978]. machine-specific [993]. machines [1712, 552, 1656, 1467, 482]. Macro [1855, 254, 319, 1297]. Macro-by-example [1855]. Made [2].

Madison [1969, 1978]. Magpie [730]. Mahler [1270]. Maintained [530, 847]. Maintaining [1025]. maintenance [672, 1134]. Make [213, 1171, 1911]. making [1352, 1411]. Mammography [1845]. man [277, 630]. management [1532, 1743, 196, 1608, 1117, 910, 1609, 1596, 1610, 387, 527, 844, 1261, 1535, 1787, 54, 224, 1103, 1526, 227, 1601]. manager [680, 1099]. managers [1123]. Managing [1704, 1232]. Mandatory [1781]. Maniplicons [1321]. manipulating [99]. Manipulation [1945, 1079, 1944, 884, 999, 302]. manual [1809, 111, 889, 890, 803]. many [1922, 1273, 303, 346, 1172, 1366]. many-sorted [303, 346, 1172, 1366]. mapping [381, 1762, 892, 1497]. March [1950, 1952]. markers [935]. marking [1693]. markup [290]. Massachusetts [1974, 1968]. massively [1671]. master [1116]. match [1491]. matching [1863, 1711, 316, 1876, 1812]. Matchmaker [1059]. material [302]. materialistic [1166]. Mathematical [163]. Matroshka [1642]. matter [564]. maturity [1502]. May [39, 1969, 1978, 311]. McCabe’s [1744]. mean [580]. means [1502]. measurable [234]. measure [818, 253, 461, 1744, 1502]. measurement [94, 1472, 1746]. Measurements [762, 936, 1156, 1203, 1204]. measures [627]. Mechanical [850, 533]. Mechanism [1446, 917, 589, 1788, 1650, 1800, 380, 685, 1801, 1574, 1028, 244, 278, 792, 1686, 490, 491]. mechanisms [1537, 194, 549, 867, 1257, 457, 1517, 558, 1002]. Mediating [203]. Meeting [274, 231, 18, 19, 20, 25, 245]. MELD [1655]. MELDing [1316]. melpomenes [274]. memories [1498]. Memory [1233, 1647, 680, 1264, 1061, 1494, 653, 387, 1261, 381, 1787, 1847, 386, 1742]. Mentat [1297]. Mentor [710]. menus [1174]. merging [1053, 1054]. Mesa [374, 392, 391, 547, 865]. Message [1641, 848, 1677, 531, 775, 1658]. message-passing [775]. Messages [525, 1679]. meta [1236, 902]. meta-cleanliness [1236]. Metaclass [1791]. Metaclasses [1308, 1803]. metalanguage [366]. Method [3, 4, 498, 413, 1908, 1782, 122, 836, 781, 812, 325, 426, 141, 1785, 276]. methodologies [716]. Methodology [583, 343, 500, 1723, 582, 1350, 1409, 716, 1138]. methods [1065, 493, 1896, 1246, 839, 1158]. metric [1685, 307, 1008]. Metrics [1834, 1242, 1156, 667, 1357, 1746]. Michael [962]. micro [814, 774]. microcomputer [937]. microprocessor [396, 375, 1278, 132, 609]. microprogramming [394, 62]. middle [140]. might [397]. migratability [936]. migration [394, 1385]. MIMD [1494, 1658]. Mimic [1202]. mind [1774]. mini [814]. minicomputer [376]. Minimal [2, 3, 458, 4, 6, 923, 368, 1824]. minimization [1735, 1777]. Minimizing [1441, 1004]. minimum [1186]. minimum-period [1186]. Minnesota [1965]. MINO [670]. Miranda [1114]. misconceptions [120]. MITRE [1648]. mixed [802, 1700]. mixture [1711]. ML [999, 1914, 1882, 890, 1880, 1912]. modcap [939, 1010, 803]. Moded [1915, 1194]. MODEF [676]. model [1602, 1529, 1542, 1373, 1030, 1034, 1649, 1626, 101, 1173, 1665, 1642, 1627, 1827, 198, 201, 1609, 1919, 544, 862, 207, 152, 308, 1021, 1041, 594, 1790, 1795, 1823, 1113, 1606, 665, 1308]. model-view-controller [1795]. Modeling [1666, 211, 1780, 230]. modelled [180]. Modelling [206, 186, 1080, 199, 209, 1130, 214, 216, 1789, 219, 226, 228]. models [1544, 569, 156, 184, 185, 1562, 196, 556, 1925, 1672, 223, 269]. modernizing [1520]. modes [1273]. modest [916]. modification [618, 1303, 789, 316]. Modified [1702, 944, 1009]. ModPascal [900, 1098]. Modula [876, 1181, 1005, 960, 1003, 941, 1382, 918, 1190, 1499, 1923, 1192, 919, 920, 949, 1144, 1195, 950, 1375, 1157, 1476, 837, 95, 800, 782, 913, 748, 1154, 1148, 832, 791, 790, 1553, 452, 634, 459, 1182, 688, 842, 966, 916]. Modula-2 [876, 1181, 1005, 960, 832, 1003, 941, 1382, 918, 1190, 1499, 1192, 919, 920, 949, 1144, 1195, 950, 1375, 1157, 837, 800, 782, 913, 748, 1154, 1148, 791, 790, 1553, 452, 634, 459, 1182, 688, 842, 966]. Modula-3 [1923, 1476]. Modular [510, 678, 817, 343, 1377, 711, 1551, 1525]. modularity [710, 1839, 540, 857]. Module [766, 299, 1875, 1800, 897, 1136, 1921, 688]. Modules [1317, 874, 1157, 1634]. modulo [1169, 893, 821]. monitor [1556, 146, 1177, 791, 1567]. monitor-based [1556]. Monitoring [708, 1585, 1058]. monitors [520, 669, 394, 776]. Montréal [1955, 1957]. mostly [952]. move [414]. movement [1512]. Moving [1101]. M’Pal [840]. MS [1367, 1826]. MS-DOS [1367, 1826]. Mul [1719]. Mul-T [1719]. mulitasker [1838]. Multi [906, 1194, 825, 396, 151, 907, 1875, 835, 1832, 1122, 737]. multi-exit [907, 810]. Multi-language [906, 1122, 737]. multi-microprocessor [396]. Multi-moded [1194]. multi-module [1875]. multi-processing [151]. Multi-sequential [825]. multi-tasking [835, 1832]. multicomputer [1483]. multilanguage [1260]. Multilevel [807, 1184]. Multilingual [615, 614]. multimethods [1785]. Multiple [219, 1788, 1782, 1023, 1087, 1448, 1289, 1823, 1755]. multiple-inheritance [1023]. multiplication [1269]. multiprocess [929]. multiprocessing [1645]. Multiprocessor [1460, 1259, 1266, 1261, 1535, 1787, 1279]. multiprocessor-based [1460]. multiprocessors [1432, 1433, 1576, 1494]. Multitasking [1837, 1568, 147, 1645, 1846]. multitasking/multiprocessing [1645]. Multitasking/multiuser [1837]. multiuser [1837]. Munich [1963]. music [1533]. mutation [1226]. mutual [875]. My [955].

NAKL [1821]. Naming [347, 1757]. NARPL [1618]. narrow [1061]. National [39, 1141]. native [1221, 1222]. natural [644, 671, 470, 1912, 840]. natural-language [644]. navigator [628]. need [1382, 564, 517]. needs [564]. neither [511, 1178]. NEMO [944]. Nestable [637]. Nested [513, 107, 657, 894, 453, 1177, 522, 1147, 633, 654, 834]. Nesting [144, 253, 461]. net [1080, 1831]. net-based [1080]. Neta [894]. nets [1666, 1842]. network [200, 59, 1851, 786]. network-structured [200]. networks [1919, 556, 1489, 1866]. Neumann [552]. Neural [1774, 1831]. newly [1110]. news [1769]. Newton [315, 1758]. next [1390, 517, 1621]. Nial [1417]. nicely [944]. NIL [557, 833]. No [564, 1872, 1706]. Non [1480, 928, 1889, 724, 1146, 712, 70]. non-circular [724]. non-disruptive [712]. non-equivalent [70]. non-intering [1889]. Non-intrusive [1480]. non-recursion [1146]. Non-technical [928]. noncorrecting [1726]. nondeterminacy [1569]. nondeterminism [1171]. nonrecursion [1250]. nonrecursive [887, 674, 777]. nor [511]. notation [114, 1174, 1822, 470, 1819, 1168]. notational [215, 939]. notations [749, 750, 672]. note [1186, 874, 821, 869, 1242, 1815, 447, 487, 1169, 930, 661, 752, 523, 799, 98, 1474, 820, 1252, 1701, 1357, 310, 357, 1181, 834, 655, 689, 786]. Notes [476, 1502, 893, 518]. Notices [1013]. notion [1883, 576, 539, 856]. Nov [1949]. novel [450]. November [1968, 19, 25]. NSLR [1811, 1727]. nucleus [395]. Number [964, 1810]. Numbers [12, 78, 99, 1879]. numeric [914]. numerical [1230, 1709, 341]. NY [1960]. NYU [150].

O [1747, 82, 837, 1287, 267, 268, 914]. O-computable [259, 260]. Oaklisp [1055]. ob [79]. OBERON [1748, 1614]. Object [1959, 1975, 1060, 1355, 1976, 1640, 1035, 1628, 1830, 1977, 1310, 1094, 1531, 1543, 1596, 1778, 1051, 1684, 1610, 1092, 1814, 462, 1312, 1196, 1767, 1077, 1544, 72, 1069, 1334, 1417, 1761, 1529, 1542, 1382, 1330, 629, 1082, 1532, 1335, 1053, 1054, 1078, 1190, 1024, 1030, 1034, 1080, 1625, 1806, 1088, 1649, 1626, 1725, 1766, 658, 1788, 1665, 493, 1627, 1089, 1058, 1099, 1678, 1299, 1528, 1057, 1773, 1792, 1073, 1117, 1071, 1029, 1037, 200, 1297, 1816, 1320, 1176, 989, 1769, 1038, 1652, 1653, 1680, 1674, 1074, 1090, 1059, 1061, 1681, 1316, 1660, 1295, 1294, 1304, 1527, 1223]. object [526, 843, 1055, 1682, 1763, 1072, 1669, 1563, 1296, 1244, 1683, 1027, 1040, 1801, 1102, 1697, 1772, 1538, 1632, 1513, 1178, 1070, 1328, 1338, 1021, 1041, 1026, 1042, 1098, 1028, 1043, 1789, 1303, 1018, 1762, 1337, 1535, 1787, 1533, 1658, 1780, 1540, 1673, 1796, 1941, 1103, 1663, 1068, 1336, 1045, 1056, 1022, 1046, 1781, 1017, 1048, 1646, 1185, 1676, 1770, 1541, 1031, 1309, 1356, 1634, 1519, 1118, 1601, 1635, 1536, 1300, 1050, 1251, 1664, 1960, 1520, 1013, 1636]. Object-Based [1975, 1977, 1684, 1680, 526, 843, 1540, 1309, 1634, 1635]. object-bases [1118]. Object-Orientated [1959, 1960]. Object-Oriented [1531, 1976, 1035, 1628, 1310, 1543, 1778, 1051, 1092, 1312, 1196, 1767, 1077, 1069, 1334, 1417, 1542, 1330, 629, 1082, 1335, 1053, 1054, 1078, 1024, 1030, 1034, 1080, 1625, 1806, 1088, 1649, 1626, 1725, 1766, 658, 1089, 1058, 1299, 1057, 1773, 1792, 1073, 1071, 1029, 1037, 200, 1297, 1816, 1320, 1769, 1038, 1653, 1674, 1074, 1090, 1059, 1061, 1681, 1316, 1660, 1295, 1294, 1304, 1527, 1223, 1055, 1682, 1763, 1072, 1669, 1244, 1683, 1801, 1102, 1697, 1772, 1538, 1513, 1070, 1328, 1338, 1098, 1028, 1043, 1789, 1303, 1018, 1762, 1337, 1535, 1787, 1533, 1658, 1796, 1941, 1103, 1663]. object-oriented [1068, 1336, 1045, 1056, 1022, 1046, 1781, 1017, 1646, 1676, 1541, 1031, 1356, 1536, 1300, 1050, 1664, 1013]. object-relationship [1780]. object/action [1563]. objective [1257, 1543, 1333]. Objective-C [1333]. objectives [880]. Objects [1654, 1076, 1016, 1039, 1075, 1647, 1311, 1639, 1305, 1650, 1768, 1324, 1651, 1857, 1268, 1195, 1850, 1659, 1329, 1351, 1410, 1224, 1225, 1891, 1656, 1668, 1072, 1671, 489, 545, 863, 1631, 1302, 1287, 1315, 1672, 1758, 1902, 1148, 1518, 1181, 1100, 966, 1036]. OBJEKT [1176]. ObjVlisp [1308]. observation [605]. observations [1110, 625]. obtain [662]. obviously [165]. OCCAM [772, 1868, 515, 1548]. OCCAM-like [1868]. occur [1691]. Oct. [1976]. October [1959]. od [79]. odd [165]. odder [79]. odds [826]. offensive [578]. O’Leary [962]. On-the-fly [1738]. one [810, 735]. one-entry [810]. one-entry/multi-exit [810]. one-pass [735]. only [1032, 1, 1033, 1034, 1035, 1037, 1039, 1040, 1041, 1042, 1043, 1044, 1046, 1047, 1048, 1049, 1050, 1036]. onto [892]. OODB [1344, 1403, 1349, 1408, 1348, 1407]. OOPSLA [1959, 1976]. opaque [1545]. open [1478, 1318, 204]. operand [382]. operands [233, 234]. Operating [1974, 375, 1556, 1058, 1057, 1832, 395, 1682, 1683, 1772, 364, 1535, 1787, 1744, 1567]. operation [821, 1169, 1249]. operational [1853]. operations [1872, 1421, 1458, 1539]. Operator [65, 447, 115, 108, 1534, 893, 965]. Operators [1815]. OPG [1420]. Opportunistic [1554, 1764]. OPS5. [902]. Optimal [1466, 968, 1903]. optimisation [1512]. Optimization [456, 418, 537, 854, 974, 1464, 1465, 992, 1504, 1381, 1503, 1165]. Optimizations [727, 1431, 728, 1904]. optimized [719, 1445, 613]. optimizer [1468, 1368, 908, 1448, 978, 726, 1450, 1508]. Optimizing [763, 952, 1066, 1735, 1364, 432, 1725, 1440, 378, 1346, 1516, 1405, 980, 990, 611, 748, 1420, 915]. option [1162]. Optional [358]. OPUS [797, 1327]. ORBIT [990]. order [1233, 1877, 1928, 1452, 1453, 1918, 1724]. ordered [1547]. ordering [421]. orders [1578]. ordinary [1802]. Oregon [1979]. organization [212, 841]. Organizing [550, 1118]. Orient84 [1048, 1074, 1017]. Orient84/K [1048, 1074, 1017]. Orientated [1959, 1960]. Oriented [809, 458, 1531, 1824, 1544, 1069, 1334, 1417, 1761, 151, 1542, 1382, 1330, 629, 1082, 1335, 1053, 1054, 1078, 1024, 1976, 1030, 1034, 1080, 1625, 1806, 1088, 1649, 1640, 1626, 1725, 1766, 481, 658, 493, 1089, 1035, 1628, 1058, 1299, 1137, 1057, 1773, 1792, 1830, 1073, 682, 1071, 1029, 1037, 200, 1297, 1816, 1320, 1769, 1038, 1653, 1579, 1583, 1674, 1074, 1090, 1310, 1059, 1061, 1681, 1316, 1660, 1295, 1294, 1304, 1527, 1223, 1055, 1682, 1543, 1763, 1072, 1669, 946, 1244, 1683, 1027, 1040, 1801, 1630, 1102, 1697, 1772, 1538, 1778, 1513, 1178, 1070, 1328, 1338, 1051, 1026, 1042]. oriented [1098, 1028, 1043, 1789, 1303, 1092, 1018, 462, 1762, 1337, 1535, 1787, 1533, 1658, 1312, 496, 116, 1796, 1941, 1103, 1663, 1068, 1336, 1045, 1056, 1134, 1022, 1046, 284, 1781, 1017, 1048, 1646, 1185, 1676, 504, 1541, 443, 1031, 1356, 1519, 1196, 1767, 1077, 1536, 1300, 1050, 1664, 1814, 1013]. Orlando [1343, 1402]. Orleans [1976]. OROS [1790]. orthogonality [370]. Orwell [1526]. other [927, 1351, 1410, 457, 931]. OTM [1329]. output [639, 351, 919, 1698, 104, 38, 625, 1423]. overhead [368]. overloaded [236]. Overloading [134]. OVERSEE [1125]. Overview [507, 783, 1019, 409, 1612, 110, 486, 374, 1611, 174, 900, 1015, 1014, 1047, 1114, 1525]. Owicki [824]. OWL [565, 1052].

P [940]. P754 [1935]. PA [1954]. Pacific [1950, 1952]. Package [435, 1128, 1092, 172]. Packages [855, 538, 998, 356, 141]. packet [1173]. packet-based [1173]. packets [1765]. packing [1158]. page [509]. Paged [104, 1261]. Painless [1319]. Palo [1961, 1964, 1962]. Pancode [1247, 1820, 1560]. Panel [1354, 1413, 1350, 1353, 1409, 1412, 1352, 1411, 1341, 1400, 1348, 1407, 1332]. panes [1319]. paper [1629, 1018, 1811, 226]. papers [1958]. paradigm [1600, 1417, 1769, 1652, 1178, 1550, 1086, 275, 1796, 1049]. paradigms [1022, 1046]. Parallel [720, 1720, 1656, 1282, 482, 1296, 1446, 1969, 1489, 1978, 1970, 1854, 1647, 1872, 1478, 1260, 1585, 1173, 1627, 1495, 1569, 1578, 1573, 1576, 1498, 1579, 1680, 1484, 1719, 1492, 661, 752, 1671, 97, 1574, 434, 89, 1572, 1490, 1581, 970, 1658, 1580, 904, 1770]. Parallelism [1236, 1716, 1485, 1665, 1717, 1496, 325, 130, 1491]. parallelization [1466, 1457, 985, 1852, 1486, 986]. parameter [457, 355]. parameters [1373, 1703, 358, 1383, 85, 1201, 838, 1150, 636, 1011]. parametric [1805]. Parasight [1480, 1575]. Paris [1356]. Parkinson [67]. Parlog [1194, 1752]. parser [1726, 683, 105, 1420, 318, 1559]. parsers [313, 1728, 1705, 961, 414, 1813, 357, 1378]. parsing [982, 1246, 661, 752, 983, 1727, 1477, 634, 1115, 630, 1811]. part [1427, 1156, 734]. Partial [1578, 38, 1928, 399, 1224, 1225, 1891, 879, 1371]. Partial-record [38]. partially [1862]. partitioning [1905, 1484, 970]. parts [1071, 14]. Pascal [473, 1386, 167, 1357, 76, 77, 63, 249, 520, 1749, 351, 82, 105, 401, 89, 364, 327, 317, 677, 149, 755, 757, 519, 1475, 501, 1159, 639, 1360, 484, 574, 110, 47, 810, 1242, 1708, 645, 348, 1696, 1245, 1109, 1377, 646, 94, 1188, 682, 513, 951, 1241, 1555, 660, 823, 911, 1111, 1183, 1709, 50, 111, 514, 479, 88, 1146, 743, 407, 266, 1706, 774, 624, 126, 241, 308, 1000, 1387, 816, 782, 1240, 1551, 787, 1369, 1694, 469, 1250, 747, 841, 1557, 931, 459, 664, 636, 828, 881, 1700, 758, 779, 1756, 771, 675]. Pascal-FC [1360, 1708]. Pascal-I [47]. Pascal-like [308, 664]. Pascal-S [47]. Pascal-SC [1475]. Pascal-XT [1245]. PASES [317]. pass [438, 451, 1448, 735]. passed [1373]. Passing [848, 1931, 1677, 531, 457, 775, 1658, 355]. past [224]. patches [837, 1621]. path [589, 1384, 1583, 110, 111]. pattern [1863, 1579, 1711, 316, 1876, 1536, 1812]. pattern-oriented [1579]. patterns [1677]. Paul [1965]. PC [945, 826, 914]. PCI [820]. PCLOS [1784]. PCOL [1251]. PCTE [1612, 1117]. PDAs [495]. PDP [390]. PDP-11 [390]. PECAN [697]. pedagogic [1415]. pedagogical [764]. peek [1644]. Peep [726]. Peephole [727, 728, 726, 1508]. peeves [939]. PegaSys [541, 858]. PEN [291, 299]. penalty [1441]. Pennsylvania [1956]. perception [119]. Perfect [2, 3, 458, 4, 6, 1824, 923]. Performance [94, 1275, 1856, 1280, 1264, 1528, 1576, 971, 372, 1719, 556, 901, 747, 712, 940]. period [1186]. permutation [335]. persistence [1236]. persistent [441, 1268, 1176, 1302, 1388, 1100, 1601]. Personal [1164, 23]. personnel [1164]. perspective [816, 216, 218]. perspectives [1156]. pet [939]. Petri [1080, 1666]. Pgraphite [1601]. phase [421]. phases [1124]. Phoan [705]. phrase [935]. PHYSCAL [1109]. physical [1243, 1109, 938]. physics [314]. Pi [1088]. PIC [293]. piggy [1462, 1463]. piggy-back [1462, 1463]. PIP [172]. pipelined [969, 373, 1271]. Pipelining [1280, 1467]. Pitfalls [1108, 1613, 947, 1150]. Pittsburgh [1954, 1956]. PL [1004, 763, 350, 265, 96, 1317]. PL.8 [409]. PL/I [1004, 763, 350, 265, 96]. PL/LL [1317]. PLAIN [279]. plan [1597]. plan-based [1597]. Plasma [1643]. Plasma-II [1643]. Playground [1773].

Please [1397]. pLucid [1212]. plus [1520, 61, 1369]. Poe [696]. Point [10, 7, 9, 11, 1937, 1938, 1141, 1940, 8, 977, 1939, 99, 1935, 1700]. Pointer [27, 794, 1715, 32, 1429, 692]. Pointers [109]. Points [467, 219]. poker [1479]. policies [1515]. Poly [889, 890]. Polylith [527, 844]. Polymorphic [1914, 1881, 1871, 1730]. Polymorphism [1816, 896, 1087, 1925, 1911]. Polynomial [6, 505]. polynomials [797]. poor [277, 630]. POPL [1963, 1966, 1972, 57, 957]. Portability [1418, 60]. Portable [1749, 1551, 1687, 1741, 605, 1468, 1695, 433, 1511, 1325, 748, 1818, 664]. PORTAL [81, 54]. porting [408, 774, 931]. portions [246]. Portland [1945, 1959, 1944, 1979]. Position [1671, 226, 1629]. positions [498]. possibilities [648]. postorder [334]. potential [233, 234, 901]. potpourri [939]. Power [1815, 232, 1883, 1314]. powerdomains [1867]. powerful [961, 1801]. PPEALS [1970]. PQCC [140]. Practical [1954, 1956, 1964, 1968, 778, 982, 413, 1593, 1445, 1852, 426, 141, 1171, 57, 1121, 1755]. practice [1893]. pragmas [1620]. pragmatic [1302]. Praxis [950]. pre [824, 242]. pre-processor [824, 242]. Precision [1269, 1700]. precompiler [493, 1377]. Predicting [901]. Prediction [269]. prefab [1071]. preferred [1157]. prefix [1115, 328]. Preliminary [616, 711, 707, 601, 617, 102, 608, 1934, 134, 1909, 587, 320, 338, 247]. premature [666]. Prentice [1238]. preorder [334]. preparing [122, 287]. preprocessor [273, 1422, 173, 630]. presence [1729]. present [224]. Presentation [183]. presented [1958]. Preserved [852, 535]. Press [940]. PRESTO [1416]. prettyprinter [249, 1756]. Prettyprinting [232, 673, 296]. priced [945]. primitive [1065, 884]. primitives [169]. Principles [1963, 1966, 1972]. print [502, 446]. priority [669, 739]. priority-based [739]. PRISM [884, 917]. probabilistic [1934]. probe [1058]. probe-based [1058]. Problem [1762, 297, 1618, 894, 92, 1107, 997, 765, 1144, 1395, 1424, 1555, 1145, 85, 1691, 1706, 830, 316, 467, 666]. Problem-solution [1762]. problem-specific [1145]. Problems [1376, 329, 398, 13, 1692, 870, 563, 459, 523]. Procedural [252, 1758, 672, 1165]. Procedure [93, 1703, 1747, 1441, 251, 1698, 1383, 1201, 379, 1459, 1150, 172, 1011]. procedures [1373, 351, 810, 1359, 478, 921, 337, 327, 525]. Proceedings [1959, 1966, 1967, 1971, 1972, 1944, 1975, 1958, 1977, 1956, 1964, 1968, 1952, 1969, 1957, 1978, 1946, 1953, 1962, 1965, 1970, 1979, 1974, 1976, 1949, 246, 1942, 1945, 1947, 1951, 1954, 1955, 1961]. Process [1630, 1385, 1718, 1535, 1553, 1499, 101, 811, 463, 1375, 1839, 1598, 563, 790, 1866, 1022, 1046, 1742, 100]. process-oriented [1046]. Processes [1447, 1922, 326, 820, 1380, 1582, 786]. Processing [201, 151, 805, 720, 450, 1843, 89, 359, 1700, 117, 1750]. processor [1265, 1280, 644, 1688, 374, 824, 1753, 1492, 1064, 242, 385, 1283]. processors [373]. PROCOL [1770, 1664]. produce [41]. Producer [1333]. producers [1032]. production [428, 300, 1327, 1491]. products [1352, 1411]. Professor [962]. profile [1164]. profiler [419]. profiles [827]. profiling [1480, 1267]. PROGEN [508]. Program [1481, 809, 695, 1674, 850, 282, 443, 233, 1906, 1648, 762, 1082, 1577, 1436, 505, 1714, 640, 1363, 494, 1264, 1571, 1576, 114, 1890, 1145, 672, 1930, 542, 859, 52, 464, 1486, 1161, 242, 307, 362, 1008, 1274, 96, 651, 533, 595, 541, 858, 335, 753, 714, 1907, 1221, 1222, 697, 1744, 1701, 1739, 496, 116, 310, 827, 1372, 119, 955, 1713, 173, 281, 128, 628, 731, 1361, 770]. programme [264]. programmer [554, 1164, 628]. programmers [907, 119]. Programming [1942, 1951, 1959, 1960, 1967, 1973, 1974, 1975, 1817, 1803, 39, 1108, 493, 1958, 1109, 798, 1110, 1977, 849, 1500, 1189, 760, 1943, 827, 466, 1238, 1013, 1979, 1963, 1966, 1972, 656, 668, 189, 179, 1807, 796, 1638, 1929, 690, 1382, 500, 1478, 1239, 1260, 1335, 1053, 1054, 1976, 805, 1625, 679, 1360, 162, 1806, 1088, 1649, 1864, 195, 1725, 1875, 719, 943, 1948, 78, 658, 670, 1642, 1627, 83, 48, 454, 1628, 1227, 699, 1120, 1290, 1278, 817, 1427, 451, 1257, 1390, 146, 262, 511, 447, 543, 861, 573, 1759, 1379, 715, 701, 486, 822, 123, 1029, 1037, 200]. programming [474, 648, 532, 305, 239, 1262, 951, 1038, 1653, 1721, 1583, 933, 1680, 1530, 1858, 164, 823, 1247, 1757, 1076, 1016, 1039, 1075, 1316, 1660, 1675, 1295, 1294, 1527, 593, 1656, 544, 862, 325, 67, 825, 526, 843, 1643, 1492, 68, 1822, 1605, 1394, 1543, 1763, 1669, 482, 946, 1296, 1244, 489, 507, 1801, 1630, 243, 1662, 1778, 1513, 1631, 1149, 296, 1328, 1338, 1122, 816, 1051, 1388, 53, 947, 106, 1684, 579, 870, 1026, 1042, 1220, 1301, 528, 845, 1098, 1550, 563, 1460, 1928, 913, 309, 562, 1455, 551, 1760, 1794, 1018, 462, 1140, 1210, 840, 1121, 178, 220]. programming [663, 1727, 1505, 1853, 81, 558, 1312, 317, 560, 1941, 1353, 1412, 564, 1045, 1056, 497, 700, 676, 61, 471, 557, 833, 547, 865, 127, 355, 56, 1397, 1526, 1017, 1048, 1185, 1676, 508, 802, 279, 1356, 597, 1946, 1970, 246, 1196, 996, 1258, 1509, 1915, 1077, 1754, 566, 449, 369, 529, 846, 1814, 1811, 768, 1953]. programming-language [563, 560]. Programs [1179, 1446, 852, 872, 1854, 1878, 1568, 63, 442, 1219, 1126, 1230, 763, 1585, 101, 1216, 144, 1127, 1384, 454, 1089, 1242, 1495, 1900, 1464, 1465, 1569, 711, 1917, 23, 910, 263, 1890, 1889, 660, 1734, 1674, 105, 1362, 1206, 1207, 350, 1753, 143, 265, 1904, 1487, 349, 1563, 988, 709, 141, 1365, 899, 624, 1429, 1574, 1849, 1028, 1043, 1572, 1913, 480, 1483, 1604, 469, 219, 1581, 436, 970, 1357, 367, 1580, 44, 612, 602, 535, 1566, 883, 613, 71]. progress [756, 757, 829, 1874]. project [1618, 702, 1119, 696, 103, 1596, 1116, 1358]. Prolog [1108, 161, 1175, 947, 524, 1001, 460, 490, 491, 1219, 925, 1367, 1276, 926, 1254, 1396, 806, 869, 1073, 1379, 1163, 932, 1395, 1424, 888, 766, 1217, 1218, 1691, 899, 870, 1474, 871, 1415, 1193, 580, 1819, 965, 937, 1462, 1463, 1754]. Prolog-based [1073]. Promises [1459]. prompts [1893]. pronoun [787]. proof [1851]. Proofs [651, 494, 1913, 581, 786]. Propagation [848, 984, 531, 1539]. proper [1882]. properties [234, 1746]. property [381]. PROPHET [258]. proposal [76, 906, 1368, 475, 516, 787, 881, 1700, 1002]. Proposals [10, 32]. Proposed [7, 9, 11, 891, 8, 521, 1611, 789, 902, 1935]. proprietary [171]. Protecting [1181, 966]. Protection [171, 380, 386, 996]. proto [772]. proto-OCCAM [772]. protocol [723, 1251, 1664]. protocol-constrained [1251, 1664]. protocols [1770]. prototype [873, 650, 700, 1756]. prototypes [1797, 1766, 655]. prototypical [1072]. Prototyping [1786, 1116, 1622, 1128, 1623, 1590, 725, 1776, 738, 1332]. proving [255]. provision [672]. PS [441, 1423]. PS-Algol [1423, 441]. PSAIL [1511]. pseudo [664, 172]. pseudo-code [664]. pseudo-library [172]. PSG [529, 846]. PSI [1275]. PSM-E [1491]. purpose [249]. purposes [107]. Puzzles [1175, 1254].

QL [891]. QL/Bn [891]. Qlisp [1488]. Quality [1688, 1368, 1504, 1381, 1503]. quasi [89]. quasi-parallel [89]. Qué [1955]. Queens [1144, 997, 1107, 1010]. questions [794]. queueing [1489]. QUICKTALK [1065].

radar [1311]. railway [467]. ramifications [882]. Random [1810]. range [418].

RAPID [1332]. Rapidly [1125]. rate [480, 1358]. Rated [1744]. rationale [1751, 1472, 127]. re [1206, 1207]. re-execution [1206, 1207]. reactive [1921]. READ [327]. readability [657, 251, 328, 956]. Readable [1808]. Reading [1613]. Real [1432, 1433, 1843, 81, 1786, 1230, 1499, 1864, 835, 1832, 1834, 609, 1868, 1091, 1681, 306, 145, 243, 1684, 1844, 1199, 1567, 1566]. Real-time [1432, 1433, 81, 1786, 1499, 1864, 835, 1832, 1834, 609, 1868, 1681, 306, 243, 1684, 1844, 1199, 1567]. Realistic [1930, 1874, 1455]. realization [277]. really [679]. Reasonably [945]. Reasoning [1731, 45, 194, 1920, 1896, 206]. Rebus [808]. receiving [337]. reclamation [1740, 712, 1515]. recognition [1536]. recompilation [974, 1136]. reconsidered [506]. reconstruction [1730]. record [38, 1158]. records [1875, 684, 1912, 642]. recovery [1, 1105, 413, 1726, 935, 357, 792]. recurring [467]. Recursion [1146, 1250, 956, 875, 1200, 73]. Recursive [1812, 1470, 277, 1359, 512, 989, 724, 1220, 357]. RED [344]. redisplay [298]. redoing [713]. reduce [683, 1813]. reduced [1740]. reduced-communication [1740]. reduction [1004, 1722, 1393]. reductionistic [83]. redundant [63, 1879]. referee [957]. reference [250, 1740, 124, 1809, 523, 1718, 803, 329]. references [1429, 1693]. Refined [1461]. Reflection [1541, 1792, 1307]. Reflective [1793, 1626, 1635]. reflectively [164]. regarding [113]. region [824]. Register [416, 739, 377, 1736, 994, 1439, 1737, 1441, 995, 70]. registers [1743, 1414]. Regular [1712, 1427, 734, 307, 1557]. Reid [18, 19, 20, 35, 34]. rejoinder [1250]. Relating [1925]. Relation [1636]. Relational [640, 708, 1106, 1528, 208, 709, 1365, 507, 1336]. relational/functional [1106]. Relations [536, 853, 1337, 1194, 545, 863, 1539]. Relationship [1292, 1641, 1598, 1780]. Relationships [185, 647, 234]. relative [1870]. Relativism [207]. relaxing [75]. RELFUN [1106]. reliability [13]. Remarks [1172, 463, 1170, 1366, 1501, 1110, 684, 1248]. remote [1840, 525]. removing [64]. rendezvous [146]. reorganizer [991]. Repair [778, 415]. repeat [764, 270]. repeatable [358]. repetition [465]. replace [341]. replication [1848]. reply [1357, 1566]. Report [1104, 768, 1622, 135, 1255, 829, 1623, 1934, 1874, 1610, 833, 279]. reporting [1845]. Reports [34, 18, 19, 20]. repository [1100]. represent [1383]. representation [872, 258, 138, 1074, 1362, 1167, 1582, 1017, 174]. representation-free [1167]. representations [753]. representing [1890]. reprogramming [271]. Requirements [588, 1622, 492, 1623, 389, 585]. Requirements/design [588]. Research [1941, 175, 177, 179, 1807, 1651, 813, 897, 1681, 216, 1550]. reserved [923]. Resetting [877]. resilient [555]. resolution [135, 939]. Resolving [1909, 1477]. Resource [54]. respect [1366]. Response [1248, 949, 1191, 876]. responsibility [1767]. responsibility-driven [1767]. Responsive [1380]. restructurers [1376]. Restructuring [1487, 1486, 1198]. result [237]. results [683, 1503, 1450]. retargetable [991, 330, 980, 427, 1508]. retargeting [814]. Retrieval [888]. Retrofitting [1324]. return [1617]. RETURNs [785]. reusability [745, 1149, 626]. reverse [1572]. reversible [1571]. Review [1808, 1613, 940, 1238, 1784]. Revised [1104, 279, 256]. Revision [39, 22, 1611]. revisions [1558]. revisited [1367, 1751, 1247, 674, 1148]. Rewrite [1927]. rewriting [1907]. REXX [1142]. Richard [940]. right [1427, 734]. right-part [734]. Ripken [115]. RISC [1272, 1231, 1227, 1277]. RISCs [1276]. robot [788]. role [1545, 596, 1391]. Rosette [1646]. Rothon [643]. router [1490]. routines [743, 1252]. RTX [1832, 1843, 1846]. Rule [1797, 1253, 1711, 1168, 546, 864]. Rule-based [1797, 546, 864]. rules [1773, 902, 1472, 1679]. Run [681, 502, 1038, 973, 665]. Run-time [681, 502, 1038, 973, 665]. runtime [1496, 138]. RUSSELL [987].

S [523, 257, 47, 439, 257]. S8 [39]. SAC [795]. Sacrificing [1892]. sad [1501]. safe [1625]. Safety [1431, 1842, 883]. SAGA [702]. sage [518]. SAIL [1511]. salesman [955]. Salvadori [113]. San [1951, 1966, 1953, 1977]. Sans [27]. Sarek [1676]. SARTEX. [805]. SASL [783]. Saving [1768]. scalar [1271]. scale [1576, 1492]. scales [1746]. Scanner [954, 1112]. Scannerless [1727, 1811]. Scavenging [712]. Scheduling [1872, 396, 1281, 969, 1375, 1467, 970, 1553]. schema [225]. scheme [1105, 1058, 1740, 824, 1055, 108, 1693, 1104, 990, 1449]. schemes [505, 910, 1802, 1540]. scheming [1160]. schlonked [484]. science [1162, 1187, 365, 1358, 518]. Scientific [1700, 1854, 1638, 1090, 1385, 721, 1776]. scope [1003, 1008]. scoped [432]. scopes [513]. scoping [570]. Scratchpad [1209]. Screen [682]. Screen-oriented [682]. SDF [1809]. search [426, 1372, 1509]. Seasons [1971]. Seattle [1958]. security [1781]. seeding [1710]. SELECT [506, 635]. Selected [456]. selection [236, 1509]. Selective [1216]. self [301, 1725, 1766, 1314]. self-documenting [301]. SEMANOL [151]. Semantic [415, 1852, 1456, 569, 534, 851, 735, 1177, 1422, 439, 1252, 1337, 136, 738]. Semantics [1867, 429, 93, 1714, 1344, 1403, 1363, 1666, 1804, 1917, 425, 1868, 1874, 1894, 1895, 1907, 725, 1374, 1430, 430, 975, 1853, 431, 1866]. semantics-directed [425]. Semaphores [780, 499]. semi [1113]. semi-syntactic [1113]. sense [1543]. sensitive [1436, 590, 503, 1497]. sentences [478, 41]. separate [401, 745, 1240, 436]. separately [874, 622]. separators [1736]. September [1959, 1977]. sequence [1703]. sequential [1922, 825, 1380]. Serial [97]. series [58, 1294, 1211]. seriously [1629]. Session [583, 599, 588, 582, 598, 586, 603]. set [1264, 773, 981, 1590, 392, 1611, 687, 391, 393]. SETL [921]. sets [836, 1248, 1873, 1170, 406, 1771, 70]. Seven [367]. shared [1494, 1072, 1662, 1302, 1421]. shared-memory [1494]. Sharing [744, 886, 812]. SHD [682]. shell [1299, 84, 1415, 572]. Shift [1813]. Shift-reduce [1813]. short [404, 106]. Should [496, 180, 1500, 443]. Side [849, 1437, 1438, 532]. side-effect [1437, 1438]. SIGACT [1963]. SIGACT-SIGPLAN [1963]. SIGForth [1971]. signal [1843]. Significant [13]. SIGOA [1945, 1944]. SIGOPS [1969, 1978]. SIGPLAN [1942, 1945, 1947, 1951, 1954, 1955, 1961, 231, 1953, 1950, 1963, 1967, 1973, 1944, 1975, 1958, 1977, 1956, 1964, 1968, 1952, 1969, 760, 1943, 1957, 1013, 1978, 1946, 1962, 1965, 1970, 1979, 246]. SIGPLAN/SIGOA [1944]. SIGSOFT [1954, 231, 1950, 1956, 1964, 1968, 1952]. SIGSOFT/SIGPLAN [1954, 1950, 1964, 1956, 1968, 1952]. SIGSPACE [231]. SIGSPACE/SIGSOFT/SIGPLAN [231]. SILs [928]. similar [1922]. similarity [762]. SIMOB [605]. Simple [2, 837, 954, 873, 917, 101, 818, 511, 478, 1087, 470, 545, 863, 363, 1112, 89, 1165, 328, 924, 1686]. simpler [765]. simplest [511, 497]. Simplicity [328, 690, 1314, 1892]. simplified [819, 108]. simplify [627]. Simplifying [64, 787].

Simula [605]. Simulating [1417]. simulation [1330, 1128, 1773, 1489, 1326]. simulator [1203, 1204, 1202, 788]. SINA [1537]. SINA/ST [1537]. single [438, 1908, 699]. single-pass [438]. sixteenth [1972]. size [1264]. skeletons [559]. sketch [996]. Skinny [1870]. Slade [1238]. SLCSE [1591]. SLD [1690]. slicing [1435]. small [898, 362, 1694, 903, 693]. small-X [898]. smallest [1274]. Smalltalk [1097, 1884, 1347, 1325, 839, 1079, 1319, 1066, 1065, 1786, 1322, 1639, 1803, 1272, 1063, 1666, 493, 1333, 1099, 1406, 1522, 1793, 1800, 1110, 1083, 1516, 1084, 1524, 1327, 1064, 1232, 1460, 1086, 1794, 1534, 1062, 1505, 1312, 1795, 1523, 580, 1645, 1101, 1776, 1521, 1525, 1085, 1331, 1326]. smalltalk-80 [1884, 1065, 1639, 1803, 1272, 1063, 1666, 1333, 1099, 1793, 1110, 1064, 1086, 1523, 1521]. smart [720]. SN [1750]. SN-2 [1750]. SNOBOL4 [477, 808, 316]. SNOBOL4/Icon [808]. SOAR [1062, 1491]. Soar/PSM [1491]. Soar/PSM-E [1491]. sociobiology [579]. SODOS [40]. soft [835]. Software [1950, 1951, 1954, 703, 667, 813, 1956, 1964, 1968, 814, 1952, 1467, 760, 14, 708, 1746, 437, 1602, 323, 680, 702, 1267, 1135, 1593, 1128, 537, 854, 1257, 797, 818, 1156, 647, 148, 641, 1117, 807, 812, 1769, 372, 40, 1597, 622, 360, 389, 979, 1418, 1166, 73, 1592, 1685, 42, 1123, 550, 1605, 1596, 1244, 545, 863, 1607, 1130, 651, 1187, 704, 1594, 1124, 1136, 596, 1551, 1598, 714, 1086, 314, 1116, 1610, 364, 371, 1790, 365, 1358, 223, 1263, 1591, 171, 1606, 1125, 1131, 1953, 460, 271, 770, 1615, 1139, 706, 1591]. software-evolution [704]. Solution [830, 657, 1618, 997, 887, 765, 870, 563, 1167, 1762, 316, 459]. Solutions [1107]. Solving [1144]. Some [1501, 120, 1330, 1108, 113, 199, 876, 684, 389, 243, 266, 800, 625, 106, 1386, 794, 882, 758, 779, 905, 1246, 104, 339, 935, 35, 635, 832]. sorted [303, 346, 1172, 1366]. sorting [688]. sorts [367]. SOS [1556]. sound [1533]. Source [149, 63, 1445, 610, 774]. source-level [1445]. Source-to-source [149]. space [1004, 1549, 1296, 1538, 1221, 1222]. space-time [1221, 1222]. special [1971, 1013]. Specialization [880]. specific [1213, 1214, 1145, 993, 422]. Specification [1849, 555, 1600, 873, 190, 1239, 411, 1355, 197, 1035, 645, 1667, 65, 1339, 1398, 73, 735, 514, 98, 430, 222, 170, 460]. Specifications [163, 644, 252, 781, 591, 455, 1855, 276, 1713, 221]. specified [1862]. Specifying [1668, 1043, 463, 1174, 1028]. speech [1351, 1410]. Speed [1749, 1217, 1218, 1838, 1474, 956]. Spelling [925, 288]. Sperry [58]. Spill [1735]. spilling [1872, 416]. SPLOS [1974]. SPLOS-III [1974]. Spreadsheet [1425, 1685, 1092]. spreadsheets [1092]. SPUR [994]. SQURL [240]. Srinivasan [71]. ST [1537]. St. [1965]. stack [377, 1232, 1616]. stacks [49]. stage [687]. stand [1523]. stand-alone [1523]. Standard [39, 10, 7, 9, 11, 1937, 1938, 1141, 1940, 8, 22, 577, 521, 1939, 1362, 306, 28, 890, 21, 1386, 366, 1935, 473]. standardisation [77]. standardization [1347, 1406]. standards [68, 1558, 1391, 771]. STAR [1615]. StarLite [1128]. state [1712, 261, 1469]. statement [894, 80, 1671, 516, 1423, 882]. statements [1500, 242, 53, 986]. Static [746, 1570, 988, 1805, 612, 570, 1908, 899, 1746]. statically [1924, 1625, 1788]. statically-typed [1924]. station [1840]. statistical [1753]. status [433, 1284, 833]. STD [1611, 1141]. Steelman [247]. step [584]. step-wise [584]. Stephen [1238]. steps [1323, 1796]. still [564]. stochastic [1489]. stock [1432, 1433, 1517, 1693]. storage [442, 398, 46, 1431, 1740, 910, 1787, 712, 1515, 1510]. store [1176]. Stores [1891]. story [274]. Strategies [1824, 1776, 590, 1288, 365]. strategy [1187, 436]. Stream [1636]. streams [1860, 172]. strength [1592, 1509]. strict [167]. Strictness [1859, 1888, 1860, 1861]. String [767, 1501, 884, 501, 1699, 450, 363, 166]. String-handling [767]. strings [399, 878]. Strong [938, 1032, 1428]. Strongly [724, 1806, 1197]. strongly-typed [1806]. struck [518]. Structural [1883, 73, 1134]. structure [657, 234, 1060, 481, 710, 797, 1137, 1292, 701, 421, 522, 1434, 1027, 1040, 394, 1028, 1043, 368, 654, 1134, 284, 548, 866, 517, 1118, 628]. structure-oriented [1137, 1134, 284]. Structured [540, 857, 83, 860, 265, 691, 277, 784, 449, 668, 483, 78, 273, 48, 340, 122, 200, 568, 1247, 1702, 285, 518, 731, 795]. structuredness [44, 71]. Structures [53, 259, 260, 1642, 512, 283, 5, 453, 1383, 339, 880, 215, 1000, 1862, 1147, 1291, 891, 627, 90, 1101, 842, 642]. structuring [349, 255]. student [1618]. Study [1808, 1447, 233, 1276, 1639, 1088, 1839, 1124, 1460, 1158, 1336, 1370, 62, 1271, 393, 1196]. Style [157, 624, 1931, 718, 1186, 483, 719, 446, 1543, 266, 1550, 1151]. styles [673]. Sub [723]. Sub-protocol-evaluators [723]. subclassing [1096, 1097]. subject [316]. subjects [1351, 1410]. sublanguages [107]. subprograms [347]. subrange [488]. subset [495, 1009, 257]. substitution [92, 316]. substrate [1068]. substring [1726]. subtype [1034]. Subtypes [488, 1030, 1885]. subtyping [1883]. succeeded [103]. suggested [127]. suitability [1161]. suitable [519, 1748]. summarizing [1716]. Summary [37, 583, 599, 620, 1587, 36, 431, 245, 588, 1436, 741, 1389, 582, 598, 1930, 586, 1018, 603, 187]. Sun [1528]. super [814]. super-mini [814]. supercomputers [1271]. superimposition [1898]. Supernode [1905]. Superoptimizer [1274]. supplemental [753]. supplying [166]. Support [1974, 1197, 1591, 396, 437, 1602, 375, 1382, 1260, 162, 1267, 1603, 1133, 1278, 1119, 131, 1293, 40, 1091, 389, 1059, 1304, 1685, 1524, 1590, 1459, 1596, 394, 1772, 1873, 527, 844, 384, 626, 287, 1847, 386, 540, 857]. Supporting [1572, 387]. supports [917, 1597, 1178]. surprise [674]. Survey [466, 813, 471]. Sussman [1160]. SW [526, 843]. Swamp [1064]. swarms [579]. SWAT [615]. Swing [261]. symbiosis [1639]. symbiotic [536, 853]. symbol [258]. symbolic [614, 610, 1277]. symmetric [1236]. Symposium [1942, 1945, 1947, 1950, 1951, 1954, 1955, 1961, 1963, 1958, 1956, 1964, 1968, 1952, 476, 760, 768, 1962, 1965, 1972, 1944, 1957, 1953]. synchronizable [1211]. Synchronization [326, 1570, 1495, 1650, 463, 244, 1771]. Synchronous [1458]. syntactic [72, 413, 342, 1855, 332, 938, 935, 366, 1113, 792]. Syntactical [778]. Syntax [473, 561, 475, 668, 72, 438, 1451, 690, 322, 1231, 250, 1726, 659, 251, 102, 324, 124, 1809, 166, 1452, 1453, 788, 1252, 801, 693, 283]. syntax-directed [668, 438, 788, 283]. syntaxes [1161]. Synthesis [1922, 1082, 705, 1921, 1533]. Synthesizer [698, 282]. System [529, 846, 888, 913, 1079, 188, 1482, 1069, 1638, 437, 190, 1622, 258, 1786, 375, 629, 1532, 1060, 1556, 1589, 1318, 1585, 702, 1305, 1923, 1396, 195, 292, 1593, 618, 1058, 705, 1099, 1650, 1299, 1256, 1416, 1522, 711, 1057, 424, 1571, 1773, 58, 707, 715, 1623, 1117, 1832, 1297, 300, 204, 660, 1674, 897, 979, 395, 953, 1295, 1294, 1304, 1527, 1132, 1327, 1590, 1682, 1683, 332, 212, 1772, 1130, 1302, 1070, 775, 541, 858, 1633, 1599, 653, 309, 1604, 1044, 737, 1415, 892, 1845, 1866, 557, 1491, 1526, 1567, 1676, 738, 1508, 1521, 905, 1750, 1615]. System [1355, 546, 864, 695, 1202, 1081]. System/370 [1202]. Systematic [1825, 1004, 380, 607]. Systems [1951, 1959, 1974, 39, 1531, 940, 1936, 1922, 1213, 1214, 1311, 601, 1562, 1478, 1260, 1078, 1033, 1976, 1080, 574, 492, 640, 1129, 1257, 1667, 1573, 822, 1834, 1653, 929, 1091, 1654, 356, 1059, 999, 1681, 1660, 898, 1837, 1072, 1459, 946, 1881, 145, 650, 1798, 1594, 1253, 1171, 697, 1199, 1461, 1535, 1787, 1673, 224, 1336, 225, 1591, 833, 1918, 1781, 1048, 1646, 228, 1031, 1634, 1847, 1953, 1970, 1851, 1848, 1915, 1536, 760]. systolic [1481, 1577, 971, 1836, 1904].

T [1238, 1719]. table [258, 411, 410, 1702]. Tables [448, 683, 66, 691, 381, 834]. TAE [1520]. Tags [1263]. tail [1722]. Tailoring [422]. take [816]. Taking [1629]. TAO [580]. Task [1935, 762]. Tasking [158, 85, 1473, 1107, 835, 1832, 235, 152, 127]. tasks [1329]. TCOL [140]. Teaching [1295, 1761, 907, 1360, 880, 479, 1763, 816, 107, 924, 518]. team [1526, 1603]. tears [486]. Technical [17, 1012, 3, 4, 928, 287]. technique [1, 1716, 442, 398, 982, 1216, 609, 1087, 73, 1467, 662, 935, 903, 931]. Techniques [694, 1965, 186, 1735, 678, 1498, 736, 408, 1535, 1787, 1658, 355, 1713, 1271]. technology [1725, 1292, 1799, 611]. teledebugging [1564]. tell [397]. Temporal [1929, 237, 1920, 1899]. Tenuring [1515]. termination [1699]. test [500, 1504, 1381, 406]. testbed [1651, 1224, 1225]. Testing [49, 590, 1603, 348, 131, 591, 1934, 52, 464, 422]. Text [1945, 1944, 481, 297, 837, 302, 496, 284, 288, 443]. Text-oriented [481]. textual [1675, 90]. their [959, 5, 724, 1855, 1461, 1739, 75, 939]. them [13, 255, 564]. Theoretical [234]. Theory [855, 538, 927, 1893, 402, 274, 362, 1098, 1903, 1744, 1746]. there [1143, 146, 769]. ThingLab [1799]. Things [1841]. ThinkerToy [1321]. thinking [1761, 1830, 1165]. Third [1974, 1063]. Thoughts [1241, 905]. Threaded [822, 405]. Three [1796, 1824, 492, 1257, 263, 165, 1753, 215]. three-level [215]. three-processor [1753]. TICLOS [1783]. TIGRE [1722]. Time [1549, 6, 1432, 1433, 1786, 1499, 1856, 1864, 1437, 1438, 835, 681, 502, 1832, 1498, 1834, 1038, 609, 1835, 1868, 1091, 623, 1681, 973, 1294, 306, 145, 243, 1843, 1886, 1684, 653, 1844, 1221, 1222, 1199, 970, 81, 1567, 1888, 995, 1566, 665]. timeline [1071]. times [1739]. timeshared [715]. timestamp [1765]. Timing [1503]. tips [1473]. tiring [1509]. TM [772]. Token [809]. Token-Oriented [809]. tokens [1705]. Tompkins [1186, 631, 632]. too [961, 15]. tool [1485, 1333, 699, 148, 622, 479, 527, 844]. toolbox [605]. toolkit [1518]. tools [1568, 1602, 1345, 1404, 641, 1590, 472, 222, 1223]. toolset [1607]. top [322]. top-down [322]. Topaz [1564]. Topics [456]. topological [263]. tower [949, 830]. Towers [887, 765, 922, 674, 1167, 777]. trace [1281, 1919, 1851]. trace-based [1851]. traced [1897]. Tracing [578]. tradeoffs [1549, 1278, 372]. transaction [559]. Transactions [1660, 1661, 559, 1663]. ’transfer [516, 784]. ’transfer/junction [516]. transform [1828, 1609, 1361]. Transformation [1544, 1673, 736, 1930, 976, 369]. Transformational [1604, 1754, 1540]. Transformations [975, 1716, 1464, 1465, 1855, 332]. Transition [159, 148, 1866]. translating [1333, 147]. translation [149, 1477, 369]. Translator [1749, 150, 424, 472, 1694, 1168]. transmit [1383]. Transparent [1323]. transportability [936]. transportable [1377, 1520]. trapping [1842]. Traps [1613]. traversable [49]. TRC [953]. treatment [335]. Treaty [1343, 1402]. Tree [736, 1863, 1697, 830, 627]. tree-like [627]. trees [968, 98, 1903]. Trellis [1301, 1052]. Trellis/Owl [1052]. Trenches [950, 918]. trouble [875]. troubles [1473]. truly [425]. truths [397]. TS [1516]. tuple [1538]. Turbo [1163]. Turing [1228, 1369]. Tutorial [175, 177, 179, 1511, 1286]. tutoring [1078]. Two [1672, 1221, 1222, 119, 842, 678, 451, 576, 41]. Two-level [1221, 1222, 41]. two-pass [451]. TX [1971, 1972]. Type [569, 72, 135, 1729, 648, 1083, 1157, 745, 1730, 1912, 1885, 1625, 193, 194, 1883, 1923, 644, 358, 1428, 402, 576, 1816, 1882, 514, 363, 1805, 1098, 276, 1461, 1790, 167, 27, 1263, 352, 1209, 1317, 905, 1915]. type-assignment [1882]. Type-checking [1083]. type-safe [1625]. Typed [933, 1924, 1806, 1725, 1766, 1788, 1650, 1197, 1886, 1601]. Types [181, 344, 230, 258, 678, 1545, 920, 1180, 252, 1896, 960, 781, 445, 202, 138, 1812, 880, 208, 1055, 759, 274, 213, 1871, 1926, 1421, 217, 1461, 663, 223, 1205, 69, 1103, 228, 229, 555, 655, 1025, 675, 665]. typesetting [293, 294, 295]. typing [1924, 926, 1914, 938, 308]. typographical [288].

Ubik [1679]. UCSD [520, 1000]. undecidable [457]. Underflow [12]. understandability [114]. Understanding [1531]. Unfair [1375]. Unfold [1464, 1465]. Unfold/fold [1464, 1465]. UNH [932]. unicorn [1001]. unification [1914]. Unified [1743]. uniform [1627, 198]. Unifying [981, 1223, 1606]. UNIMAC [51]. Uninitialized [1190, 1148, 1195, 1181, 966]. uniprocessor [1261]. units [1243, 1109, 938]. Univac [58]. Universal [965, 228, 619]. University [1969]. unIX [571, 465, 1687, 139]. unnecessary [974]. unquote [447]. unquote-operator [447]. until [270]. unwanted [806]. unwarranted [832]. Up-down [1115]. update [1901]. US&R [713]. USA [1942, 1945, 1947, 1951, 1954, 1959, 1960, 1961, 1950]. usage [1441, 1124, 653, 393]. Use [160, 1294, 212, 188, 1716, 343, 193, 960, 622, 1785, 1461, 627, 581, 795, 369]. used [274, 1841, 405, 955]. useful [1819, 1011]. User [1602, 309, 1069, 1318, 1133, 286, 1132, 111, 394, 1799, 166, 562, 55, 1520, 1518]. user-defined [166]. user-interface [1799]. users [203]. Using [1311, 719, 132, 811, 478, 1383, 1484, 1531, 1685, 1072, 826, 1538, 238, 1747, 1562, 942, 1743, 1107, 1528, 772, 447, 1698, 1769, 1435, 1777, 999, 423, 1503, 830, 392, 725, 1604, 1154, 1539, 1535, 1787, 1010, 827, 172, 1270, 1286, 642, 1746]. USSR [1829]. utilization [1498]. Utopia84 [304]. UW [1443].

V [257]. validation [129]. Value [337, 1387, 1879, 1249, 1617]. value-yielding [1249]. valued [1106]. Values [489, 1869, 502, 578, 972, 1730, 1458]. Vamp [1779]. Vanilla [1687]. variable [1496, 878, 172]. Variables [794, 1878, 621, 138, 1715, 1894, 109, 118, 692, 675]. variance [1739]. variant [125]. variants [684, 1912]. various [673, 1503]. varying [358]. VAX [614, 1280, 1122, 139, 393, 774]. VAX-11 [393]. VAXELN [743]. vector [1549, 262, 930]. vectorization [1457, 1906]. verifiability [247]. verifiable [257]. Verification [925, 1849, 142]. verifying [141]. versatility [161]. Version [1608, 465, 1909, 1472]. versions [1889]. versus [488, 1146, 1093, 1370, 128]. vertical [394]. Very [983, 436, 1818]. VHDL [1364, 1723]. via [1368, 1585, 1901, 416, 1571, 1736, 611, 1933]. viable [1162]. video [299]. View [950, 918, 896, 202, 981, 1166, 219, 1795]. Viewing [699]. viewpoint [922, 178]. Views [1531, 1293, 203, 207, 709, 1796, 1580, 1876]. Virtual [1061, 1801, 1067, 1787, 1847, 1266, 153, 1296, 1261, 405, 1082]. visibility [742]. Visionaries [962]. Visual [1239, 1390, 551, 1128, 1530, 1353, 1412, 90]. visualization [1674]. visualize [648]. VLIW [1281, 1467]. VLSI [1259]. Voyeur [1580]. vs [918, 1276, 949, 950, 1662, 1250, 525, 1439]. vu [1193].

W [1963]. walk [649]. walking [565]. want [769]. Washington [1958]. way [1699, 648, 1876]. weather [1840]. well [977]. well-defined [977]. Wesley [1613]. Western [261]. Wetherell [523]. Where [915, 1892]. wherefore [282]. which [1640, 103]. while [1701, 764]. Who [564]. whole [1127]. WI [1978]. window [1522, 1785, 1676]. windows [1319, 1439]. winning [826]. Wisconsin [1969, 1969]. wise [584]. wish [916]. Within [852, 596, 495, 535]. without [486, 1524, 326, 1867, 1062, 1847]. word [923]. words [1828]. Work [878, 756, 757]. Working [1044]. Workshop [1960, 1975, 187, 1255, 1344, 1403, 1345, 1404, 1977, 1346, 1405, 1339, 1398, 1587, 1969, 1349, 1408, 1356, 1978, 1971, 1593, 1610, 1013]. workstation [1123, 706, 1124, 1279]. world [788]. world-builder [788]. Worlds [1118]. write [769, 957, 327]. writers [289]. writing [424, 1733, 798, 660, 350, 472, 1808].

X [898]. X3.160 [1387]. X3.160-198x [1387]. X3.9 [39, 39]. X3.9-1978 [39]. X3.9-198x [39]. X3J13 [1355]. X3J3 [33, 18, 19, 20, 25]. XC [1253]. XE [1751]. XT [1245].

YABBER [472]. YABBER-yet [472]. yacc [1422, 944, 630]. YALE [967]. yielding [1249]. York [1949]. Yorktown [1960].

Z [281]. zero [1699]. zero-termination [1699]. ZS [1283]. ZS-1 [1283].

References

Anderson:1982:AII

[1] S. O. Anderson and R. C. Backhouse. An alternative implementation of an insertion-only recovery technique. Acta Informatica, 18(3):289–298, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cichelli:1980:MPH

[2] R. J. Cichelli. Minimal perfect hash functions made simple. Communications of the ACM, 23(1):17–19, January 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks in [4], and the author’s response in [3], and also later extension [6]. A simple improvement giving dramatic speedups is described in [1824].

Cichelli:1980:TCC

[3] R. J. Cichelli. Technical correspondence: On Cichelli’s minimal perfect hash functions method. Communications of the ACM, 23(12):728–729, December 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This is the author’s response to the comments in [4] about [2]. See also [6].

Jaeschke:1980:TCC

[4] G. Jaeschke and G. Osterburg. Technical correspondence: On Cichelli’s minimal perfect hash functions method. Communications of the ACM, 23(12): 728–729, December 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This letter to the editor contains comments on [2], together with a response from R. J. Cichelli [3].

Gonnet:1983:CAD

[5] G. Gonnet and F. W. Tompa. A constructive approach to the design of algorithms and their data structures. Communications of the ACM, 26(11): 912–920, November 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sager:1985:PTG

[6] Thomas J. Sager. A polynomial time generator for minimal perfect hash functions. Communications of the ACM, 28(5):523–532, May 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/3538.html. See [2].

Coonen:1980:IGP

[7] Jerome T. Coonen. An implementation guide to a proposed standard for floating point arithmetic. Computer, 13 (1):68–79, January 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See errata in [9].

Stevenson:1981:PSBa

[8] David Stevenson. A proposed standard for binary floating-point arithmetic. Computer, 14(3):51–62, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [1937, 1938].

Coonen:1981:EIG

[9] Jerome T. Coonen. Errata: An implementation guide to a proposed standard for floating point arithmetic. Computer, 14(3):62–??, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [7].

Cody:1981:APF

[10] William J. Cody, Jr. Analysis of proposals for the floating-point standard. Computer, 14(3):63–68, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:1981:API

[11] David Hough. Applications of the proposed IEEE-754 standard for floating point arithmetic. Computer, 14 (3):70–74, March 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1981:UDN

[12] Jerome T. Coonen. Underflow and the denormalized numbers. Computer, 14 (3):75–87, March 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berns:1985:SFR

[13] G. M. Berns. Significant Fortran reliability problems and what Fortran 8X can do about them. ACM FORTRAN Forum, 4(1):25–32, March 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Smith:1985:SPF

[14] B. T. Smith. Software parts in Fortran 8X. ACM FORTRAN Forum, 4 (1):33–38, March 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1985:IFG

[15] J. L. Wagener. Is Fortran getting too big? ACM FORTRAN Forum, 4(2): 16, June 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Aharonian:1985:LKF

[16] G. Aharonian. Let’s kill future Fortrans. ACM FORTRAN Forum, 4(3): 21–23, October 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Aharonian:1985:TCA

[17] Gregory Aharonian. Technical contributions: Automated Fortran conversion. ACM FORTRAN Forum, 4(4): 13–19, December 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRa

[18] John Reid. John Reid reports: X3J3 meeting, August 1986. ACM FOR-TRAN Forum, 6(1):4–77, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRb

[19] John Reid. John Reid reports: X3J3 meeting, November 1986. ACM FOR-TRAN Forum, 6(1):8–11, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRc

[20] John Reid. John Reid reports: X3J3 meeting, February 1987. ACM FOR-TRAN Forum, 6(1):12–27, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1987:FES

[21] M. Metcalf. Fortran 8X–the emerging standard. ACM FORTRAN Forum, 6 (1):28–47, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Adams:1987:FDF

[22] J. Adams and J. Reid. Fortran 8X– the draft Fortran standard revision. ACM FORTRAN Forum, 6(2):27–38, October 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Guinier:1987:FPP

[23] D. Guinier. Fortran programs on a personal computer. ACM FOR-TRAN Forum, 6(2):39–47, October 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1987:CFD

[24] Lorer P. Meissner. Concerning the Fortran 8X draft. ACM FORTRAN Forum, 6(3):1–31, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:XMN

[25] John Reid. X3J3 meeting, November 1987. ACM FORTRAN Forum, 6(3): 4–96, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Hybl:1987:C

[26] Albert Hybl. Correspondence. ACM FORTRAN Forum, 6(3):10–11, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:1987:PFS

[27] L. Schonfelder and S. Morgan. Pointer functionality sans data-type. ACM FORTRAN Forum, 6(3):12–26, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Lahey:1987:FS

[28] T. Lahey. The Fortran 8X standard. ACM FORTRAN Forum, 6(3): 27–30, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Fullerton:1987:ADF

[29] J. Fullerton. An alternate design for Fortran 8X. ACM FORTRAN Forum, 6(3):31–40, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Miyawaki:1987:FA

[30] S. Miyawaki. Fortran 8X abridgement. ACM FORTRAN Forum, 6(3): 41–78, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Dolan:1988:FD

[31] Kent P. Dolan. Fortran 8X discussion. ACM FORTRAN Forum, 7(1):17–27, April 1988. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1988:MFP

[32] Loren Meissner. More on Fortran 8X pointer proposals. ACM FORTRAN Forum, 7(1):28–36, April 1988. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:XAD

[33] Loren P. Meissner. X3J3 avoids deadlock. ACM FORTRAN Forum, 8(1): 1–53, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:JRR

[34] John Reid. John Reid reports. ACM FORTRAN Forum, 8(1):6–32, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:ERC

[35] Edward Reid. Edward Reid: some comments on Fortran 8X. ACM FOR-TRAN Forum, 8(1):33–36, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:SF

[36] Loren P. Meissner. Summary of Fortran 88. ACM FORTRAN Forum, 8(2):15– 31, May 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Adams:1989:SCF

[37] Jeanne C. Adams and Jerrold L. Wagener. Summary of changes to Fortran draft. ACM FORTRAN Forum, 8(3): 13–19, August 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:PIO

[38] Loren P. Meissner. Partial-record input and output. ACM FORTRAN Forum, 8(3):20–23, August 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

CBEMA:1989:FD

[39] Computer and Business Equipment Manufacturers Association (CBEMA). Fortran 8X draft, May 1989: American National Standard for Information Systems Programming Language Fortran, S8 (X3.9-198x), Revision of X3.9-1978. ACM FORTRAN Forum, 8(4): various, December 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Horowitz:1986:SSD

[40] Ellis Horowitz and Ronald C. Williamson. SODOS: a software documentation support environment — its definition.

IEEE Transactions on Software Engineering, SE-12(8):849–859, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Grune:1984:HPA

[41] D. Grune. How to produce all sentences from a two-level grammar. Information Processing Letters, 19(4):181–185, ?? 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kral:1984:SE

[42] J. Kral. On software equations. Information Processing Letters, 19(4): 191–196, ?? 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Smith:1984:HDI

[43] S. Smith, D. Barnard, and I. MacLeod. Holophrasted displays in an interactive environment. Int. J. Man-Mach. Stud., 20(4):343–355, April 1984.

Srinivasan:1979:CFC

[44] B. Srinivasan and V. Gopalakrishna. Control flow complexity and structuredness of programs. ACM SIGPLAN Notices, 14(11):110–115, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [71].

Reynolds:1980:RAA

[45] John C. Reynolds. Reasoning about arrays. ACM SIGPLAN Notices, 15 (1):23, January 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beidler:1980:ILD

[46] John Beidler and John Meinke. An intermediate level dynamic storage capability. ACM SIGPLAN Notices, 15(1): 24–33, January 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cichelli:1980:PIC

[47] Richard J. Cichelli. Pascal-I: interactive, conversational Pascal-S. ACM SIGPLAN Notices, 15(1):34–44, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culik:1980:WFL

[48] Karel Culik. What is a flowchart loop and about structured programming. ACM SIGPLAN Notices, 15(1): 45–57, January 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gannon:1980:TTS

[49] John Gannon, Paul McMullin, Richard Hamlet, and Mark Ardis. Testing traversable stacks. ACM SIG-PLAN Notices, 15(1):58–65, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaye:1980:IPI

[50] Douglas R. Kaye. Interactive Pascal input. ACM SIGPLAN Notices, 15(1): 66–68, January 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaPlant:1980:ULA

[51] William P. LaPlant, Jr. and Matthew A. Jaro. The UNIMAC lexical analyzer. ACM SIGPLAN Notices, 15(1):69–76, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laski:1980:HAP

[52] Janusz W. Laski. A hierarchical approach to program testing. ACM SIG-PLAN Notices, 15(1):77–85, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moss:1980:SPL

[53] C. D. S. Moss. Structures programming with LOOP statements. ACM SIG-PLAN Notices, 15(1):86–94, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schoenberger:1980:RMP

[54] Roger Schoenberger. Resource management in PORTAL. ACM SIGPLAN Notices, 15(1):95–102, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sutton:1980:AUI

[55] Stephen A. Sutton. An algorithm for user interaction. ACM SIGPLAN Notices, 15(1):103–104, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1980:APL

[56] Richard N. Taylor. Assertions in programming languages. ACM SIGPLAN Notices, 15(1):105–114, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosen:1980:HPP

[57] Barry K. Rosen. How practical is POPL? ACM SIGPLAN Notices, 15 (1):115–116, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Firestone:1980:ELS

[58] Roger M. Firestone. An experimental LISP system for the Sperry Univac 1100 series. ACM SIGPLAN Notices, 15(1):117–129, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mayor:1980:LNA

[59] Michael A. Mayor. A language for network analysis and definition. ACM SIGPLAN Notices, 15(1):130–138, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehrlich:1980:PHL

[60] F. P. Mehrlich and S. M. Tanner. Portability: high level language implementation. ACM SIGPLAN Notices, 15(1):139–145, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stodola:1980:PPL

[61] Frank W. Stodola. The PLUS programming language. ACM SIGPLAN Notices, 15(1):146–155, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weidner:1980:CCS

[62] Thomas G. Weidner. CHAMIL: a case study in microprogramming language design. ACM SIGPLAN Notices, 15(1): 156–166, January 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1980:SRI

[63] Henry G. Baker, Jr. A source of redundant identifiers in PASCAL programs. ACM SIGPLAN Notices, 15(2): 14–16, February 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boute:1980:SAR

[64] Raymond T. Boute. Simplifying ADA by removing limitations. ACM SIG-PLAN Notices, 15(2):17–29, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganzinger:1980:OIA

[65] Harald Ganzinger and Knut Ripken. Operator identification in ADA: formal specification, complexity, and concrete implementation. ACM SIGPLAN Notices, 15(2):30–42, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [115].

Ince:1980:IIL

[66] D. C. Ince. An interpretative implementation of limited entry decision tables in ALGOL 68. ACM SIGPLAN Notices, 15(2):43–45, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kral:1980:PP

[67] Jaroslav Král. Parkinson programming. ACM SIGPLAN Notices, 15(2): 46–50, February 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1980:FPS

[68] Godfrey Lee, Tim Boreham, Bob Minns, Fred Smith, and Rick Soderstrom. FORTRAN programming standards. ACM SIGPLAN Notices, 15(2): 51–63, February 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silberschatz:1980:ECA

[69] A. Silberschatz and R. B. Kieburtz. The external consistency of abstract data types. ACM SIGPLAN Notices, 15(2):64–73, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weber:1980:GRA

[70] Manfred Weber and Susan L. Bernstein. Global register allocation for non-equivalent register sets. ACM SIG-PLAN Notices, 15(2):74–81, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zucconi:1980:CSF

[71] Linda Zucconi. A comment on “Control flow complexity and structuredness of programs” by Srinivasan and Gopalakrishna. ACM SIGPLAN Notices, 15(2):82–84, February 1980. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [44].

Anderson:1980:TSL

[72] Bruce Anderson. Type syntax in the language ‘C’, an object lesson in syntactic innovation. ACM SIG-PLAN Notices, 15(3):21–27, March 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klaeren:1980:ASS

[73] Herbert A. Klaeren. An abstract software specification technique based on structural recursion. ACM SIG-PLAN Notices, 15(3):28–34, March 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raiha:1980:BAG

[74] K.-J. Raiha. Bibliography on attribute grammars. ACM SIGPLAN Notices, 15 (3):35–44, March 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1980:TRA

[75] Mary Shaw and Wm. A. Wulf. Toward relaxing assumptions in languages and their implementations. ACM SIG-PLAN Notices, 15(3):45–61, March 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Addyman:1980:DPP

[76] A. M. Addyman. A draft proposal for PASCAL. ACM SIGPLAN Notices, 15 (4):1–66, April 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Addyman:1980:PS

[77] A. M. Addyman. PASCAL standardisation. ACM SIGPLAN Notices, 15 (4):67–69, April 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1980:GNN

[78] Norman H. Cohen. Gödel numbers: a new approach to structured programming. ACM SIGPLAN Notices, 15 (4):70–74, April 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harel:1980:DCO

[79] David Harel. “do considered od” considered odder than “do considered ob”. ACM SIGPLAN Notices, 15 (4):75, April 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hetrick:1980:CCH

[80] Brian A. Hetrick and H. Eisenberg. Confinement of a class of harmful effects of the goto statement. ACM SIGPLAN Notices, 15(4):76–78, April 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schild:1980:RTP

[81] R. Schild and H. Lienhard. Realtime programming in PORTAL. ACM SIGPLAN Notices, 15(4):79–92, April 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cichelli:1980:FP

[82] Richard J. Cichelli. Fixing PASCAL’s I/O. ACM SIGPLAN Notices, 15(5): 19, May 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowley:1980:SPR

[83] Charles Crowley. Structured programming is reductionistic! ACM SIG-PLAN Notices, 15(5):20–23, May 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ellis:1980:LS

[84] John R. Ellis. A Lisp shell. ACM SIGPLAN Notices, 15(5):24–34, May 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1980:TPP

[85] Douglas W. Jones. Tasking and parameters: a problem area in Ada. ACM SIGPLAN Notices, 15(5):37–40, May 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leinbaugh:1980:IC

[86] Dennis W. Leinbaugh. Indenting for the compiler. ACM SIGPLAN Notices, 15(5):41–48, May 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levine:1980:WLB

[87] John Levine. Why a Lisp-based command language? ACM SIG-PLAN Notices, 15(5):49–53, May 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ljungkvist:1980:PEF

[88] Sten Ljungkvist. Pascal and existing Fortran files. ACM SIGPLAN Notices, 15(5):54–55, May 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noodt:1980:SEP

[89] Terje Noodt and Dag Belsnes. A simple extension of PASCAL for quasi-parallel processing. ACM SIGPLAN Notices, 15(5):56–65, May 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanZandt:1980:VTS

[90] John Van Zandt. On visual and textual structures. ACM SIGPLAN Notices, 15(5):78–80, May 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Galkowski:1980:CDC

[91] J. T. Galkowski. A critique of the DOD common language effort. ACM SIGPLAN Notices, 15(6):15–18, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blatt:1980:GBS

[92] D. W. E. Blatt. On the great big substitution problem. ACM SIG-PLAN Notices, 15(6):19–27, June 1980.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bradshaw:1980:PSL

[93] Franklyn T. Bradshaw, George W. Ernst, Raymond J. Hookway, and William F. Ogden. Procedure semantics and language definition. ACM SIG-PLAN Notices, 15(6):28–33, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:1980:PMP

[94] Victor S. Foster. Performance measurement of a Pascal compiler. ACM SIGPLAN Notices, 15(6):34–38, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahjoub:1980:NMC

[95] Ahmed Mahjoub. A new Modula compiler for the LSI-11. ACM SIG-PLAN Notices, 15(6):39–45, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNeil:1980:PPE

[96] Michael McNeil and William Tracz. PL/I program efficiency. ACM SIG-PLAN Notices, 15(6):46–60, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1980:SAP

[97] Brian Meek. Serial attitudes, parallel attitudes. ACM SIGPLAN Notices, 15 (6):61–63, June 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moitra:1980:NAS

[98] Abha Moitra. A note on algebraic specification of binary trees. ACM SIG-PLAN Notices, 15(6):64–67, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:1980:FMF

[99] J. K. Reid. Functions for manipulating floating-point numbers. ACM SIG-PLAN Notices, 15(6):68–76, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBos:1980:CAP

[100] Jan van den Bos. Comments on ADA process communication. ACM SIG-PLAN Notices, 15(6):77–81, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandy:1980:SMD

[101] K. M. Chandy and J. Misra. A simple model of distributed programs based on implementation-hiding and process autonomy. ACM SIGPLAN Notices, 15(7–8):26–35, July/August 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeRemer:1980:SDP

[102] Frank DeRemer, Tom Pennello, and Richard Meyers. A syntax diagram for (preliminary) Ada. ACM SIG-PLAN Notices, 15(7–8):36–47, July/ August 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Glass:1980:PWF

[103] Robert L. Glass. The “project which failed” which succeeded. ACM SIG-PLAN Notices, 15(7–8):48–51, July/ August 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1980:PIO

[104] D. C. Ince. Paged input/output in some high level languages. ACM SIG-PLAN Notices, 15(7–8):52–57, July/ August 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jackel:1980:FPP

[105] Manfred Jackel. A formatting parser for PASCAL programs. ACM SIG-PLAN Notices, 15(7–8):58–63, July/ August 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nicolescu:1980:SSC

[106] Radu Nicolescu. Some short comments on the definition and the documentation of the Ada programming language. ACM SIGPLAN Notices, 15(7–8):64– 71, July/August 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pagan:1980:NSA

[107] Frank G. Pagan. Nested sublanguages of ALGOL 68 for teaching purposes. ACM SIGPLAN Notices, 15(7–8):72– 81, July/August 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pennello:1980:SOI

[108] Tom Pennello, Frank DeRemer, and Richard Meyers. A simplified operator identification scheme for Ada. ACM SIGPLAN Notices, 15(7–8):82– 87, July/August 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smart:1980:PLV

[109] Robert Smart. Pointers to local variables. ACM SIGPLAN Notices, 15(7– 8):88–89, July/August 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Campbell:1980:OPP

[110] Roy H. Campbell and Robert B. Kolstad. An overview of Path Pascal’s design. ACM SIGPLAN Notices, 15(9): 13–14, September 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kolstad:1980:PPU

[111] Robert B. Kolstad and Roy H. Campbell. Path Pascal user manual. ACM SIGPLAN Notices, 15(9):15– 24, September 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chiarini:1980:FLC

[112] A. Chiarini. On FP languages combining forms. ACM SIGPLAN Notices, 15(9):25–27, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dodd:1980:SCR

[113] W. P. Dodd. Some comments on a recent article by salvadori and dumont regarding the evaluation of compound conditional expressions. ACM SIG-PLAN Notices, 15(9):28–31, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanata:1980:CCP

[114] Syuetsu Hanata and Tadamasa Satoh. Compact Chart — a program logic notation with high describability and understandability. ACM SIGPLAN Notices, 15(9):32–38, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Janas:1980:COI

[115] Jürgen M. Janas. A comment on ‘operator identification in ADA’ by Ganzinger and Ripken. ACM SIG-PLAN Notices, 15(9):39–43, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [65].

Sharp:1980:DOP

[116] John A. Sharp. Data oriented program design. ACM SIGPLAN Notices, 15(9): 44–57, September 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yucel:1980:BPF

[117] M. Nadir Yücel and Boğos Pinar. Bit processing with FORTRAN. ACM SIGPLAN Notices, 15(9):58– 60, September 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Takaoka:1980:IHV

[118] Tadao Takaoka. Introduction of history to variables. ACM SIGPLAN Notices, 15(9):61–63, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanik:1980:TEP

[119] Murat M. Tanik. Two experiments on a program complexity perception by programmers. ACM SIGPLAN Notices, 15(9):64–66, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashcroft:1980:SCM

[120] Ed Ashcroft and Bill Wadge. Some common misconceptions about Lucid. ACM SIGPLAN Notices, 15(10):15– 26, October 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brender:1980:CAA

[121] Ronald F. Brender. The case against Ada as an APSE command language. ACM SIGPLAN Notices, 15(10):27– 34, October 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gimpel:1980:CMP

[122] James F. Gimpel. Contour: a method of preparing structured flowcharts. ACM SIGPLAN Notices, 15(10):35– 41, October 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gobin:1980:FHP

[123] Marc A. Gobin. File handling in programming languages. ACM SIGPLAN Notices, 15(10):42–47, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goodenough:1980:AJS

[124] J. B. Goodenough. Ada (July 1980) syntax cross reference listing. ACM SIGPLAN Notices, 15(10):48–56, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledgard:1980:HEV

[125] Henry F. Ledgard. A human engineered variant of BNF. ACM SIG-PLAN Notices, 15(10):57–62, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moffat:1980:CPB

[126] D. V. Moffat. A categorized Pascal bibliography (June, 1980). ACM SIG-PLAN Notices, 15(10):63–75, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tai:1980:CSI

[127] Kuo-Chung Tai and Ken Garrard. Comments on the suggested implementation of tasking facilities in the ’Rationale for the design of the Ada programming language’. ACM SIGPLAN Notices, 15(10):76–84, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yehudai:1980:AIV

[128] Amiram Yehudai. Automatic indention versus program formatting. ACM SIG-PLAN Notices, 15(10):85–87, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goodenough:1980:ACV

[129] John B. Goodenough. The Ada compiler validation capability. ACM SIG-PLAN Notices, 15(11):1–8, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Notkin:1980:EPA

[130] David S. Notkin. An experience with parallelism in Ada. ACM SIGPLAN Notices, 15(11):9–15, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fairley:1980:ADT

[131] Richard E. Fairley. Ada debugging and testing support environments. ACM SIGPLAN Notices, 15(11):16– 25, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duncan:1980:UAI

[132] A. G. Duncan and J. S. Hutchison. Using Ada for industrial embedded microprocessor applications. ACM SIG-PLAN Notices, 15(11):26–35, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goos:1980:TCF

[133] Gerhard Goos and Georg Winterstein. Towards a compiler front-end for Ada. ACM SIGPLAN Notices, 15(11):36– 46, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Persch:1980:OPA

[134] Guido Persch, Georg Winterstein, Manfred Dausmann, and Sophia Drossopoulou. Overloading in preliminary Ada. ACM SIGPLAN Notices, 15(11):47–56, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Belmont:1980:TRA

[135] Peter A. Belmont. Type resolution in Ada: an implementation report. ACM SIGPLAN Notices, 15(11):57– 61, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sherman:1980:FSA

[136] Mark S. Sherman and Martha S. Borkan. A flexible semantic analyzer for Ada. ACM SIGPLAN Notices, 15(11):62–71, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenberg:1980:CAC

[137] Jonathan Rosenberg, David Alex Lamb, Andy Hisgen, and Mark Sherman. The Charrette Ada compiler. ACM SIGPLAN Notices, 15(11):72– 81, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hisgen:1980:RRA

[138] Andy Hisgen, David Alex Lamb, Jonathan Rosenberg, and Mark Sherman. A runtime representation for Ada variables and types. ACM SIGPLAN Notices, 15(11):82–90, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sherman:1980:ACG

[139] Mark Sherman, Andy Hisgen, David Alex Lamb, and Jonathan Rosenberg. An Ada code generator for VAX 11/780 with Unix. ACM SIGPLAN Notices, 15(11):91–100, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brosgol:1980:TME

[140] Benjamin M. Brosgol. TCOL_{Ada} and the “middle end” of the PQCC Ada compiler. ACM SIGPLAN Notices, 15(11):101–112, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Luckham:1980:PMD

[141] David C. Luckham and Wolfgang Polak. A practical method of documenting and verifying Ada programs with packages. ACM SIGPLAN Notices, 15(11):113–122, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Young:1980:GVA

[142] William D. Young and Donald I. Good. Generics and verification in Ada.

ACM SIGPLAN Notices, 15(11):123– 127, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krieg-Bruckner:1980:ATL

[143] Bernd Krieg-Brückner and David C. Luckham. ANNA: towards a language for annotating Ada programs. ACM SIGPLAN Notices, 15(11):128– 138, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1980:NAP

[144] Lori A. Clarke, Jack C. Wileden, and Alexander L. Wolf. Nesting in Ada programs is for the birds. ACM SIGPLAN Notices, 15(11):139– 145, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1980:ETA

[145] Lee MacLaren. Evolving toward Ada in real time systems. ACM SIGPLAN Notices, 15(11):146–155, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eventoff:1980:RMC

[146] W. Eventoff, D. Harvey, and R. J. Price. The rendezvous and monitor concepts: is there an efficiency difference? (concurrent programming language). ACM SIGPLAN Notices, 15 (11):156–165, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stevenson:1980:ATA

[147] D. R. Stevenson. Algorithms for translating Ada multitasking. ACM SIGPLAN Notices, 15(11):166–175, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Filipski:1980:AST

[148] Gary L. Filipski, Donald R. Moore, and John E. Newton. Ada as a software transition tool. ACM SIGPLAN Notices, 15(11):176–182, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albrecht:1980:SST

[149] Paul F. Albrecht, Philip E. Garrison, Susan L. Graham, Robert H. Hyerle, Patricia Ip, and Bernd Krieg-Brückner. Source-to-source translation: Ada to Pascal and Pascal to Ada. ACM SIGPLAN Notices, 15(11):183– 193, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dewar:1980:NAT

[150] Robert B. K. Dewar, Gerald A. Fisher, Jr., Edmond Schonberg, Robert Froehlich, Stephen Bryant, Clinton F. Goss, and Michael Burke. The NYU Ada translator and interpreter. ACM SIGPLAN Notices, 15(11):194– 201, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Belz:1980:MPI

[151] F. C. Belz, E. K. Blum, and D. Heimbigner. A multi-processing implementation-oriented formal definition of Ada in SEMANOL. ACM SIGPLAN Notices, 15(11):202–212, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lovengreen:1980:FMT

[152] Hans Henrik Løvengreen and Dines Bjørner. On a formal model of the tasking concept in Ada. ACM SIGPLAN Notices, 15(11):213–222, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Groves:1980:DVM

[153] L. J. Groves and W. J. Rogers. The design of a virtual machine for Ada. ACM SIGPLAN Notices, 15(11):223– 234, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bishop:1980:EMD

[154] Judy M. Bishop. Effective machine descriptors for Ada. ACM SIGPLAN Notices, 15(11):235–242, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1980:C

[155] Anonymous. Compilers. ACM SIG-PLAN Notices, 15(11):??, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1980:EMA

[156] Anonymous. Execution models and architecture. ACM SIGPLAN Notices, 15 (11):??, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1980:S

[157] Anonymous. Style. ACM SIGPLAN Notices, 15(11):??, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1980:Ta

[158] Anonymous. Tasking. ACM SIG-PLAN Notices, 15(11):??, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1980:Tb

[159] Anonymous. Transition. ACM SIG-PLAN Notices, 15(11):??, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1980:U

[160] Anonymous. Use. ACM SIGPLAN Notices, 15(11):??, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baxter:1980:VP

[161] L. Baxter. The versatility of PRO-LOG. ACM SIGPLAN Notices, 15(12): 15–16, December 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buxton:1980:IBP

[162] J. N. Buxton. An informal bibliography on programming support environments. ACM SIGPLAN Notices, 15(12):17–30, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cleaveland:1980:MS

[163] J. Craig Cleaveland. Mathematical specifications. ACM SIGPLAN Notices, 15(12):31–42, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1980:DCR

[164] LeRoy Johnson. do considered reflectively: a contribution to group programming. ACM SIGPLAN Notices, 15(12):43, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1980:UDU

[165] LeRoy Johnson. \underline{do} considered \underline{ob}viously \underline{odd} in three dimensions. ACM SIGPLAN Notices, 15(12):44, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pentzlin:1980:SCS

[166] Karl L. Pentzlin. A syntax for character and string constants supplying userdefined character codes. ACM SIG-PLAN Notices, 15(12):45–52, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sale:1980:CFS

[167] Arthur Sale. Counterview in favour of strict type compatibility (pascal).

ACM SIGPLAN Notices, 15(12):53– 55, December 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siero:1980:AAC

[168] P. L. J. Siero. APL and Algol68, the correspondence and the differences, especially in applications of graph-analysis. ACM SIGPLAN Notices, 15(12):56–61, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stroet:1980:ACP

[169] Jan Stroet. An alternative to the communication primitives in ADA. ACM SIGPLAN Notices, 15(12):62–74, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sumpter:1980:CSH

[170] Anthony G. Sumpter and Gerry E. Quick. Concurrency specification in high level languages. ACM SIGPLAN Notices, 15(12):75–81, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1980:PPS

[171] Nora M. Taylor. Protection of proprietary software. ACM SIGPLAN Notices, 15(12):82–84, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tutleman:1980:HLC

[172] Jonathan S. Tutleman. Handling large, complex and variable job-control streams using a procedure invocation package (‘PIP’) and a pseudo-library. ACM SIGPLAN Notices, 15(12):85– 91, December 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wagner:1980:FPL

[173] Neal R. Wagner. A FORTRAN preprocessor for the large program environment. ACM SIGPLAN Notices, 15(12): 92–103, December 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mylopoulos:1981:OKR

[174] John Mylopoulos. An overview of Knowledge Representation. ACM SIG-PLAN Notices, 16(1):5–12, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:TAI

[175] Anonymous. Tutorial on artificial intelligence research. ACM SIGPLAN Notices, 16(1):13–18, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLeod:1981:AD

[176] Dennis McLeod and John Miles Smith. Abstraction in databases. ACM SIG-PLAN Notices, 16(1):19–25, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:TDR

[177] Anonymous. Tutorial on database research. ACM SIGPLAN Notices, 16(1):

26–28, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rowe:1981:DAP

[178] Lawrence A. Rowe. Data abstraction from a programming language viewpoint. ACM SIGPLAN Notices, 16(1): 29–35, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:TPL

[179] Anonymous. Tutorial on programming language research. ACM SIG-PLAN Notices, 16(1):36–39, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:WSM

[180] Anonymous. What should be modelled? ACM SIGPLAN Notices, 16 (1):40–42, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:T

[181] Anonymous. Types. ACM SIGPLAN Notices, 16(1):43–52, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:B

[182] Anonymous. Behaviour. ACM SIG-PLAN Notices, 16(1):53–61, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:P

[183] Anonymous. Presentation. ACM SIG-PLAN Notices, 16(1):62–71, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:CM

[184] Anonymous. Consistency of models. ACM SIGPLAN Notices, 16(1):72–76, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:RBA

[185] Anonymous. Relationships between and among models. ACM SIGPLAN Notices, 16(1):77–82, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:AMT

[186] Anonymous. Application of modelling techniques. ACM SIGPLAN Notices, 16(1):83–87, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:WS

[187] Anonymous. Workshop Summary. ACM SIGPLAN Notices, 16(1):88–90, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albano:1981:IIS

[188] Antonio Albano and Renzo Orsini. An interactive integrated system to design and use data bases. ACM SIG-PLAN Notices, 16(1):91–93, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amble:1981:IDB

[189] Tore Amble. Integration of data base design in programming languages. ACM SIGPLAN Notices, 16(1):94, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balzer:1981:DSS

[190] Robert M. Balzer. Dynamic system specification. ACM SIGPLAN Notices, 16(1):95–97, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borgida:1981:DAE

[191] Alexander T. Borgida and Sol Greenspan. Data and activities: Exploiting hierarchies of classes. ACM SIGPLAN Notices, 16(1):98–100, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brodie:1981:DAD

[192] Michael L. Brodie. Data abstraction for designing database-intensive applications. ACM SIGPLAN Notices, 16(1): 101–103, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buneman:1981:UDT

[193] Peter Buneman and Ira Winston. The use of data type information in an interactive database environment. ACM SIGPLAN Notices, 16(1):104–106, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carbonell:1981:DRI

[194] Jaime G. Carbonell. Default reasoning and inheritance mechanisms on type hierarchies. ACM SIGPLAN Notices, 16(1):107–109, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cattell:1981:IDS

[195] R. G. G. Cattell. Integrating a database system and programming / information environment. ACM SIG-PLAN Notices, 16(1):110–111, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Codd:1981:DMD

[196] E. F. Codd. Data models in database management. ACM SIGPLAN Notices, 16(1):112–114, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cristian:1981:SBC

[197] Flaviu Cristian. On the specification of behavioural constraints. ACM SIG-PLAN Notices, 16(1):115–117, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deutsch:1981:CUM

[198] L. Peter Deutsch. Constraints: a uniform model for data and control. ACM SIGPLAN Notices, 16(1):118–120, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feather:1981:SCC

[199] Martin S. Feather. Some contrasts and considerations of an approach to modelling. ACM SIGPLAN Notices, 16(1): 121–123, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldstein:1981:INS

[200] Ira Goldstein. Integrating a networkstructured database into an objectoriented programming language. ACM SIGPLAN Notices, 16(1):124–125, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardgrave:1981:PDM

[201] W. Terry Hardgrave and Donald R. Deutsch. Processing data model abstractions. ACM SIGPLAN Notices, 16 (1):126–127, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hayes:1981:LVT

[202] Patrick J. Hayes and Gary G. Hendrix. A logical view of types. ACM SIG-PLAN Notices, 16(1):128–130, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hendrix:1981:MVD

[203] Gary G. Hendrix. Mediating the views of databases and database users. ACM SIGPLAN Notices, 16(1):131–132, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hitchcock:1981:DDO

[204] Peter Hitchcock. Data dictionaries in open system communication. ACM SIGPLAN Notices, 16(1):133–134, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katz:1981:HDH

[205] Randy H. Katz. Heterogeneous databases and high level abstraction. ACM SIGPLAN Notices, 16(1):135– 137, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

King:1981:MCR

[206] Jonathan J. King. Modelling concepts for reasoning about access to knowledge. ACM SIGPLAN Notices, 16(1): 138–140, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kreps:1981:RVC

[207] Peter Kreps. Relativism and views in a conceptual data base model. ACM SIGPLAN Notices, 16(1):141–143, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lacroix:1981:ATD

[208] Michel Lacroix and Alain Pirotte. Associating types with domains of relational data bases. ACM SIGPLAN Notices, 16(1):144–146, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavenworth:1981:DAA

[209] B. Leavenworth. A data abstraction approach to database modelling. ACM SIGPLAN Notices, 16(1):147–149, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levesque:1981:IKB

[210] Hector J. Levesque. Incompleteness in knowledge bases. ACM SIGPLAN Notices, 16(1):150–152, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lundberg:1981:AIM

[211] Bengt Lundberg and Janis A. Bubenko, Jr. Axiomatic information modeling. ACM SIGPLAN Notices, 16(1):153– 154, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mark:1981:UDO

[212] William Mark. Use of database organization in the Consul system. ACM SIGPLAN Notices, 16(1):155–157, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mayr:1981:MMD

[213] Heinrich C. Mayr. Make more of data types. ACM SIGPLAN Notices, 16(1): 158–160, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLeod:1981:CDM

[214] Dennis McLeod. On conceptual database modelling. ACM SIGPLAN Notices, 16(1):161–163, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyer:1981:TLA

[215] Bertrand Meyer. A three-level approach to the description of data structures, and notational framework. ACM SIGPLAN Notices, 16(1):164–166, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mylopoulos:1981:PRC

[216] John Mylopoulos. A perspective for research on conceptual modelling. ACM SIGPLAN Notices, 16(1):167–170, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paolini:1981:ADT

[217] Paolo Paolini. Abstract data types and Data Bases. ACM SIGPLAN Notices, 16(1):171–173, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiter:1981:DBL

[218] Raymond Reiter. Data bases: a logical perspective. ACM SIGPLAN Notices, 16(1):174–176, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rich:1981:MPV

[219] Charles Rich. Multiple points of view in modelling programs. ACM SIG-PLAN Notices, 16(1):177–179, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rowe:1981:IDD

[220] Lawrence A. Rowe. Issues in the design of database programming languages. ACM SIGPLAN Notices, 16(1):180– 182, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

dosSantos:1981:TCA

[221] C. S. dos Santos, A. L. Furtado, J. M. V. de Castilho, and S. E. R. de Carvalho. Towards constructive axiomatic specifications. ACM SIGPLAN Notices, 16(1):183–185, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:1981:DAT

[222] Joachim W. Schmidt. Data abstraction tools: Design, specification and application. ACM SIGPLAN Notices, 16(1): 186–188, January 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1981:ADT

[223] Mary Shaw. Abstraction, data types, and models for software. ACM SIG-PLAN Notices, 16(1):189–191, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sibley:1981:DMS

[224] E. H. Sibley. Database management systems past and present. ACM SIGPLAN Notices, 16(1):192, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sowa:1981:CSK

[225] John F. Sowa. A conceptual schema for Knowledge-based systems. ACM SIG-PLAN Notices, 16(1):193–195, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thatcher:1981:DAD

[226] J. W. Thatcher. Data abstraction, data bases and conceptual modelling (position Paper). ACM SIGPLAN Notices, 16(1):196–197, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserman:1981:EDA

[227] Anthony I. Wasserman. The extension of data abstraction to database management. ACM SIGPLAN Notices, 16(1):198–200, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weber:1981:DTU

[228] Herbert Weber. Are data types universal modelling concepts for data base systems? ACM SIGPLAN Notices, 16(1):201–202, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wedekind:1981:CAD

[229] Hartmut H. Wedekind. Constructive Abstract Data types (CAD). ACM SIGPLAN Notices, 16(1):203–206, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zilles:1981:TAM

[230] Stephen N. Zilles. Types, algebras and modeling. ACM SIGPLAN Notices, 16(1):207–209, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bixler:1981:ASS

[231] David C. Bixler. Los Angeles SIGSPACE/SIGSOFT/SIGPLAN December meeting. ACM SIGPLAN Notices, 16(2):12, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashworth:1981:PP

[232] Robert Ashworth. Prettyprinting for power (FORTRAN). ACM SIGPLAN Notices, 16(2):16–17, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bandyopadhyay:1981:SPL

[233] S. K. Bandyopadhyay. A study on program level dependency of implemented algorithms on its potential operands. ACM SIGPLAN Notices, 16(2):18– 25, February 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bandyopadhyay:1981:TRB

[234] S. K. Bandyopadhyay. Theoretical relationships between potential operands and basic measurable properties of algorithm structure. ACM SIGPLAN Notices, 16(2):26–34, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haridi:1981:IEE

[235] Seif Haridi, John-Olof Bauner, and Gert Svensson. An implementation and empirical evaluation of the tasking facilities in ADA. ACM SIGPLAN Notices, 16(2):35–47, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cormack:1981:ASO

[236] G. V. Cormack. An algorithm for the selection of overloaded functions in ADA. ACM SIGPLAN Notices, 16(2): 48–52, February 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dewhurst:1981:ERT

[237] Stephen C. Dewhurst. An equivalence result for temporal logic. ACM SIG-PLAN Notices, 16(2):53–55, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rutter:1981:UHL

[238] Paul Rutter. Using a high level language as a cross assembler. ACM SIG-PLAN Notices, 16(2):56–58, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanson:1981:PL

[239] David R. Hanson. The Y programming language. ACM SIGPLAN Notices, 16(2):59–68, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLeanJones:1981:IS

[240] Douglas McLean Jones. Introduction to SQURL. ACM SIGPLAN Notices, 16(2):69–76, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moffat:1981:EPA

[241] David V. Moffat. Enumerations in Pascal, Ada, and beyond. ACM SIG-PLAN Notices, 16(2):77–82, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Magee:1981:ICC

[242] Patrick W. Magee. Improving the COBOL CALLed and CALLing program interface with a pre-processor or CALL statements are harmful, but can be controlled. ACM SIGPLAN Notices, 16(2):83–88, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahjoub:1981:SCA

[243] Ahmed Mahjoub. Some comments on Ada as a real-time programming language. ACM SIGPLAN Notices, 16(2): 89–95, February 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silberschatz:1981:SMA

[244] Abraham Silberschatz. On the synchronization mechanism of the Ada language. ACM SIGPLAN Notices, 16(2): 96–103, February 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitaker:1981:SAI

[245] William A. Whitaker. Summary of the ADA implementor’s meeting: December 1980. ACM SIGPLAN Notices, 16(2):104–109, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitaker:1981:CPA

[246] William A. Whitaker. Comments on portions of the ACM SIGPLAN conference on the ADA programming language not available in the proceedings. ACM SIGPLAN Notices, 16(2): 110–112, February 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Young:1981:SVP

[247] William D. Young and Donald I. Good. Steelman and the verifiability of (preliminary) ADA. ACM SIGPLAN Notices, 16(2):113–119, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zimmer:1981:CFC

[248] J. A. Zimmer. A control form for complex iterations. ACM SIGPLAN Notices, 16(2):120–131, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bates:1981:PPD

[249] Rodney M. Bates. A PASCAL prettyprinter with a different purpose. ACM SIGPLAN Notices, 16(3):10–17, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cole:1981:ASC

[250] S. N. Cole. Ada syntax cross reference. ACM SIGPLAN Notices, 16(3):18–47, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dailey:1981:PIS

[251] Thomas M. Dailey III. A procedure invocation syntax for enhanced readability. ACM SIGPLAN Notices, 16 (3):48–52, March 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furtado:1981:PSI

[252] A. L. Furtado and P. A. S. Veloso. Procedural specifications and implementations for abstract data types. ACM SIGPLAN Notices, 16(3):53–62, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1981:CMB

[253] Warren A. Harrison and Kenneth I. Magel. A complexity measure based on nesting level. ACM SIGPLAN Notices, 16(3):63–74, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mengarini:1981:MFA

[254] Bill Mengarini. Macro facilities in ADA. ACM SIGPLAN Notices, 16 (3):75–81, March 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Orejas:1981:EMA

[255] Fernando Orejas. Even more on advice on structuring compilers and proving them correct: changing an arrow. ACM SIGPLAN Notices, 16(3):82–84, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Persch:1981:LGR

[256] G. Persch, G. Winterstein, S. Drossopoulou, and M. Dausmann. An LALR (1) grammar for (revised) Ada. ACM SIGPLAN Notices, 16(3):85–98, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tripathi:1981:HVV

[257] Anand R. Tripathi, William D. Young, Donald I. Good, and James C. Browne. HAL/S/V: a verifiable subset for HAL/S. ACM SIGPLAN Notices, 16 (3):102–113, March 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barach:1981:STR

[258] David R. Barach and Frederick N. Webb. The symbol table representation of types in the PROPHET system. ACM SIGPLAN Notices, 16(4):17–26, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergstra:1981:CDS

[259] J. A. Bergstra and J. J. Ch. Meyer. I/O-computable data structures. ACM SIGPLAN Notices, 16(4):27–32, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergstra:1981:ICD

[260] J. A. Bergstra and J. J. Ch. Meyer. I/O-computable data structures. ACM SIGPLAN Notices, 16(4):27–32, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowley:1981:FSM

[261] Charles Crowley. A finite state machine for Western Swing. ACM SIGPLAN Notices, 16(4):33–35, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Falcone:1981:VPL

[262] Joseph R. Falcone. A vector programming language. ACM SIGPLAN Notices, 16(4):36–50, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1981:TAC

[263] Warren Harrison and Kenneth Magel. A topological analysis of the complexity of computer programs with less than three binary branches. ACM SIGPLAN Notices, 16(4):51–63, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunt:1981:BPC

[264] J. G. Hunt. Bracketing programme constructs. ACM SIGPLAN Notices, 16(4):64–67, April 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kylov:1981:SPP

[265] Haim I. Kylov and S. Nudelman. Structured programs in PL/I. ACM SIGPLAN Notices, 16(4):68–69, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marca:1981:SPS

[266] David Marca. Some Pascal style guidelines. ACM SIGPLAN Notices, 16 (4):70–80, April 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perkins:1981:LA

[267] Hal Perkins. Lazy I/O is not the answer. ACM SIGPLAN Notices, 16 (4):81–88, April 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perkins:1981:LIA

[268] H. Perkins. Lazy I/O is not the answer. ACM SIGPLAN Notices, 16(4):81–88, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanik:1981:PMC

[269] Murat M. Tanik. Prediction models for cyclomatic complexity. ACM SIGPLAN Notices, 16(4):89–97, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sale:1981:ARU

[270] A. H. J. Sale. Addition of repeat & until as identifiers. ACM SIG-PLAN Notices, 16(4):98–103, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolberg:1981:CCS

[271] John R. Wolberg. Comparing the cost of software conversion to the cost of reprogramming. ACM SIGPLAN Notices, 16(4):104–110, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1981:NB

[272] R. L. Wexelblat. New books. ACM SIGPLAN Notices, 16(5):16–18, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1981:SAP

[273] Michael L. Cook and Mark G. Arnold. A structured APL preprocessor. ACM SIGPLAN Notices, 16(5):22–31, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1981:MMS

[274] Michael Levy. Meeting on melpomenes — a story used to introduce the theory of data types. ACM SIGPLAN Notices, 16(5):32–35, May 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Redwine:1981:PDE

[275] Samuel T. Redwine, Jr. A paradigm for dealing with exceptions. ACM SIG-PLAN Notices, 16(5):36–38, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Remy:1981:EMC

[276] J. L. Remy and P. A. S. Veloso. An economical method for comparing data type specifications. ACM SIG-PLAN Notices, 16(5):39–42, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sassa:1981:PMR

[277] Masataka Sassa. A poor man’s realization of recursive structured FOR-TRAN. ACM SIGPLAN Notices, 16 (5):43–53, May 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stepoway:1981:ELB

[278] Stephen L. Stepoway and Abraham Silberschatz. An extension to the language-based access-control mechanism of Jones and Liskov. ACM SIG-PLAN Notices, 16(5):54–58, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserman:1981:RRP

[279] A. I. Wasserman, D. D. Sheretz, M. L. Kersten, van de Riet, R. P., and M. D. Dippe. Revised report on the programming language PLAIN. ACM SIG-PLAN Notices, 16(5):59–80, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:NBa

[280] Anonymous. New books. ACM SIG-PLAN Notices, 16(5):??, May 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wood:1981:ZPE

[281] Steven R. Wood. Z: the 95% program editor. ACM SIGPLAN Notices, 16(6):1–7, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Teitelbaum:1981:WWC

[282] Tim Teitelbaum, Thomas Reps, and Susan Horwitz. The why and wherefore of the Cornell Program Synthesizer. ACM SIGPLAN Notices, 16(6):8–16, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1981:SDE

[283] Christopher W. Fraser. SYNTAXdirected editing of general data structures. ACM SIGPLAN Notices, 16 (6):17–21, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stromfors:1981:IES

[284] O. Strömfors and L. Jonesjö. The implementation and experiences of a structure-oriented text editor. ACM SIGPLAN Notices, 16(6):22–27, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1981:DLD

[285] Joseph M. Morris and Mayer D. Schwartz. The design of a languagedirected editor for block-structured languages. ACM SIGPLAN Notices, 16 (6):28–33, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Good:1981:EFU

[286] Michael Good. Etude and the folklore of user interface design. ACM SIGPLAN Notices, 16(6):34–43, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walker:1981:DES

[287] Janet H. Walker. The document editor: a support environment for preparing technical documents. ACM SIG-PLAN Notices, 16(6):44–50, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turba:1981:CST

[288] Thomas N. Turba. Checking for spelling and typographical errors in computer-based text. ACM SIG-PLAN Notices, 16(6):51–60, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cherry:1981:CAW

[289] Lorinda Cherry. Computer aids for writers. ACM SIGPLAN Notices, 16 (6):61–67, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldfarb:1981:GAD

[290] C. F. Goldfarb. A generalized approach to document markup. ACM SIGPLAN Notices, 16(6):68–73, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:1981:PHD

[291] Todd Allen, Robert Nix, and Alan Perlis. PEN: a hierarchical document editor. ACM SIGPLAN Notices, 16 (6):74–81, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chamberlin:1981:JIS

[292] Donald D. Chamberlin, James C. King, Donald R. Slutz, Stephen J. P. Todd, and Bradford W. Wade. JANUS: An interactive system for document composition. ACM SIGPLAN Notices, 16 (6):82–91, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kernighan:1981:PLT

[293] Brian W. Kernighan. PIC: a language for typesetting graphics. ACM SIG-PLAN Notices, 16(6):92–98, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanWyde:1981:GTL

[294] C. T. Van Wyde. A graphics typesetting language. ACM SIGPLAN Notices, 16(6):97–107, June 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanWyk:1981:GTL

[295] Christopher J. Van Wyk. A graphics typesetting language. ACM SIG-PLAN Notices, 16(6):99–107, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mikelsons:1981:PIP

[296] Martin Mikelsons. Prettyprinting in an interactive programming environment. ACM SIGPLAN Notices, 16(6): 108–116, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Achugbue:1981:LBP

[297] James O. Achugbue. On the line breaking problem in text formatting. ACM SIGPLAN Notices, 16(6):117– 122, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gosling:1981:RA

[298] James Gosling. A redisplay algorithm. ACM SIGPLAN Notices, 16 (6):123–129, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barach:1981:DPV

[299] David R. Barach, David H. Taenzer, and Robert E. Wells. Design of the PEN video editor display module. ACM SIGPLAN Notices, 16(6): 130–136, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammer:1981:IEI

[300] Michael Hammer, Richard Ilson, Tim Anderson, Edward Gilbert, Michael Good, Bahram Niamir, Larry Rosentein, and Sandor Schoichet. The implementation of Etude, an integrated and interactive document production system. ACM SIGPLAN Notices, 16 (6):137–146, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stallman:1981:EEC

[301] Richard M. Stallman. EMACS: the extensible, customizable selfdocumenting display editor. ACM SIG-PLAN Notices, 16(6):147–156, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:1981:ABB

[302] Brian K. Reid and David Hanson. An annotated bibliography of background material on text manipulation. ACM SIGPLAN Notices, 16(6): 157–160, June 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goguen:1981:CMS

[303] J. A. Goguen and J. Meseguer. Completeness of many-sorted equational logic. ACM SIGPLAN Notices, 16 (7):24–32, July 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gonnet:1981:CBU

[304] Gaston H. Gonnet. Coding benchmark for Utopia84. ACM SIGPLAN Notices, 16(7):33–34, July 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hankin:1981:DFP

[305] C. L. Hankin and H. W. Glaser. The data flow programming language CAJOLE-an informal introduction. ACM SIGPLAN Notices, 16(7):35–44, July 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kneis:1981:DSI

[306] Wilfried Kneis. Draft standard industrial real-time FORTRAN. ACM SIG-PLAN Notices, 16(7):45–60, July 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Magel:1981:REP

[307] Kenneth Magel. Regular expressions in a program complexity metric. ACM SIGPLAN Notices, 16(7):61–65, July 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moffat:1981:MPL

[308] David V. Moffat. A model for Pascallike typing. ACM SIGPLAN Notices, 16(7):66–74, July 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perrott:1981:UEI

[309] R. H. Perrott and D. K. Stevenson. Users’ experiences with the ILLIAC IV system and its programming languages. ACM SIGPLAN Notices, 16(7):75–88, July 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silberschatz:1981:NDP

[310] Abraham Silberschatz. A note on the distributed program component cell. ACM SIGPLAN Notices, 16(7):89–96, July 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1981:NBM

[311] R. L. Wexelblat. New books: May 1981. ACM SIGPLAN Notices, 16 (8):6–9, August 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1981:NBJ

[312] R. L. Wexelblat. New books: June 1981. ACM SIGPLAN Notices, 16(8): 10–13, August 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burgess:1981:IBL

[313] C. Burgess and L. James. An indexed bibliography for LR grammars and parsers. ACM SIGPLAN Notices, 16 (8):14–26, August 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Payne:1981:BSP

[314] A. J. Payne. A basis for software physics? ACM SIGPLAN Notices, 16 (8):27–30, August 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rapin:1981:NL

[315] Ch. Rapin and J. Menu. The Newton language. ACM SIGPLAN Notices, 16 (8):31–40, August 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schilit:1981:SGB

[316] Bill N. Schilit. A solution to the great big substitution problem: subject modification during pattern matching in SNOBOL4. ACM SIGPLAN Notices, 16(8):41–49, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shapiro:1981:PPE

[317] Ehud Shapiro, Gregg Collins, Lewis Johnson, and John Ruttenberg. PASES: a programming environment for PASCAL. ACM SIGPLAN Notices, 16(8):50–57, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1981:FLP

[318] David Spector. Full LR(1) parser generation. ACM SIGPLAN Notices, 16 (8):58–66, August 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderLinden:1981:MFA

[319] Peter van der Linden and Leonard K. Nicholson. Macro facilities in the Ada environment. ACM SIGPLAN Notices, 16(8):67–68, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winkler:1981:DBPa

[320] J. F. H. Winkler. Differences between preliminary and final Ada. ACM SIG-PLAN Notices, 16(8):69–81, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1981:NBb

[321] Anonymous. New books. ACM SIG-PLAN Notices, 16(8):??, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonet:1981:ASD

[322] Rafael Bonet, Antonio Kung, Knut Ripken, Robert K. Yates, Manfred Sommer, and Jürgen Winkler. Ada syntax diagrams for top-down analysis. ACM SIGPLAN Notices, 16(9): 29–41, September 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Booch:1981:DSD

[323] Grady Booch. Describing software design in Ada. ACM SIGPLAN Notices, 16(9):42–47, September 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeRemer:1981:ASC

[324] Frank DeRemer, Tom Penncllo, and W. M. McKeeman. Ada syntax chart. ACM SIGPLAN Notices, 16(9):48– 59, September 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Konstam:1981:MCP

[325] Aaron H. Konstam. A method for controlling parallelism in programming languages. ACM SIGPLAN Notices, 16 (9):60–65, September 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lauer:1981:SCP

[326] P. E. Lauer. Synchronization of concurrent processes without globality assumptions. ACM SIGPLAN Notices, 16(9):66–80, September 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rowland:1981:EPR

[327] David A. Rowland. An extension to PASCAL READ and WRITE procedures. ACM SIGPLAN Notices, 16(9): 81–82, September 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shimon:1981:SER

[328] Cohen Shimon. Simplicity = efficiency = readability: a simple INFIX to PREFIX algorithm. ACM SIGPLAN Notices, 16(9):83–89, September 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wetherell:1981:PAR

[329] C. S. Wetherell. Problems with the Ada reference grammar. ACM SIG-PLAN Notices, 16(9):90–104, September 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See note [523].

Ganapathi:1981:BAR

[330] M. Ganapathi and C. N. Fischer. Bibliography on automated retargetable code generation. ACM SIGPLAN Notices, 16(10):9–12, October 1981.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gustafson:1981:CFD

[331] David A. Gustafson. Control flow, data flow and data independence. ACM SIGPLAN Notices, 16(10):13–19, October 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maher:1981:DDS

[332] B. Maher and D. H. Sleeman. A data driven system for syntactic transformations. ACM SIGPLAN Notices, 16(10): 50–52, October 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malcolm:1981:BCC

[333] James A. Malcolm. Brevity and clarity in command languages. ACM SIG-PLAN Notices, 16(10):53–59, October 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maurer:1981:ACP

[334] W. D. Maurer. An alternate characterization of preorder, inorder, and postorder. ACM SIGPLAN Notices, 16(10):60–61, October 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narayana:1981:FTP

[335] K. T. Narayana and M. Clint. On the formal treatment of a permutation finding program. ACM SIGPLAN Notices, 16(10):62–73, October 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jenks:1981:LCA

[336] Richard D. Jenks and Barry M. Trager. A language for computational algebra. ACM SIGPLAN Notices, 16(11):22– 29, November 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mayer:1981:VRP

[337] Alastair J. W. Mayer. Value receiving procedures. ACM SIGPLAN Notices, 16(11):30–34, November 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winkler:1981:DBPb

[338] J. F. H. Winkler. Differences between preliminary and final Ada. ACM SIG-PLAN Notices, 16(11):35–48, November 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1981:IFS

[339] Darrel C. Ince. The implementation of file structures in some high level languages. ACM SIGPLAN Notices, 16(11):49–55, November 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1981:FSD

[340] D. G. Elliott. FORTRAN 77 and structured design. ACM SIGPLAN Notices, 16(12):7–9, December 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1981:CAR

[341] Alfred H. Morris, Jr. Can Ada replace FORTRAN for numerical computation? ACM SIGPLAN Notices, 16(12): 10–13, December 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fitzhorn:1981:CTC

[342] Patrick A. Fitzhorn and Gearold R. Johnson. C: toward a concise syntactic description. ACM SIGPLAN Notices, 16(12):14–21, December 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ben-Ari:1981:MMU

[343] Mordechai Ben-Ari and Amiram Yehudai. A methodology for modular use of Ada. ACM SIGPLAN Notices, 16(12): 22–26, December 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanDeusen:1981:TR

[344] Mary Van Deusen. Types in RED. ACM SIGPLAN Notices, 16(12):27– 38, December 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1982:NBD

[345] R. L. Wexelblat. New books: December 1981. ACM SIGPLAN Notices, 17 (1):4–8, January 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goguen:1982:CMS

[346] J. A. Goguen and J. Meseguer. Completeness of many-sorted equational logic. ACM SIGPLAN Notices, 17(1): 9–17, January 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Booch:1982:NSC

[347] Grady Booch. Naming subprograms with clarity. ACM SIGPLAN Notices, 17(1):18–22, January 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dixon:1982:PCT

[348] Douglas F. Dixon. A Pascal compiler testing facility. ACM SIGPLAN Notices, 17(1):23–26, January 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laski:1982:AGI

[349] Janusz W. Laski. Application of the goal invariant to the structuring of programs. ACM SIGPLAN Notices, 17(1): 27–36, January 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kilov:1982:GWP

[350] Haim Kilov and S. Nudelman. Guidelines for writing PL/I programs. ACM SIGPLAN Notices, 17(1):37–39, January 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bykat:1982:EPI

[351] Alex Bykat. An extension to PASCAL input-output procedures. ACM SIG-PLAN Notices, 17(1):40–41, January 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stroustrup:1982:CAD

[352] Bjarne Stroustrup. Classes: an abstract data type facility for the C language. ACM SIGPLAN Notices, 17(1): 42–51, January 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1982:NBa

[353] Anonymous. New books. ACM SIG-PLAN Notices, 17(1):??, January 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1982:NBF

[354] R. L. Wexelblat. New books: February 1981. ACM SIGPLAN Notices, 17(2): 17–19, February 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tai:1982:CPP

[355] Kuo-Chung Tai. Comments on parameter passing techniques in programming languages. ACM SIGPLAN Notices, 17(2):24–27, February 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jessop:1982:APD

[356] Warren H. Jessop. Ada packages and distributed systems. ACM SIGPLAN Notices, 17(2):28–36, February 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Topor:1982:NER

[357] Rodney W. Topor. A note on error recovery in recursive descent parsers. ACM SIGPLAN Notices, 17(2):37– 40, February 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ford:1982:ORV

[358] Gary Ford and Brian Hansche. Optional, repeatable, and varying type parameters. ACM SIGPLAN Notices, 17(2):41–48, February 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1982:MBP

[359] Tim N. Roberts. More on bit processing with FORTRAN. ACM SIG-PLAN Notices, 17(2):49–52, February 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1982:SDG

[360] Mark Scott Johnson. A software debugging glossary. ACM SIGPLAN Notices, 17(2):53–70, February 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1982:NBb

[361] Anonymous. New books. ACM SIGPLAN Notices, 17(2):??, February 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Magel:1982:TSP

[362] Kenneth Magel. A theory of small program complexity. ACM SIGPLAN Notices, 17(3):37–45, March 1982. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morrison:1982:SSD

[363] Ronald Morrison. The string as a simple data type. ACM SIGPLAN Notices, 17(3):46–52, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prael:1982:POS

[364] C. E. Prael. PASCAL for operating software? A critical examination. ACM SIGPLAN Notices, 17(3):53–57, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salt:1982:DSS

[365] Norman F. Salt. Defining software science counting strategies. ACM SIG-PLAN Notices, 17(3):58–67, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scowen:1982:SSM

[366] R. S. Scowen. A standard syntactic metalanguage. ACM SIGPLAN Notices, 17(3):68–73, March 1982. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scowen:1982:SSP

[367] R. S. Scowen and Z. J. Ciechanowicz. Seven sorts of programs. ACM SIGPLAN Notices, 17(3):74–79, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1982:MOG

[368] David Spector. Minimal overhead garbage collection of complex list structure. ACM SIGPLAN Notices, 17(3): 80–82, March 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBosch:1982:TPL

[369] Peter van den Bosch. The translation of programming languages through the use of a graph transformation language. ACM SIGPLAN Notices, 17(3):83–92, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderLinden:1982:AOA

[370] Peter van der Linden. Ambiguity and orthogonality in Ada. ACM SIG-PLAN Notices, 17(3):93–94, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rattner:1982:HSC

[371] Justin Rattner. Hardware/software cooperation in the iAPX-32. ACM SIG-PLAN Notices, 17(4):1, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hennessy:1982:HST

[372] John Hennessy, Norman Jouppi, Forest Baskett, Thomas Gross, and John Gill. Hardware/software tradeoffs for increased performance. ACM SIG-PLAN Notices, 17(4):2–11, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rymarczyk:1982:CGP

[373] James W. Rymarczyk. Coding guidelines for pipelined processors. ACM SIGPLAN Notices, 17(4):12–19, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnsson:1982:OMP

[374] Richard K. Johnsson and John D. Wick. An overview of the Mesa processor architecture. ACM SIGPLAN Notices, 17(4):20–29, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berenbaum:1982:OSL

[375] Alan D. Berenbaum, Michael W. Condry, and Priscilla M. Lu. The operating system and language support features of the BELLMACTM-32 microprocessor. ACM SIGPLAN Notices, 17(4):30–38, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radin:1982:M

[376] George Radin. The 801 minicomputer. ACM SIGPLAN Notices, 17(4):39–47, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ditzel:1982:RAF

[377] David R. Ditzel and H. R. McLellan. Register allocation for free: The C machine stack cache. ACM SIGPLAN Notices, 17(4):48–56, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harbison:1982:AAO

[378] Samuel P. Harbison. An architectural alternative to optimizing compilers. ACM SIGPLAN Notices, 17 (4):57–65, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lampson:1982:FPC

[379] Butler W. Lampson. Fast procedure calls. ACM SIGPLAN Notices, 17 (4):66–76, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1982:SPM

[380] Douglas W. Jones. Systematic protection mechanism design. ACM SIGPLAN Notices, 17(4):77–80, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reed:1982:GPM

[381] Karl Reed. On a general property of memory mapping tables. ACM SIGPLAN Notices, 17(4):81–86, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1982:EIO

[382] Robert P. Cook and Nitin Donde. An experiment to improve operand addressing. ACM SIGPLAN Notices, 17 (4):87–91, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fusaoka:1982:CCH

[383] Akira Fusaoka and Masaharu Hirayama. Compiler chip: a hardware implementation of compiler. ACM SIGPLAN Notices, 17(4):92–95, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rau:1982:ASE

[384] B. R. Rau, C. D. Glaeser, and E. M. Greenawalt. Architectural support for the efficient generation of code for horizontal architectures. ACM SIGPLAN Notices, 17(4):96–99, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLear:1982:GCD

[385] R. E. McLear, D. M. Scheibelhut, and E. Tammaru. Guidelines for creating a debuggable processor. ACM SIG-PLAN Notices, 17(4):100–106, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilkes:1982:HSM

[386] M. V. Wilkes. Hardware support for memory protection: Capability implementations. ACM SIGPLAN Notices, 17(4):107–116, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pollack:1982:SAM

[387] Fred J. Pollack, George W. Cox, Dan W. Hammerstrom, Kevin C. Kahn, Konrad K. Lai, and Justin R. Rattner. Supporting Ada memory management in the iAPX-432. ACM SIGPLAN Notices, 17(4):117– 131, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sansonnet:1982:DEL

[388] J. P. Sansonnet, M. Castan, C. Percebois, D. Botella, and J. Perez. Direct execution of LISP on a list-directed architecture. ACM SIGPLAN Notices, 17 (4):132–139, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1982:SRA

[389] Mark Scott Johnson. Some requirements for architectural support of software debugging. ACM SIGPLAN Notices, 17(4):140–148, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Middelburg:1982:EPA

[390] C. A. Middelburg. The effect of the PDP-11 architecture on code generation for chill. ACM SIGPLAN Notices, 17(4):149–157, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sweet:1982:EAM

[391] Richard E. Sweet and James G. Sandman, Jr. Empirical analysis of the Mesa instruction set. ACM SIG-PLAN Notices, 17(4):158–166, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McDaniel:1982:AMI

[392] Gene McDaniel. An analysis of a Mesa instruction set using dynamic instruction frequencies. ACM SIG-PLAN Notices, 17(4):167–176, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiecek:1982:CSV

[393] Cheryl A. Wiecek. A case study of VAX-11 instruction set usage for compiler execution. ACM SIGPLAN Notices, 17(4):177–184, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maekawa:1982:FSA

[394] Mamoru Maekawa, Ken Sakamura, and Chiaki Ishikawa. Firmware structure and architectural support for monitors, vertical migration and user microprogramming. ACM SIGPLAN Notices, 17(4):185–194, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamibayashi:1982:HOS

[395] N. Kamibayashi, H. Ogawana, K. Nagayama, and H. Aiso. Heart: An operating system nucleus machine implemented by firmware. ACM SIG-PLAN Notices, 17(4):195–204, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahuja:1982:MMA

[396] Sudhir R. Ahuja and Abhaya Asthana. A multi-microprocessor architecture with hardware support for communication and scheduling. ACM SIG-PLAN Notices, 17(4):205–209, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1982:HDW

[397] Edsger W. Dijkstra. How do we tell truths that might hurt? ACM SIG-PLAN Notices, 17(5):13–15, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barach:1982:TFSa

[398] David R. Barach, David H. Taenzer, and Robert E. Wells. A technique for finding storage allocation errors in C-language problems. ACM SIG-PLAN Notices, 17(5):16–23, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ancona:1982:BCL

[399] M. Ancona, G. Dodero, and V. Gianuzzi. Building collections of LR(k) items with partial expansion of lookahead strings. ACM SIGPLAN Notices, 17(5):25–28, May 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guazzo:1982:AFA

[400] Mauro Guazzo. APL finds its APLOMB. ACM SIGPLAN Notices, 17(5):29–37, May 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kranc:1982:SCF

[401] Morris E. Kranc. A separate compilation facility for PASCAL. ACM SIG-PLAN Notices, 17(5):38–46, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greiter:1982:DTT

[402] Gebhard Greiter. A data type theory. ACM SIGPLAN Notices, 17(5):47–53, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carter:1982:CI

[403] Breck Carter. On choosing identifiers. ACM SIGPLAN Notices, 17(5):54–59, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holt:1982:SIC

[404] R. C. Holt. A short introduction to Concurrent Euclid. ACM SIG-PLAN Notices, 17(5):60–79, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thanh:1982:ITC

[405] Nguyen the Thanh and E. Walter Raschner. Indirect threaded code used to emulate a virtual machine. ACM SIGPLAN Notices, 17(5):80–89, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seyfer:1982:CTS

[406] H. K. Seyfer. Compiler test sets. ACM SIGPLAN Notices, 17(5):90–97, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madhavji:1982:CP

[407] N. H. Madhavji and I. R. Wilson. Cray Pascal. ACM SIGPLAN Notices, 17 (6):1–14, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kipps:1982:EPT

[408] I. M. Kipps. Experience with porting techniques on a COBOL 74 compiler. ACM SIGPLAN Notices, 17(6):15–21, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auslander:1982:OPC

[409] Marc Auslander and Martin Hopkins. An overview of the PL.8 compiler. ACM SIGPLAN Notices, 17(6):22–31, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1982:ETD

[410] Susan L. Graham, Robert R. Henry, and Robert A. Schulman. An experiment in table driven code generation. ACM SIGPLAN Notices, 17(6):32–43, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bird:1982:ICG

[411] Peter L. Bird. An implementation of a code generator specification language for table driven code generators. ACM SIGPLAN Notices, 17(6):44–55, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Landwehr:1982:EAC

[412] Rudolf Landwehr, Hans-Stephan Jansohn, and Gerhard Goos. Experience with an automatic code-generator generator. ACM SIGPLAN Notices, 17 (6):56–66, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burke:1982:PMS

[413] Michael Burke and Gerald A. Fisher, Jr. A practical method for syntactic error diagnosis and recovery. ACM SIGPLAN Notices, 17(6):67–78, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mauney:1982:FMA

[414] Jon Mauney and Charles N. Fischer. A forward move algorithm for LL and LR parsers. ACM SIGPLAN Notices, 17 (6):79–87, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1982:SED

[415] C. W. Johnson and C. Runciman. Semantic errors — diagnosis and repair. ACM SIGPLAN Notices, 17(6):88–97, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaitin:1982:RAS

[416] G. J. Chaitin. Register allocation and spilling via graph coloring. ACM SIGPLAN Notices, 17(6):98–105, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chow:1982:DDF

[417] Anita L. Chow and Andres Rudmik. The design of a data flow analyzer. ACM SIGPLAN Notices, 17 (6):106–113, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Markstein:1982:ORC

[418] Victoria Markstein, John Cocke, and Peter Markstein. Optimization of range checking. ACM SIGPLAN Notices, 17 (6):114–119, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1982:GCG

[419] Susan L. Graham, Peter B. Kessler, and Marshall K. Mckusick. Gprof: a call graph execution profiler. ACM SIGPLAN Notices, 17(6):120–126, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turba:1982:FDE

[420] Thomas N. Turba. A facility for the downward extension of a high-level language. ACM SIGPLAN Notices, 17 (6):127–133, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaLonde:1982:FCS

[421] Wilf R. LaLonde and Jim des Rivieres. A flexible compiler structure that allows dynamic phase ordering. ACM SIGPLAN Notices, 17(6): 134–139, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seyfer:1982:TTS

[422] Harlan K. Seyfer. Tailoring testing to a specific compiler — experiences. ACM SIGPLAN Notices, 17 (6):140–152, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koskimies:1982:CCU

[423] Kai Koskimies, Kari-Jouko Räihä, and Matti Sarjakoski. Compiler construction using attribute grammars. ACM SIGPLAN Notices, 17(6):153– 159, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farrow:1982:LYA

[424] Rodney Farrow. LINGUIST-86: Yet another translator writing system based on attribute grammars. ACM SIGPLAN Notices, 17(6):160– 171, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganzinger:1982:TGS

[425] Harald Ganzinger, Robert Giegerich, Ulrich Möncke, and Reinhard Wilhelm. A truly generative semantics-directed compiler generator. ACM SIGPLAN Notices, 17(6):172–184, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krumme:1982:PMC

[426] David W. Krumme and David H. Ackley. A practical method for code generation based on exhaustive search. ACM SIGPLAN Notices, 17(6):185– 196, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morgan:1982:AEI

[427] Thomas M. Morgan and Lawrence A. Rowe. Analyzing exotic instructions for a retargetable code generator. ACM SIGPLAN Notices, 17(6): 197–204, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crawford:1982:EPC

[428] John Crawford. Engineering a production code generator. ACM SIG-PLAN Notices, 17(6):205–215, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bodwin:1982:EEC

[429] James Bodwin, Laurette Bradley, Kohji Kanda, Diane Litle, and Uwe Pleban. Experience with an experimental compiler generator based on denotational semantics. ACM SIG-PLAN Notices, 17(6):216–229, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raskovsky:1982:DSS

[430] Martin R. Raskovsky. Denotational semantics as a specification of code generators. ACM SIGPLAN Notices, 17 (6):230–244, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sethi:1982:CFA

[431] Ravi Sethi. Control flow aspects of semantics directed compiling (summary). ACM SIGPLAN Notices, 17 (6):245–260, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brooks:1982:OCL

[432] Rodney A. Brooks, Richard P. Gabriel, and Guy L. Steele, Jr. An optimizing compiler for lexically scoped LISP. ACM SIGPLAN Notices, 17(6): 261–275, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griss:1982:CSP

[433] Martin L. Griss, Eric Benson, and Anthony C. Hearn. Current status of a portable LISP compiler. ACM SIG-PLAN Notices, 17(6):276–283, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moor:1982:ACP

[434] Ian W. Moor. An applicative compiler for a parallel machine. ACM SIG-PLAN Notices, 17(6):284–293, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marshall:1982:LGP

[435] Howard Z. Marshall. The linear graph package, a compiler building environment. ACM SIGPLAN Notices, 17(6):

294–300, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rudmik:1982:ESC

[436] Andres Rudmik and Barbara G. Moore. An efficient separate compilation strategy for very large programs. ACM SIGPLAN Notices, 17 (6):301–307, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avakian:1982:DIS

[437] Arra Avakian, Sam Haradhvala, Julian Horn, and Bruce Knobe. The design of an integrated support software system. ACM SIGPLAN Notices, 17 (6):308–317, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1982:SPS

[438] T. P. Baker. A single-pass syntaxdirected front end for Ada. ACM SIG-PLAN Notices, 17(6):318–326, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pratt:1982:EFS

[439] Terrence W. Pratt and George D. Maydwell. Experience with the formal semantic definition of HAL/S. ACM SIG-PLAN Notices, 17(6):327–333, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drossopoulou:1982:AGA

[440] S. Drossopoulou, J. Uhl, G. Persch, G. Goos, M. Dausmann, and G. Winterstein. An attribute grammar for Ada. ACM SIGPLAN Notices, 17(6): 334–339, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atkinson:1982:PAA

[441] Malcolm Atkinson, Ken Chisholm, and Paul Cockshott. PS-algol: an Algol with a persistent heap. ACM SIG-PLAN Notices, 17(7):24–31, July 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barach:1982:TFS

[442] David R. Barach, David H. Taenzer, and Robert E. Wells. A technique for finding storage allocation errors in C-language programs. ACM SIG-PLAN Notices, 17(7):32–38, July 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waters:1982:PES

[443] Richard C. Waters. Program editors should not abandon text oriented commands. ACM SIGPLAN Notices, 17 (7):39–46, July 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:1982:NLF

[444] J. C. H. Park. A new LALR formalism. ACM SIGPLAN Notices, 17(7):47–61, July 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harland:1982:ALE

[445] David M. Harland and Hamish I. E. Gunn. Another look at enumerated types. ACM SIGPLAN Notices, 17 (7):62–71, July 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gladney:1982:HPS

[446] H. M. Gladney. How to print with style. ACM SIGPLAN Notices, 17(7):72–79, July 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:1982:NGL

[447] Ronald Fischer. A note on gotoless programming in APL using the unquote-operator. ACM SIGPLAN Notices, 17(7):80–87, July 1982. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lesk:1982:EDT

[448] Michael Lesk. English for decision tables considered harmful. ACM SIG-PLAN Notices, 17(7):88, July 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See note [834].

deJong:1982:SDF

[449] M. D. de Jong and C. L. Hankin. Structured data flow programming. ACM SIGPLAN Notices, 17(8):18–27, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCaig:1982:TNS

[450] Martin McCaig. Towards a novel string processing language. ACM SIG-PLAN Notices, 17(8):28–33, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Early:1982:FPT

[451] Grady Early. Functional programming and the two-pass assembler. ACM SIG-PLAN Notices, 17(8):34–42, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1982:AIM

[452] David Spector. Ambiguities and insecurities in Modula-2. ACM SIG-PLAN Notices, 17(8):43–51, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1982:IDN

[453] Gregory P. Hill. An improvement over deeply nested IF-THEN-ELSE control structures. ACM SIGPLAN Notices, 17 (8):52–56, August 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [522].

Culik:1982:APL

[454] Karel Culik. An advanced programming language definition and domain correctness of programs. ACM SIG-PLAN Notices, 17(8):57–66, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1982:CLS

[455] Mark A. Jones. A comparison of LISP specifications of function definition and argument handling. ACM SIGPLAN Notices, 17(8):67–73, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansche:1982:SBC

[456] B. Hansche, S. Hudson, B., and Huey. Selected bibliography of compiler optimization topics. ACM SIGPLAN Notices, 17(8):74–83, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jokinen:1982:EPP

[457] Matti O. Jokinen. The effect of parameter passing and other implementation dependent mechanisms is undecidable. ACM SIGPLAN Notices, 17(9): 16–17, September 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1982:LOM

[458] Curtis R. Cook and R. R. Oldehoeft. A letter oriented minimal perfect hashing function. ACM SIGPLAN Notices, 17(9):18–27, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sumner:1982:MSP

[459] Roger T. Sumner and R. E. Gleaves. Modula-2 — a solution to Pascal’s problems. ACM SIGPLAN Notices, 17 (9):28–33, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1982:BPS

[460] Walter W. Wilson. Beyond PRO-LOG: software specification by grammar. ACM SIGPLAN Notices, 17(9): 34–43, September 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Piwowarski:1982:NLC

[461] Paul Piwowarski. A nesting level complexity measure. ACM SIGPLAN Notices, 17(9):44–50, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rentsch:1982:OOP

[462] Tim Rentsch. Object oriented programming. ACM SIGPLAN Notices, 17(9):51–57, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greiter:1982:RLC

[463] G. Greiter. Remarks on language concepts for specifying process synchronization. ACM SIGPLAN Notices, 17(9):58–61, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laski:1982:DFG

[464] Janusz Laski. On data flow guided program testing. ACM SIGPLAN Notices, 17(9):62–71, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marca:1982:RCU

[465] David Marca. A repetition construct for UNIX version 6. ACM SIGPLAN Notices, 17(9):72–75, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stotts:1982:CSCa

[466] Paul David Stotts, Jr. A comparative survey of concurrent programming languages. ACM SIGPLAN Notices, 17(9):76–87, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Identical to [471].

Stone:1982:PRH

[467] Roger G. Stone. Points recurring. the history of a railway problem. ACM SIGPLAN Notices, 17(9):88– 94, September 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowe:1982:IC

[468] M. K. Crowe. An incremental compiler. ACM SIGPLAN Notices, 17(10): 13–22, October 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rees:1982:AAA

[469] Michael J. Rees. Automatic assessment aids for Pascal programs. ACM SIG-PLAN Notices, 17(10):33–42, October 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1982:SNN

[470] B. J. MacLennan. A simple, natural notation for applicative languages. ACM SIGPLAN Notices, 17(10):43–49, October 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stotts:1982:CSCb

[471] Paul David Stotts, Jr. A comparative survey of concurrent programming languages. ACM SIGPLAN Notices, 17(10):50–61, October 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Identical to [466].

Meijer:1982:YYA

[472] H. Meijer and A. Nijholt. YABBERyet another bibliography: translator writing tools. ACM SIGPLAN Notices, 17(10):62–72, October 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aretz:1982:SDI

[473] F. E. J. Kruseman Aretz. Syntax diagrams for ISO Pascal Standard. ACM SIGPLAN Notices, 17(10):73–78, October 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gootherts:1982:CPL

[474] Paul D. Gootherts and James W. Davis. Common programming language ambiguity. ACM SIGPLAN Notices, 17(11):7–10, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pentzlin:1982:SCD

[475] Karl L. Pentzlin. Syntax of comments: discussion and a proposal. ACM SIG-PLAN Notices, 17(11):11–13, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karlsson:1982:NAS

[476] Kent Karlsson and Kent Petersson. Notes from: the Aspenäs Symposium of Functional Languages and Computer Architecture. ACM SIGPLAN Notices, 17(11):14–22, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duff:1982:CBS

[477] Steven G. Duff. The case for BUFFERS in SNOBOL4. ACM SIGPLAN Notices, 17(11):24–30, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Groves:1982:USE

[478] Lindsay J. Groves. Using simple English sentences to call procedures. ACM SIGPLAN Notices, 17(11):31– 38, November 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1982:CAP

[479] Eric B. Levy. The case against Pascal as a teaching tool. ACM SIGPLAN Notices, 17(11):39–41, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Payne:1982:BRC

[480] A. J. Payne. A basis for the rate of change in programs. ACM SIGPLAN Notices, 17(11):42–44, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1982:TOS

[481] Ellis Cohen. Text-oriented structure commands for structure editors. ACM SIGPLAN Notices, 17(11):45– 49, November 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Litvin:1982:PEP

[482] Yury Litvin. Parallel evolution programming language for data flow machines. ACM SIGPLAN Notices, 17 (11):50–58, November 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boddy:1982:SSA

[483] David E. Boddy. A style for structured APL. ACM SIGPLAN Notices, 17(12): 27–30, December 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cailliau:1982:HAG

[484] R. Cailliau. How to avoid getting schlonked by Pascal. ACM SIGPLAN Notices, 17(12):31–40, December 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corbett:1982:EAF

[485] Robert Paul Corbett. Enhanced arithmetic for Fortran. ACM SIGPLAN Notices, 17(12):41–48, December 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geurts:1982:OBP

[486] Leo Geurts. An overview of the B programming language or B without tears. ACM SIGPLAN Notices, 17(12): 49–58, December 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gladney:1982:NE

[487] H. M. Gladney. A note on enumerations. ACM SIGPLAN Notices, 17(12): 59–64, December 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harland:1982:SVC

[488] David M. Harland. Subtypes versus cell constancy with subrange constraints. ACM SIGPLAN Notices, 17(12):65– 69, December 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1982:VOP

[489] B. J. MacLennan. Values and objects in programming languages. ACM SIG-PLAN Notices, 17(12):70–79, December 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1982:EFA

[490] M. J. Wise. EPILOG=PROLOG+data flow: Arguments for combining PRO-LOG with a data driven mechanism. ACM SIGPLAN Notices, 17(12):80– 86, December 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1982:EPD

[491] Michael J. Wise. EPILOG = PROLOG + Data Flow: arguments for combining PROLOG with a data driven mechanism. ACM SIGPLAN Notices, 17(12): 80–86, December 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Celko:1983:DTR

[492] Joe Celko, John S. Davis, and John Mitchell. A demonstration of three requirements language systems. ACM SIGPLAN Notices, 18(1):9–14, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cox:1983:OOP

[493] Brad J. Cox. The object oriented precompiler: Programming Smalltalk 80 methods in C language. ACM SIG-PLAN Notices, 18(1):15–22, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culik:1983:FIP

[494] K. Culik. On formal and informal proofs for program correctness. ACM SIGPLAN Notices, 18(1):23–28, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olsen:1983:SAD

[495] Dan R. Olsen, Jr. A subset algorithm for deterministic FSMs within deterministic PDAs. ACM SIGPLAN Notices, 18(1):29–34, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shani:1983:SPE

[496] Uri Shani. Should program editors not abandon text oriented commands? ACM SIGPLAN Notices, 18(1):35–41, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1983:SFP

[497] David Spector. The simplest functional programming language. ACM SIG-PLAN Notices, 18(1):42–46, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See rebuttal [511].

Ambler:1983:CGC

[498] Arol Ambler and Robert Trawick. Chatin’s graph coloring algorithm as a method for assigning positions to Diana attributes. ACM SIGPLAN Notices, 18(2):37–38, February 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barz:1983:ISB

[499] Hans W. Barz. Implementing semaphores by binary semaphores. ACM SIGPLAN Notices, 18(2):39– 45, February 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1983:NMG

[500] Daniel M. Berry. A new methodology for generating test cases for a programming language compiler. ACM SIG-PLAN Notices, 18(2):46–56, February 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boswell:1983:SEP

[501] F. D. Boswell, M. J. Carmody, and T. R. Grove. A string extension for Pascal. ACM SIGPLAN Notices, 18(2): 57–61, February 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Finkel:1983:RTP

[502] Raphael Finkel. Run-time print values. ACM SIGPLAN Notices, 18(2): 62–64, February 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jackel:1983:CSF

[503] Manfred Jackel. Context sensitive formatting. ACM SIGPLAN Notices, 18 (2):65–68, February 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanOost:1983:DDO

[504] E. M. J. C. Van Oost. DUIF: a dataoriented flowchart environment. ACM SIGPLAN Notices, 18(2):69–75, February 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cantone:1983:CPS

[505] Giovanni Cantone, Aniello Cimitile, and Lucio Sansone. Complexity in program schemes: the characteristic polynomial. ACM SIGPLAN Notices, 18 (3):22–31, March 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

denHertog:1983:DSR

[506] E. H. den Hertog, H. J. C. Gerbscheid, and M. L. Kersten. DO-SELECT reconsidered. ACM SIG-PLAN Notices, 18(3):32–35, March 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1983:ORP

[507] B. J. MacLennan. Overview of relational programming. ACM SIG-PLAN Notices, 18(3):36–45, March 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vermaak:1983:PPE

[508] Desiré Vermaak. PROGEN: a programming environment. ACM SIG-PLAN Notices, 18(3):46–47, March 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1983:CPI

[509] Richard L. Wexelblat. Correction to page 27 of 1/83 issue. ACM SIG-PLAN Notices, 18(4):2, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyd:1983:MC

[510] Stowe Boyd. Modular C. ACM SIGPLAN Notices, 18(4):45–54, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fehr:1983:SFP

[511] Elfriede Fehr. The ‘simplest functional programming language’ is neither simple nor functional. ACM SIGPLAN Notices, 18(4):55–57, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [497].

Francis:1983:CDC

[512] Rhys S. Francis. Containment defines a class of recursive data structures. ACM SIGPLAN Notices, 18(4):58–64, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardy:1983:NSP

[513] Norman Hardy. Nested scopes in Pascal and Algol 68. ACM SIGPLAN Notices, 18(4):65, April 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laut:1983:ASP

[514] Alfred Laut. An algebraic specification of Pascal’s file type. ACM SIGPLAN Notices, 18(4):66–68, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

May:1983:O

[515] David May. OCCAM (language). ACM SIGPLAN Notices, 18(4):69–79, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pentzlin:1983:DSL

[516] Karl L. Pentzlin. Discussion of statement labels; proposal of a ’transfer/ junction’ construct as alternative to ’GOTO’. ACM SIGPLAN Notices, 18 (4):80–81, April 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

TeixeiradeQueiroz:1983:NCS

[517] F. Teixeira de Queiroz. The next control structure that I need. ACM SIGPLAN Notices, 18(4):82–85, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tompkins:1983:DTS

[518] Howard E. Tompkins. In defense of teaching structured COBOL as computer science (or, notes on being sage struck). ACM SIGPLAN Notices, 18 (4):86–94, April 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Antoy:1983:PSG

[519] Sergio Antoy and Giuliana Dettori. Is Pascal suitable for GKS? ACM SIG-PLAN Notices, 18(5):11–14, May 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boddy:1983:IDA

[520] David E. Boddy. Implementing data abstractions and monitors in UCSD PASCAL. ACM SIGPLAN Notices, 18 (5):15–24, May 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harle:1983:PSB

[521] James Harle. The proposed standard for Basic. ACM SIGPLAN Notices, 18 (5):25–40, May 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lakhotia:1983:IID

[522] Arun Lakhotia. An improvement over: “An improvement over deeply nested IF-THEN-ELSE control structure”. ACM SIGPLAN Notices, 18 (5):41–43, May 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [453].

Machanick:1983:NCW

[523] Philip Machanick. A note on C. S. Wetherell’s ‘Problems with the Ada reference grammar’. ACM SIGPLAN Notices, 18(5):44–45, May 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [329].

OKeefe:1983:PCL

[524] R. A. O’Keefe. PROLOG compared with LISP? ACM SIGPLAN Notices, 18(5):46–56, May 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scott:1983:MVR

[525] Michael Lee Scott. Messages vs. remote procedures is a false dichotomy. ACM SIGPLAN Notices, 18(5):57–62, May 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laff:1983:SOB

[526] Mark R. Laff and Brent Hailpern. SW 2: An object-based programming environment. ACM SIGPLAN Notices, 18(6):1–11, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Purtilo:1983:PES

[527] James Purtilo. Polylith: An environment to support management of tool interfaces. ACM SIGPLAN Notices, 18 (6):12–18, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ODonnell:1983:DBC

[528] John T. O’Donnell. Dialogues: a basis for constructing programming environments. ACM SIGPLAN Notices, 18 (6):19–27, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bahlke:1983:PPS

[529] Rolf Bahlke and Gregor Snelting. The PSG — Programming System Generator. ACM SIGPLAN Notices, 18 (6):28–33, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beshers:1983:MCA

[530] George McArthur Beshers and Roy Harold Campbell. Maintained and constructor attributes. ACM SIGPLAN Notices, 18(6):34–42, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Demers:1983:APM

[531] Alan Demers, Anne Rogers, and Frank Kenneth Zadeck. Attribute propagation by message passing. ACM SIGPLAN Notices, 18(6):43–59, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hall:1983:DSE

[532] Cordelia V. Hall and John T. O’Donnell. Debugging in a side effect free programming environment. ACM SIGPLAN Notices, 18(6):60–68, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Metayer:1983:MAP

[533] D. Le Metayer. Mechanical analysis of program complexity. ACM SIG-PLAN Notices, 18(6):69–73, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dybvig:1983:SE

[534] R. Kent Dybvig and Bruce T. Smith. A semantic editor. ACM SIGPLAN Notices, 18(6):74–82, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1983:LPP

[535] Mark B. Wells, Margaret A. Hug, and Rollo Silver. Libraries as programs preserved within compiler continuations. ACM SIGPLAN Notices, 18(6):83–91, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horwitz:1983:RAS

[536] Susan Horwitz and Tim Teitelbaum. Relations and attributes: a symbiotic basis for editing environments. ACM SIGPLAN Notices, 18(6):93–106, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1983:IIA

[537] Keith D. Cooper, Ken Kennedy, and Linda Torczon. The impact of interprocedural analysis and optimization on the design of a software development environment. ACM SIG-PLAN Notices, 18(6):107–116, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agnarsson:1983:TTP

[538] Snorri Agnarsson and M. S. Krishnamoorthy. Towards a theory of packages. ACM SIGPLAN Notices, 18(6): 117–130, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gries:1983:NNE

[539] David Gries and Jan Prins. A new notion of encapsulation. ACM SIG-PLAN Notices, 18(6):131–139, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caplinger:1983:SES

[540] Michael Caplinger. Structured Editor Support for Modularity and Data Abstraction. ACM SIGPLAN Notices, 18 (6):140–147, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moriconi:1983:PSG

[541] Mark Moriconi and Dwight F. Hare. PegaSys: a system for graphical explanation of program designs. ACM SIG-PLAN Notices, 18(6):148–160, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1983:APF

[542] Bent Bruun Kristensen, Ole Lehrmann Madsen, Birger Møller-Pedersen, and Kristen Nygaard. An algebra for program fragments. ACM SIGPLAN Notices, 18(6):161–170, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Francez:1983:ELP

[543] Nissim Francez, Shalom Goldenberg, Ron Y. Pinter, Michael Tiomkin, and Shalom Tsur. An environment for logic programming. ACM SIGPLAN Notices, 18(6):179–190, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Komorowski:1983:MIL

[544] Henryk Jan Komorowski and Shigeo Omori. A model and an implementation of a logic programming environment. ACM SIGPLAN Notices, 18(6): 191–198, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1983:SSE

[545] Bruce J. MacLennan. A simple software environment based on objects and relations. ACM SIGPLAN Notices, 18 (6):199–207, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fickas:1983:DIR

[546] Stephen Fickas. Design issues in a Rule-Based System. ACM SIG-PLAN Notices, 18(6):208–215, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sweet:1983:MPE

[547] Richard E. Sweet. The Mesa programming environment. ACM SIG-PLAN Notices, 18(6):216–229, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swinehart:1983:SC

[548] Daniel C. Swinehart, Polle T. Zellweger, and Robert B. Hagmann. The structure of Cedar. ACM SIGPLAN Notices, 18(6):230–244, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donahue:1983:IMC

[549] James Donahue. Integration mechanisms in Cedar. ACM SIGPLAN Notices, 18(6):245–251, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lampson:1983:OSD

[550] B. W. Lampson and E. E. Schmidt. Organizing software in a distributed environment. ACM SIGPLAN Notices, 18(6S):1–13, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Powell:1983:VAI

[551] M. L. Powell and M. A. Linton. Visual abstraction in an interactive programming environment. ACM SIG-PLAN Notices, 18(6S):14–21, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Honschopp:1983:CFL

[552] U. Honschopp, W.-M. Lippe, and F. Simon. Compiling functional languages for von Neumann machines. ACM SIGPLAN Notices, 18(6S):22–27, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farrow:1983:AGD

[553] R. Farrow. Attribute grammars and data-flow languages. ACM SIGPLAN Notices, 18(6S):28–40, June 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heinanen:1983:PCA

[554] J. Heinanen. A programmer controlled approach to data and control abstraction. ACM SIGPLAN Notices, 18 (6S):41–52, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weihl:1983:SIR

[555] W. Weihl and B. Liskov. Specification and implementation of resilient, atomic data types. ACM SIGPLAN Notices, 18(6S):53–64, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeBlanc:1983:ALM

[556] T. J. LeBlanc and R. P. Cook. An analysis of language models for highperformance communication in localarea networks. ACM SIGPLAN Notices, 18(6S):65–72, June 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strom:1983:NIL

[557] R. E. Strom and S. Yemini. NIL: an integrated language and system for distributed programming. ACM SIGPLAN Notices, 18(6S):73–82, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:1983:AMD

[558] J. W. Schmidt and M. Mall. Abstraction mechanisms for database programming. ACM SIGPLAN Notices, 18 (6S):83–93, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ridjanovic:1983:ATS

[559] D. Ridjanovic and M. L. Brodie. Action and transaction skeletons: high level language constructs for database transactions. ACM SIGPLAN Notices, 18 (6S):94–99, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1983:DPL

[560] M. Shaw, E. Borison, M. Horowitz, T. Lane, D. Nichols, and R. Pausch. Descartes: a programming-language approach to interactive display interfaces. ACM SIGPLAN Notices, 18(6S):

100–111, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olsen:1983:SDG

[561] D. R. Olsen, Jr. and E. P. Dempsey. Syntax directed graphical interaction. ACM SIGPLAN Notices, 18(6S):112– 117, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pilote:1983:PLF

[562] M. Pilote. A programming language framework for designing user interfaces. ACM SIGPLAN Notices, 18(6S): 118–136, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossher:1983:FPL

[563] H. L. Ossher and B. K. Reid. Fable: a programming-language solution to IC process automation problems. ACM SIGPLAN Notices, 18(6S): 137–148, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smoliar:1983:WNL

[564] S. W. Smoliar and D. Barstow. Who needs languages, and why do they need them? or no matter how high the level, it’s still programming. ACM SIG-PLAN Notices, 18(6S):149–157, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donner:1983:DOL

[565] M. D. Donner. The design of OWL, a language for walking. ACM SIG-PLAN Notices, 18(6S):158–165, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zippel:1983:CP

[566] R. Zippel. Capsules (programming). ACM SIGPLAN Notices, 18(6S):166– 169, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Novak:1983:DAG

[567] G. S. Novak, Jr. Data abstraction in GLISP. ACM SIGPLAN Notices, 18 (6S):170–177, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hood:1983:EAI

[568] Robert Hood. Efficient abstractions for the implementation of structured editors. ACM SIGPLAN Notices, 18(6S): 171–178, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albano:1983:THS

[569] A. Albano. Type hierarchies and semantic data models. ACM SIGPLAN Notices, 18(6S):178–186, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cormack:1983:ESS

[570] G. V. Cormack. Extensions to static scoping. ACM SIGPLAN Notices, 18 (6S):187–191, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lima:1983:DCF

[571] Isabel Gouveia Lima, Richard Hopkins, Lindsay Marshall, David Mundy, and Philip Treleaven. Decentralised control flow — BASed on unIX BASIX. ACM SIGPLAN Notices, 18(6S):192– 201, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shultis:1983:FS

[572] J. Shultis. A functional shell. ACM SIGPLAN Notices, 18(6S):202–211, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1983:HLP

[573] C. W. Fraser and D. R. Hanson. A high-level programming and command language. ACM SIGPLAN Notices, 18 (6S):212–219, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calmet:1983:PHL

[574] J. Calmet and R.-M. di Scala. Pascal as host language of computer algebra systems. ACM SIGPLAN Notices, 18(7):15–24, July 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldmann:1983:AG

[575] Harry Feldmann. Acceptive grammars. ACM SIGPLAN Notices, 18(7):25–44, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greiter:1983:CTD

[576] Gebh Greiter. Comparing two definitions for the notion ‘data type’. ACM SIGPLAN Notices, 18(7):45–49, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guntheroth:1983:NAB

[577] Kurt Guntheroth. The new ANSI BA-SIC standard. ACM SIGPLAN Notices, 18(7):50–59, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gutfreund:1983:TOV

[578] Steven Gutfreund. Tracing offensive values from exceptions. ACM SIG-PLAN Notices, 18(7):60–61, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nilges:1983:BST

[579] Edward G. Nilges. On bugs and swarms: towards a sociobiology of programming? ACM SIGPLAN Notices, 18(7):62–64, July 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Takeuchi:1983:THM

[580] Ikuo Takeuchi, Hiroshi Okuno, and Nobuyasu Ohsato. TAO: a harmonic mean of Lisp, Prolog and Smalltalk. ACM SIGPLAN Notices, 18(7):65–74, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thompson:1983:UAF

[581] John R. Thompson. The use and abuse of formal proofs. ACM SIG-PLAN Notices, 18(7):75–79, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gramlich:1983:DMS

[582] Wayne C. Gramlich. Debugging methodology: session summary. ACM SIGPLAN Notices, 18(8):1–3, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gramlich:1983:SSD

[583] Wayne C. Gramlich. Session summary — debugging methodology. ACM SIGPLAN Notices, 18(8):1–3, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamlet:1983:DLS

[584] Dick Hamlet. Debugging ‘Level’: stepwise debugging. ACM SIGPLAN Notices, 18(8):4–8, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seidner:1983:IDR

[585] Rich Seidner and Nick Tindall. Interactive debug requirements. ACM SIGPLAN Notices, 18(8):9–22, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Linton:1983:KBD

[586] Mark A. Linton, Robert Balzer, R. J. Cunningham, Cordell Green, Robert Sedlmeyer, Elliott Solloway, William Swartout, and Pamela Zave. Knowledge-based debugging: session summary. ACM SIGPLAN Notices, 18 (8):23–24, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sedlmeyer:1983:KBF

[587] Robert L. Sedlmeyer, William B. Thompson, and Paul E. Johnson. Knowledge-based fault localization in debugging: preliminary draft. ACM SIGPLAN Notices, 18(8):25–31, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bates:1983:RDD

[588] Peter Bates. Requirements/design debugging: session summary. ACM SIG-PLAN Notices, 18(8):32–33, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruegge:1983:GPE

[589] Bernd Bruegge and Peter Hibbard. Generalized path expressions: a high level debugging mechanism. ACM SIG-PLAN Notices, 18(8):34–44, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1983:AES

[590] Lori A. Clarke and Debra J. Richardson. The application of error-sensitive testing strategies to debugging. ACM SIGPLAN Notices, 18(8):45–52, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jard:1983:ATS

[591] C. Jard and G. von Bochmann. An approach to testing specifications. ACM SIGPLAN Notices, 18(8):53–59, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1983:IE

[592] Insup Lee. Integrated environments. ACM SIGPLAN Notices, 18(8):60–62, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kishimoto:1983:EDL

[593] Zen Kishimoto. An experimental debugger in a limited programming environment. ACM SIGPLAN Notices, 18 (8):63–66, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Powell:1983:DMD

[594] Michael L. Powell and Mark A. Linton. A database model of debugging. ACM SIGPLAN Notices, 18(8):67–70, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mikelsons:1983:IPE

[595] Martin Mikelsons. Interactive program execution in Lispedit. ACM SIG-PLAN Notices, 18(8):71–80, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mullerburg:1983:RDW

[596] Monika A. F. Müllerburg. The role of debugging within software engineering environments. ACM SIGPLAN Notices, 18(8):81–90, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wertz:1983:ILP

[597] Harald Wertz. An integrated Lisp programming environment. ACM SIG-PLAN Notices, 18(8):91–95, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gross:1983:DDS

[598] Thomas Gross. Distributed debugging: session summary. ACM SIG-PLAN Notices, 18(8):96–97, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gross:1983:SSD

[599] Thomas Gross. Session summary — distributed debugging. ACM SIG-PLAN Notices, 18(8):96–97, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baiardi:1983:DDC

[600] F. Baiardi, N. De Francesco, E. Matteoli, S. Stefanini, and G. Vaglini. Development of a debugger for a concurrent language. ACM SIGPLAN Notices, 18 (8):98–106, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bates:1983:AHL

[601] Peter Bates and Jack C. Wileden. An approach to high-level debugging of distributed systems: preliminary draft. ACM SIGPLAN Notices, 18(8):107– 111, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weber:1983:IDC

[602] Janice Cynthia Weber. Interactive debugging of concurrent programs. ACM SIGPLAN Notices, 18(8):112–113, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weber:1983:IIS

[603] Janice Cynthia Weber. Implementation issues: session summary. ACM SIG-PLAN Notices, 18(8):114–116, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kenniston:1983:III

[604] Michael S. Kenniston. Implementation issues II. ACM SIGPLAN Notices, 18(8):117–120, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barra:1983:SPT

[605] Knut Barra and Hans Petter Dahle. SIMOB: a portable toolbox for observation of Simula executions. ACM SIG-PLAN Notices, 18(8):121–122, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrante:1983:HLL

[606] Jeanne Ferrante. High level language debugging with a compiler. ACM SIG-PLAN Notices, 18(8):123–129, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fritzson:1983:SAA

[607] Peter Fritzson. A systematic approach to advanced debugging through incremental compilation. ACM SIGPLAN Notices, 18(8):130–139, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gentleman:1983:HAH

[608] W. Morven Gentleman and Henry Hoeksma. Hardware assisted high level debugging: preliminary draft. ACM SIGPLAN Notices, 18(8):140–144, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1983:RTM

[609] Charles R. Hill. A real-time microprocessor debugging technique. ACM SIG-PLAN Notices, 18(8):145–148, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1983:IIS

[610] John D. Johnson and Gary W. Kenney. Implementation issues for a source level symbolic debugger. ACM SIG-PLAN Notices, 18(8):149–151, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ottenstein:1983:HLD

[611] Karl J. Ottenstein and Linda M. Ottenstein. High-level debugging assistance via optimizing compiler technology. ACM SIGPLAN Notices, 18(8): 152–154, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tischler:1983:SAP

[612] Ron Tischler, Robin Schaufler, and Charlotte Payne. Static analysis of programs as an aid to debugging. ACM SIGPLAN Notices, 18(8):155–158, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zellweger:1983:IHL

[613] Polle T. Zellweger. An interactive high-level debugger for control-flow optimized programs. ACM SIGPLAN Notices, 18(8):159–172, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beander:1983:VDI

[614] Bert Beander. VAX DEBUG: an interactive, symbolic, multilingual debugger. ACM SIGPLAN Notices, 18(8): 173–179, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cardell:1983:MDS

[615] James R. Cardell. Multilingual debugging with the SWAT high-level debugger. ACM SIGPLAN Notices, 18(8): 180–189, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargill:1983:BD

[616] T. A. Cargill. The Blit debugger (preliminary draft). ACM SIGPLAN Notices, 18(8):190–200, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargill:1983:BDP

[617] Thomas A. Cargill. The Blit debugger: preliminary draft. ACM SIG-PLAN Notices, 18(8):190–200, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1983:DDM

[618] Robert P. Cook and Insup Lee. DY-MOS: a dynamic modification system. ACM SIGPLAN Notices, 18(8):201– 202, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walter:1983:DUD

[619] Carol K. Walter. DELTA: the Universal Debugger for CP-6. ACM SIG-PLAN Notices, 18(8):203–205, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1983:SEC

[620] Mark Scott Johnson. Summary of evaluation and comments. ACM SIGPLAN Notices, 18(8):206, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casimir:1983:CV

[621] Rommert J. Casimir. Collections of variables. ACM SIGPLAN Notices, 18(9):26–30, September 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1983:CST

[622] Darrel C. Ince. A compatibility software tool for use with separately compiled languages. ACM SIGPLAN Notices, 18(9):31–34, September 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jalili:1983:GLT

[623] Fahimeh Jalili. A general lineartime evaluator for attribute grammars. ACM SIGPLAN Notices, 18(9): 35–44, September 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meekings:1983:SAP

[624] B. A. E. Meekings. Style analysis of Pascal programs. ACM SIGPLAN Notices, 18(9):45–54, September 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mulders:1983:SOO

[625] Huub Mulders. Some observations on the in- and output in high-level languages. ACM SIGPLAN Notices, 18 (9):55–58, September 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1983:LFS

[626] David Spector. Language features to support reusability. ACM SIGPLAN Notices, 18(9):59–61, September 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tamine:1983:UTL

[627] Jacques J. Tamine. On the use of tree-like structures to simplify measures of complexity. ACM SIGPLAN Notices, 18(9):62–69, September 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1983:PND

[628] C. K. Yuen. The programmer as navigator: a discourse on program structure. ACM SIGPLAN Notices, 18(9): 70–78, September 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhaskar:1983:HOO

[629] K. S. Bhaskar. How object-oriented is your system? ACM SIGPLAN Notices, 18(10):8–11, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanKatwijk:1983:PYP

[630] J. van Katwijk. A preprocessor for YACC or a poor man’s approach to parsing attributed grammars. ACM SIGPLAN Notices, 18(10):12–15, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:1983:FDM

[631] Edward Reid. Fighting the disease: more comments after Dijkstra and Tompkins. ACM SIGPLAN Notices, 18(10):16–21, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:1983:MCA

[632] E. Reid. More comments after Dijkstra and Tompkins. ACM SIGPLAN Notices, 18(10):16–21, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenbloom:1983:DNI

[633] Michael H. Rosenbloom. Deeply nested IF-THEN-ELSE’s. ACM SIGPLAN Notices, 18(10):22–24, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1983:LPM

[634] David Spector. Lexing and parsing Modula-2. ACM SIGPLAN Notices, 18(10):25–32, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1983:IIT

[635] David Taylor. An improvement to the

‘if-then-else’ and ‘select’ clauses,

and some comments on Arun Lakhotias’ an improvement over . . .. ACM SIGPLAN Notices, 18(10):33–37, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tennent:1983:ACA

[636] R. D. Tennent. An alternative to conformant-array parameters in Pascal. ACM SIGPLAN Notices, 18(10): 38–43, October 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weinman:1983:NBC

[637] Joseph B. Weinman, Jr. Nestable bracketed comments. ACM SIGPLAN Notices, 18(10):44–47, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chow:1983:ILC

[638] F. C. Chow and M. Ganapathi. Intermediate languages in compiler construction–A bibliography. ACM SIGPLAN Notices, 18(11):21–23, November 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burns:1983:EIO

[639] A. Burns. Enhanced input/output on Pascal. ACM SIGPLAN Notices, 18(11):24–33, November 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ceri:1983:RDB

[640] S. Ceri and S. Crespi-Reghizzi. Relational data bases in the design of program construction systems. ACM SIG-PLAN Notices, 18(11):34–44, November 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gait:1983:CLS

[641] Jason Gait. Command level software tools. ACM SIGPLAN Notices, 18(11): 45–53, November 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woitok:1983:ALD

[642] Rainer Woitok. Abstracting linked data structures using incremental records. ACM SIGPLAN Notices, 18 (11):54–63, November 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brown:1983:ARD

[643] E. J. Brown. On the application of Rothon diagrams to data abstraction. ACM SIGPLAN Notices, 18(12):17– 24, December 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Comer:1983:ENL

[644] James R. Comer. An experimental natural-language processor for generating data type specifications. ACM SIG-PLAN Notices, 18(12):25–33, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Despeyroux:1983:ASP

[645] Joëlle Despeyroux. An algebraic specification of a Pascal compiler. ACM SIGPLAN Notices, 18(12):34–48, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanders:1983:CP

[646] Harley Flanders. Coroutines in Pascal. ACM SIGPLAN Notices, 18(12): 49–56, December 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Evangelist:1983:RAC

[647] W. M. Evangelist. Relationships among computational, software, and intuitive complexity. ACM SIGPLAN Notices, 18(12):57–59, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grosch:1983:TDG

[648] Josef Grosch. Type derivation graphs: a way to visualize the type building possibilities of programming languages. ACM SIGPLAN Notices, 18(12):60– 68, December 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1983:WTA

[649] Leon S. Levy. A walk through AWK. ACM SIGPLAN Notices, 18(12):69– 85, December 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1983:AII

[650] B. J. MacLennan. Abstraction in the Intel iAPX-432 prototype systems implementation language. ACM SIG-PLAN Notices, 18(12):86–95, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merrill:1983:PPC

[651] Gary Merrill. Proofs, program correctness, and software engineering. ACM SIGPLAN Notices, 18(12):96–105, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newman:1983:ITE

[652] P. S. Newman. IF-THEN-ELSE, again. ACM SIGPLAN Notices, 18(12):106– 111, December 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olsson:1983:MUL

[653] Olle Olsson. The memory usage of a LISP system: the Belady lifetime function. ACM SIGPLAN Notices, 18(12):112–119, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivasan:1983:NCS

[654] B. Srinivasan. Is nested control structure more complex? ACM SIG-PLAN Notices, 18(12):120–121, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wharton:1983:NTP

[655] R. Michael Wharton. A note on types and prototypes. ACM SIGPLAN Notices, 18(12):122–126, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albizuri-Romero:1984:GAP

[656] Miren Begona Albizuri-Romero. A graphical abstract programming language. ACM SIGPLAN Notices, 19(1): 14–23, January 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amit:1984:DSI

[657] Neta Amit. A different solution for improving the readability of deeply nested IF-THEN-ELSE control structure. ACM SIGPLAN Notices, 19(1): 24–30, January 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1984:DAD

[658] A. Toni Cohen. Data abstraction, data encapsulation and object-oriented programming. ACM SIGPLAN Notices, 19(1):31–35, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Craeynest:1984:CES

[659] Dirk Craeynest, Geert Vansteenkiste, and Johan Lewi. Construction of an ELL(1) syntax analyser for Ada with the compiler-generator LILA. ACM SIGPLAN Notices, 19(1):36–45, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunter:1984:SWP

[660] R. B. Hunter. A system for writing Pascal programs interactively. ACM SIG-PLAN Notices, 19(1):46–56, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loka:1984:NPPa

[661] Raghavendra Rao Loka. A note on parallel parsing. ACM SIGPLAN Notices, 19(1):57–59, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Majocchi:1984:TOE

[662] Luca Majocchi, Luca Rodda, and Giacomo R. Sechi. A technique to obtain effective locality of computation in FORTRAN IV. ACM SIGPLAN Notices, 19(1):60–62, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruane:1984:ADT

[663] Lawrence M. Ruane. Abstract data types in assembly language programming. ACM SIGPLAN Notices, 19(1): 63–67, January 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1984:PPC

[664] David Taylor. A portable pseudo-code for Pascal-like languages. ACM SIG-PLAN Notices, 19(1):68–77, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanKatwijk:1984:DMR

[665] J. van Katwijk and J. van Someren. The doublet model: run-time model and implementation of Ada types. ACM SIGPLAN Notices, 19(1):78–92, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1984:FCP

[666] C. K. Yuen. Further comments on the premature loop exit problem. ACM SIGPLAN Notices, 19(1):93–94, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1984:SCM

[667] W. Harrison. Software complexity metrics: a bibliography and category index. ACM SIGPLAN Notices, 19(2): 17–27, February 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albizuri-Romero:1984:GGS

[668] Miren Begona Albizuri-Romero. GRASE: a graphical syntax-directed editor for structured programming. ACM SIG-PLAN Notices, 19(2):28–37, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boddy:1984:DMP

[669] David E. Boddy. On the design of monitors with priority conditions. ACM SIGPLAN Notices, 19(2):38–46, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooke:1984:PLM

[670] D. J. Cooke and M. H. J. Al-Noufaly. The programming language MINO and its formal definition. ACM SIGPLAN Notices, 19(2):47–57, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldmann:1984:AGN

[671] Harry Feldmann. An acceptive grammar for the natural language English. ACM SIGPLAN Notices, 19(2): 58–67, February 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1984:PPF

[672] D. C. Ince. The provision of procedural and functional interfaces for the maintenance of program design language and program language notations. ACM SIGPLAN Notices, 19(2):68–74, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavens:1984:PSV

[673] Gary T. Leavens. Prettyprinting styles for various languages. ACM SIGPLAN Notices, 19(2):75–79, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mayer:1984:THR

[674] Herbert Mayer and Don Perkins. Towers of Hanoi revisited. A nonrecursive surprise. ACM SIGPLAN Notices, 19 (2):80–84, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deAPrice:1984:DDV

[675] Ana M. de A. Price. Defining dynamic variables and abstract data types in Pascal. ACM SIGPLAN Notices, 19(2): 85–91, February 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steensgaard-Madsen:1984:DPL

[676] J. Steensgaard-Madsen and L. M. Olsen. Definition of the programming language MODEF. ACM SIGPLAN Notices, 19(2):92–110, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agrawal:1984:DAP

[677] Mukul Babu Agrawal and Vijay Kumar Garg. Dimensional analysis in Pascal. ACM SIGPLAN Notices, 19 (3):7–11, March 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyd:1984:FBG

[678] Stowe Boyd. Free and bound generics: two techniques for abstract data types in modular C. ACM SIG-PLAN Notices, 19(3):12–20, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bulman:1984:WPR

[679] Edward Bulman. What is programming really all about? ACM SIG-PLAN Notices, 19(3):21–23, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callaghan:1984:SIM

[680] John L. Callaghan. A software implemented memory manager. ACM SIGPLAN Notices, 19(3):24–31, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Faust:1984:RTD

[681] Don Faust. Run-time declaration elaboration. ACM SIGPLAN Notices, 19 (3):32–38, March 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garg:1984:SOH

[682] Vijay Kumar Garg. Screen-oriented highlevel debugger (SHD) for Pascal. ACM SIGPLAN Notices, 19(3):39–41, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Groening:1984:CAR

[683] Klaus Groening. “Combined actions to reduce LR-parser tables”: (experimental results). ACM SIGPLAN Notices, 19(3):42–45, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoyer:1984:SRR

[684] Wolfgang Hoyer. Some remarks on records with variants. ACM SIG-PLAN Notices, 19(3):46–50, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lau:1984:YAH

[685] Francis C. M. Lau and Atul Asthana. Yet another history mechanism for command interpreters. ACM SIG-PLAN Notices, 19(3):51–56, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mamone:1984:EFC

[686] Salvatore Mamone. Error free code. is it attainable? ACM SIGPLAN Notices, 19(3):57–60, March 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pugh:1984:ASS

[687] John R. Pugh. Actors: the stage is set. ACM SIGPLAN Notices, 19(3):61–65, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiener:1984:GSM

[688] Richard S. Wiener. A generic sorting module in Modula-2. ACM SIG-PLAN Notices, 19(3):66–72, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wingerter:1984:NDC

[689] Richard Wingerter. A note on determining the complexity of algorithms. ACM SIGPLAN Notices, 19(3):73–78, March 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benepe:1984:DSC

[690] David B. Benepe, Jr. In defense of simplicity of comment syntax (in programming languages). ACM SIGPLAN Notices, 19(4):32–33, April 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McMullen:1984:SDT

[691] Bill McMullen. Structured decision tables. ACM SIGPLAN Notices, 19 (4):34–43, April 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thacher:1984:EPV

[692] Henry C. Thacher, Jr. On the elimination of pointer variables and dynamic allocation in higher level languages. ACM SIGPLAN Notices, 19 (4):44–46, April 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zelkowitz:1984:SCE

[693] Marvin V. Zelkowitz. A small contribution to editing with a syntax directed editor. ACM SIGPLAN Notices, 19(5):1–6, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horgan:1984:TIL

[694] J. R. Horgan and D. J. Moore. Techniques for improving language-based editors. ACM SIGPLAN Notices, 19 (5):7–14, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giacalone:1984:III

[695] Alessandro Giacalone, Martin C. Rinard, and Thomas W. Doeppner, Jr. IDEOSY: An Ideographic and Interactive Program Description System. ACM SIGPLAN Notices, 19(5):15–20, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:1984:PLB

[696] C. N. Fischer, Gregory F. Johnson, Jon Mauney, Anil Pal, and Daniel L. Stock. The Poe language-based editor project. ACM SIGPLAN Notices, 19(5):21–29, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiss:1984:GPD

[697] Steven P. Reiss. Graphical program development with PECAN program development systems. ACM SIG-PLAN Notices, 19(5):30–41, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reps:1984:SG

[698] Thomas Reps and Tim Teitelbaum. The Synthesizer Generator. ACM SIG-PLAN Notices, 19(5):42–48, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Delisle:1984:VPE

[699] Norman M. Delisle, David E. Menicosy, and Mayer D. Schwartz. Viewing a programming environment as a single tool. ACM SIGPLAN Notices, 19(5):49–56, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Standish:1984:APA

[700] Thomas A. Standish and Richard N. Taylor. Arcturus: a prototype advanced Ada programming environment. ACM SIGPLAN Notices, 19 (5):57–64, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garlan:1984:GIP

[701] David B. Garlan and Philip L. Miller. GNOME: An introductory programming environment based on a family of structure editors. ACM SIG-PLAN Notices, 19(5):65–72, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Campbell:1984:SPS

[702] Roy H. Campbell and Peter A. Kirslis. The SAGA project: a system for software development. ACM SIG-PLAN Notices, 19(5):73–80, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chesi:1984:IIS

[703] M. Chesi, E. Dameri, M. P. Franceschi, M. G. Gatti, and C. Simonelli. ISDE: An Interactive Software Development Environment. ACM SIGPLAN Notices, 19(5):81–88, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minsky:1984:DSE

[704] N. Minsky and A. Borgida. The Darwin software-evolution environment. ACM SIGPLAN Notices, 19(5):89–95, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeTreville:1984:PIS

[705] John DeTreville. Phoan: An intelligent system for distributed control synthesis. ACM SIGPLAN Notices, 19 (5):96–103, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leblang:1984:CAS

[706] David B. Leblang and Robert P. Chase, Jr. Computer-aided Software Engineering in a distributed workstation environment. ACM SIGPLAN Notices, 19 (5):104–112, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fritzson:1984:PED

[707] Peter Fritzson. Preliminary experience from the DICE system a distributed incremental compiling environment. ACM SIGPLAN Notices, 19 (5):113–123, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snodgrass:1984:MSD

[708] Richard Snodgrass. Monitoring in a software development environment: a relational approach. ACM SIG-PLAN Notices, 19(5):124–131, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Linton:1984:IRV

[709] Mark A. Linton. Implementing relational views of programs. ACM SIG-PLAN Notices, 19(5):132–140, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donzeau-Gouge:1984:DSM

[710] V. Donzeau-Gouge, G. Kahn, B. Lang, and B. Mélèse. Document structure and modularity in Mentor. ACM SIG-PLAN Notices, 19(5):141–148, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Estublier:1984:PEC

[711] J. Estublier, S. Ghoul, and S. Krakowiak. Preliminary experience with a configuration control system for modular programs. ACM SIGPLAN Notices, 19 (5):149–156, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungar:1984:GSN

[712] David Ungar. Generation Scavenging: a non-disruptive high performance storage reclamation algorithm. ACM SIGPLAN Notices, 19(5):157– 167, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vitter:1984:UNF

[713] Jeffrey Scott Vitter. US&R: a new framework for redoing (extended abstract). ACM SIGPLAN Notices, 19 (5):168–176, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ottenstein:1984:PDG

[714] Karl J. Ottenstein and Linda M. Ottenstein. The program dependence graph in a software development environment. ACM SIGPLAN Notices, 19 (5):177–184, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:1984:PET

[715] Richard P. Gabriel and Martin E. Frost. A programming environment for a timeshared system. ACM SIG-PLAN Notices, 19(5):185–192, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soloway:1984:CBM

[716] Elliot Soloway. A cognitively-based methodology for designing languages/ environments/methodologies. ACM SIGPLAN Notices, 19(5):193–196, May 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karr:1984:CGC

[717] Michael Karr. Code generation by coagulation. ACM SIGPLAN Notices, 19(6):1–12, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aigrain:1984:EGG

[718] Philippe Aigrain, Susan L. Graham, Robert R. Henry, Marshall Kirk McKusick, and Eduardo Pelegri-Llopart. Experience with a Graham-Glanville style code generator. ACM SIGPLAN Notices, 19(6):13–24, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:1984:UDP

[719] Thomas W. Christopher, Philip J. Hatcher, and Ronald C. Kukuk. Using dynamic programming to generate optimized code in a Graham-Glanville style code generator. ACM SIG-PLAN Notices, 19(6):25–36, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:1984:PPS

[720] Joseph A. Fisher, John R. Ellis, John C. Ruttenberg, and Alexandru Nicolau. Parallel processing: a smart compiler and a dumb machine. ACM SIGPLAN Notices, 19(6):37–47, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Touzeau:1984:FCF

[721] Roy F. Touzeau. A Fortran compiler for the FPS-164 scientific computer. ACM SIGPLAN Notices, 19(6):48–57, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnsson:1984:ECL

[722] Thomas Johnsson. Efficient compilation of lazy evaluation. ACM SIG-PLAN Notices, 19(6):58–69, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farrow:1984:SPE

[723] Rodney Farrow. Sub-protocol-evaluators for attribute grammars. ACM SIGPLAN Notices, 19(6):70–80, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jourdan:1984:SNC

[724] Martin Jourdan. Strongly non-circular attribute grammars and their recursive evaluation. ACM SIGPLAN Notices, 19(6):81–93, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pleban:1984:CPU

[725] Uwe F. Pleban. Compiler prototyping using formal semantics. ACM SIGPLAN Notices, 19(6):94–105, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kessler:1984:PAD

[726] Robert R. Kessler. Peep: an architectural description driven peephole optimizer. ACM SIGPLAN Notices, 19 (6):106–110, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davidson:1984:AGP

[727] Jack W. Davidson and Christopher W. Fraser. Automatic generation of peephole optimizations. ACM SIGPLAN Notices, 19(6):111–116, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1984:AGP

[728] C. W. Fraser. Automatic generation of peephole optimizations. ACM SIG-PLAN Notices, 19(6):111–116, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1984:ACA

[729] Christopher W. Fraser, Eugene W. Myers, and Alan L. Wendt. Analyzing and compressing assembly code. ACM SIG-PLAN Notices, 19(6):117–121, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwartz:1984:ICM

[730] Mayer D. Schwartz, Norman M. Delisle, and Vimal S. Begwani. Incremental compilation in Magpie. ACM SIGPLAN Notices, 19(6):122– 131, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zadeck:1984:IDF

[731] Frank Kenneth Zadeck. Incremental data flow analysis in a structured program editor. ACM SIG-PLAN Notices, 19(6):132–143, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiss:1984:AIC

[732] Steven P. Reiss. An approach to incremental compilation. ACM SIG-PLAN Notices, 19(6):144–156, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganzinger:1984:ACG

[733] Harald Ganzinger and Robert Giegerich. Attribute coupled grammars. ACM SIGPLAN Notices, 19(6):157–170, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jullig:1984:RRP

[734] Richard K. Jullig and Frank DeRemer. Regular right-part attribute grammars. ACM SIGPLAN Notices, 19 (6):171–178, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koskimies:1984:SLO

[735] Kai Koskimies. A specification language for one-pass semantic analysis. ACM SIGPLAN Notices, 19(6): 179–189, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1984:TTT

[736] S. E. Keller, J. A. Perkins, T. F. Payton, and S. P. Mardinly. Tree transformation techniques and experiences. ACM SIGPLAN Notices, 19 (6):190–201, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:1984:MLC

[737] Uwe Schmidt and Reinhard Völler. A multi-language compiler system with automatically generated code generators. ACM SIGPLAN Notices, 19(6): 202–212, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wand:1984:SPS

[738] Mitchell Wand. A semantic prototyping system. ACM SIGPLAN Notices, 19(6):213–221, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chow:1984:RAP

[739] Frederick Chow and John Hennessy. Register allocation by priority-based coloring. ACM SIGPLAN Notices, 19 (6):222–232, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:1984:ALI

[740] John R. Allen and Ken Kennedy. Automatic loop interchange. ACM SIG-PLAN Notices, 19(6):233–246, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1984:ECF

[741] Keith D. Cooper and Ken Kennedy. Efficient computation of flow insensitive interprocedural summary information. ACM SIGPLAN Notices, 19(6): 247–258, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blower:1984:EIV

[742] M. I. Blower. An efficient implementation of visibility in Ada. ACM SIG-PLAN Notices, 19(6):259–265, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1984:IRV

[743] M. Donald MacLaren. Inline routines in VAXELN Pascal. ACM SIG-PLAN Notices, 19(6):266–275, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bray:1984:SCA

[744] Gary Bray. Sharing code among instances of Ada generics. ACM SIG-PLAN Notices, 19(6):276–284, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1984:TCS

[745] Michael R. Levy. Type checking, separate compilation and reusability. ACM SIGPLAN Notices, 19(6):285– 289, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baiardi:1984:SCI

[746] F. Baiardi, L. Ricci, and M. Vanneschi. Static checking of interprocess communication in ECSP. ACM SIG-PLAN Notices, 19(6):290–299, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robbins:1984:EHC

[747] David C. Robbins. Engineering a highcapacity Pascal compiler for high performance. ACM SIGPLAN Notices, 19 (6):300–309, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Powell:1984:POC

[748] Michael L. Powell. A portable optimizing compiler for Modula-2. ACM SIG-PLAN Notices, 19(6):310–318, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliot:1984:EEN

[749] I. B. Elliot. The EPN and ESN notations. ACM SIGPLAN Notices, 19 (7):9–17, July 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1984:EEN

[750] Irving B. Elliott. The EPN and ESN notations. ACM SIGPLAN Notices, 19(7):9–17, July 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kao:1984:DFA

[751] H. Kao and T. Y. Chen. Data flow analysis for Cobol. ACM SIG-PLAN Notices, 19(7):18–21, July 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loka:1984:NPPb

[752] Raghavendra Rao Loka. A note on parallel parsing. ACM SIGPLAN Notices, 19(7):22–24, July 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ottenstein:1984:IPR

[753] Karl J. Ottenstein. Intermediate program representations in compiler construction: a supplemental bibliography. ACM SIGPLAN Notices, 19(7):25–27, July 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1984:FCE

[754] Anonymous. Foreword to the candidate extension library. ACM SIG-PLAN Notices, 19(7):28–44, July 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1984:PFC

[755] Anonymous. Pascal: forward to the candidate extension library. ACM SIG-PLAN Notices, 19(7):28–44, July 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1984:FWP

[756] Anonymous. Foreword to work in progress. ACM SIGPLAN Notices, 19 (7):45–64, July 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1984:PFW

[757] Anonymous. Pascal: foreword to work in progress. ACM SIGPLAN Notices, 19(7):45–64, July 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winkler:1984:SIIa

[758] J. F. H. Winkler. Some improvements of ISO-Pascal. ACM SIG-PLAN Notices, 19(7):65–78, July 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavens:1984:BDT

[759] G. T. Leavens. Bibliography on data types. ACM SIGPLAN Notices, 19(8): 41–50, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rowe:1984:PLI

[760] Lawrence A. Rowe, Peter Deutsch, Stu Feldman, Butler Lampson, Barbara Liskov, and Terry Winograd. Programming language issues for the 1980’s: SIGPLAN ’83: Symposium on Programming Languages Issues in Software Systems. ACM SIGPLAN Notices, 19(8):51–61, August 1984.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baldwin:1984:BDG

[761] Roger R. Baldwin. The backwarddirected GO TO in FORTRAN. ACM SIGPLAN Notices, 19(8):62–64, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berghel:1984:MPS

[762] H. L. Berghel and D. L. Sallach. Measurements of program similarity in identical task environments. ACM SIGPLAN Notices, 19(8):65–76, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brownsmith:1984:OLP

[763] Joseph D. Brownsmith and Leslie H. Oliver. Optimizing loops in programs compiled with the IBM PL/I optimizing compiler. ACM SIGPLAN Notices, 19(8):77–84, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Covington:1984:PDR

[764] Michael A. Covington. A pedagogical disadvantage of repeat and while. ACM SIGPLAN Notices, 19(8):85–86, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Franklin:1984:SIS

[765] Wm. Randolph Franklin. A simpler iterative solution to the Towers of Hanoi problem. ACM SIGPLAN Notices, 19 (8):87–88, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1984:MIL

[766] D. C. Ince. Module interconnection languages and Prolog. ACM SIG-PLAN Notices, 19(8):89–93, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noyelle:1984:SHF

[767] Yves Noyelle. String-handling facilities of the LSE language. ACM SIG-PLAN Notices, 19(8):94–103, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smolka:1984:RIS

[768] Scott A. Smolka. Report on the 6th International Symposium on Programming. ACM SIGPLAN Notices, 19(8): 104–105, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanenbaum:1984:DAT

[769] Andrew S. Tanenbaum, Ed Keizer, and Hans van Staveren. Does anybody out there want to write HALF of a compiler? ACM SIGPLAN Notices, 19(8): 106–108, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woodman:1984:PDL

[770] Mark Woodman. A Program Design Language for software engineering. ACM SIGPLAN Notices, 19(8): 109–118, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zemrowski:1984:DBA

[771] Kenneth M. Zemrowski. Differences between ANS and ISO standards for Pascal: ambiguities clarified. ACM SIG-PLAN Notices, 19(8):119–126, August 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fay:1984:EUI

[772] D. Q. M. Fay. Experiences using Inmos proto-OCCAM (TM). ACM SIG-PLAN Notices, 19(9):5–11, September 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furtado:1984:GSC

[773] A. L. Furtado. Generalized set comparison. ACM SIGPLAN Notices, 19(9): 12–15, September 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mayer:1984:EPP

[774] Herbert Mayer. Experiences porting Pascal source from a micro computer to a Vax. ACM SIGPLAN Notices, 19(9): 16–20, September 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milewski:1984:LIA

[775] Jaroslaw Milewski. Loslan implementation of the AMPL message-passing system. ACM SIGPLAN Notices, 19(9): 21–29, September 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pramanik:1984:IKM

[776] Sakti Pramanik and Bernhard Weinberg. The implementation kit with monitors. ACM SIGPLAN Notices, 19 (9):30–33, September 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stadel:1984:ANA

[777] Manfred Stadel. Another nonrecursive algorithm for the Towers of Hanoi. ACM SIGPLAN Notices, 19(9):34– 36, September 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanKatwijk:1984:PEA

[778] J. van Katwijk. Practical experiences with automatic repair of syntactical errors, or, syntactical error repair in the DAS compiler. ACM SIGPLAN Notices, 19(9):37–48, September 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winkler:1984:SIIb

[779] J. F. H. Winkler. Some improvements of ISO-Pascal. ACM SIGPLAN Notices, 19(9):49–62, September 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gait:1984:SOK

[780] Jason Gait. Semaphores outside the kernel. ACM SIGPLAN Notices, 19 (10):12–21, October 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greiter:1984:ASF

[781] Gebh. Greiter. acce-specifications: a formal method to define data types. ACM SIGPLAN Notices, 19(10):22– 31, October 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1984:DBM

[782] Hausi A. Muller. Differences between Modula-2 and Pascal. ACM SIGPLAN Notices, 19(10):32–39, October 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richards:1984:OAS

[783] H. Richards. An overview of ARC SASL. ACM SIGPLAN Notices, 19 (10):40–45, October 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1984:STC

[784] David O. Williams. Structured transfer of control. ACM SIGPLAN Notices, 19(10):46–51, October 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1984:AR

[785] David O. Williams. Alternate RE-TURNs. ACM SIGPLAN Notices, 19 (10):52–57, October 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zedan:1984:NDF

[786] H. Zedan. A note on deadlock-free proofs of network of processes. ACM SIGPLAN Notices, 19(10):58–62, October 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pentzlin:1984:EPP

[787] Karl Ludwig Pentzlin. An extension proposal for Pascal: simplifying assignments like “a := a + 1;” by a “pronoun character”. ACM SIG-PLAN Notices, 19(11):17, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rising:1984:SDE

[788] Linda Rising. A syntax-directed editor, world-builder and simulator for the language of KAREL the robot. ACM SIG-PLAN Notices, 19(11):18–21, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenthal:1984:PMA

[789] David Rosenthal. A proposed modification to the absolute coordinates of Logo. ACM SIGPLAN Notices, 19 (11):22, November 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewry:1984:MPF

[790] D. A. Sewry. Modula-2 process facilities. ACM SIGPLAN Notices, 19(11): 23–32, November 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewry:1984:MMC

[791] D. A. Sewry. Modula-2 and the monitor concept. ACM SIGPLAN Notices, 19(11):33–41, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turba:1984:EBM

[792] Thomas N. Turba. An exception-based mechanism for syntactic error recovery. ACM SIGPLAN Notices, 19(11): 42–51, November 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanTuyl:1984:EF

[793] Robert R. Van Tuyl. On evolution of FORTRAN. ACM SIGPLAN Notices, 19(11):52–55, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:1984:SQA

[794] ShouHan Wang. Some questions about “On the Elimination of Pointer Variables and Dynamic Allocation in Higher Level Languages”. ACM SIG-PLAN Notices, 19(11):56–58, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zethmayr:1984:CSD

[795] David Zethmayr. CUL-de-SAC and dramatic EXIT for use in structured languages. ACM SIGPLAN Notices, 19 (11):59–62, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anselmi:1984:NPL

[796] L. Anselmi. A new programming language: GALILEI. ACM SIGPLAN Notices, 19(12):12–26, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ejiogu:1984:OSS

[797] Lem O. Ejiogu. OPUS #16G software structure: its characteristic polynomials. ACM SIGPLAN Notices, 19(12): 27–34, December 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geller:1984:PW

[798] Dennis P. Geller. Programming and writing. ACM SIGPLAN Notices, 19 (12):35–37, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyer:1984:NIH

[799] Bertrand Meyer. A note on iterative Hanoi. ACM SIGPLAN Notices, 19(12):38–40, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moffat:1984:SCA

[800] David V. Moffat. Some concerns about Modula-2. ACM SIGPLAN Notices, 19(12):41–47, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [832, 876].

Taylor:1984:ACL

[801] David Taylor. An alternative to current looping syntax. ACM SIGPLAN Notices, 19(12):48–53, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vouk:1984:CML

[802] Mladen A. Vouk. On the cost of mixed language programming. ACM SIG-PLAN Notices, 19(12):54–60, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1984:MRM

[803] Mark B. Wells, Margaret A. Hug, Roland Silver, and John Brockmeyer. Modcap reference manual. ACM SIG-PLAN Notices, 19(12):61–79, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Birtwistle:1985:CH

[804] Graham Birtwistle. The coroutines of Hanoi. ACM SIGPLAN Notices, 20 (1):9–10, January 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Breguet:1985:SPL

[805] Pierre Breguet, François Grize, and Alfred Strohmeier. SARTEX. A programming language for graph processing. ACM SIGPLAN Notices, 20(1): 11–19, January 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Covington:1985:EUL

[806] Michael A. Covington. Eliminating unwanted loops in Prolog. ACM SIG-PLAN Notices, 20(1):20–26, January 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammons:1985:MFS

[807] Charles Hammons and Paul Dobbs. Multilevel files for software development. ACM SIGPLAN Notices, 20(1): 27–30, January 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1985:RSI

[808] Ralph E. Griswold. Rebus — a SNOBOL4/Icon hybrid. ACM SIG-PLAN Notices, 20(2):7–16, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brun:1985:TOA

[809] G. Brun, A. Businger, and R. Schoenberger. The token-oriented approach to program editing. ACM SIGPLAN Notices, 20(2):17–20, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cottam:1985:EPO

[810] Ian D. Cottam. Extending Pascal with one-entry/multi-exit procedures. ACM SIGPLAN Notices, 20(2):21–31, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1985:UBI

[811] Peter C. J. Graham. Using BINS for inter-process communication. ACM SIGPLAN Notices, 20(2):32–41, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1985:MSI

[812] Warren Harrison and Curtis Cook. A method of sharing industrial software complexity data. ACM SIGPLAN Notices, 20(2):42–51, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1985:SCR

[813] Warren Harrison and Curtis Cook. Software complexity research: a survey. ACM SIGPLAN Notices, 20(2): 52–53, February 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heyman:1985:SCR

[814] Jerrold Heyman. Software conversion: retargeting from a super-mini to a micro. ACM SIGPLAN Notices, 20(2): 54–57, February 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meertens:1985:DB

[815] Lambert Meertens and Steven Pemberton. Description of B. ACM SIG-PLAN Notices, 20(2):58–76, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Molnar:1985:PLP

[816] Ludovít Molnár, Pavol Návrat, and Jiří Šafařík. The programming language Pascal (in teaching perspective): take it or leave it. ACM SIGPLAN Notices, 20(2):77–79, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dutta:1985:MPC

[817] Kalyan Dutta. Modular programming in C: an approach and an example. ACM SIGPLAN Notices, 20(3):9–15, March 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ejiogu:1985:SMS

[818] Lem O. Ejiogu. A simple measure of software complexity. ACM SIG-PLAN Notices, 20(3):16–31, March 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Machanick:1985:CSL

[819] Philip Machanick. The case for simplified languages. ACM SIGPLAN Notices, 20(3):32–36, March 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Petermann:1985:NPD

[820] Uwe Petermann and Andrzej Szalas. A note on PCI: distributed processes communicating by interrupts. ACM SIGPLAN Notices, 20(3):37–46, March 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:1985:NMO

[821] A. P. Chang. A note on the MODULO operation. ACM SIGPLAN Notices, 20 (4):19–23, April 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Glass:1985:TIS

[822] Harvey Glass. Threaded interpretive systems and functional programming environments. ACM SIGPLAN Notices, 20(4):24–32, April 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonak:1985:PCP

[823] Jan E. Jonak. Pascal and comms programming. ACM SIGPLAN Notices, 20(4):33–41, April 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kilian:1985:CCR

[824] Michael F. Kilian. A conditional critical region pre-processor for C based on the Owicki and Gries scheme. ACM SIGPLAN Notices, 20(4):42–56, April 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1985:MSE

[825] Bent Bruun Kristensen, Ole Lehrmann Madsen, Birger Møller-Pedersen, and Kristen Nygaard. Multi-sequential execution in the BETA programming language. ACM SIGPLAN Notices, 20 (4):57–70, April 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mamone:1985:UPI

[826] Salvatore Mamone. Using a PC to increase your odds of winning the lottery. ACM SIGPLAN Notices, 20(4):71–79, April 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sklenar:1985:PLA

[827] Ivan Sklenar. Programming language assessments using the program profiles. ACM SIGPLAN Notices, 20(4):80–87, April 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tomek:1985:CIP

[828] Ivan Tomek and Tomasz Muldner. A CAI implementation of Pascal. ACM SIGPLAN Notices, 20(4):88–95, April 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dunlavey:1985:PRD

[829] Michael R. Dunlavey. A progress report on D: a compiled language featuring continuations. ACM SIG-PLAN Notices, 20(5):8–15, May 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maziar:1985:STH

[830] Stepan Maziar. Solution of the Tower of Hanoi problem using a binary tree. ACM SIGPLAN Notices, 20(5):16–20, May 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ragan:1985:CCI

[831] Richard R. Ragan. CYBIL: Cyber Implementation Language. ACM SIG-PLAN Notices, 20(5):21–30, May 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Segre:1985:SCA

[832] Lidia Segre and Michael Stanton. ‘Some concerns about Modula-2’ considered unwarranted. ACM SIG-PLAN Notices, 20(5):31–35, May 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [800, 876].

Strom:1985:NDS

[833] Rob Strom and Shaula Yemini. The NIL distributed systems programming language: a status report. ACM SIG-PLAN Notices, 20(5):36–44, May 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanthienen:1985:NED

[834] J. Vanthienen. A note on English for decision tables considered harmful and the nested IF-THEN-ELSE. ACM SIG-PLAN Notices, 20(5):45–47, May 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [448].

Dasiewicz:1985:SRT

[835] P. Dasiewicz. A soft real-time multitasking BASIC interpreter. ACM SIG-PLAN Notices, 20(6):10–15, June 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gough:1985:NMG

[836] K. John Gough. A new method of generating LL (1) lookahead sets. ACM SIGPLAN Notices, 20(6):16–19, June 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lo:1985:SPM

[837] Chingmin Jim Lo. Simple patches to Modula-2 text I/O. ACM SIG-PLAN Notices, 20(6):20–25, June 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OKeefe:1985:AKP

[838] R. A. O’Keefe. Alternatives to keyword parameters. ACM SIGPLAN Notices, 20(6):26–32, June 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OKeefe:1985:FSM

[839] R. A. O’Keefe. Finding SmallTalk methods. ACM SIGPLAN Notices, 20 (6):33–38, June 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenbloom:1985:MPL

[840] Michael H. Rosenbloom. M’Pal: a programming language based on natural English definitions. ACM SIG-PLAN Notices, 20(6):39–46, June 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rudnicki:1985:OPB

[841] Piotr Rudnicki. On the organization of Pascal blocks. ACM SIG-PLAN Notices, 20(6):47–55, June 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiener:1985:TAI

[842] Richard S. Wiener and Richard F. Sincovec. Two approaches to implementing generic data structures in Modula-2. ACM SIGPLAN Notices, 20(6): 56–64, June 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laff:1985:SOB

[843] Mark R. Laff and Brent Hailpern. SW 2: An object-based programming environment. ACM SIGPLAN Notices, 20(7):1–11, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Purtilo:1985:PES

[844] James Purtilo. Polylith: An environment to support management of tool interfaces. ACM SIGPLAN Notices, 20 (7):12–18, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ODonnell:1985:DBC

[845] John T. O’Donnell. Dialogues: a basis for constructing programming environments. ACM SIGPLAN Notices, 20(7):19–27, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bahlke:1985:PPS

[846] Rolf Bahlke and Gregor Snelting. The PSG — Programming System Generator. ACM SIGPLAN Notices, 20(7): 28–33, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beshers:1985:MCA

[847] George McArthur Beshers and Roy Harold Campbell. Maintained and constructor attributes. ACM SIGPLAN Notices, 20(7):34–42, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Demers:1985:APM

[848] Alan Demers, Anne Rogers, and Frank Kenneth Zadeck. Attribute propagation by message passing. ACM SIGPLAN Notices, 20(7):43–59, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hall:1985:DSE

[849] Cordelia V. Hall and John T. O’Donnell. Debugging in a side effect free programming environment. ACM SIGPLAN Notices, 20(7):60– 68, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeMetayer:1985:MAP

[850] D. Le Metayer. Mechanical analysis of program complexity. ACM SIG-PLAN Notices, 20(7):69–73, July 1985.

CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dybvig:1985:SE

[851] R. Kent Dybvig and Bruce T. Smith. A semantic editor. ACM SIGPLAN Notices, 20(7):74–82, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1985:LPP

[852] Mark B. Wells, Margaret A. Hug, and Rollo Silver. Libraries as programs preserved within compiler continuations. ACM SIGPLAN Notices, 20(7):83–92, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horwitz:1985:RAS

[853] Susan Horwitz and Tim Teitelbaum. Relations and attributes: a symbiotic basis for editing environments. ACM SIGPLAN Notices, 20(7):93– 106, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1985:IIA

[854] Keith D. Cooper, Ken Kennedy, and Linda Torczon. The impact of interprocedural analysis and optimization on the design of a software development environment. ACM SIGPLAN Notices, 20(7):107–116, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agnarsson:1985:TTP

[855] Snorri Agnarsson and M. S. Krishnamoorthy. Towards a theory of packages. ACM SIGPLAN Notices, 20(7): 117–130, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gries:1985:NNE

[856] David Gries and Jan Prins. A new notion of encapsulation. ACM SIGPLAN Notices, 20(7):131–139, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caplinger:1985:SES

[857] Michael Caplinger. Structured Editor Support for Modularity and Data Abstraction. ACM SIGPLAN Notices, 20 (7):140–147, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moriconi:1985:PSG

[858] Mark Moriconi and Dwight F. Hare. PegaSys: a system for graphical explanation of program designs. ACM SIGPLAN Notices, 20(7):148–160, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1985:APF

[859] Bent Bruun Kristensen, Ole Lehrmann Madsen, Birger Møller-Pedersen, and Kristen Nygaard. An algebra for program fragments. ACM SIGPLAN Notices, 20(7):161–170, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hood:1985:EAI

[860] Robert Hood. Efficient abstractions for the implementation of structured editors. ACM SIGPLAN Notices, 20(7): 171–178, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Francez:1985:ELP

[861] Nissim Francez, Shalom Goldenberg, Ron Y. Pinter, Michael Tiomkin, and Shalom Tsur. An environment for logic programming. ACM SIGPLAN Notices, 20(7):179–190, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Komorowski:1985:MIL

[862] Henryk Jan Komorowski and Shigeo Omori. A model and an implementation of a logic programming environment. ACM SIGPLAN Notices, 20(7): 191–198, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1985:SSE

[863] Bruce J. MacLennan. A simple software environment based on objects and relations. ACM SIGPLAN Notices, 20 (7):199–207, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fickas:1985:DIR

[864] Stephen Fickas. Design issues in a Rule-Based System. ACM SIGPLAN Notices, 20(7):208–215, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sweet:1985:MPE

[865] Richard E. Sweet. The Mesa programming environment. ACM SIGPLAN Notices, 20(7):216–229, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swinehart:1985:SC

[866] Daniel C. Swinehart, Polle T. Zellweger, and Robert B. Hagmann. The structure of Cedar. ACM SIGPLAN Notices, 20(7):230–244, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donahue:1985:IMC

[867] James Donahue. Integration mechanisms in Cedar. ACM SIGPLAN Notices, 20(7):245–251, July 1985. CO-DEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seyfer:1985:BCA

[868] H. K. Seyfer. A bibliography on compiler assessment. II. ACM SIG-PLAN Notices, 20(8):20–27, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Covington:1985:FNL

[869] Michael A. Covington. A further note on looping in Prolog. ACM SIG-PLAN Notices, 20(8):28–31, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nute:1985:PSC

[870] Donald Nute. A programming solution to certain problems with loops in Prolog. ACM SIGPLAN Notices, 20 (8):32–37, August 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Poole:1985:ELP

[871] David Poole and Randy Goebel. On eliminating loops in Prolog. ACM SIG-PLAN Notices, 20(8):38–40, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albizuri-Romero:1985:IRP

[872] Miren Begoña Albizuri-Romero. Internal representation of programs in GRASE. ACM SIGPLAN Notices, 20 (8):41–50, August 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amoroso:1985:FSP

[873] Edward G. Amoroso. The formal specification and prototype implementation of a simple editor. ACM SIG-PLAN Notices, 20(8):51–59, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bron:1985:NCI

[874] C. Bron, E. J. Dijkstra, and T. J. Rossingh. A note on the checking of interfaces between separately compiled modules. ACM SIGPLAN Notices, 20 (8):60–63, August 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gerber:1985:TMR

[875] A. J. Gerber. The trouble with mutual recursion in Concurrent Euclid. ACM SIGPLAN Notices, 20(8):64–70, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldberg:1985:RSC

[876] Morton Goldberg. A response to “Some concerns about Modula-2”. ACM SIG-PLAN Notices, 20(8):71–72, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [800, 832].

Harris:1985:RD

[877] Lawrence A. Harris and Yannich Wets. Resetting displays. ACM SIGPLAN Notices, 20(8):73–77, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jares:1985:WSV

[878] Vladimir Jareš. Work with strings of variable length in COBOL. ACM SIG-PLAN Notices, 20(8):78–81, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1985:EPE

[879] Neil D. Jones, Peter Sestoft, and Harald S{øndergaard. An experiment in partial evaluation: the generation of a compiler generator. ACM SIG-PLAN Notices, 20(8):82–87, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaLonde:1985:SGI

[880] Wilf R. LaLonde and John R. Pugh. Specialization, generalization and inheritance, teaching objectives beyond data structures and data types. ACM SIGPLAN Notices, 20(8):88–92, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turba:1985:PEH

[881] Thomas N. Turba. The Pascal exception handling proposal. ACM SIG-PLAN Notices, 20(8):93–98, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitelaw:1985:SRE

[882] M. W. Whitelaw. Some ramifications of the EXIT statement in loop control. ACM SIGPLAN Notices, 20(8):99–106, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zedan:1985:SDD

[883] H. Zedan. Safety decomposition of distributed programs. ACM SIGPLAN Notices, 20(8):107–112, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnett:1985:PSM

[884] Michael P. Barnett. The primitive string manipulation (PRISM) language. ACM SIGPLAN Notices, 20(9): 11–20, September 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Binding:1985:CCC

[885] Carl Binding. Cheap concurrency in C. ACM SIGPLAN Notices, 20(9): 21–26, September 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1985:IS

[886] A. Toni Cohen and Thomas J. Myers. Information sharing. ACM SIG-PLAN Notices, 20(9):27–31, September 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eggers:1985:THY

[887] Bleicke Eggers. The Towers of Hanoi: yet another nonrecursive solution. ACM SIGPLAN Notices, 20(9): 32–42, September 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hornsby:1985:DIF

[888] C. Hornsby and C. H. C. Leung. The design and implementation of a flexible retrieval language for a Prolog database system. ACM SIGPLAN Notices, 20(9):43–51, September 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matthews:1985:PM

[889] David C. J. Matthews. Poly manual. ACM SIGPLAN Notices, 20(9):52–76, 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matthews:1985:PMP

[890] D. C. J. Matthews. Poly manual: Poly and standard ML. ACM SIGPLAN Notices, 20(9):52–76, September 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Postma:1985:PDS

[891] Stef. W. Postma. Proposed data structures for QL/Bn. ACM SIGPLAN Notices, 20(9):77–85, September 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shao:1985:ICM

[892] J. Shao and G. E. Quick. Implementation considerations of mapping LISP onto a cellular computer system. ACM SIGPLAN Notices, 20(9): 86–93, September 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thornburg:1985:FNM

[893] Jonathan Thornburg. Further notes on the modulo operator. ACM SIG-PLAN Notices, 20(9):94, September 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baldwin:1985:GNA

[894] Roger R. Baldwin. A generalization of Neta Amit’s statement of the general problem of deeply nested IF-THEN-ELSEs. ACM SIGPLAN Notices, 20 (10):9–13, October 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowe:1985:CCI

[895] Malcolm Crowe, Clark Nicol, Michael Hughes, and David Mackay. On converting a compiler into an incremental compiler. ACM SIGPLAN Notices, 20(10):14–22, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harland:1985:AVP

[896] David M. Harland, Martyn W. Szyplewski, and John B. Wainwright. An alternative view of polymorphism. ACM SIGPLAN Notices, 20(10):23– 35, October 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1985:ISD

[897] D. C. Ince, Walton Hall, and Milton Keynes. The influence of system design complexity research on the design of module interconnection languages. ACM SIGPLAN Notices, 20(10):36– 43, October 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaplan:1985:SXL

[898] Randy M. Kaplan. small-X — a language and interpreter for building expert systems. ACM SIGPLAN Notices, 20(10):44–47, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matsumoto:1985:SAP

[899] Hidekazu Matsumoto. A static analysis of Prolog programs. ACM SIG-PLAN Notices, 20(10):48–59, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olthoff:1985:OM

[900] Walter Olthoff. An overview on Mod-Pascal. ACM SIGPLAN Notices, 20 (10):60–71, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Otto:1985:PPC

[901] Jerome A. Otto. Predicting potential COBOL performance on low level machine architectures. ACM SIGPLAN Notices, 20(10):72–78, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenthal:1985:AMR

[902] Don Rosenthal. Adding meta rules to OPS5. A proposed extension. ACM SIGPLAN Notices, 20(10):79–86, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sager:1985:TCS

[903] Thomas J. Sager. A technique for creating small fast compiler frontends. ACM SIGPLAN Notices, 20(10):87– 94, October 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Szalas:1985:EHP

[904] Andrzej Szalas and Danuta Szczepanska. Exception handling in parallel computations. ACM SIGPLAN Notices, 20(10):95–104, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yamane:1985:STT

[905] Yoshito Yamane. Some thoughts on the type system of B. ACM SIGPLAN Notices, 20(10):105–110, October 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aharonian:1985:MLB

[906] G. Aharonian. Multi-language BLAS: a proposal. ACM SIGPLAN Notices, 20(11):11–13, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buhr:1985:CTM

[907] P. A. Buhr. A case for teaching multiexit loops to beginning programmers. ACM SIGPLAN Notices, 20(11):14– 22, November 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burkhard:1985:MIC

[908] Neil A. Burkhard. Machine-independent C optimizer. ACM SIGPLAN Notices, 20(11):23–26, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dwyer:1985:IGL

[909] Barry Dwyer. Improving Gough’s LL(1) lookahead generator. ACM SIG-PLAN Notices, 20(11):27–29, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:1985:ESM

[910] Rajiv Gupta and Mary Lou Soffa. The efficiency of storage management schemes for Ada programs. ACM SIG-PLAN Notices, 20(11):30–38, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joslin:1985:EPI

[911] David Joslin. Extensions in Pascal implementations. ACM SIGPLAN Notices, 20(11):39–46, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marcus:1985:GI

[912] Robert Marcus. Generalized inheritance. ACM SIGPLAN Notices, 20 (11):47–48, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pase:1985:SPM

[913] Douglas M. Pase. System programming in Modula-2. ACM SIGPLAN Notices, 20(11):49–53, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silvestri:1985:IPN

[914] Antonio C. Silvestri. An IBM-PC numeric I/O interrupt. ACM SIGPLAN Notices, 20(11):54–63, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolfe:1985:WOC

[915] Michael Wolfe and Tom Macke. Where are the optimizing compilers? ACM SIGPLAN Notices, 20(11):64– 68, November 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zimmer:1985:MMW

[916] J. A. Zimmer. A modest Modula wish list. ACM SIGPLAN Notices, 20(11): 69–77, November 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnett:1985:SMS

[917] Michael P. Barnett. A simple mechanism that supports the PRISM language. ACM SIGPLAN Notices, 20 (12):5–12, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bielak:1985:AVM

[918] Richard Bielak. ADA(*) vs. Modula-2: a view from the trenches. ACM SIG-PLAN Notices, 20(12):13–17, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [950].

Crawford:1985:HLI

[919] Albert L. Crawford. High level input/ output in Modula-2. ACM SIGPLAN Notices, 20(12):18–25, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Czyzowicz:1985:IGT

[920] Jurek Czyzowicz and Michal Iglewski. Implementing generic types in Modula-2. ACM SIGPLAN Notices, 20(12): 26–32, December 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hyun:1985:IES

[921] Kio C. Hyun and Ernst-Erich Doberkat. Inline expansion of SETL procedures. ACM SIGPLAN Notices, 20 (12):33–38, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Konopasek:1985:THD

[922] Milos Konopasek. The Towers of Hanoi from a different viewpoint. ACM SIG-PLAN Notices, 20(12):39–46, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sebesta:1985:MPH

[923] Robert W. Sebesta and Mark A. Taylor. Minimal perfect hash functions for reserved word lists. ACM SIGPLAN Notices, 20(12):47–53, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terry:1985:CST

[924] P. D. Terry. CLANG: a simple teaching language. ACM SIGPLAN Notices, 20(12):54–63, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berghel:1986:SVP

[925] Hal Berghel and Eric Traudt. Spelling verification in Prolog. ACM SIG-PLAN Notices, 21(1):19–27, January 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brand:1986:TP

[926] Daniel Brand. On typing in Prolog. ACM SIGPLAN Notices, 21(1):28–30, January 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeKerf:1986:ACO

[927] Joseph L. F. De Kerf. APL compared with other languages according to Halstead’s theory. ACM SIG-PLAN Notices, 21(1):31–39, January 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dunn:1986:NTI

[928] Dick Dunn. Non-technical issues in acceptance of SILs. ACM SIGPLAN Notices, 21(1):40–45, January 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunt:1986:DDM

[929] J. G. Hunt. Detection of deadlocks in multiprocess systems. ACM SIG-PLAN Notices, 21(1):46–48, January 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1986:NVC

[930] Kuo-Cheng Li. A note on the vector C language. ACM SIGPLAN Notices, 21 (1):49–57, January 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sorens:1986:TAP

[931] Michael J. Sorens. A technique for automatically porting dialects of Pascal to each other. ACM SIGPLAN Notices, 21(1):58–63, January 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heering:1986:EEI

[932] J. Heering and P. Klint. The efficiency of the Equation Interpreter compared with the UNH Prolog interpreter. ACM SIGPLAN Notices, 21(2):18–21, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:1986:TFP

[933] W. T. Huang and D. C. You. Typed functional programming. ACM SIG-PLAN Notices, 21(2):22–26, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonak:1986:EFL

[934] J. E. Jonak. Experience with a FORTH-like language. ACM SIG-PLAN Notices, 21(2):27–36, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muchnick:1986:SES

[935] T. G. Muchnick. On some extensions of syntactic error recovery technique based on phrase markers. ACM SIG-PLAN Notices, 21(2):37–39, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skelton:1986:MMT

[936] S. Skelton. Measurements of migratability and transportability. ACM SIG-PLAN Notices, 21(2):40–45, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weeks:1986:CFA

[937] J. Weeks and H. Berghel. A comparative feature-analysis of microcomputer Prolog implementations. ACM SIG-PLAN Notices, 21(2):46–61, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manner:1986:STP

[938] R. Männer. Strong typing and physical units (syntactic extensions of high level languages). ACM SIGPLAN Notices, 21(3):11–20, March 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1986:PNP

[939] Mark B. Wells. A potpourri of notational pet peeves (and their resolution in modcap). ACM SIGPLAN Notices, 21(3):21–30, March 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rutter:1986:BRP

[940] Paul Rutter. Book review: Performance and Evaluation of Lisp Systems by Richard P. Gabriel (MIT Press, 1985). ACM SIGPLAN Notices, 21 (4):21, April 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beidler:1986:CGM

[941] J. Beidler and P. Jackowitz. Consistent generics in Modula-2. ACM SIGPLAN Notices, 21(4):32–41, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:1986:ALE

[942] T. Y. Chen and S. C. Kwan. An analysis of length equation using a dynamic approach. ACM SIGPLAN Notices, 21 (4):42–47, April 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cisneros:1986:IPL

[943] G. Cisneros and H. V. McIntosh. Introduction to the programming language Convert. ACM SIGPLAN Notices, 21 (4):48–57, April 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Groening:1986:NNM

[944] K. Groening and C. Ohsendoth. NEMO: a nicely modified YACC. ACM SIGPLAN Notices, 21(4):58–66, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kelsh:1986:RPC

[945] J. P. Kelsh and J. C. Hansen. Reasonably priced compilers and interpreters for the IBM PC and compatible computers. ACM SIGPLAN Notices, 21 (4):67–69, April 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Longe:1986:LOA

[946] M. R. Longe. The line-oriented approach (programming systems). ACM SIGPLAN Notices, 21(4):70–74, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ng:1986:PPP

[947] K. W. Ng and W. Y. Ma. Pitfalls in PROLOG programming. ACM SIGPLAN Notices, 21(4):75–79, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [1108].

Nielson:1986:BAI

[948] F. Nielson. A bibliography on abstract interpretation. ACM SIG-PLAN Notices, 21(5):31–38, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1986:AVM

[949] M. B. Feldman. Ada vs. Modula-2: a response from the ivory tower. ACM SIGPLAN Notices, 21(5):39–44, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenwood:1986:CVT

[950] J. R. Greenwood. Comments on “A View from the Trenches”. Ada vs. Modula-2 vs. Praxis. ACM SIG-PLAN Notices, 21(5):45–49, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [918].

Hayward:1986:CAP

[951] V. Hayward. Compared anatomy of the programming languages Pascal and C. ACM SIGPLAN Notices, 21(5):50–60, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jazayeri:1986:OCH

[952] M. Jazayeri and M. Haden. Optimizing compilers are here (mostly). ACM SIGPLAN Notices, 21(5):61–63, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kary:1986:TES

[953] D. D. Kary and P. L. Juell. TRC: An expert system compiler. ACM SIG-PLAN Notices, 21(5):64–68, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mossenbock:1986:ASEa

[954] H. Mössenböck. Alex: a simple and efficient scanner generator. ACM SIG-PLAN Notices, 21(5):69–78, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tracz:1986:CUP

[955] W. Tracz. Confessions of a used program salesman — The loves in My Life. ACM SIGPLAN Notices, 21(5):79–81, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watson:1986:RCB

[956] J. F. Watson. Recursion: a choice between readability and execution speed. ACM SIGPLAN Notices, 21(5):82–90, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegman:1986:WLP

[957] M. N. Wegman. What it’s like to be a POPL referee; or how to write an extended abstract so that it is more likely to be accepted. ACM SIG-PLAN Notices, 21(5):91–95, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1986:E

[958] Richard L. Wexelblat. Editorial. ACM SIGPLAN Notices, 21(6):1–2, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gfeller:1986:AAT

[959] Martin Gfeller. APL arrays and their editor. ACM SIGPLAN Notices, 21 (6):18–27, June 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsby:1986:CUG

[960] M. E. Goldsby. Concurrent use of generic types in MODULA-2. ACM SIGPLAN Notices, 21(6):28–34, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Machanick:1986:LPT

[961] P. Machanick. Are LR parsers too powerful? ACM SIGPLAN Notices, 21 (6):35–40, June 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1986:VIP

[962] D. Taylor. Visionaries: An interview with Professor Michael O’Leary. ACM SIGPLAN Notices, 21(6):41–54, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thornburg:1986:KIC

[963] J. Thornburg. Keyword input for C. ACM SIGPLAN Notices, 21(6):55–58, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thornburg:1986:NCC

[964] J. Thornburg. Number crunching in C. ACM SIGPLAN Notices, 21(6):59–60, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vasak:1986:UCO

[965] T. Vasak. Universal closure operator for Prolog. ACM SIGPLAN Notices, 21(6):61–62, June 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiener:1986:PAU

[966] R. S. Wiener. Protecting against uninitialized abstract objects in Modula-2. ACM SIGPLAN Notices, 21(6):63–69, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See note [1181].

Zavodnik:1986:YDY

[967] R. J. Zavodnik and M. D. Middleton. YALE: The design of yet another language-based editor. ACM SIGPLAN Notices, 21(6):70–78, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bernstein:1986:OCE

[968] David Bernstein, Haran Boral, and Ron Y. Pinter. Optimal chaining in expression trees. ACM SIGPLAN Notices, 21(7):1–10, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p1-bernstein/

.

Gibbons:1986:EIS

[969] Philip B. Gibbons and Steven S. Muchnick. Efficient instruction scheduling for a pipelined architecture. ACM SIGPLAN Notices, 21(7):11–16, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p11-gibbons/

.

Sarkar:1986:CTP

[970] Vivek Sarkar and John Hennessy. Compile-time partitioning and scheduling of parallel programs. ACM SIGPLAN Notices, 21(7):17–26, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p17-sarkar/

.

Gross:1986:CHP

[971] Thomas Gross and Monica S. Lam. Compilation for a high-performance systolic array. ACM SIGPLAN Notices, 21(7):27–38, July 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/12276/p27-gross/

.

Hoover:1986:EIE

[972] Roger Hoover and Tim Teitelbaum. Efficient incremental evaluation of aggregate values in attribute grammars. ACM SIGPLAN Notices, 21(7):39–50, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p39-hoover/

.

Kaiser:1986:GRT

[973] Gail E. Kaiser. Generation of runtime environments. ACM SIG-PLAN Notices, 21(7):51–57, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p51-kaiser/

.

Cooper:1986:IOE

[974] Keith D. Cooper, Ken Kennedy, and Linda Torczon. Interprocedural optimization: eliminating unnecessary recompilation. ACM SIG-PLAN Notices, 21(7):58–67, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p58-cooper/

.

Royer:1986:TDS

[975] Veronique Royer. Transformations of denotational semantics in semantics directed compiler generation. ACM SIGPLAN Notices, 21(7):68–73, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p68-royer/

.

Raiha:1986:GTA

[976] K. J. Räihä and Jorma Tarhio. A globalizing transformation for attribute grammars. ACM SIGPLAN Notices, 21(7):74–84, July 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p74-raiha/

.

Farrow:1986:AGF

[977] Rodney Farrow. Automatic generation of fixed-point-finding evaluators for circular, but well-defined, attribute grammars. ACM SIG-PLAN Notices, 21(7):85–98, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p85-farrow/

.

Johnson:1986:EML

[978] Mark S. Johnson and Terrence C. Miller. Effectiveness of a machinelevel, global optimizer. ACM SIG-PLAN Notices, 21(7):99–108, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p99-johnson/

.

Johnson:1986:DFA

[979] Harold Johnson. Data flow analysis for ‘intractable’ imbedded system software. ACM SIGPLAN Notices, 21(7):109–117, July 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p109-johnson/

.

Kessler:1986:ERH

[980] R. R. Kessler, J. C. Peterson, H. Carr, G. P. Duggan, J. Knell, and J. J. Krohnfeldt. EPIC — a retargetable, highly optimizing Lisp compiler. ACM SIGPLAN Notices, 21(7):118–130, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p118-kessler/

.

Ives:1986:UVR

[981] Fred Ives. Unifying view of recent LALR(1) lookahead set algorithms. ACM SIGPLAN Notices, 21(7):131–135, July 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/12276/p131-ives/

.

Bermudez:1986:PAL

[982] Manuel E. Bermudez and Karl M. Schimpf. A practical arbitrary lookahead LR parsing technique. ACM SIGPLAN Notices, 21(7):136–144, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p136-bermudez/

.

Pennello:1986:VFL

[983] Thomas J. Pennello. Very fast LR parsing. ACM SIGPLAN Notices, 21(7):145–151, July 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p145-pernnello/

.

Callahan:1986:ICP

[984] David Callahan, Keith D. Cooper, Ken Kennedy, and Linda Torczon. Interprocedural constant propagation (compilers). ACM SIG-PLAN Notices, 21(7):152–161, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p152-callahan/

.

Burke:1986:IDA

[985] Michael Burke and Ron Cytron. Interprocedural dependence analysis and parallelization. ACM SIG-PLAN Notices, 21(7):162–175, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p162-burke/

.

Triolet:1986:DPC

[986] Rémi Triolet, François Irigoin, and Paul Feautrier. Direct parallelization of CALL statements. ACM SIG-PLAN Notices, 21(7):176–185, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p176-triolet/

.

Boehm:1986:IR

[987] Hans-Juergen Boehm and Alan Demers. Implementing RUSSELL. ACM SIGPLAN Notices, 21(7):186–195, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p186-boehm/

.

Lindstrom:1986:SEF

[988] Gary Lindstrom. Static evaluation of functional programs. ACM SIG-PLAN Notices, 21(7):196–206, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p196-lindstrom/

.

Harrison:1986:ECL

[989] Peter G. Harrison and Hessam Khoshnevisan. Efficient compilation of linear recursive functions into object level loops. ACM SIGPLAN Notices, 21(7):207–218, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/12276/p207-harrison/

.

Kranz:1986:OOC

[990] David Kranz, Richard Kelsey, Jonathan Rees, Paul Hudak, James Philbin, and Norman Adams. ORBIT: an optimizing compiler for Scheme. ACM SIGPLAN Notices, 21(7):219–233, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p219-kranz/

.

Davidson:1986:RIR

[991] Jack W. Davidson. A retargetable instruction reorganizer. ACM SIG-PLAN Notices, 21(7):234–241, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p234-davidson/

.

Fraser:1986:ICG

[992] Christopher W. Fraser and Alan L. Wendt. Integrating code generation and optimization. ACM SIG-PLAN Notices, 21(7):242–248, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p242-fraser/

.

Kessler:1986:DMS

[993] Peter B. Kessler. Discovering machinespecific code improvements. ACM SIG-PLAN Notices, 21(7):249–254, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/12276/p249-kessler/

.

Larus:1986:RAS

[994] James R. Larus and Paul N. Hilfinger. Register allocation in the SPUR Lisp compiler. ACM SIG-PLAN Notices, 21(7):255–263, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p255-larus/

.

Wall:1986:GRA

[995] David W. Wall. Global register allocation at link time. ACM SIG-PLAN Notices, 21(7):264–275, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/12276/p264-wall/

.

Wilkes:1986:ESP

[996] Maurice V. Wilkes. EMPIRIC: a sketch of a programming language designed to facilitate a fine grain of protection. ACM SIGPLAN Notices, 21(8):16–27, August 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clay:1986:NSQ

[997] C. Clay. A new solution to the N ≤ 8 queens problem. ACM SIG-PLAN Notices, 21(8):28–30, August 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grune:1986:GPC

[998] Dick Grune. Generic packages in C. ACM SIGPLAN Notices, 21(8):31–39, August 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jouvelot:1986:DNL

[999] P. Jouvelot. Designing new languages or new language manipulation systems using ML. ACM SIGPLAN Notices, 21 (8):40–52, August 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moffat:1986:GDS

[1000] David V. Moffat. Generic data structures in UCSD Pascal. ACM SIG-PLAN Notices, 21(8):53–61, August 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsey:1986:LUM

[1001] Bruce D. Ramsey. The lion and the unicorn met PROLOG. ACM SIG-PLAN Notices, 21(8):62–70, August 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zang:1986:PIC

[1002] Xiaolin Zang. A proposal for implementing the concurrent mechanisms of Ada. ACM SIGPLAN Notices, 21(8): 71–79, August 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1986:SII

[1003] Terry L. Anderson. The scope of imported identifiers in Modula-2. ACM SIGPLAN Notices, 21(9):17– 21, September 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baldwin:1986:SIP

[1004] Roger R. Baldwin. Systematic indentation in PL/I: minimizing the reduction in horizontal space. ACM SIG-PLAN Notices, 21(9):22–26, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christiansen:1986:MLC

[1005] H. P. Christiansen, A. Lacroix, K. Lundberg, A. McKeeman, and P. D. V. van der Stok. More or less concerns about MODULA-2. ACM SIG-PLAN Notices, 21(9):27–31, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gisselquist:1986:ECC

[1006] Richard Gisselquist. An experimental C compiler for the Cray 2 computer. ACM SIGPLAN Notices, 21(9): 32–41, September 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hekmatpour:1986:FLF

[1007] S. Hekmatpour and D. C. Ince. Forms as a language facility. ACM SIGPLAN Notices, 21(9):42–48, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Magel:1986:ECS

[1008] Kenneth Magel. Efficient calculation of the scope program complexity metric. ACM SIGPLAN Notices, 21(9): 49–56, September 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parker:1986:SFC

[1009] J. R. Parker. A subset FORTRAN compiler for a modified Harvard architecture. ACM SIGPLAN Notices, 21 (9):57–62, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silver:1986:CQA

[1010] Rollo Silver, Mark Wells, Shaun-Inn Wu, and Margaret Hug. A concurrent N ≤ 8 Queens’ algorithm using Modcap. ACM SIGPLAN Notices, 21(9): 63–76, September 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanEijk:1986:UAF

[1011] Peter van Eijk. A useful application of formal procedure parameters. ACM SIGPLAN Notices, 21(9): 77–78, September 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1986:TC

[1012] Anonymous. Technical contributions. ACM SIGPLAN Notices, 21(9):??, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1986:ISI

[1013] Peter Wegner and Bruce Shriver. Introduction to the special issue of the SIG-PLAN Notices on the Object-Oriented Programming Workshop. ACM SIG-PLAN Notices, 21(10):1–6, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stroustrup:1986:OC

[1014] Bjarne Stroustrup. An overview of C++. ACM SIGPLAN Notices, 21 (10):7–18, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1986:COC

[1015] Alan Snyder. CommonObjects: an overview (Common Lisp extension). ACM SIGPLAN Notices, 21(10):19– 28, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahn:1986:OCLa

[1016] Kenneth Kahn, Eric Dean Tribble, Mark S. Miller, and Daniel G. Bobrow. Objects in concurrent logic programming languages. ACM SIGPLAN Notices, 21(10):29–38, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This is an abridged version of the paper of the same name published in the OOPSLA 86 Proceedings.

Tokoro:1986:CPO

[1017] Mario Tokoro and Yutaka Ishikawa.

Concurrent programming in Orient84/K: an object-oriented knowledge representation language. ACM SIG-PLAN Notices, 21(10):39–48, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiss:1986:OOF

[1018] Steven P. Reiss. An object-oriented framework for graphical programming (summary paper). ACM SIGPLAN Notices, 21(10):49–57, October 1986.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agha:1986:OAL

[1019] Gul Agha. An overview of actor languages. ACM SIGPLAN Notices, 21 (10):58–67, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deJong:1986:CA

[1020] Peter de Jong. Compilation into AC-TORS. ACM SIGPLAN Notices, 21 (10):68–77, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nguyen:1986:GOMa

[1021] Van Nguyen and Brent Hailpern. A generalized object model. ACM SIG-PLAN Notices, 21(10):78–87, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strom:1986:COOa

[1022] Rob Strom. A comparison of the object-oriented and process paradigms. ACM SIGPLAN Notices, 21(10):88– 97, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hendler:1986:EMI

[1023] James Hendler. Enhancement for multiple-inheritance. ACM SIGPLAN Notices, 21(10):98–106, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borgida:1986:EOO

[1024] Alexander Borgida. Exceptions in object-oriented languages. ACM SIG-PLAN Notices, 21(10):107–119, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zdonik:1986:MCD

[1025] Stanley B. Zdonik. Maintaining consistency in a database with changing types. ACM SIGPLAN Notices, 21(10):120–127, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nygaard:1986:BCOa

[1026] Kristen Nygaard. Basic concepts in object oriented programming. ACM SIG-PLAN Notices, 21(10):128–132, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madsen:1986:BSOa

[1027] Ole Lehrmann Madsen. Block structure and object oriented languages. ACM SIGPLAN Notices, 21(10):133– 142, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossher:1986:MSS

[1028] Harold L. Ossher. A mechanism for specifying the structure of large, layered, object-oriented programs. ACM SIGPLAN Notices, 21(10):143–152, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goguen:1986:EFO

[1029] Joseph A. Goguen and José Meseguer. Extensions and foundations of objectoriented programming. ACM SIG-PLAN Notices, 21(10):153–162, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:1986:AMSa

[1030] Kim B. Bruce and Peter Wegner. An algebraic model of subtypes in objectoriented languages (draft). ACM SIG-PLAN Notices, 21(10):163–172, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1986:COO

[1031] Peter Wegner. Classification in objectoriented systems. ACM SIGPLAN Notices, 21(10):173–182, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agha:1986:FCA

[1032] Gul Agha. Fair concurrency in actors (abstract only): eager evaluation producers strong convergence. ACM SIGPLAN Notices, 21(10):183, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borgida:1986:CEB

[1033] Alex Borgida. Can entity-based information systems live with exceptions? (abstract only). ACM SIG-PLAN Notices, 21(10):183, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:1986:AMSb

[1034] Kim Bruce and Peter Wegner. An algebraic model of subtype in objectoriented languages (abstract only). ACM SIGPLAN Notices, 21(10):183– 184, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dahl:1986:OOS

[1035] Ole-Johann Dahl. Object-oriented specification (abstract only). ACM SIGPLAN Notices, 21(10):184, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deJong:1986:COA

[1036] Peter de Jong. Compiling objects into actors (abstract only). ACM SIG-PLAN Notices, 21(10):184, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goguen:1986:FEO

[1037] J. A. Goguen and J. Meseguer. Foundations and extensions of object-oriented programming (abstract only). ACM SIGPLAN Notices, 21(10):185, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hendler:1986:RTE

[1038] Jim Hendler. Run-time enhancements for object-oriented programming. ACM SIGPLAN Notices, 21(10):185, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahn:1986:OCLb

[1039] Ken Kahn. Objects in concurrent logic programming languages (abstract only). ACM SIGPLAN Notices, 21 (10):186, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madsen:1986:BSOb

[1040] Ole Madsen. Block structure and object oriented languages (abstract only). ACM SIGPLAN Notices, 21 (10):186, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nguyen:1986:GOMb

[1041] Van Nguyen and Brent Hailpern. A generalized object model (abstract only). ACM SIGPLAN Notices, 21 (10):187, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nygaard:1986:BCOb

[1042] Kristen Nygaard. Basic concepts in object oriented programming (abstract only). ACM SIGPLAN Notices, 21(10): 187–188, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossher:1986:SSL

[1043] Harold Ossher. Specifying the structure of large, layered, object-oriented programs (abstract only). ACM SIG-PLAN Notices, 21(10):188, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiss:1986:WGS

[1044] Steven Reiss. Working in the garden system (abstract only). ACM SIGPLAN Notices, 21(10):188, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1986:COO

[1045] Alan Snyder. CommonObjects: objectoriented programming in Common Lisp. ACM SIGPLAN Notices, 21 (10):189, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strom:1986:COOb

[1046] Rob Strom. A comparison of the object-oriented and process-oriented paradigms (abstract only). ACM SIG-PLAN Notices, 21(10):189, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stroustrup:1986:OCA

[1047] Bjarne Stroustrup. An overview of C++ (abstract only). ACM SIG-PLAN Notices, 21(10):189, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tokoro:1986:OKA

[1048] Mario Tokoro and Yutaka Ishikawa.

Orient84/K (abstract only): an object oriented concurrent programming language for knowledge systems. ACM SIGPLAN Notices, 21(10):190–191, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1986:CPC

[1049] Peter Wegner. Classification as a paradigm for computing (abstract only). ACM SIGPLAN Notices, 21 (10):191, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zdonik:1986:ACO

[1050] Stanley Zdonik. Approaches to change in an object-oriented database (abstract only). ACM SIGPLAN Notices, 21(10):191–192, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moon:1986:OOP

[1051] David A. Moon. Object-oriented programming with Flavors. ACM SIG-PLAN Notices, 21(11):1–8, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schaffert:1986:ITO

[1052] Craig Schaffert, Topher Cooper, Bruce Bullis, Mike Kilian, and Carrie Wilpolt. An introduction to Trellis/ Owl. ACM SIGPLAN Notices, 21(11): 9–16, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bobrow:1986:CLM

[1053] D. G. Bobrow, K. Kahn, G. Kiczales, L. Masinter, Stefik, M., and F. Zdybel. Common Loops, merging Lisp and object-oriented programming. ACM SIGPLAN Notices, 21(11): 17–29, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bobrow:1986:CML

[1054] Daniel G. Bobrow, Kenneth Kahn, Gregor Kiczales, Larry Masinter, Mark Stefik, and Frank Zdybel. Common-Loops: merging Lisp and objectoriented programming. ACM SIG-PLAN Notices, 21(11):17–29, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lang:1986:OOO

[1055] Kevin J. Lang and Barak A. Pearlmutter. Oaklisp: an object-oriented scheme with first class types. ACM SIGPLAN Notices, 21(11):30–37, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1986:EIO

[1056] Alan Snyder. Encapsulation and inheritance in object-oriented programming languages. ACM SIGPLAN Notices, 21(11):38–45, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ewing:1986:OOO

[1057] Juanita J. Ewing. An object-oriented operating system interface. ACM SIG-PLAN Notices, 21(11):46–56, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dasgupta:1986:PBM

[1058] Partha Dasgupta. A probe-based monitoring scheme for an objectoriented distributed operating system.

ACM SIGPLAN Notices, 21(11):57– 66, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1986:MMK

[1059] Michael B. Jones and Richard F. Rashid. Mach and Matchmaker: kernel and language support for object-oriented distributed systems. ACM SIGPLAN Notices, 21(11):67– 77, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Black:1986:OSE

[1060] Andrew Black, Norman Hutchinson, Eric Jul, and Henry Levy. Object structure in the Emerald system. ACM SIG-PLAN Notices, 21(11):78–86, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaehler:1986:VMN

[1061] Ted Kaehler. Virtual memory on a narrow machine for an object-oriented language. ACM SIGPLAN Notices, 21 (11):87–106, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Samples:1986:SSB

[1062] A. Dain Samples, David Ungar, and Paul Hilfinger. SOAR: Smalltalk without bytecodes. ACM SIGPLAN Notices, 21(11):107–118, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caudill:1986:TGS

[1063] Patrick J. Caudill and Allen Wirfs-Brock. A third generation Smalltalk-80 implementation. ACM SIGPLAN Notices, 21(11):119–130, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:1986:SFP

[1064] David M. Lewis, David R. Galloway, Robert J. Francis, and Brian W. Thomson. Swamp: a fast processor for Smalltalk-80. ACM SIGPLAN Notices, 21(11):131–139, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ballard:1986:QSD

[1065] Mark B. Ballard, David Maier, and Allen Wirfs-Brock. QUICKTALK: a Smalltalk-80 dialect for defining primitive methods. ACM SIGPLAN Notices, 21(11):140–150, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atkinson:1986:HOC

[1066] Robert G. Atkinson. Hurricane:

an optimizing compiler for Smalltalk. ACM SIGPLAN Notices, 21(11):151– 158, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mittal:1986:VCB

[1067] Sanja Mittal, Daniel G. Bobrow, and Kenneth M. Kahn. Virtual copies: at the boundary between classes and instances. ACM SIGPLAN Notices, 21 (11):159–166, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1986:ISO

[1068] Reid G. Smith, Rich Dinitz, and Paul Barth. Impulse-86. A substrate for object-oriented interface design2. ACM SIGPLAN Notices, 21(11):167– 176, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1986:EFD

[1069] David B. Anderson. Experience with Flamingo: a distributed, objectoriented user interface system. ACM SIGPLAN Notices, 21(11):177–185, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyrowitz:1986:IAC

[1070] Norman Meyrowitz. Intermedia: The architecture and construction of an object-oriented hypermedia system and applications framework. ACM SIGPLAN Notices, 21(11):186–201, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garrett:1986:BTE

[1071] L. Nancy Garrett and Karen E. Smith. Building a timeline editor from prefab parts: the architecture of an objectoriented application. ACM SIGPLAN Notices, 21(11):202–213, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberman:1986:UPO

[1072] Henry Lieberman. Using prototypical objects to implement shared behavior in object-oriented systems. ACM SIGPLAN Notices, 21(11):214– 223, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fukunaga:1986:EPB

[1073] Koichi Fukunaga and Shin ichi Hirose. An experience with a Prologbased object-oriented language. ACM SIGPLAN Notices, 21(11):224–231, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ishikawa:1986:COO

[1074] Yutaka Ishikawa and Mario Tokoro. A concurrent object-oriented knowledge representation language Orient84/K: its features and implementation. ACM SIGPLAN Notices, 21(11):232–241, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahn:1986:OCLc

[1075] K. Kahn, E. D. Tribble, M. S. Miller, and D. G. Bobrow. Objects in concurrent logic programming languages. ACM SIGPLAN Notices, 21(11):242, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahn:1986:OCL

[1076] Kenneth Kahn, Eric Dean Tribble, Mark S. Miller, and Daniel G. Bobrow. Objects in concurrent logic programming languages. ACM SIGPLAN Notices, 21(11):242–257, November 1986.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yonezawa:1986:OOC

[1077] Akinori Yonezawa, Jean-Pierre Briot, and Etsuya Shibayama. Objectoriented concurrent programming ABCL/1. ACM SIGPLAN Notices, 21 (11):258–268, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonar:1986:OOA

[1078] Jeffrey Bonar, Robert Cunningham, and Jamie Schultz. An object-oriented architecture for intelligent tutoring systems. ACM SIGPLAN Notices, 21 (11):269–276, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abdali:1986:SSA

[1079] S. Kamal Abdali, Guy W. Cherry, and Neil Soffer. A Smalltalk system for algebraic manipulation. ACM SIGPLAN Notices, 21(11):277–283, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruno:1986:PNB

[1080] Giorgio Bruno and Alessandra Balsamo. Petri net-based object-oriented modelling of distributed systems. ACM SIGPLAN Notices, 21(11):284–293, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1986:AAI

[1081] Michael S. Miller, Howard Cunningham, Chan Lee, and Steven R. Vegdahl. The Application Accelerator Illustration System. ACM SIGPLAN Notices, 21(11):294–302, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhaskar:1986:VIO

[1082] K. S. Bhaskar, J. K. Peckol, and J. L.

Beug. Virtual Instruments: objectoriented program synthesis. ACM SIGPLAN Notices, 21(11):303–314, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1986:TCS

[1083] Ralph E. Johnson. Type-checking Smalltalk. ACM SIGPLAN Notices, 21 (11):315–321, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaLonde:1986:EBS

[1084] Wilf R. LaLonde, Dave A. Thomas, and John R. Pugh. An exemplar based Smalltalk. ACM SIGPLAN Notices, 21(11):322–330, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yokote:1986:DIC

[1085] Yasuhiko Yokote and Mario Tokoro.

The design and implementation of Concurrent Smalltalk. ACM SIGPLAN Notices, 21(11):331–340, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pascoe:1986:ENS

[1086] Geoffrey A. Pascoe. Encapsulators: a new software paradigm in Smalltalk-80. ACM SIGPLAN Notices, 21(11):341– 346, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ingalls:1986:STH

[1087] Daniel H. H. Ingalls. A simple technique for handling multiple polymorphism. ACM SIGPLAN Notices, 21 (11):347–349, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargill:1986:PCS

[1088] T. A. Cargill. Pi: a case study in object-oriented programming. ACM SIGPLAN Notices, 21(11):350–360, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cunningham:1986:DOO

[1089] Ward Cunningham and Kent Beck. A diagram for object-oriented programs. ACM SIGPLAN Notices, 21(11):361– 367, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacky:1986:OOA

[1090] Jonathan Jacky and Ira Kalet. An object-oriented approach to a large scientific application. ACM SIGPLAN Notices, 21(11):368–376, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacobson:1986:LSC

[1091] Ivar Jacobson. Language support for changeable large real time systems. ACM SIGPLAN Notices, 21(11):377– 384, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Piersol:1986:OOS

[1092] Kurt W. Piersol. Object-oriented spreadsheets: the analytic spreadsheet package. ACM SIGPLAN Notices, 21(11):385–390, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyer:1986:GVI

[1093] Bertrand Meyer. Genericity versus inheritance. ACM SIGPLAN Notices, 21(11):391–405, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khoshafian:1986:OI

[1094] Setrag N. Khoshafian and George P.

Copeland. Object identity. ACM SIGPLAN Notices, 21(11):406–416, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McAllester:1986:BC

[1095] David McAllester and Ramin Zabih.

Boolean classes. ACM SIGPLAN Notices, 21(11):417–423, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandberg:1986:AS

[1096] David Sandberg. An alternative to subclassing. ACM SIGPLAN Notices, 21(11):424–428, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandberg:1986:ASS

[1097] D. Sandberg. An alternative to subclassing (smalltalk). ACM SIGPLAN Notices, 21(11):424–428, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olthoff:1986:AOO

[1098] Walter G. Olthoff. Augmentation of object-oriented programming by concepts of abstract data type theory: the ModPascal experience. ACM SIGPLAN Notices, 21(11):429–443, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Decouchant:1986:DDO

[1099] Dominique Decouchant. Design of a distributed object manager for the Smalltalk-80 system. ACM SIGPLAN Notices, 21(11):444–452, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiebe:1986:DRI

[1100] Douglas Wiebe. A distributed repository for immutable persistent objects. ACM SIGPLAN Notices, 21(11):453– 465, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vegdahl:1986:MSB

[1101] Steven R. Vegdahl. Moving structures between Smalltalk images. ACM SIGPLAN Notices, 21(11):466–471, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maier:1986:DOO

[1102] David Maier, Jacob Stein, Allen Otis, and Alan Purdy. Development of an object-oriented DBMS. ACM SIGPLAN Notices, 21(11):472– 482, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skarra:1986:MCT

[1103] Andrea H. Skarra and Stanley B.

Zdonik. The management of changing types in an object-oriented database. ACM SIGPLAN Notices, 21(11):483– 495, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rees:1986:RRA

[1104] Jonathan Rees, William Clinger, H. Abelson, N. I. Adams, IV, D. H. Bartley, G. Brooks, R. K. Dybvig, D. P. Friedman, R. Halstead, C. Hanson, C. T. Haynes, E. Kohlbecker, D. Oxley, K. M. Pitman, G. J. Rozas, G. J. Sussman, and M. Wand. Revised report^{3} on the algorithmic language SCHEME. ACM SIGPLAN Notices, 21(12):37– 79, December 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). MIT A. I. MEMO 848a.

Balanescu:1986:HER

[1105] T. Balanescu, S. Gavrila, M. Gheorghe, R. Nicolescu, and L. Sofonea. On Hartmann’s error-recovery scheme. ACM SIGPLAN Notices, 21(12):80–86, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boley:1986:RRF

[1106] H. Boley. RELFUN: a relational/ functional integration with valued clauses. ACM SIGPLAN Notices, 21 (12):87–98, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clapp:1986:SQP

[1107] R. M. Clapp, T. N. Mudge, and R. A. Volz. Solutions to the n Queens problem using tasking in Ada. ACM SIG-PLAN Notices, 21(12):99–110, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clocksin:1986:SCP

[1108] W. F. Clocksin. Some comments on “Pitfalls in Prolog Programming”. ACM SIGPLAN Notices, 21(12):111– 113, December 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [947].

Dreiheller:1986:PPP

[1109] A. Dreiheller, M. Moerschbacher, and B. Mohr. PHYSCAL: Programming Pascal with physical units. ACM SIG-PLAN Notices, 21(12):114–123, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gonsalves:1986:PSO

[1110] G. Gonsalves and A. C. Silvestri. Programming in Smalltalk-80: observations and remarks from the newly initiated. ACM SIGPLAN Notices, 21 (12):124–130, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joslin:1986:EPI

[1111] D. A. Joslin. Extended Pascal — illustrative features. ACM SIGPLAN Notices, 21(12):131–138, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mossenbock:1986:ASEb

[1112] H. Mössenböck. Alex: a simple and efficient scanner generator. ACM SIG-PLAN Notices, 21(12):139–148, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shyu:1986:SSL

[1113] Yuh-Huei H. Shyu. From semisyntactic lexical analyzer to a new compiler model. ACM SIGPLAN Notices, 21(12):149–157, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turner:1986:OM

[1114] D. Turner. An overview of Miranda.

ACM SIGPLAN Notices, 21(12):158– 166, December 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turner:1986:PPG

[1115] P. K. Turner. Up-down parsing with prefix grammars. ACM SIGPLAN Notices, 21(12):167–174, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Penedo:1987:PPM

[1116] Maria H. Penedo. Prototyping a project master database for software engineering environments. ACM SIG-PLAN Notices, 22(1):1–11, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gallo:1987:OMS

[1117] Ferdinando Gallo, Regis Minot, and Ian Thomas. The object management system of PCTE as a software engineering database management system. ACM SIGPLAN Notices, 22(1):12–15, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wile:1987:WOS

[1118] David S. Wile and Dennis G. Allard. Worlds: an organizing structure for object-bases. ACM SIGPLAN Notices, 22(1):16–26, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dowson:1987:IIP

[1119] Mark Dowson. ISTAR: an integrated project support environment. ACM SIGPLAN Notices, 22(1):27–33, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Delisle:1987:PEC

[1120] Norman Delisle and Mayer Schwartz.

A programming environment for CSP. ACM SIGPLAN Notices, 22(1):34–41, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ross:1987:ICP

[1121] Graham Ross. Integral-C: a practical environment for C programming. ACM SIGPLAN Notices, 22(1):42–48, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:1987:EVA

[1122] Charles Z. Mitchell. Engineering VAX Ada for a multi-language programming environment. ACM SIGPLAN Notices, 22(1):49–58, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kret:1987:IWS

[1123] M. A. Kret. An information workstation for software managers. ACM SIG-PLAN Notices, 22(1):59–69, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mital:1987:CSW

[1124] Rakesh M. Mital, Marcia M. Kim, and Richard A. Berg. A case study of workstation usage during the early phases of the software development life cycle.

ACM SIGPLAN Notices, 22(1):70–76, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wartik:1987:RES

[1125] Steven Wartik. Rapidly evolving software and the OVERSEE environment. ACM SIGPLAN Notices, 22(1):77–83, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Belkatir:1987:EDB

[1126] N. Belkatir and J. Estublier. Experience with a data base of programs. ACM SIGPLAN Notices, 22(1):84–91, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Copper:1987:ECW

[1127] Keith D. Copper, Ken Kennedy, and Linda Torczon. Editing and compiling whole programs. ACM SIGPLAN Notices, 22(1):92–101, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1987:SVS

[1128] Robert P. Cook and Rihard J. Auletta. StarLite, a visual simulation package for software prototyping. ACM SIG-PLAN Notices, 22(1):102–110, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Childs:1987:AAA

[1129] Carolyn Childs and Filippos I. Vokolos. AWB-ADE: an application development environment for interactive, integrated systems. ACM SIGPLAN Notices, 22(1):111–120, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marzullo:1987:JSS

[1130] Keith Marzullo and Douglas Wiebe.

Jasmine: a software system modelling facility. ACM SIGPLAN Notices, 22(1):121–130, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserman:1987:GEI

[1131] Anthony I. Wasserman and Peter A.

Pircher. A graphical, extensible integrated environment for software development. ACM SIGPLAN Notices, 22 (1):131–142, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

King:1987:IUI

[1132] Roger King and Scott Hudson. Implementing a user interface as a system of attributes. ACM SIGPLAN Notices, 22(1):143–149, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dewan:1987:DES

[1133] Prasun Dewan and Marvin Solomon.

Dost: an environment to support automatic generation of user interfaces. ACM SIGPLAN Notices, 22(1):150– 159, January 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Staudt:1987:SAM

[1134] Barbara J. Staudt, Charles W. Krueger, and David Garlan. A structural approach to the maintenance of structure-oriented environments. ACM SIGPLAN Notices, 22(1):160–170, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carroll:1987:IAS

[1135] Martin Carroll and Barbara G. Ryder. An incremental algorithm for software analysis. ACM SIGPLAN Notices, 22(1):171–179, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1987:ERM

[1136] Hausi A. Muller, Robert Hood, and Ken Kennedy. Efficient recompilation of module interfaces in a software development environment. ACM SIG-PLAN Notices, 22(1):180–189, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Engels:1987:SSO

[1137] G. Engels, M. Nagl, and W. Schafer. On the structure of structure-oriented editors for different applications. ACM SIGPLAN Notices, 22(1):190–198, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weiderman:1987:MEE

[1138] Nelson H. Weiderman, A. Nico Habermann, Mark W. Borger, and Mark H.

Klein. A methodology for evaluating environments. ACM SIGPLAN Notices, 22(1):199–207, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanLamsweerde:1987:KGS

[1139] A. van Lamsweerde, M. Buyse, B. Delcourt, E. Delor, M. Ervier, and M. C. Schayes. The kernel of a generic software development environment. ACM SIGPLAN Notices, 22(1):208–217, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reppy:1987:FPE

[1140] J. H. Reppy and E. R. Gansner. A foundation for programming environments. ACM SIGPLAN Notices, 22(1): 218–227, January 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

IEEE:1987:AIS

[1141] ANSI/IEEE Std 754-1985. an American National Standard: IEEE Standard for Binary Floating-Point Arithmetic. ACM SIGPLAN Notices, 22(2): 9–25, February 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://portalparts.acm.org/
30000/24686/fm/frontmatter.pdf.

See [1937].

Cowlishaw:1987:DRL

[1142] M. Cowlishaw. The design of the REXX language. ACM SIGPLAN Notices, 22(2):26–35, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:1987:WGG

[1143] G. Alan Creak. When GOTO goes to, how does it get there? ACM SIG-PLAN Notices, 22(2):36–42, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsby:1987:SQP

[1144] Michael E. Goldsby. Solving the “N ≤ 8 Queens” problem with CSP and Modula-2. ACM SIGPLAN Notices, 22(2):43–52, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hrycej:1987:KBP

[1145] Tomas Hrycej. A knowledge-based problem-specific program generator. ACM SIGPLAN Notices, 22(2):53– 61, February 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Louden:1987:RVN

[1146] Kenneth Louden. Recursion versus non-recursion in Pascal: recursion can be faster. ACM SIGPLAN Notices, 22(2):62–67, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perkins:1987:CDN

[1147] G. R. Perkins, R. W. Norman, and S. Danicic. Coping with deeply nested control structures. ACM SIGPLAN Notices, 22(2):68–77, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Savit:1987:UMA

[1148] Jeffrey Savit. Uninitialized Modula-2 abstract objects, revisited. ACM SIG-PLAN Notices, 22(2):78–84, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyer:1987:EPR

[1149] Bertrand Meyer. Eiffel: programming for reusability and extendibility. ACM SIGPLAN Notices, 22(2):85–94, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Slonneger:1987:PPP

[1150] Ken Slonneger. Pitfalls with procedure parameters. ACM SIGPLAN Notices, 22(2):95–99, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1987:LGG

[1151] David Spector and Prescott K. Turner.

Limitations of Graham-Glanville style code generation. ACM SIGPLAN Notices, 22(2):100–107, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harel:1987:LDC

[1152] D. Harel. Logic and databases: a critique. ACM SIGPLAN Notices, 22 (3):14–20, March 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pagan:1987:GFL

[1153] F. G. Pagan. A graphical FP language. ACM SIGPLAN Notices, 22(3):21–39, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rehmer:1987:EHU

[1154] K. Rehmer. Error handling using Modula-2. ACM SIGPLAN Notices, 22 (3):40–48, March 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kennedy:1987:EDC

[1155] S. C. Kennedy. Engineering development of computer languages. ACM SIGPLAN Notices, 22(3):49–58, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ejiogu:1987:CIS

[1156] L. O. Ejiogu. The critical issues of software metrics part 0: perspectives on software measurements. ACM SIGPLAN Notices, 22(3):59–64, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leeson:1987:TIM

[1157] J. J. Leeson and M. L. Spear. Type independent modules: the preferred approach to generic ADTs in Modula-2. ACM SIGPLAN Notices, 22(3):65–70, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parker:1987:SRP

[1158] J. R. Parker. A study of record packing methods. ACM SIGPLAN Notices, 22 (3):71–77, March 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Braunschober:1987:CCP

[1159] W. Braunschober. COMPAS: compatible Pascal. ACM SIGPLAN Notices, 22 (3):78–82, March 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1987:CAS

[1160] P. Wadler. A critique of Abelson and Sussman or why calculating is better than scheming. ACM SIG-PLAN Notices, 22(3):83–94, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lomax:1987:SLS

[1161] A. Lomax. The suitability of language syntaxes for program generation. ACM SIGPLAN Notices, 22(3):95–101, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LePera:1987:FVO

[1162] T. A. LePera and H. E. Donley. FOR-TRAN: a viable option for the introductory computer science course. ACM SIGPLAN Notices, 22(3):102– 110, March 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hankley:1987:FAT

[1163] W. Hankley. Feature analysis of Turbo Prolog. ACM SIGPLAN Notices, 22 (3):111–118, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tracz:1987:PPP

[1164] W. Tracz. Personal personnel programmer profile. ACM SIGPLAN Notices, 22(3):119–120, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schooler:1987:KBT

[1165] R. Schooler. Knowing is better than thinking: a simple approach to interprocedural optimization. ACM SIG-PLAN Notices, 22(3):121–122, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kerr:1987:MVS

[1166] R. Kerr. A materialistic view of the software “engineering” analogy. ACM SIGPLAN Notices, 22(3):123– 125, March 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rohl:1987:THR

[1167] J. S. Rohl. The Towers of Hanoi: a representation-free solution. ACM SIGPLAN Notices, 22(3):126, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watson:1987:GRN

[1168] J. F. Watson. A grammar rule notation translator. ACM SIGPLAN Notices, 22 (4):16–27, April 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1987:NMO

[1169] G. A. Hill. A note on the modulo operation in Edison. ACM SIGPLAN Notices, 22(4):28–29, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:1987:RRA

[1170] J. C. H. Park. Remarks on recent algorithms for LALR lookahead sets. ACM SIGPLAN Notices, 22(4):30–32, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radensky:1987:LEN

[1171] A. Radensky. Lazy evaluation and nondeterminism make Backus’ FP-systems more practical. ACM SIGPLAN Notices, 22(4):33–40, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goguen:1987:RRM

[1172] J. A. Goguen and J. Meseguer. Remarks on remarks on many-sorted equational logic. ACM SIGPLAN Notices, 22(4):41–48, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheese:1987:CCP

[1173] A. Cheese. Combinatory code and a parallel packet-based computational model. ACM SIGPLAN Notices, 22 (4):49–58, April 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hekmatpour:1987:NSM

[1174] S. Hekmatpour. A notation for specifying menus. ACM SIGPLAN Notices, 22(4):59–62, April 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jerardi:1987:PPL

[1175] T. W. Jerardi. Puzzles, PROLOG and logic. ACM SIGPLAN Notices, 22 (4):63–69, April 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harland:1987:OPO

[1176] D. M. Harland and B. Beloff. OB-JEKT: a persistent object store with an integrated garbage collector. ACM SIGPLAN Notices, 22(4):70–79, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kotulski:1987:ASN

[1177] L. Kotulski. About the semantic nested monitor calls. ACM SIGPLAN Notices, 22(4):80–82, April 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Methfessel:1987:IAO

[1178] R. Methfessel. Implementing an access and object oriented paradigm in a language that supports neither. ACM SIGPLAN Notices, 22(4):83–93, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ion:1987:PCC

[1179] I. Ion, R. Arhire, and M. Macesanu. Programs complexity: comparative analysis hierarchy, classification. ACM SIGPLAN Notices, 22(4):94–102, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eckart:1987:IAD

[1180] J. D. Eckart. Iteration and abstract data types. ACM SIGPLAN Notices, 22(4):103–110, April 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Torbett:1987:NPA

[1181] M. A. Torbett. A note on “Protecting against uninitialized abstract objects in Modula-2”. ACM SIGPLAN Notices, 22(5):8–10, May 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [966].

Torbett:1987:MAI

[1182] M. A. Torbett. More ambiguities and insecurities in Modula-2. ACM SIG-PLAN Notices, 22(5):11–17, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joslin:1987:EPI

[1183] D. A. Joslin. Extended Pascal — illustrative features. ACM SIGPLAN Notices, 22(5):18–19, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [1241].

Snyder:1987:MEC

[1184] W. Van Snyder. Multilevel EXIT and CYCLE aren’t so bad. ACM SIG-PLAN Notices, 22(5):20–22, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Touati:1987:AOO

[1185] H. Touati. Is Ada an object oriented programming language? ACM SIG-PLAN Notices, 22(5):23–26, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baldwin:1987:NHT

[1186] R. R. Baldwin. A note on H. E. Tompkins’s minimum-period COBOL style. ACM SIGPLAN Notices, 22(5):27–31, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1987:SSC

[1187] D. M. Miller, R. S. Maness, J. W. Howatt, and W. H. Shaw. A software science counting strategy for the full Ada language. ACM SIG-PLAN Notices, 22(5):32–41, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furnari:1987:PIL

[1188] M. Furnari. Pascal implementation of a LISP interpreter. ACM SIG-PLAN Notices, 22(5):42–46, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaplan:1987:PLL

[1189] I. Kaplan. Programming the Loral LDF 100 dataflow machine. ACM SIG-PLAN Notices, 22(5):47–57, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bondy:1987:UMA

[1190] J. Bondy. Uninitialized Modula-2 abstract object instances, yet again. ACM SIGPLAN Notices, 22(5):58–63, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gallaire:1987:LDR

[1191] H. Gallaire, J. Minker, and J. Nicolas.

Logic and databases: a response. ACM SIGPLAN Notices, 22(6):20–24, June 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Collado:1987:MIC

[1192] M. Collado, R. Morales, and J. J.

Moreno. A Modula-2 implementation of CSP. ACM SIGPLAN Notices, 22 (6):25–38, June 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sosnowski:1987:PDD

[1193] R. A. Sosnowski. Prolog dialects:

a déjà vu of BASICs. ACM SIG-PLAN Notices, 22(6):39–48, June 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheese:1987:MMR

[1194] A. Cheese. Multi-moded relations in Parlog. ACM SIGPLAN Notices, 22 (6):49–51, June 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenfield:1987:DUM

[1195] S. Greenfield and R. Norton. Detecting uninitialized Modula-2 abstract objects. ACM SIGPLAN Notices, 22 (6):52–58, June 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiener:1987:OOP

[1196] R. S. Wiener. Object-oriented programming in C++: a case study. ACM SIGPLAN Notices, 22(6):59–68, June 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gantenbein:1987:SDB

[1197] R. E. Gantenbein. Support for dynamic binding in strongly typed languages. ACM SIGPLAN Notices, 22 (6):69–75, June 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1987:IAR

[1198] J. C. Miller and B. M. Strauss III. Implications of automated restructuring of COBOL. ACM SIGPLAN Notices, 22(6):76–82, June 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rizk:1987:DIC

[1199] A. Rizk and F. Halsall. Design and implementation of a C-based language for distributed real-time systems. ACM SIGPLAN Notices, 22(6):83–100, June 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Indermark:1987:CFL

[1200] K. Indermark and H. Klaeren. Compiling Fibonacci-like recursion. ACM SIGPLAN Notices, 22(6):101–108, June 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krumme:1987:CEP

[1201] D. W. Krumme. Comments on an example for procedure parameters. ACM SIGPLAN Notices, 22(6): 109–111, June 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

May:1987:MFS

[1202] C. May. Mimic: a fast System/ 370 simulator. ACM SIGPLAN Notices, 22(7):1–13, July 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/29650/p1-may/.

Huguet:1987:BAG

[1203] M. Huguet, T. Lang, and Y. Tamir. A block-and-actions generator as an alternative to a simulator for collecting architecture measurements. ACM SIG-PLAN Notices, 22(7):14–25, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huguet:1987:BGA

[1204] M. Huguet, T. Lang, and Y. Tamir. A block-and-actions generator as an alternative to a simulator for collecting architecture measurements. ACM SIGPLAN Notices, 22(7):14–25, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p14-huguet/

.

Shebs:1987:ADI

[1205] S. T. Shebs and R. R. Kessler. Automatic design and implementation of language data types. ACM SIGPLAN Notices, 22(7):26–37, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p26-shebs/

.

Karinthi:1987:IRE

[1206] R. R. Karinthi and M. Weiser. Incremental re-execution of programs. ACM SIGPLAN Notices, 22(7):38–44, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karinthi:1987:IRP

[1207] R. R. Karinthi and M. Weiser. Incremental re-execution of programs. ACM SIGPLAN Notices, 22(7):38–44, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p38-karinthi/

.

Notkin:1987:ETE

[1208] D. Notkin and W. G. Griswold. Enhancement through extension: the extension interpreter. ACM SIG-PLAN Notices, 22(7):45–55, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p45-notkin/

.

Sutor:1987:TIC

[1209] R. S. Sutor and R. D. Jenks. The type inference and coercion facilities in the Scratchpad II interpreter. ACM SIGPLAN Notices, 22(7):56–63, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p56-sutor/

.

Robison:1987:IFP

[1210] A. D. Robison. The Illinois functional programming interpreter. ACM SIGPLAN Notices, 22(7):64–73, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p64-robison/

.

Waters:1987:EIS

[1211] R. C. Waters. Efficient interpretation of synchronizable series expressions. ACM SIGPLAN Notices, 22(7):74–85, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/plan/29650/p74-waters/

.

Faustini:1987:EIL

[1212] A. A. Faustini and W. W. Wadge. An eductive interpreter for the language pLucid. ACM SIGPLAN Notices, 22(7):86–91, July 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/29650/p86-faustini/

.

Bahlke:1987:GLD

[1213] R. Bahlke, B. Moritz, and G. Snelting. A generator for languagespecific debugging systems. ACM SIGPLAN Notices, 22(7):92–101, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/29650/p92-bahlke/

.

Bahlke:1987:GLS

[1214] R. Bahlke, B. Moritz, and G. Snelting. A generator for language-specific debugging systems. ACM SIGPLAN Notices, 22(7):92–101, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skedzielewski:1987:IDI

[1215] S. K. Skedzielewski, R. K. Yates, and R. R. Oldehoeft. DI: an interactive debugging interpreter for applicative languages. ACM SIG-PLAN Notices, 22(7):102–112, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p102-skedzielewski/

.

Chase:1987:SIT

[1216] B. B. Chase and R. T. Hood.

Selective interpretation as a technique for debugging computationally intensive programs. ACM SIG-PLAN Notices, 22(7):113–124, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p113-chase/

.

Krall:1987:IHP

[1217] A. Krall. Implementation of a highspeed Prolog interpreter. ACM SIG-PLAN Notices, 22(7):125–131, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p125-krall/

.

Krall:1987:IHS

[1218] A. Krall. Implementation of a highspeed Prolog interpreter. ACM SIG-PLAN Notices, 22(7):125–131, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barklund:1987:EIP

[1219] J. Barklund. Efficient interpretation of Prolog programs. ACM SIG-PLAN Notices, 22(7):132–137, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p132-barklund/

.

OBagy:1987:RII

[1220] J. O’Bagy and R. E. Griswold. A recursive interpreter for the Icon programming language. ACM SIG-PLAN Notices, 22(7):138–149, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p138-o_bagy/

.

Pittman:1987:THI

[1221] T. Pittman. Two-level hybrid interpreter/native code execution for combined space-time program efficiency. ACM SIGPLAN Notices, 22(7):150–152, July 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/29650/p150-pittman/

.

Pittman:1987:TLH

[1222] T. Pittman. Two-level hybrid interpreter/native code execution for combined space-time program efficiency. ACM SIGPLAN Notices, 22(7):150– 152, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koskimies:1987:TUA

[1223] K. Koskimies and J. Paakki. TOOLS: a unifying approach to object-oriented language interpretation. ACM SIG-PLAN Notices, 22(7):153–164, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p153-koskimies/

.

Johnson:1987:DTC

[1224] G. F. Johnson. GL — a denotational testbed with continuations and partial continuations as first-class objects. ACM SIGPLAN Notices, 22(7): 165–176, July 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1987:GDT

[1225] G. F. Johnson. GL — a denotational testbed with continuations and partial continuations as firstclass objects. ACM SIGPLAN Notices, 22(7):165–176, July 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/29650/p165-johnson/

.

Offutt:1987:FIM

[1226] A. J. Offutt, VI and K. N. King.

A Fortran 77 interpreter for mutation analysis. ACM SIGPLAN Notices, 22(7):177–188, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/29650/p177-offutt/

.

Davidson:1987:CRI

[1227] J. W. Davidson and J. V. Gresh.

Cint: a RISC interpreter for the C programming language. ACM SIG-PLAN Notices, 22(7):189–198, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p189-davidson/

.

Cordy:1987:DIE

[1228] James R. Cordy and T. C. N. Graham. Design of an interpretive environment for Turing. ACM SIG-PLAN Notices, 22(7):199–204, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p199-cordy/

.

Harr:1987:IAL

[1229] H. Harr, M. W. Evens, and J. Sprowl. Interpreting ABF — a language for document construction. ACM SIG-PLAN Notices, 22(7):205–213, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p205-harr/

.

Boehm:1987:CRI

[1230] H. J. Boehm. Constructive real interpretation of numerical programs. ACM SIGPLAN Notices, 22(7):214–221, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p214-boehm/

.

Clark:1987:JIR

[1231] C. F. Clark. The JADE interpreter: a RISC interpreter for syntax directed editing. ACM SIG-PLAN Notices, 22(7):222–228, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p222-clark/

.

Moss:1987:MSF

[1232] J. E. B. Moss. Managing stack frames in Smalltalk. ACM SIG-PLAN Notices, 22(7):229–240, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p229-moss/

.

Danvy:1987:MAH

[1233] O. Danvy. Memory allocation and higher-order functions. ACM SIG-PLAN Notices, 22(7):241–252, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/29650/p241-danvy/

.

Lang:1987:IIC

[1234] B. Lang and F. Dupont. Incremental incrementally compacting garbage collection. ACM SIG-PLAN Notices, 22(7):253–263, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/plan/29650/p253-lang/

.

Eckart:1987:DGC

[1235] J. D. Eckart and R. J. LeBlanc. Distributed garbage collection. ACM SIG-PLAN Notices, 22(7):264–273, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/29650/p264-eckart/

.

Gelernter:1987:PPM

[1236] D. Gelernter, S. Jagannathan, and T. London. Parallelism, persistence and meta-cleanliness in the symmetric Lisp interpreter. ACM SIG-PLAN Notices, 22(7):274–282, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/29650/p274-gelernter/

.

Kearns:1987:CIL

[1237] P. Kearns and M. Freeman. CCAL: an interpreted language for experimentation in concurrent control. ACM SIGPLAN Notices, 22(7):283–291, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/29650/p283-kearns/

.

Urdang:1987:BRP

[1238] Erik G. Urdang. Book review: The T Programming Language: A Dialect of Lisp by Stephen Slade, Prentice Hall, Inc. 1987. ACM SIGPLAN Notices, 22(8):13, August 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhujade:1987:VSB

[1239] Moreshwar R. Bhujade. Visual specification of blocks in programming languages. ACM SIGPLAN Notices, 22 (8):24–26, August 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nani:1987:ISC

[1240] G. Nani. Implementing separate compilations in Pascal. ACM SIGPLAN Notices, 22(8):27–32, August 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

House:1987:TEP

[1241] Ronald T. House. Thoughts on “extended Pascal — illustrative examples”. ACM SIGPLAN Notices, 22(8): 33–38, August 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1183].

Davies:1987:NMP

[1242] Gordon Davies and A. Tan. A note on metrics of Pascal programs. ACM SIG-PLAN Notices, 22(8):39–44, August 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See reply [1357].

Baldwin:1987:IPU

[1243] Geoff Baldwin. Implementation of physical units. ACM SIGPLAN Notices, 22(8):45–50, August 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lujun:1987:OOP

[1244] Shang Lujun and Sun Zhongxiu. An object-oriented programming language for developing distributed software. ACM SIGPLAN Notices, 22(8):51–56, August 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drechsler:1987:PXC

[1245] K. H. Drechsler and M. P. Stadel. The Pascal-XT code generator. ACM SIG-PLAN Notices, 22(8):57–78, August 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gerardy:1987:ECS

[1246] Robert Gerardy. Experimental comparison of some parsing methods. ACM SIGPLAN Notices, 22(8):79–88, August 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonsson:1987:PBS

[1247] Dan Jonsson. Pancode and boxcharts: structured programming revisited. ACM SIGPLAN Notices, 22(8): 89–98, August 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ives:1987:RRR

[1248] Fred Ives. Response to remarks on recent algorithms for LALR lookahead sets. ACM SIGPLAN Notices, 22(8): 99–104, August 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakkinen:1987:CVY

[1249] Markku Sakkinen. Comparison as a value-yielding operation. ACM SIG-PLAN Notices, 22(8):105–110, August 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rickert:1987:RVN

[1250] N. W. Rickert. Recursion vs. nonrecursion in Pascal: a rejoinder. ACM SIG-PLAN Notices, 22(9):12–13, September 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBos:1987:PPC

[1251] Jan van den Bos. PCOL: a protocol-constrained object language. ACM SIGPLAN Notices, 22(9):14– 19, September 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruf:1987:NIS

[1252] J. A. Ruf. A note introducing syntax through semantic routines. ACM SIG-PLAN Notices, 22(9):20–22, September 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nuutila:1987:XLE

[1253] E. Nuutila, J. Kuusela, M. Tamminen, J. Veilhti, and J. Arkko. XC: a language for embedded rule based systems. ACM SIGPLAN Notices, 22(9): 23–32, September 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casimir:1987:PP

[1254] R. J. Casimir. Prolog puzzles. ACM SIGPLAN Notices, 22(9):33– 37, September 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1987:WR

[1255] D. B. Benson and M. G. Main. Workshop report. ACM SIGPLAN Notices, 22(9):38–43, September 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1987:AS

[1256] I. B. Elliott. The astro system. ACM SIGPLAN Notices, 22(9):44– 53, September 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edelson:1987:HOM

[1257] D. R. Edelson. How objective mechanisms facilitate the development of large software systems in three programming languages. ACM SIGPLAN Notices, 22(9):54–63, September 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1987:HAP

[1258] Niklaus Wirth. Hardware architectures for programming languages and programming languages for hardware architectures. ACM SIGPLAN Notices, 22(10):2–8, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beck:1987:VAM

[1259] Bob Beck, Bob Kasten, and Shreekant Thakkar. VLSI assist for a multiprocessor. ACM SIGPLAN Notices, 22(10): 10–20, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bisiani:1987:ASM

[1260] Roberto Bisiani and Alessandro Forin.

Architectural support for multilanguage parallel programming on heterogeneous systems. ACM SIGPLAN Notices, 22(10):21–30, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rashid:1987:MIV

[1261] Richard Rashid, Avadis Tevanian, Michael Young, David Golub, and Robert Baron. Machine-independent virtual memory management for paged uniprocessor and multiprocessor architectures. ACM SIGPLAN Notices, 22 (10):31–39, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hayes:1987:ADE

[1262] John R. Hayes, Martin E. Fraeman, Robert L. Williams, and Thomas Zaremba. An architecture for the direct execution of the Forth programming language. ACM SIGPLAN Notices, 22(10):42–49, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steenkiste:1987:TTC

[1263] Peter Steenkiste and John Hennessy. Tags and type checking in LISP: hardware and software approaches. ACM SIGPLAN Notices, 22(10):50–59, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davidson:1987:EIS

[1264] Jack W. Davidson and Richard A. Vaughan. The effect of instruction set complexity on program size and memory performance. ACM SIGPLAN Notices, 22(10):60–64, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atkinson:1987:DP

[1265] Russell R. Atkinson and Edward M. McCreight. The dragon processor. ACM SIGPLAN Notices, 22(10):65– 69, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goodman:1987:CMV

[1266] James R. Goodman. Coherency for multiprocessor virtual address caches. ACM SIGPLAN Notices, 22(10):72– 81, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargill:1987:CHS

[1267] T. A. Cargill and B. N. Locanthi. Cheap hardware support for software debugging and profiling. ACM SIG-PLAN Notices, 22(10):82–83, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Georgiou:1987:ECI

[1268] C. J. Georgiou, S. L. Palmer, and P. L.

Rosenfeld. An experimental coprocessor for implementing persistent objects on an IBM 4381. ACM SIGPLAN Notices, 22(10):84–87, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Magenheimer:1987:IMD

[1269] Daniel J. Magenheimer, Liz Peters, Karl Pettis, and Dan Zuras. Integer multiplication and division on the HP Precision Architecture. ACM SIG-PLAN Notices, 22(10):90–99, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:1987:MEU

[1270] David W. Wall and Michael L. Powell. The Mahler experience: using an intermediate language as the machine description. ACM SIGPLAN Notices, 22(10):100–104, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weiss:1987:SSC

[1271] Shlomo Weiss and James E. Smith.

A study of scalar compilation techniques for pipelined supercomputers. ACM SIGPLAN Notices, 22(10):105– 109, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bush:1987:CSR

[1272] William R. Bush, A. Dain Samples, David Ungar, and Paul N. Hilfinger.

Compiling Smalltalk-80 to a RISC. ACM SIGPLAN Notices, 22(10):112– 116, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chow:1987:HMA

[1273] F. Chow, S. Correll, M. Himelstein, E. Killian, and L. Weber. How many addressing modes are enough? ACM SIGPLAN Notices, 22(10):117– 121, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Massalin:1987:SLS

[1274] Henry Massalin. Superoptimizer: a look at the smallest program. ACM SIGPLAN Notices, 22(10):122– 126, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taki:1987:PAE

[1275] Kazuo Taki, Katzuto Nakajima, Hiroshi Nakashima, and Morihiro Ikeda. Performance and architectural evaluation of the PSI machine. ACM SIG-PLAN Notices, 22(10):128–135, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borriello:1987:RVC

[1276] Gaetano Borriello, Andrew R. Cherenson, Peter B. Danzig, and Michael N. Nelson. RISCs vs. CISCs for Prolog: a case study. ACM SIGPLAN Notices, 22(10):136–145, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kieburtz:1987:RAS

[1277] Richard B. Kieburtz. A RISC architecture for symbolic computation. ACM SIGPLAN Notices, 22(10):146– 155, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ditzel:1987:DTS

[1278] David R. Ditzel, Hubert R. McLellan, and Alan D. Berenbaum. Design tradeoffs to support the C programming language in the CRISP microprocessor. ACM SIGPLAN Notices, 22(10): 158–163, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thacker:1987:FMW

[1279] Charles P. Thacker and Lawrence C.

Stewart. Firefly: a multiprocessor workstation. ACM SIGPLAN Notices, 22(10):164–172, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1987:PPV

[1280] Douglas W. Clark. Pipelining and performance in the VAX 8800 processor. ACM SIGPLAN Notices, 22(10): 173–177, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Colwell:1987:VAT

[1281] Robert P. Colwell, Robert P. Nix, John J. O’Donnell, David B. Papworth, and Paul K. Rodman. A VLIW architecture for a trace scheduling compiler. ACM SIGPLAN Notices, 22(10):

180–192, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levinthal:1987:PCG

[1282] Adam Levinthal, Pat Hanrahan, Mike Paquette, and Jim Lawson. Parallel computers for graphics applications. ACM SIGPLAN Notices, 22(10):193– 198, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1987:ZCP

[1283] J. E. Smith, G. E. Dermer, B. D. Vanderwarn, S. D. Klinger, and C. M. Rozewski. The ZS-1 central processor. ACM SIGPLAN Notices, 22(10): 199–204, October 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stone:1987:IBS

[1284] D. L. Stone and J. R. Nestor. IDL: Background and status. ACM SIG-PLAN Notices, 22(11):5–9, November 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snodgrass:1987:DII

[1285] R. Snodgrass. Displaying IDL instances. ACM SIGPLAN Notices, 22 (11):10–17, November 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Warren:1987:TIU

[1286] W. B. Warren, J. Kickenson, and R. Snodgrass. A tutorial introduction to using IDL. ACM SIGPLAN Notices, 22(11):18–34, November 1987.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newcomer:1987:EBI

[1287] J. M. Newcomer. Efficient binary I/O of IDL objects. ACM SIGPLAN Notices, 22(11):35–43, November 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lamb:1987:ISD

[1288] D. A. Lamb. Implementation strategies for DIANA attributes. ACM SIG-PLAN Notices, 22(11):44–54, November 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rohrich:1987:GAM

[1289] J. Rohrich. Graph attribution with multiple attribute grammars. ACM SIGPLAN Notices, 22(11):55– 70, November 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Didriksen:1987:IDD

[1290] T. Didriksen, A. Lie, and R. Conradi.

IDL as a data description language for a programming environment database. ACM SIGPLAN Notices, 22(11):71– 78, November 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Persch:1987:EID

[1291] G. Persch. Editing IDL data structures. ACM SIGPLAN Notices, 22(11): 79–86, November 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feiler:1987:RBI

[1292] P. H. Feiler. Relationship between IDL and structure editor generation technology. ACM SIGPLAN Notices, 22(11):87–94, November 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garlan:1987:EIS

[1293] D. Garlan. Extending IDL to support concurrent views. ACM SIGPLAN Notices, 22(11):95–110, November 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kerr:1987:UOO

[1294] R. K. Kerr and D. B. Percival. Use of object-oriented programming in a time series analysis system. ACM SIG-PLAN Notices, 22(12):1–10, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kempf:1987:TOO

[1295] Renate Kempf and Marilyn Stelzner. Teaching object-oriented programming with the KEE system. ACM SIGPLAN Notices, 22(12):11–25, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lucco:1987:PPV

[1296] Steven E. Lucco. Parallel programming in a virtual object space. ACM SIG-PLAN Notices, 22(12):26–34, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grimshaw:1987:MOO

[1297] Andrew S. Grimshaw and Jane W. S.

Liu. Mentat: An object-oriented macro data flow system. ACM SIGPLAN Notices, 22(12):35–47, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borning:1987:CH

[1298] Alan Borning, Robert Duisberg, Bjorn Freeman-Benson, Axel Kramer, and Michael Woolf. Constraint hierarchies. ACM SIGPLAN Notices, 22(12):48– 60, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Diederich:1987:OOD

[1299] Jim Diederich and Jack Milton. An object-oriented design system shell. ACM SIGPLAN Notices, 22(12):61– 77, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Young:1987:OOF

[1300] Robert L. Young. An object-oriented framework for interactive data graphics. ACM SIGPLAN Notices, 22(12): 78–90, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OBrien:1987:TPE

[1301] Patrick D. O’Brien, Daniel C. Halbert, and Michael F. Kilian. The Trellis programming environment. ACM SIG-PLAN Notices, 22(12):91–102, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merrow:1987:PSS

[1302] Thomas Merrow and Jame Laursen. A pragmatic system for shared persistent objects. ACM SIGPLAN Notices, 22(12):103–110, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Penney:1987:CMG

[1303] D. Jason Penney and Jacob Stein. Class modification in the GemStone object-oriented DBMS. ACM SIG-PLAN Notices, 22(12):111–117, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:1987:COS

[1304] Won Kim, Jay Banerjee, Hong-Tai Chou, Jorge F. Garza, and Darrel Woelk. Composite object support in an object-oriented database system. ACM SIGPLAN Notices, 22(12):118– 125, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caplinger:1987:ISB

[1305] Michael Caplinger. An information system based on distributed objects. ACM SIGPLAN Notices, 22(12):126– 137, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stein:1987:DI

[1306] Lynn Andrea Stein. Delegation is inheritance. ACM SIGPLAN Notices, 22(12):138–146, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maes:1987:CEC

[1307] Pattie Maes. Concepts and experiments in computational reflection. ACM SIGPLAN Notices, 22(12):147– 155, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cointe:1987:MFC

[1308] Pierre Cointe. Metaclasses are first class: The ObjVlisp Model. ACM SIG-PLAN Notices, 22(12):156–162, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1987:DOB

[1309] Peter Wegner. Dimensions of objectbased language design. ACM SIG-PLAN Notices, 22(12):168–182, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacobson:1987:OOD

[1310] Ivar Jacobson. Object-oriented development in an industrial environment. ACM SIGPLAN Notices, 22(12):183– 191, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barry:1987:UOD

[1311] Brian M. Barry and John R. Altoft. Using objects to design and build radar ESM systems. ACM SIGPLAN Notices, 22(12):192–201, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seidewitz:1987:OOP

[1312] Ed Seidewitz. Object-oriented programming in Smalltalk and ADA. ACM SIGPLAN Notices, 22(12):202– 213, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kempf:1987:EC

[1313] James Kempf, Warren Harris, Roy D’Souza, and Alan Snyder. Experience with CommonLoops. ACM SIG-PLAN Notices, 22(12):214–226, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungar:1987:SPS

[1314] David Ungar and Randall B. Smith. Self: The power of simplicity. ACM SIGPLAN Notices, 22(12):227–242, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nierstrasz:1987:AOH

[1315] O. M. Nierstrasz. Active objects in hybrid. ACM SIGPLAN Notices, 22 (12):243–253, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaiser:1987:MDF

[1316] Gail E. Kaiser and David Garlan. MELDing data flow and objectoriented programming. ACM SIG-PLAN Notices, 22(12):254–267, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thorelli:1987:MTC

[1317] Lars-Erik Thorelli. Modules and type checking in PL/LL. ACM SIG-PLAN Notices, 22(12):268–276, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Call:1987:COS

[1318] Lisa A. Call, David L. Cohrs, and Barton P. Miller. CLAM: an open system for graphical user interfaces. ACM SIGPLAN Notices, 22(12):277– 286, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alexander:1987:PPS

[1319] James H. Alexander. Painless panes for Smalltalk windows. ACM SIG-PLAN Notices, 22(12):287–294, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grossman:1987:LCO

[1320] Mark Grossman and Raimund K. Ege.

Logical composition of object-oriented interfaces. ACM SIGPLAN Notices, 22 (12):295–306, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gutfreund:1987:MT

[1321] Steven H. Gutfreund. Maniplicons in ThinkerToy. ACM SIGPLAN Notices, 22(12):307–317, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bennett:1987:DID

[1322] John K. Bennett. The design and implementation of distributed Smalltalk. ACM SIGPLAN Notices, 22(12):318– 330, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCullough:1987:TFF

[1323] Paul L. McCullough. Transparent forwarding: First steps. ACM SIG-PLAN Notices, 22(12):331–341, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Filman:1987:RO

[1324] Robert E. Filman. Retrofitting objects. ACM SIGPLAN Notices, 22(12):342– 353, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miranda:1987:BPS

[1325] Eliot Miranda. BrouHaHa: a portable SMALLTALK interpreter. ACM SIG-PLAN Notices, 22(12):354–365, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderMeulen:1987:IIS

[1326] Pieter S. van der Meulen. INSIST: Interactive Simulation in Smalltalk. ACM SIGPLAN Notices, 22(12):366– 376, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laursen:1987:OSP

[1327] Jane Laursen and Robert Atkinson. Opus: A Smalltalk production system.

ACM SIGPLAN Notices, 22(12):377– 387, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minsky:1987:LAO

[1328] Naftaly H. Minsky and David Rozenshtein. A law-based approach to objectoriented programming. ACM SIG-PLAN Notices, 22(12):382–493, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hogg:1987:OAO

[1329] John Hogg and Steven Weiser. OTM:

Applying objects to tasks. ACM SIG-PLAN Notices, 22(12):388–393, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bezivin:1987:SEO

[1330] Jean Bezivin. Some experiments in object-oriented simulation. ACM SIG-PLAN Notices, 22(12):394–405, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yokote:1987:EEC

[1331] Yasuhiko Yokote and Mario Tokoro.

Experience and evolution of concurrent Smalltalk. ACM SIGPLAN Notices, 22(12):406–415, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freburger:1987:RPC

[1332] Karl Freburger. RAPID: Prototyping control panel interfaces. ACM SIG-PLAN Notices, 22(12):416–422, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cox:1987:PTT

[1333] Brad J. Cox and Kurt J. Schmucker. Producer: a tool for translating Smalltalk-80 to Objective-C. ACM SIGPLAN Notices, 22(12):423–429, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andrews:1987:CLD

[1334] Timothy Andrews and Craig Harris. Combining language and database advances in an object-oriented development environment. ACM SIG-PLAN Notices, 22(12):430–440, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bloom:1987:IDO

[1335] Toby Bloom and Stanley B. Zdonik. Issues in the design of object-oriented database programming languages. ACM SIGPLAN Notices, 22(12):441– 451, December 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1987:ICS

[1336] Karen E. Smith and Stanley B. Zdonik. Intermedia: a case study of the differences between relational and objectoriented database systems. ACM SIG-PLAN Notices, 22(12):452–465, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rumbaugh:1987:RSC

[1337] James Rumbaugh. Relations as semantic constructs in an object-oriented language. ACM SIGPLAN Notices, 22 (12):466–481, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minsky:1987:LBA

[1338] Naftaly H. Minsky and David Rozenshtein. A law-based approach to objectoriented programming. ACM SIG-PLAN Notices, 22(12):482–493, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kerth:1987:WSD

[1339] Norman Kerth. Workshop on specification and design. ACM SIG-PLAN Notices, 23(5):9–16, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1987:KAD

[1340] Barbara Liskov. Keynote address — data abstraction and hierarchy. ACM SIGPLAN Notices, 23(5):17–34, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1987:PI

[1341] Peter Wegner. Panel on inheritance.

ACM SIGPLAN Notices, 23(5):35–40, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1987:BI

[1342] William Cook. BOF on inheritance. ACM SIGPLAN Notices, 23(5):41–42, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberman:1987:TO

[1343] Henry Lieberman, Lynn Stein, and David Ungar. Treaty of Orlando. ACM SIGPLAN Notices, 23(5):43–44, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caruso:1987:WOS

[1344] Michael Caruso and Stan Zdonik. Workshop on OODB semantics. ACM SIGPLAN Notices, 23(5):45–50, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deutsch:1987:WT

[1345] Peter Deutsch. Workshop on tools. ACM SIGPLAN Notices, 23(5):53–58, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1987:WCO

[1346] Ralph Johnson. Workshop on compiling and optimizing. ACM SIG-PLAN Notices, 23(5):59–65, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deutsch:1987:SSE

[1347] Peter Deutsch. SMALLTALK standardization efforts. ACM SIGPLAN Notices, 23(5):67–68, May 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zdonik:1987:PO

[1348] Stan Zdonik. Panel on OODB. ACM SIGPLAN Notices, 23(5):69–71, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thatte:1987:WOI

[1349] Satish Thatte. Workshop on OODB implementation. ACM SIGPLAN Notices, 23(5):73–87, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1987:PDM

[1350] Reid Smith. Panel on design methodology. ACM SIGPLAN Notices, 23 (5):91–95, May 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jackson:1987:BSO

[1351] Michael Jackson. Banquet speech — objects and other subjects. ACM SIGPLAN Notices, 23(5):97–104, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Uebbing:1987:PMP

[1352] John Uebbing. Panel on making products. ACM SIGPLAN Notices, 23(5): 105–111, May 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1987:PVP

[1353] David Smith. Panel on visual programming. ACM SIGPLAN Notices, 23 (5):113–118, May 1987. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moss:1987:PC

[1354] Eliot Moss. Panel on concurrency. ACM SIGPLAN Notices, 23(5):119– 127, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bobrow:1988:CLO

[1355] Daniel G. Bobrow, Linda G. DeMichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. Common LISP Object System specification X3J13 Document 88-002R. ACM SIGPLAN Notices, 23(SI): 1–142, September 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1988:WOO

[1356] Peter Wegner. Workshop on objectoriented programming ECOOP 1987, Paris, June 18, 1987. ACM SIG-PLAN Notices, 23(1):16–37, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scheider:1988:RNM

[1357] Victor Scheider. A reply to “A note on metrics of Pascal programs”. ACM SIGPLAN Notices, 23(1):38–39, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1242].

Schneider:1988:AHS

[1358] Victor Schneider. Approximations for the Halstead software science software error rate and project effort estimators. ACM SIGPLAN Notices, 23(1): 40–47, January 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fairfield:1988:DFA

[1359] P. Fairfield and M. A. Hennell. Data flow analysis of recursive procedures. ACM SIGPLAN Notices, 23(1):48–57, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burns:1988:PFL

[1360] A. Burns and G. Davies. Pascal-FC: a language for teaching concurrent programming. ACM SIGPLAN Notices, 23(1):58–66, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deVries:1988:FPF

[1361] Fer-Jan de Vries. A functional program for the fast Fourier transform. ACM SIGPLAN Notices, 23(1):67–74, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joy:1988:SGR

[1362] Mike Joy and Tom Axford. A standard for a graph representation for functional programs. ACM SIGPLAN Notices, 23(1):75–82, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chao:1988:DSP

[1363] Shan-Jon Chao and Barrett R. Bryant. Denotational semantics for program analysis. ACM SIGPLAN Notices, 23 (1):83–91, January 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhasker:1988:IOC

[1364] J. Bhasker. Implementation of an optimizing compiler for VHDL. ACM SIG-PLAN Notices, 23(1):92–108, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLenna:1988:FRP

[1365] B. J. MacLenna. Four relational programs. ACM SIGPLAN Notices, 23(1): 109–119, January 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanGestel:1988:RMS

[1366] Eric van Gestel and Ivo van Horebeek. Remarks on many-sorted equational logic with respect to algebraic implementations. ACM SIGPLAN Notices, 23(1):120–126, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berghel:1988:MPI

[1367] Hal Berghel, Richard Rankin, and Sanjiv Bhatia. MS-DOS Prolog implementations revisited. ACM SIGPLAN Notices, 23(1):127–134, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bierman:1988:PIO

[1368] Keith H. Bierman. A proposal for improving optimizer quality via dynamic analysis. ACM SIGPLAN Notices, 23(1):135–136, January 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perelgut:1988:TPC

[1369] S. Perelgut and J. R. Cordy. Turing Plus: a comparison with C and Pascal. ACM SIGPLAN Notices, 23(1): 137–143, January 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trickey:1988:CVL

[1370] Howard Trickey. C++ versus LISP: a case study. ACM SIGPLAN Notices, 23 (2):9–18, February 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sestoft:1988:BPE

[1371] Peter Sestoft and Harald Søndergaard.

A bibliography on partial evaluation. ACM SIGPLAN Notices, 23(2):19– 26, February 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sorensen:1988:SPC

[1372] P. F. Sorensen. In search of program complexity. ACM SIGPLAN Notices, 23(2):28–35, February 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Birtwistle:1988:MPP

[1373] Graham Birtwistle and Ken Loose. A model for procedures passed as parameters. ACM SIGPLAN Notices, 23(2): 36–43, February 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ponder:1988:BSa

[1374] Carl G. Ponder. Benchmark semantics. ACM SIGPLAN Notices, 23(2): 44–48, February 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hemmendinger:1988:UPS

[1375] David Hemmendinger. Unfair process scheduling in Modula-2. ACM SIGPLAN Notices, 23(3):7–12, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calliss:1988:PAR

[1376] Frank W. Calliss. Problems with automatic restructurers. ACM SIG-PLAN Notices, 23(3):13–21, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Egenhofer:1988:PMT

[1377] Max J. Egenhofer and Andrew U. Frank. A precompiler for modular, transportable Pascal. ACM SIG-PLAN Notices, 23(3):22–32, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yeh:1988:ACI

[1378] Dashing Yeh and Uwe Kastens. Automatic construction of incremental LR(1) — parsers. ACM SIGPLAN Notices, 23(3):33–42, March 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furtado:1988:TFP

[1379] Antonio L. Furtado. Towards functional programming in Prolog. ACM SIGPLAN Notices, 23(3):43–52, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soundararajan:1988:RSP

[1380] Neelam Soundararajan and Roger L.

Costello. Responsive sequential processes. ACM SIGPLAN Notices, 23 (3):53–62, March 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klerer:1988:NBT

[1381] Melvin Klerer and Hong Liu. A new benchmark test to estimate optimization quality of compilers. ACM SIGPLAN Notices, 23(3):63–72, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1988:WDM

[1382] Joseph Bergin and Stuart Greenfield.

What does Modula-2 need to fully support object oriented programming? ACM SIGPLAN Notices, 23(3):73–82, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holsti:1988:UFP

[1383] Niklas Holsti. Using formal procedure parameters to represent and transmit complex data structures. ACM SIGPLAN Notices, 23(3):83–92, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coward:1988:DPF

[1384] P. David Coward. Determining path feasibility for commercial programs. ACM SIGPLAN Notices, 23(3):93–101, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maguire:1988:PME

[1385] Gerald Q. Maguire, Jr. and Jonathan M. Smith. Process migration: effects on scientific computation. ACM SIG-PLAN Notices, 23(3):102–106, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ranger:1988:SCF

[1386] Phillippe Ranger. Some comments on the forthcoming “extended Pascal” standard. ACM SIGPLAN Notices, 23(3):107–110, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Molloy:1988:CDX

[1387] Mark Molloy. Comment on dpANS X3.160-198x, extended Pascal concerning the initial value clause. ACM SIGPLAN Notices, 23(4):22–26, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morrison:1988:BPP

[1388] R. Morrison, M. P. Atkinson, A. L.

Brown, and A. Dearle. Bindings persistent programming languages. ACM SIGPLAN Notices, 23(4):27–34, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1988:ECF

[1389] Keith D. Cooper and Ken Kennedy. Efficient computation of flow-insensitive interprocedural summary information — a correction. ACM SIGPLAN Notices, 23(4):35–42, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:1988:VPL

[1390] Alistair D. N. Edwards. Visual programming languages: the next generation. ACM SIGPLAN Notices, 23 (4):43–50, April 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rickert:1988:RLS

[1391] Neil W. Rickert. The role of the language standards committee. ACM SIGPLAN Notices, 23(4):51–55, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhatti:1988:IEE

[1392] M. A. Bhatti. Incremental execution environment. ACM SIGPLAN Notices, 23(4):56–64, April 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarwar:1988:IFL

[1393] S. M. Sarwar, S. J. Hahn, and J. A. Davis. Implementing functional languages on a combinator-based reduction machine. ACM SIGPLAN Notices, 23(4):65–70, April 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1988:CPL

[1394] Chang Li. Concurrent programming language — LISPTALK. ACM SIGPLAN Notices, 23(4):71–80, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hemmendinger:1988:HPPa

[1395] David Hemmendinger. The “Hamming problem” in Prolog. ACM SIGPLAN Notices, 23(4):81–86, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cattaneo:1988:CLI

[1396] Giuseppe Cattaneo and Vincenzo Loia. A Common-LISP implementation of an extended Prolog system. ACM SIGPLAN Notices, 23(4):87–102, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terwilliger:1988:PLC

[1397] Robert B. Terwilliger. Please: a language combining imperative and logic programming. ACM SIGPLAN Notices, 23(4):103–110, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kerth:1988:WSD

[1398] Norman Kerth. Workshop on specification and design. ACM SIG-PLAN Notices, 23(5):9–16, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1988:KAD

[1399] Barbara Liskov. Keynote address — data abstraction and hierarchy. ACM SIGPLAN Notices, 23(5):17–34, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1988:PI

[1400] Peter Wegner. Panel on inheritance.

ACM SIGPLAN Notices, 23(5):35–40, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1988:BI

[1401] William Cook. BOF on inheritance.

ACM SIGPLAN Notices, 23(5):41–42, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberman:1988:TO

[1402] Henry Lieberman, Lynn Stein, and David Ungar. Treaty of Orlando. ACM SIGPLAN Notices, 23(5):43–44, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caruso:1988:WOS

[1403] Michael Caruso and Stan Zdonik.

Workshop on OODB semantics. ACM SIGPLAN Notices, 23(5):45–50, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deutsch:1988:WT

[1404] Peter Deutsch. Workshop on tools. ACM SIGPLAN Notices, 23(5):53–58, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1988:WCO

[1405] Ralph Johnson. Workshop on compiling and optimizing. ACM SIG-PLAN Notices, 23(5):59–65, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deutsch:1988:SSE

[1406] Peter Deutsch. Smalltalk standardization efforts. ACM SIGPLAN Notices, 23(5):67–68, May 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zdonik:1988:PO

[1407] Stan Zdonik. Panel on OODB. ACM SIGPLAN Notices, 23(5):69–71, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thatte:1988:WOI

[1408] Satish Thatte. Workshop on OODB implementation. ACM SIGPLAN Notices, 23(5):73–87, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1988:PDM

[1409] Reid Smith. Panel on design methodology. ACM SIGPLAN Notices, 23 (5):91–95, May 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jackson:1988:BSO

[1410] Michael Jackson. Banquet speech — objects and other subjects. ACM SIGPLAN Notices, 23(5):97–104, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Uebbing:1988:PMP

[1411] John Uebbing. Panel on making products. ACM SIGPLAN Notices, 23(5): 105–111, May 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1988:PVP

[1412] David Smith. Panel on visual programming. ACM SIGPLAN Notices, 23 (5):113–118, May 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moss:1988:PC

[1413] Eliot Moss. Panel on concurrency.

ACM SIGPLAN Notices, 23(5):119– 127, May 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1988:AR

[1414] P. E. Cohen. An abundance of registers. ACM SIGPLAN Notices, 23 (6):24–34, June 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1988:PES

[1415] Victor Schneider. A pedagogic expert system shell in Prolog. ACM SIG-PLAN Notices, 23(6):35–38, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1988:PS

[1416] Irving B. Elliott. The PRESTO system. ACM SIGPLAN Notices, 23 (6):39–48, June 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beaudet:1988:SOO

[1417] Philip J. Beaudet and Michael A. Jenkins. Simulating the object-oriented paradigm to Nial. ACM SIGPLAN Notices, 23(6):49–58, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaindl:1988:PS

[1418] Hermann Kaindl. Portability of software. ACM SIGPLAN Notices, 23 (6):59–68, June 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anand:1988:CF

[1419] N. Anand. Clarify function! ACM SIGPLAN Notices, 23(6):69–79, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1988:OOP

[1420] George H. Roberts. OPG: an optimizing parser generator. ACM SIG-PLAN Notices, 23(6):80–90, June 1988.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Osterbye:1988:ADT

[1421] Kasper Osterbye. Abstract data types with shared operations. ACM SIG-PLAN Notices, 23(6):91–96, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:1988:YPC

[1422] Joseph C. H. Park. y+: a yacc preprocessor for certain semantic actions. ACM SIGPLAN Notices, 23(6):97–106, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philbrow:1988:PAD

[1423] P. C. Philbrow, I. Armour, M. P. Atkinson, and J. Livingstone. PS-Algol’s device-independent output statement. ACM SIGPLAN Notices, 23(6):107– 114, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hemmendinger:1988:HPPb

[1424] David Hemmendinger. The ‘Hamming’ problem in Prolog. ACM SIG-PLAN Notices, 23(6):115–120, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:1988:SLL

[1425] Peter Kokol. Spreadsheet language level: how high is it? ACM SIG-PLAN Notices, 23(6):121–134, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ponder:1988:ALI

[1426] Carl G. Ponder, Patrick McGeer, and Anthony P-C. Ng. Are applicative languages inefficient? ACM SIG-PLAN Notices, 23(6):135–139, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dwyer:1988:RRP

[1427] Barry Dwyer. Regular right part programming languages. ACM SIG-PLAN Notices, 23(6):140–144, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gerard:1988:AAD

[1428] Scott N. Gerard. Adding an authorization dimension to strong type checking. ACM SIGPLAN Notices, 23(6): 145–151, June 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1988:FDP

[1429] Barton P. Miller. The frequency of dynamic pointer references in “C” programs. ACM SIGPLAN Notices, 23 (6):152–156, June 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ponder:1988:BSb

[1430] Carl G. Ponder. Benchmark semantics. ACM SIGPLAN Notices, 23(6): 157–161, June 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chase:1988:SCS

[1431] D. R. Chase. Safety consideration for storage allocation optimizations. ACM SIGPLAN Notices, 23(7):1–10, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p1-chase/

.

Appel:1988:RCC

[1432] A. W. Appel, J. R. Ellis, and K. Li. Real-time concurrent collection on stock multiprocessors. ACM SIGPLAN Notices, 23(7):11–20, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p11-appel/

.

Appel:1988:RTC

[1433] A. W. Appel, J. R. Ellis, and K. Li. Real-time concurrent collection on stock multiprocessors. ACM SIG-PLAN Notices, 23(7):11–20, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Larus:1988:DCB

[1434] J. R. Larus and P. N. Hilfinger.

Detecting conflicts between structure accesses. ACM SIGPLAN Notices, 23(7):24–31, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p24-larus/

.

Horwitz:1988:ISU

[1435] S. Horwitz, T. Reps, and D. Binkley. Interprocedural slicing using dependence graphs. ACM SIG-PLAN Notices, 23(7):35–46, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/53990/p35-horwitz/

.

Callahan:1988:PSG

[1436] D. Callahan. The program summary graph and flow-sensitive interprocedural data flow analysis. ACM SIGPLAN Notices, 23(7):47–56, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/53990/p47-callahan/

.

Cooper:1988:ISA

[1437] K. D. Cooper and K. Kennedy. Interprocedural side-effect analysis in linear time. ACM SIGPLAN Notices, 23(7):57–66, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/53990/p57-cooper/

.

Cooper:1988:ISE

[1438] K. D. Cooper and K. Kennedy. Interprocedural side-effect analysis in linear time. ACM SIGPLAN Notices, 23 (7):57–66, July 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:1988:RWV

[1439] D. W. Wall. Register windows vs. register allocation. ACM SIG-PLAN Notices, 23(7):67–78, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p67-wall/

.

Fraser:1988:AGF

[1440] C. W. Fraser and A. L. Wendt. Automatic generation of fast optimizing code generators. ACM SIG-PLAN Notices, 23(7):79–84, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p79-fraser/

.

Chow:1988:MRU

[1441] F. C. Chow. Minimizing register usage penalty at procedure calls. ACM SIGPLAN Notices, 23(7):85–94, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p85-chow/

.

Keutzer:1988:AHC

[1442] K. Keutzer and W. Wolf. Anatomy of a hardware compiler. ACM SIG-PLAN Notices, 23(7):95–104, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/pldi/53990/p95-keutzer/

.

Andrews:1988:DIU

[1443] K. Andrews, R. R. Henry, and W. K. Yamamoto. Design and implementation of the UW Illustrated compiler. ACM SIGPLAN Notices, 23(7):105–114, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p105-andrews/

.

Yellin:1988:ILI

[1444] D. Yellin and R. Strom. INC: a language for incremental computations. ACM SIGPLAN Notices, 23(7):115–124, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p115-yellin/

.

Coutant:1988:DPA

[1445] D. S. Coutant, S. Meloy, and M. Ruscetta. DOC: a practical approach to source-level debugging of globally optimized code. ACM SIG-PLAN Notices, 23(7):125–134, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p125-coutant/

.

Miller:1988:MED

[1446] Barton P. Miller and Jong-Deok Choi. A mechanism for efficient debugging of parallel programs. ACM SIG-PLAN Notices, 23(7):135–144, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p135-miller/

.

Stone:1988:DCP

[1447] Janice M. Stone. Debugging concurrent processes: a case study. ACM SIGPLAN Notices, 23(7):145–153, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p145-stone/

.

Jain:1988:EAD

[1448] S. Jain and C. Thompson. An efficient approach to data flow analysis in a multiple pass global optimizer. ACM SIGPLAN Notices, 23(7):154–163, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p154-jain/

.

Shivers:1988:CFA

[1449] O. Shivers. Control flow analysis in Scheme. ACM SIGPLAN Notices, 23(7):164–174, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p164-shivers/

.

Schwarz:1988:OAD

[1450] B. Schwarz, W. Kirchgässner, and R. Landwehr. An optimizer for Ada — design, experiences and results. ACM SIGPLAN Notices, 23(7):175–184, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/53990/p175-schwarz/

.

Ballance:1988:GAI

[1451] R. A. Ballance, J. Butcher, and S. L. Graham. Grammatical abstraction and incremental syntax analysis in a language-based editor. ACM SIG-PLAN Notices, 23(7):185–198, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/53990/p185-ballance/

.

Pfenning:1988:HAS

[1452] F. Pfenning and C. Elliot. Higherorder abstract syntax. ACM SIG-PLAN Notices, 23(7):199–208, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/53990/p199-pfenning/

.

Pfenning:1988:HOA

[1453] F. Pfenning and C. Elliot. Higherorder abstract syntax. ACM SIG-PLAN Notices, 23(7):199–208, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walz:1988:IEG

[1454] J. A. Walz and G. F. Johnson. Incremental evaluation for a general class of circular attribute grammars. ACM SIGPLAN Notices, 23(7):209–221, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p209-walz/

.

Pleban:1988:AGR

[1455] U. F. Pleban and P. Lee. An automatically generated, realistic compiler for imperative programming language. ACM SIGPLAN Notices, 23(7):222–232, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p222-pleban/

.

Seshadri:1988:SAC

[1456] V. Seshadri, S. Weber, D. B. Wortman, C. P. Yu, and I. Small. Semantic analysis in a concurrent compiler. ACM SIGPLAN Notices, 23(7):233–240, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p233-seshadri/

.

Allen:1988:CCV

[1457] R. Allen and S. Johnson. Compiling C for vectorization, parallelization, and inline expansion. ACM SIG-PLAN Notices, 23(7):241–249, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p241-allen/

.

Reppy:1988:SOF

[1458] J. H. Reppy. Synchronous operations as first-class values. ACM SIG-PLAN Notices, 23(7):250–259, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p250-reppy/

.

Liskov:1988:PLS

[1459] B. Liskov and L. Shrira. Promises: Linguistic support for efficient asynchronous procedure calls in distributed systems. ACM SIGPLAN Notices, 23(7):260–267, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p260-liskov/

.

Pallas:1988:MSC

[1460] J. Pallas and D. Ungar. Multiprocessor Smalltalk: a case study of a multiprocessor-based programming environment. ACM SIG-PLAN Notices, 23(7):268–277, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p268-pallas/

.

Rose:1988:RTH

[1461] J. R. Rose. Refined types: highly differentiated type systems and their use in the design of intermediate languages. ACM SIGPLAN Notices, 23(7):278–287, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p278-rose/

.

Weiner:1988:PBC

[1462] J. L. Weiner and S. Ramakrishman. A piggy-back compiler for Prolog. ACM SIGPLAN Notices, 23(7):288–296, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weiner:1988:PCP

[1463] J. L. Weiner and S. Ramakrishman. A piggy-back compiler for Prolog. ACM SIGPLAN Notices, 23(7):288–296, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p288-weiner/

.

Debray:1988:UFT

[1464] S. K. Debray. Unfold/fold transformations and loop optimization of logic programs. ACM SIGPLAN Notices, 23 (7):297–307, July 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Debray:1988:UTL

[1465] S. K. Debray. Unfold/fold transformations and loop optimization of logic programs. ACM SIG-PLAN Notices, 23(7):297–307, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p297-debray/

.

Aiken:1988:OLP

[1466] A. Aiken and A. Nicolau. Optimal loop parallelization. ACM SIG-PLAN Notices, 23(7):308–317, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p308-aiken/

.

Lam:1988:SPE

[1467] M. Lam. Software pipelining: an effective scheduling technique for VLIW machines. ACM SIGPLAN Notices, 23(7):318–328, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p318-lam/

.

Benitez:1988:PGO

[1468] M. E. Benitez and J. W. Davidson. A portable global optimizer and linker. ACM SIGPLAN Notices, 23(7):329–338, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p329-benitez/

.

Jones:1988:HCF

[1469] D. W. Jones. How (not) to code a finite state machine. ACM SIG-PLAN Notices, 23(8):19–22, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1988:RAL

[1470] G. H. Roberts. Recursive ascent: an LR analog to recursive descent. ACM SIGPLAN Notices, 23(8):23–29, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1988:FFL

[1471] S. L. Payton Jones. FLIC: a functional language intermediate code. ACM SIG-PLAN Notices, 23(8):30–48, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weicker:1988:DBR

[1472] R. P. Weicker. Dhrystone benchmark: rationale for version 2 and measurement rules. ACM SIGPLAN Notices, 23(8):49–62, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rising:1988:TTT

[1473] L. Rising. Tasking troubles and tips.

ACM SIGPLAN Notices, 23(8):63–72, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paakki:1988:NSP

[1474] J. Paakki. A note on the speed of Prolog. ACM SIGPLAN Notices, 23 (8):73–82, August 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bohlender:1988:NDP

[1475] G. Bohlender, Ch. Ullrich, and J. Wolff von Gudenberg. New developments in Pascal-SC. ACM SIGPLAN Notices, 23(8):83–92, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1988:LLC

[1476] E. Levy. The linked list class of Modula-3. ACM SIGPLAN Notices, 23 (8):93–102, August 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Share:1988:RAP

[1477] M. Share. Resolving ambiguities in the parsing of translation grammars. ACM SIGPLAN Notices, 23(8):103–109, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bershad:1988:OEB

[1478] Brian N. Bershad, Edward D. Lazowska, Henry M. Levy, and David B. Wagner. An open environment for building parallel programming systems. ACM SIGPLAN Notices, 23(9):1–9, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

proceedings/plan/62115/p1-bershad/

.

Notkin:1988:EP

[1479] David Notkin, Lawrence Snyder, David Socha, Mary L. Bailey, Bruce Forstall, Kevin Gates, Ray Greenlaw, William G. Griswold, Thomas J. Holman, Richard Korry, Gemini Lasswell, Robert Mitchell, and Philip A. Nelson. Experiences with poker. ACM SIG-PLAN Notices, 23(9):10–20, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/62115/p10-notkin/

.

Aral:1988:NII

[1480] Ziya Aral and Ilya Gertner. Nonintrusive and interactive profiling in Parasight. ACM SIGPLAN Notices, 23(9):21–30, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/62115/p21-aral/

.

Bruegge:1988:PDS

[1481] Bernd Bruegge. Program development for a systolic array. ACM SIG-PLAN Notices, 23(9):31–41, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/62115/p31-bruegge/

.

Albert:1988:CFA

[1482] Eugene Albert, Kathleen Knobe, Joan D. Lukas, and Guy L. Steele, Jr. Compiling Fortran 8x array features for the Connection Machine computer system. ACM SIG-PLAN Notices, 23(9):42–56, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p42-albert/

.

Quinn:1988:CCP

[1483] Michael J. Quinn, Philip J. Hatcher, and Karen C. Jourdenais. Compiling C* programs for a hypercube multicomputer. ACM SIG-PLAN Notices, 23(9):57–65, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p57-quinn/

.

Katseff:1988:UDP

[1484] Howard P. Katseff. Using data partitioning to implement a parallel assembler. ACM SIGPLAN Notices, 23(9):66–76, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p66-katseff/

.

Burke:1988:ADP

[1485] Michael Burke, Ron Cytron, Jeanne Ferrante, Wilson Hsieh, Vivek Sarkar, and David Shields. Automatic discovery of parallelism: a tool and an experiment (extended abstract). ACM SIG-PLAN Notices, 23(9):77–84, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/62115/p77-burke/

.

Li:1988:EIA

[1486] Zhiyuan Li and Pen-Chung Yew. Efficient interprocedural analysis for program parallelization and restructuring. ACM SIGPLAN Notices, 23 (9):85–99, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/62115/p85-li/.

Larus:1988:RLP

[1487] James R. Larus and Paul N. Hilfinger. Restructuring Lisp programs for concurrent execution. ACM SIG-PLAN Notices, 23(9):100–110, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/62115/p100-larus/

.

Goldman:1988:QEN

[1488] Ron Goldman and Richard P. Gabriel.

Qlisp: experience and new directions. ACM SIGPLAN Notices, 23 (9):111–123, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/plan/62115/p111-goldman/

.

Nicol:1988:PDE

[1489] David M. Nicol. Parallel discreteevent simulation of FCFS stochastic queueing networks. ACM SIG-PLAN Notices, 23(9):124–137, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p124-nicol/

.

Rose:1988:PDI

[1490] Jonathan Rose. The parallel decomposition and implementation of an integrated circuit global router. ACM SIG-PLAN Notices, 23(9):138–145, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p138-rose/

.

Tambe:1988:SPI

[1491] Milind Tambe, Dirk Kalp, Anoop Gupta, Charles Forgy, Brian Milnes, and Allen Newell. Soar/PSM-E: investigating match parallelism in a learning production system. ACM SIG-PLAN Notices, 23(9):146–160, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p146-tambe/

.

LeBlanc:1988:LSP

[1492] Thomas J. LeBlanc, Michael L. Scott, and Christopher M. Brown. Largescale parallel programming: experience with BBN Butterfly parallel processor. ACM SIGPLAN Notices, 23(9):161–172, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/62115/p161-leblanc/

.

Carriero:1988:AEL

[1493] Nicholas Carriero and David Gelernter. Applications experience with Linda. ACM SIGPLAN Notices, 23 (9):173–187, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/62115/p173-carriero/

.

Lee:1988:IAL

[1494] Ching-Cheng Lee, Stephen Skedzielewski, and John Feo. On the implementation of applicative languages on shared-memory, MIMD multiprocessors. ACM SIGPLAN Notices, 23 (9):188–197, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/62115/p188-lee/

.

Davis:1988:CSB

[1495] Helen Davis and John Hennessy. Characterizing the synchronization behavior of parallel programs. ACM SIG-PLAN Notices, 23(9):198–211, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p198-davis/

.

Gupta:1988:EVG

[1496] Anoop Gupta and Andrew Tucker. Exploiting variable grain parallelism at runtime. ACM SIGPLAN Notices, 23(9):212–221, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p212-gupta/

.

Stramm:1988:CSH

[1497] Bernd Stramm and Francine Berman. Communication-sensitive heuristics and algorithms for mapping compilers. ACM SIGPLAN Notices, 23 (9):222–243, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p222-stramm/

.

Gupta:1988:CTT

[1498] Rajiv Gupta and Mary Lou Soffa. Compile-time techniques for efficient utilization of parallel memories. ACM SIGPLAN Notices, 23 (9):235–246, September 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/62115/p235-gupta/

.

Burns:1988:MIR

[1499] A. Burns, G. L. Davies, and A. J.

Wellings. A Modula-2 implementation of a real-time process abstraction. ACM SIGPLAN Notices, 23(10): 49–58, October 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaelbling:1988:PLS

[1500] M. J. Kaelbling. Programming languages should NOT have comment statements. ACM SIGPLAN Notices, 23(10):59–60, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1988:SSR

[1501] P. W. Abrahams. Some sad remarks about string handling in C. ACM SIG-PLAN Notices, 23(10):61–68, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1988:NWL

[1502] J. E. Sammet. Notes on what language maturity means, and how to measure it. ACM SIGPLAN Notices, 23(10): 69–71, October 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1988:TRV

[1503] H. Liu and M. Klerer. Timing results of various compilers using an optimization quality benchmark. ACM SIG-PLAN Notices, 23(10):72–74, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klerer:1988:BTE

[1504] M. Klerer and H. Liu. Benchmark test to estimate optimization quality of compilers. ACM SIGPLAN Notices, 23(10):75–84, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandberg:1988:SEP

[1505] D. W. Sandberg. Smalltalk and exploratory programming. ACM SIG-PLAN Notices, 23(10):85–92, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schnorf:1988:DIC

[1506] P. Schnorf. Dynamic instantiation and configuration of functionally extended, efficient lexical analysers. ACM SIG-PLAN Notices, 23(10):93–102, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Genillard:1988:G

[1507] C. Genillard and A. Strohmeier. GRAMOL. ACM SIGPLAN Notices, 23(10):103–122, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Warfield:1988:ESR

[1508] J. W. Warfield and H. R. Bauer III. An expert system for a retargetable peephole optimizer. ACM SIGPLAN Notices, 23(10):123–130, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yates:1988:DPI

[1509] J. S. Yates and R. A. Schwartz. Dynamic programming and industrialstrength instruction selection: code generation by tiring, but not exhaustive, search. ACM SIGPLAN Notices, 23(10):131–140, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weinstock:1988:EAH

[1510] C. B. Weinstock and W. A. Wulf.

An efficient algorithm for heap storage allocation. ACM SIGPLAN Notices, 23(10):141–148, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lemkin:1988:PPS

[1511] P. F. Lemkin. PSAIL: a portable SAIL to C compiler — description and tutorial. ACM SIGPLAN Notices, 23(10): 149–171, October 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dhamdhere:1988:FAC

[1512] D. M. Dhamdhere. A fast algorithm for code movement optimisation. ACM SIGPLAN Notices, 23(10):172– 180, October 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mellender:1988:ILO

[1513] F. Mellender. An integration of logic and object-oriented programming. ACM SIGPLAN Notices, 23(10): 181–185, October 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lang:1988:GF

[1514] J. E. Lang. Grammars for FFP. ACM SIGPLAN Notices, 23(10):186– 190, October 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungar:1988:TPG

[1515] David Ungar and Frank Jackson. Tenuring policies for generation-based storage reclamation. ACM SIGPLAN Notices, 23(11):1–17, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1988:TOC

[1516] Ralph E. Johnson, Justin O. Graver, and Laurance W. Zurawski. TS: an optimizing compiler for Smalltalk. ACM SIGPLAN Notices, 23(11):18– 26, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rose:1988:FDM

[1517] John R. Rose. Fast dispatch mechanisms for stock hardware. ACM SIG-PLAN Notices, 23(11):27–35, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Szekely:1988:UIT

[1518] Pedro Szekely and Brad Myers. A user interface toolkit based on graphical objects and constraints. ACM SIGPLAN Notices, 23(11):36–45, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weinand:1988:OOA

[1519] Andre Weinand, Erich Gamma, and Rudolf Marty. ET++: an object oriented application framework in C++. ACM SIGPLAN Notices, 23(11):46– 57, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Szczur:1988:TAE

[1520] Martha R. Szczur and Philip Miller.

Transportable Applications Environment (TAE) plus experiences in “Object”-ively modernizing a user interface environment. ACM SIGPLAN Notices, 23(11):58–70, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirfs-Brock:1988:ICS

[1521] Rebecca Wirfs-Brock. An integrated color Smalltalk-80 system. ACM SIG-PLAN Notices, 23(11):71–82, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Epstein:1988:SWS

[1522] Danny Epstein and Wilf R. LaLonde.

A Smalltalk window system based on constraints. ACM SIGPLAN Notices, 23(11):83–94, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridhar:1988:CSA

[1523] S. Sridhar. Configuring stand-alone Smalltalk-80 applications. ACM SIG-PLAN Notices, 23(11):95–104, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaLonde:1988:BBF

[1524] Wilf R. LaLonde and Mark Van Gulik. Building a backtracking facility in Smalltalk without kernel support. ACM SIGPLAN Notices, 23(11):105– 122, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirfs-Brock:1988:OMS

[1525] Allen Wirfs-Brock and Brian Wilkerson. A overview of modular Smalltalk. ACM SIGPLAN Notices, 23(11):123– 134, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomas:1988:OCM

[1526] Dave Thomas and Kent Johnson. Orwell: a configuration management system for team programming. ACM SIGPLAN Notices, 23(11):135–141, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:1988:IOO

[1527] Won Kim, Nat Ballou, Jay Banerjee, Hong-Tai Chou, Jorge G. Garza, and Darrell Woelk. Integrating an objectoriented programming system with a database system. ACM SIGPLAN Notices, 23(11):142–152, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duhl:1988:PCO

[1528] Joshua Duhl and Craig Damon. A performance comparison of object and relational databases using the Sun Benchmark. ACM SIGPLAN Notices, 23(11):153–163, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beech:1988:ICO

[1529] David Beech. Intensional concepts in an object database model. ACM SIGPLAN Notices, 23(11):164–175, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ingalls:1988:FVP

[1530] Dan Ingalls, Scott Wallace, Yu-Ying Chow, Frank Ludolph, and Ken Doyle. Fabrik: a visual programming environment. ACM SIGPLAN Notices, 23 (11):176–190, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kleyn:1988:GUO

[1531] Michael F. Kleyn and Paul C. Gingrich.

GraphTrace: Understanding objectoriented systems using concurrently animated views. ACM SIGPLAN Notices, 23(11):191–205, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bjornerstedt:1988:AOM

[1532] Anders Bjornerstedt and Stefan Britts.

AVANCE: an object management system. ACM SIGPLAN Notices, 23(11): 206–221, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scaletti:1988:IEO

[1533] C. A. Scaletti and R. E. Johnson. An interactive environment for objectoriented music composition and sound synthesis. ACM SIGPLAN Notices, 23 (11):222–233, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rubin:1988:SII

[1534] Kenneth S. Rubin, Patricia M. Jones, Christine M. Mitchell, and Theodore C. Goldstein. A Smalltalk implementation of an intelligent operator’s associate. ACM SIGPLAN Notices, 23(11):234– 247, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russo:1988:PME

[1535] Vincent Russo, Gary Johnston, and Roy Campbell. Process management and exception handling in multiprocessor operating systems using objectoriented design techniques. ACM SIGPLAN Notices, 23(11):248–258, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yoshida:1988:OOF

[1536] Norihiko Yoshida and Kouji Hino. An object-oriented framework of pattern recognition systems. ACM SIGPLAN Notices, 23(11):259–267, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aksit:1988:DAM

[1537] Mehmet Aksit and Anand Tripathi. Data abstraction mechanisms in SINA/ ST. ACM SIGPLAN Notices, 23(11): 267–275, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matsuoka:1988:UTS

[1538] Satoshi Matsuoka and Satoru Kawai.

Using tuple space communication in distributed object-oriented languages. ACM SIGPLAN Notices, 23(11):276– 284, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rumbaugh:1988:CPO

[1539] James Rumbaugh. Controlling propagation of operations using attributes on relations. ACM SIGPLAN Notices, 23(11):285–296, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shibayama:1988:HID

[1540] Etsuya Shibayama. How to invent distributed implementation schemes of an object-based concurrent language: a transformational approach. ACM SIGPLAN Notices, 23(11):297– 305, November 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watanabe:1988:ROO

[1541] Takuo Watanabe and Akinori Yonezawa.

Reflection in an object-oriented concurrent language. ACM SIGPLAN Notices, 23(11):306–315, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bensley:1988:EMD

[1542] Edward H. Bensley, Thomas J. Brando, and Myra Jean Prelle. An execution model for distributed object-oriented computation. ACM SIGPLAN Notices, 23(11):316–322, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberherr:1988:OOP

[1543] K. Lieberherr, I. Holland, and A. Riel. Object-oriented programming: an objective sense of style. ACM SIGPLAN Notices, 23(11):323–334, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [1546].

Alabiso:1988:TDF

[1544] Bruno Alabiso. Transformation of data flow analysis models to object oriented design. ACM SIGPLAN Notices, 23(11):335, November 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corradi:1988:ROT

[1545] Antonio Corradi and Letizia Leonardi. The role of opaque types to build abstractions. ACM SIGPLAN Notices, 23(12):24–37, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakkinen:1988:CLD

[1546] Markku Sakkinen. Comments on “The Law of Demeter” and C++. ACM SIG-PLAN Notices, 23(12):38–44, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1543].

Yeh:1988:IIE

[1547] Dashing Yeh and Uwe Kastens. Improvements of an incremental evaluation algorithm for ordered attribute grammars. ACM SIGPLAN Notices, 23 (12):45–50, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Serbedzija:1988:ACO

[1548] Nikola B. Serbedzija. Asynchronous communication in Occam. ACM SIG-PLAN Notices, 23(12):51–62, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Budd:1988:TST

[1549] Timothy A. Budd. Time space tradeoffs in vector algorithms for APL functions. ACM SIGPLAN Notices, 23(12): 63–68, December 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oman:1988:PPS

[1550] Paul W. Oman and Curtis R. Cook.

A paradigm for programming style research. ACM SIGPLAN Notices, 23 (12):69–78, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neusius:1988:PSM

[1551] Christian Neusius. Portable software in modular Pascal. ACM SIGPLAN Notices, 23(12):79–85, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furtado:1988:IAL

[1552] Antonio L. Furtado and Paulo A. S. Veloso. Iteration for applicative languages. ACM SIGPLAN Notices, 23 (12):86–94, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewry:1988:PSM

[1553] D. A. Sewry. Process scheduling in Modula-2. ACM SIGPLAN Notices, 23(12):95–97, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1988:OGC

[1554] Paul R. Wilson. Opportunistic garbage collection. ACM SIGPLAN Notices, 23 (12):98–102, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

House:1988:PEP

[1555] R. T. House. A problem with extended Pascal. ACM SIGPLAN Notices, 23(12):103–114, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boddy:1988:SMB

[1556] David E. Boddy. SOS: a monitorbased operating system for instruction. ACM SIGPLAN Notices, 23(12):115– 124, December 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sharpe:1988:ARE

[1557] Daniel Sharpe. Adding regular expressions to Pascal. ACM SIGPLAN Notices, 23(12):125–133, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1988:LSC

[1558] B. Meek. Language standards committees and revisions. ACM SIG-PLAN Notices, 23(12):134–142, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1988:EFL

[1559] David Spector. Efficient full LR(1)

parser generation. ACM SIGPLAN Notices, 23(12):143–150, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kovats:1988:CIC

[1560] Thomas A. Kovats. Comments on innovative control constructs in Pancode and EPN. ACM SIGPLAN Notices, 23(12):151–157, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elshoff:1989:DDA

[1561] I. J. P. Elshoff. A distributed debugger for Amoeba. ACM SIG-PLAN Notices, 24(1):1–10, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bates:1989:DHD

[1562] Peter Bates. Debugging heterogeneous distributed systems using event-based models of behavior. ACM SIGPLAN Notices, 24(1):11–22, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:1989:EBD

[1563] Chu-Chung Lin and Richard J. LeBlanc. Event-based debugging of object/action programs. ACM SIG-PLAN Notices, 24(1):23–34, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Redell:1989:ETT

[1564] David D. Redell. Experience with Topaz teledebugging. ACM SIGPLAN Notices, 24(1):35–44, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Black:1989:MEH

[1565] David L. Black, David B. Golub, Karl Hauth, Avadis Tevanian, and Richard Sanzi. The Mach exception handling facility. ACM SIGPLAN Notices, 24 (1):45–56, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wittie:1989:DDC

[1566] Larry D. Wittie. Debugging distributed C programs by real time reply. ACM SIGPLAN Notices, 24(1):57–67, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tokuda:1989:RTM

[1567] Hideyuki Tokuda, Makoto Kotera, and Clifford E. Mercer. A real-time monitor for a distributed real-time operating system. ACM SIGPLAN Notices, 24 (1):68–77, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appelbe:1989:ITD

[1568] William F. Appelbe and Charles E. Mc-Dowell. Integrating tools for debugging and developing multitasking programs. ACM SIGPLAN Notices, 24(1): 78–88, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Emrath:1989:ADN

[1569] Perry A. Emrath and David A. Padua.

Automatic detection of nondeterminacy in parallel programs. ACM SIG-PLAN Notices, 24(1):89–99, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callahan:1989:SAL

[1570] David Callahan and Jaspal Sublok.

Static analysis of low-level synchronization. ACM SIGPLAN Notices, 24(1): 100–111, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1989:ISP

[1571] Stuart I. Feldman and Channing B.

Brown. IGOR: a system for program debugging via reversible execution. ACM SIGPLAN Notices, 24(1): 112–123, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pan:1989:SRE

[1572] Douglas Z. Pan and Mark A. Linton. Supporting reverse execution for parallel programs. ACM SIGPLAN Notices, 24(1):124–129, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Forin:1989:DHP

[1573] Alessandro Forin. Debugging of heterogeneous parallel systems. ACM SIG-PLAN Notices, 24(1):130–140, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1989:MED

[1574] Barton P. Miller and Jong-Deok Choi. A mechanism for efficient debugging of parallel programs. ACM SIGPLAN Notices, 24(1):141–150, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aral:1989:HLD

[1575] Ziya Aral and Ilya Gertner. High-level debugging in Parasight. ACM SIG-PLAN Notices, 24(1):151–162, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fowler:1989:IAP

[1576] Robert J. Fowler, Thomas J. LeBlanc, and John M. Mellor-Crummey. An integrated approach to parallel program debugging and performance analysis on large-scale multiprocessors. ACM SIG-PLAN Notices, 24(1):163–173, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruegge:1989:PDS

[1577] Bernd Bruegge and Thomas Gross. A program debugger for a systolic array: design and implementation. ACM SIG-PLAN Notices, 24(1):174–182, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fidge:1989:POP

[1578] C. J. Fidge. Partial orders for parallel debugging. ACM SIGPLAN Notices, 24(1):183–194, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hough:1989:IEP

[1579] Alfred A. Hough and Janice E. Cuny.

Initial experiences with a patternoriented parallel debugger. ACM SIG-PLAN Notices, 24(1):195–205, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Socha:1989:VGV

[1580] David Socha, Mary L. Bailey, and David Notkin. Voyeur: graphical views of parallel programs. ACM SIGPLAN Notices, 24(1):206–215, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rubin:1989:DPP

[1581] Robert Rubin, Larry Rudolph, and Dror Zernik. Debugging parallel programs in parallel. ACM SIGPLAN Notices, 24(1):216–225, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stone:1989:GRC

[1582] Janice M. Stone. A graphical representation of concurrent processes. ACM SIGPLAN Notices, 24(1):226–235, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hseush:1989:DPD

[1583] Wenwey Hseush and Gail E. Kaiser. Data path debugging: data-oriented debugging for a concurrent programming language. ACM SIGPLAN Notices, 24(1):236–247, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lichtenstein:1989:CAD

[1584] Yossi Lichtenstein and Ehud Shapiro. Concurrent algorithmic debugging. ACM SIGPLAN Notices, 24(1):248– 260, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cameron:1989:ISD

[1585] E. J. Cameron and D. Cohen. The IC^{∗} system for debugging parallel programs via interactive monitoring and control. ACM SIGPLAN Notices, 24(1): 261–270, January 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldszmidt:1989:IBD

[1586] G. Goldszmidt, S. Katz, and S. Yemini. Interactive blackbox debugging for concurrent languages. ACM SIGPLAN Notices, 24(1):271–282, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeBlanc:1989:WS

[1587] Thomas J. LeBlanc and Barton P.

Miller. Workshop summary. ACM SIG-PLAN Notices, 24(1):ix–xxi, January 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1989:FAE

[1588] Richard N. Taylor, Frank C. Belz, Lori A. Clarke, Leon Osterweil, Richard W. Selby, Jack C. Wileden, Alexander L. Wolf, and Michael Young. Foundations for the Arcadia environment architecture. ACM SIGPLAN Notices, 24 (2):1–13, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borras:1989:CS

[1589] P. Borras, D. Clement, Th. Despeyroux, J. Incerpi, G. Kahn, B. Lang, and V. Pascual. Centaur: the system. ACM SIGPLAN Notices, 24(2): 14–24, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1989:ESS

[1590] Peter Lee, Frank Pfenning, Gene Rollins, and William Scherlis. The ergo support system: an integrated set of tools for prototyping integrated environments. ACM SIGPLAN Notices, 24(2):25–34, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strelich:1989:SLC

[1591] Tom Strelich. The Software Life Cycle Support Environment (SLCSE): a computer based framework for developing software systems. ACM SIG-PLAN Notices, 24(2):35–44, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koenig:1989:III

[1592] Shaye Koenig. (ISEF): an integrated industrial-strength software engineering framework. ACM SIGPLAN Notices, 24(2):45–54, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clemm:1989:WSP

[1593] Geoffrey M. Clemm. The workshop system: a practical knowledge-based software environment. ACM SIGPLAN Notices, 24(2):55–64, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minsky:1989:SDE

[1594] Naftaly H. Minsky and David Rozenshtein. A software development environment for law-governed systems. ACM SIGPLAN Notices, 24(2):65– 75, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodden:1989:IAI

[1595] Thomas Rodden, Pete Sawyer, and Ian Sommerville. Interacting with an active, integrated environment. ACM SIGPLAN Notices, 24(2):76–84, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1989:ODS

[1596] Lung-Chun Liu and Ellis Horowitz.

Object database support for a software project management environment. ACM SIGPLAN Notices, 24(2): 85–96, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huff:1989:PBI

[1597] Karen E. Huff and Victor R. Lesser.

A plan-based intelligent assistant that supports the software development. ACM SIGPLAN Notices, 24(2):97– 106, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Notkin:1989:RBS

[1598] David Notkin. The relationship between software development environments and the software process. ACM SIGPLAN Notices, 24(2):107– 109, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nord:1989:EAS

[1599] Robert L. Nord and Frank Pfenning.

The ergo attribute system. ACM SIG-PLAN Notices, 24(2):110–120, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alpern:1989:GAS

[1600] Bowen Alpern, Alan Carle, Barry Rosen, Peter Sweeney, and Kenneth Zadeck. Graph attribution as a specification paradigm. ACM SIGPLAN Notices, 24(2):121–129, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wileden:1989:PEP

[1601] Jack C. Wileden, Alexander L. Wolf, Charles D. Fisher, and Peri L. Tarr. Pgraphite: an experiment in persistent typed object management. ACM SIG-PLAN Notices, 24(2):130–142, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beaudouin-Lafon:1989:UIS

[1602] Michel Beaudouin-Lafon. User interface support for the integration of software tools: an iconic model of interaction. ACM SIGPLAN Notices, 24(2): 143–152, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1989:TSE

[1603] Lori A. Clarke, Debra J. Richardson, and Steven J. Zeil. TEAM: a support environment for testing, evaluation, and analysis. ACM SIGPLAN Notices, 24(2):153–162, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reddy:1989:TDP

[1604] Uday S. Reddy. Transformational derivation of programs using the focus system. ACM SIGPLAN Notices, 24(2):163–172, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewerentz:1989:EPL

[1605] Claus Lewerentz. Extended programming in the large in a software development environment. ACM SIGPLAN Notices, 24(2):173–182, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walpole:1989:UMC

[1606] J. Walpole, G. S. Blair, J. Malik, and J. R. Nicol. A unifying model for consistent distributed software development environments. ACM SIGPLAN Notices, 24(2):183–190, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahler:1989:ITE

[1607] Axel Mahler and Andreas Lampen. An integrated toolset for engineering software configurations. ACM SIGPLAN Notices, 24(2):191–200, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1989:VMG

[1608] Ellis S. Cohen, Dilip A. Soni, Raimund Gluecker, William M. Hasling, Robert W. Schwanke, and Michael E. Wagner. Version management in Gypsy. ACM SIGPLAN Notices, 24(2):201–215, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heimbigner:1989:GTM

[1609] Dennis Heimbigner and Steven Krane. A graph transform model for configuration management environments. ACM SIGPLAN Notices, 24(2):216– 225, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Penedo:1989:OMI

[1610] Maria H. Penedo, Erhard Ploedereder, and Ian Thomas. Object management issues for software engineering environments workshop report. ACM SIG-PLAN Notices, 24(2):226–234, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Munck:1989:ODS

[1611] Robert Munck, Patricia Oberndorf, Erhard Ploedereder, and Richard Thall. An overview of DOD-STD-1838 A (proposed) the common APSE interface set: revision. ACM SIGPLAN Notices, 24(2):235–247, February 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boudier:1989:OPP

[1612] Gerard Boudier, Ferdinando Gallo, Regis Minot, and Ian Thomas. An overview of PCTE and PCTE+. ACM SIGPLAN Notices, 24(2):248– 257, February 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leach:1989:BRC

[1613] George W. Leach. Book review: C Traps and Pitfalls by Andrew Koenig: (Reading, MA: Addison-Wesley, 1989, ISBN 0-201-17928-8). ACM SIG-PLAN Notices, 24(3):20–22, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanDelft:1989:CO

[1614] A. J. E. van Delft. Comments on Oberon. ACM SIGPLAN Notices, 24 (3):23–30, March 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:1989:ASS

[1615] Ming-Yuan Y. Zhu. AUTO STAR — a software development system. ACM SIGPLAN Notices, 24(3):31–45, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VelazquezIturbide:1989:FCS

[1616] J. Ángel Velázquez Iturbide. Formalization of the control stack. ACM SIGPLAN Notices, 24(3):46–54, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakkinen:1989:HBR

[1617] Markku Sakkinen. How to best return the value of a function. ACM SIGPLAN Notices, 24(3):55–56, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atteson:1989:NSS

[1618] Kevin Atteson, Michael Lorenz, and William F. Dowling. NARPL: a solution to the student compiler project problem. ACM SIGPLAN Notices, 24 (3):57–66, March 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberherr:1989:FBL

[1619] Karl J. Lieberherr and Ian Holland. Formulations and benefits of the Law of Demeter. ACM SIGPLAN Notices, 24 (3):67–78, March 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grogono:1989:CAP

[1620] Peter Grogono. Comments, assertions, and pragmas. ACM SIGPLAN Notices, 24(3):79–84, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonsson:1989:NEG

[1621] Dan Jonsson. Next: the elimination of Goto-patches? ACM SIG-PLAN Notices, 24(3):85–92, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balzer:1989:DRR

[1622] Robert Balzer, Richard P. Gabriel, Frank Belz, Robert Dewar, David Fisher, John Guttag, Paul Hudak, and Mitchell Wand. Draft report on requirements for a common prototyping system. ACM SIGPLAN Notices, 24 (3):93–161, March 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:1989:DRR

[1623] R. P. Gabriel. Draft report on requirements for a common prototyping system. ACM SIGPLAN Notices, 24(3): 93–165, March 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briot:1989:DDI

[1624] Jean-Pierre Briot and Jean de Ratuld. Design of a distributed implementation of ABCL/1. ACM SIG-PLAN Notices, 24(4):15–17, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p15-briot/

.

Buhr:1989:ACS

[1625] P. A. Buhr, Glen Ditchfield, and C. R. Zarnke. Adding concurrency to a statically type-safe object-oriented programming language. ACM SIG-PLAN Notices, 24(4):18–21, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p18-buhr/

.

Caseau:1989:MRO

[1626] Yves Caseau. A model for a reflective object-oriented language. ACM SIGPLAN Notices, 24(4):22–24, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p22-caseau/

.

Crowl:1989:UOM

[1627] Lawrence A. Crowl. A uniform object model for parallel programming. ACM SIGPLAN Notices, 24(4):25–27, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p25-crowl/

.

Dally:1989:OOC

[1628] William J. Dally and Andrew A. Chien. Object-oriented concurrent programming in CST. ACM SIG-PLAN Notices, 24(4):28–31, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p28-dally/

.

Herlihy:1989:TCS

[1629] Maurice Herlihy. Taking concurrency seriously (position paper). ACM SIGPLAN Notices, 24(4):32–33, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p32-herlihy/

.

Magnusson:1989:POP

[1630] Boris Magnusson. Process oriented programming. ACM SIG-PLAN Notices, 24(4):34–36, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/

proceedings/plan/67386/p34-magnusson/

.

Mello:1989:COL

[1631] Paola Mello. Concurrent objects in a logic programming framework. ACM SIGPLAN Notices, 24(4):37–39, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p37-mello/

.

Mehrotra:1989:COA

[1632] Piyush Mehrotra and John Van Rosendale. Concurrent object access in BLAZE 2. ACM SIG-PLAN Notices, 24(4):40–42, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p40-mehrotra/

.

Nakajima:1989:DLS

[1633] Tatsuo Nakajima, Yasuhiko Yokote, Mario Tokoro, Shinichi Ochiai, and Tatsuo Nagamatsu. DistributedConcurrentSmalltalk: a language and system for the interpersonal environment. ACM SIGPLAN Notices, 24(4):43–45, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p43-nakajima/

.

Wegner:1989:GMO

[1634] Peter Wegner. Granularity of modules in object-based concurrent systems. ACM SIGPLAN Notices, 24(4):46–49, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p46-wegner/

.

Yonezawa:1989:IOB

[1635] Akinori Yonezawa and Takuo Watanabe. An introduction to object-based reflective concurrent computation. ACM SIGPLAN Notices, 24(4):50–54, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p50-yonezawa/

.

Yoshida:1989:ASO

[1636] Kaoru Yoshida and Takashi Chikayama. A′UM = Stream + Object + Relation. ACM SIGPLAN Notices, 24(4):55–58, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p55-yoshiba/

.

Agha:1989:FIC

[1637] Gul Agha. Foundational issues in concurrent computing. ACM SIG-PLAN Notices, 24(4):60–65, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p60-agha/

.

Athas:1989:CAP

[1638] W. C. Athas and N. J. Boden. Cantor: an actor programming system for scientific computing. ACM SIG-PLAN Notices, 24(4):66–68, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p66-athas/

.

Briot:1989:OAS

[1639] Jean-Pierre Briot. From objects to actors: study of a limited symbiosis in Smalltalk-80. ACM SIG-PLAN Notices, 24(4):69–72, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p69-briot/

.

Carre:1989:OOL

[1640] Françoise Carre and Pascal Clere.

Object oriented languages and actors: which language for a distributed approach? ACM SIG-PLAN Notices, 24(4):73–75, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p73-carre/

.

Christopher:1989:MDC

[1641] Thomas W. Christopher. Message driven computing and its relationship to actors. ACM SIG-PLAN Notices, 24(4):76–78, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/plan/67386/p76-christopher/

.

Crowl:1989:CDS

[1642] Lawrence A. Crowl. Concurrent data structures and actor programming under the Matroshka model. ACM SIGPLAN Notices, 24(4):79–80, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p79-crowl/

.

Lapalme:1989:PIA

[1643] Guy Lapalme and Patrick Salle. Plasma-II: an actor approach to concurrent programming. ACM SIG-PLAN Notices, 24(4):81–83, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p81-lapaime/

.

Manning:1989:PAA

[1644] Carl Manning. A peek at Acore, an actor core language. ACM SIG-PLAN Notices, 24(4):84–86, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p84-manning/

.

Thomas:1989:AMM

[1645] David A. Thomas, Wilf R. LaLonde, John Duimovich, Michael Wilson, Jeff McAffer, and Brian Barry. Actra — a multitasking/multiprocessing Smalltalk. ACM SIGPLAN Notices, 24(4):87–90, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p87-thomas/

.

Tomlinson:1989:ROO

[1646] Chris Tomlinson, Won Kim, Mark Scheevel, Vineet Singh, Becky Will, and Gul Agha. Rosette: An object-oriented concurrent systems architecture. ACM SIGPLAN Notices, 24(4):91–93, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p91-tomlinson/

.

Bain:1989:IGO

[1647] William L. Bain. Indexed, global objects for distributed memory parallel architectures. ACM SIG-PLAN Notices, 24(4):95–98, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p95-bain/

.

Bensley:1989:MFG

[1648] E. H. Bensley, T. J. Brando, J. C.

Fohlin, M. J. Prelle, and A. M. Wollrath. MITRE’s future generation computer architectures program. ACM SIGPLAN Notices, 24(4):99–101, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p99-bensley/

.

Caromel:1989:GMC

[1649] Denis Caromel. A general model for concurrent and distributed objectoriented programming. ACM SIG-PLAN Notices, 24(4):102–104, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p102-caromel/

.

Decouchant:1989:SMT

[1650] D. Decouchant, S. Krakowiak, M. Meysembourg, M. Riveill, and X. Rousset de Pina. A synchronization mechanism for typed objects in a distributed system. ACM SIG-PLAN Notices, 24(4):105–107, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p105-decouchant/

.

Gasser:1989:MHL

[1651] Les Gasser. MACE: High-level distributed objects in a flexible testbed for distributed AI research. ACM SIG-PLAN Notices, 24(4):108–110, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p108-gasser/

.

Heuser:1989:EOP

[1652] L. Heuser, A. Schill, and M. Mühlhäuser.

Extensions to the object paradigm for the development of distributed applications. ACM SIGPLAN Notices, 24(4):111–113, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p111-heuser/

.

Hindel:1989:OOP

[1653] B. Hindel. An object-oriented programming language for distributed systems: Heraklit. ACM SIG-PLAN Notices, 24(4):114–116, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p114-hindel/

.

Jazayeri:1989:ODS

[1654] Mehdi Jazayeri. Objects for distributed systems. ACM SIGPLAN Notices, 24(4):117–119, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p117-jazayeri/

.

Kaiser:1989:CM

[1655] Gail E. Kaiser. Concurrent MELD.

ACM SIGPLAN Notices, 24(4):120– 122, April 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:80/pubs/

                                        citations/proceedings/plan/67386/
                                        p120-kaiser/.

Koegel:1989:POD

[1656] John F. Koegel. Parallel objects on distributed constraint logic programming machines. ACM SIG-PLAN Notices, 24(4):123–125, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p123-koegel/

.

Papazoglou:1989:IIE

[1657] M. P. Papazoglou, J. Bubenko, and M. Norrie. Intent: an integrated environment for distributed heterogeneous databases. ACM SIG-PLAN Notices, 24(4):126–128, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p126-papazoglou/

.

Segall:1989:COO

[1658] Edward J. Segall. Combining objectoriented techniques with data-driven parallel execution on message passing, MIMD computers. ACM SIG-PLAN Notices, 24(4):129–131, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p129-segall/

.

Herlihy:1989:LCO

[1659] Maurice P. Herlihy and Jeannette M. Wing. Linearizable concurrent objects. ACM SIGPLAN Notices, 24(4):133–135, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p133-herlihy/

.

Kaiser:1989:TCO

[1660] Gail E. Kaiser. Transactions for concurrent object-oriented programming systems. ACM SIGPLAN Notices, 24(4):136–138, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p136-kaiser/

.

Langworthy:1989:ECC

[1661] David E. Langworthy. Evaluating correctness criteria for transactions. ACM SIGPLAN Notices, 24(4):139–141, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p139-langworthy/

.

Martin:1989:CPV

[1662] Bruce Martin. Concurrent programming vs. concurrency control: shared events or shared data. ACM SIG-PLAN Notices, 24(4):142–144, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p142-martin/

.

Skarra:1989:CCC

[1663] Andrea H. Skarra. Concurrency control for cooperating transactions in an object-oriented database. ACM SIG-PLAN Notices, 24(4):145–147, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p145-skarra/

.

vandenBos:1989:PPC

[1664] Jan van den Bos. PROCOL: a protocolconstrained concurrent object-oriented language. ACM SIGPLAN Notices, 24(4):149–151, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p149-van_
                                        den_bos/.

Corradi:1989:POM

[1665] Antonio Corradi and Letizia Leonardi. PO: an object model to express parallelism. ACM SIGPLAN Notices, 24(4):152–155, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p152-corradi/

.

Christodoulakis:1989:MSS

[1666] Dimitris N. Christodoulakis. Modeling the semantics of Smalltalk-80 with Petri nets. ACM SIG-PLAN Notices, 24(4):156–158, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p156-christodoulakis/

.

Feather:1989:LIS

[1667] Martin S. Feather. Language issues in the specification and design of composite systems. ACM SIG-PLAN Notices, 24(4):159–161, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p159-feather/

.

Kramer:1989:SCO

[1668] Bernd Krämer. Specifying concurrent objects. ACM SIGPLAN Notices, 24(4):162–164, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p162-kramer/

.

Lim:1989:HOO

[1669] Jeong Lim and Ralph E. Johnson.

The heart of object-oriented concurrent programming. ACM SIG-PLAN Notices, 24(4):165–167, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p165-lim/

.

Shang:1989:CB

[1670] Lujun Shang, Changpeng Fan, and Lihui Xu. Concurrent behaviors. ACM SIGPLAN Notices, 24(4):168–170, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p168-lujun/

.

Luke:1989:PSC

[1671] Edward A. Luke, Helen C. Takacs, and William C. Welch. Position statement on concurrent objects for massively parallel architectures. ACM SIG-PLAN Notices, 24(4):171–173, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p171-luke/

.

Nierstrasz:1989:TMC

[1672] Oscar M. Nierstrasz. Two models of concurrent objects. ACM SIG-PLAN Notices, 24(4):174–176, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p174-nierstrasz/

.

Shibayama:1989:TCO

[1673] Etsuya Shibayama. Transformation of concurrent object systems. ACM SIG-PLAN Notices, 24(4):177–179, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p177-shibayama/

.

Ichikawa:1989:PDV

[1674] Itaru Ichikawa, Seiichi Aikawa, Mayumi Kamiko, Etsuo Ono, and Tomoharu Mohri. Program design visualization system for object-oriented programs. ACM SIGPLAN Notices, 24(4):181–183, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p181-ichikawa/

.

Kaplan:1989:GGT

[1675] Simon M. Kaplan and Steven K. Goering. Garp: a graphical/textual language for concurrent programming. ACM SIGPLAN Notices, 24(4):184–186, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p184-kaplan/

.

Uehara:1989:SWS

[1676] Minoru Uehara, Chisato Numaoka, Yasuhiko Yokote, and Mario Tokoro. Sarek: a window system interface for object-oriented concurrent programming languages. ACM SIG-PLAN Notices, 24(4):187–189, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p187-uehara/

.

Ball:1989:APM

[1677] Charles R. Ball, Theodore W. Leung, and Carl A. Waldspurger. Analyzing patterns of message passing. ACM SIGPLAN Notices, 24(4):191–193, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p191-ball/

.

Delagi:1989:ELA

[1678] Bruce A. Delagi and Nakul P. Saraiya. Elint in Lamina: application of a concurrent object language. ACM SIG-PLAN Notices, 24(4):194–196, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p194-delagi/

.

deJong:1989:UCF

[1679] Peter de Jong. The Ubik configurator: a fusion of messages, daemons, and rules. ACM SIG-PLAN Notices, 24(4):197–199, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/67386/p197-de_
jong/.

Hvannberg:1989:OBP

[1680] Ebba Thora Hvannberg and Mukkai S. Krishnamoorthy. An object-based parallel programming assistant. ACM SIGPLAN Notices, 24(4):200–202, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/plan/67386/p200-hvannberg/

.

Kafura:1989:COO

[1681] Dennis Kafura. Concurrent objectoriented real-time systems research. ACM SIGPLAN Notices, 24(4):203– 205, April 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:80/pubs/
citations/proceedings/plan/67386/
p203-kafura/.

Lester:1989:FOO

[1682] Martin Lester and Roger Christensen. Felix, an object-oriented operating system. ACM SIGPLAN Notices, 24(4):206–208, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p206-lester/

.

MacKay:1989:HOO

[1683] S. A. MacKay, W. M. Gentleman, D. A. Stewart, and M. Wein. Harmony as an object-oriented operating system. ACM SIGPLAN Notices, 24(4):209–211, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/67386/p209-mackay/

.

Nigro:1989:OBR

[1684] Libero Nigro and Francesco Tisato. Object-based real-time programming. ACM SIGPLAN Notices, 24(4):212–214, April 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/67386/p212-nigro/

.

Kokol:1989:USS

[1685] Peter Kokol. Using spreadsheet software to support metric’s life cycle activities. ACM SIGPLAN Notices, 24 (5):27–37, May 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1989:SBB

[1686] Paul R. Wilson. A simple bucketbrigade advancement mechanism for generation-based garbage collection. ACM SIGPLAN Notices, 24(5):38–46, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rain:1989:PIV

[1687] Mark Rain. Portable IPC on Vanilla Unix. ACM SIGPLAN Notices, 24 (5):47–56, May 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Green:1989:QAP

[1688] Collette Green, Brian Meek, and K. K. Siu. Quality assessment of processor documentation. ACM SIGPLAN Notices, 24(5):57–65, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robillard:1989:AC

[1689] Pierre-N. Robillard. Automating comments. ACM SIGPLAN Notices, 24 (5):66–70, May 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ringwood:1989:SFA

[1690] G. A. Ringwood. SLD: a folk acronym?

ACM SIGPLAN Notices, 24(5):71–75, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marriott:1989:POC

[1691] Kim Marriott and Harald Søndergaard.

On Prolog and the occur check problem. ACM SIGPLAN Notices, 24 (5):76–82, May 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casimir:1989:FGP

[1692] Rommert J. Casimir. Fourth generation problems. ACM SIGPLAN Notices, 24(5):83–86, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1989:CMS

[1693] Paul R. Wilson and Thomas G. Moher. A “card-marking” scheme for controlling intergenerational references in generation-based garbage collection on stock hardware. ACM SIGPLAN Notices, 24(5):87–92, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pintelas:1989:TSE

[1694] P. E. Pintelas, K. P. Ventouris, and M. D. Papassimakopoulou. A translator from small Euclid to Pascal. ACM SIGPLAN Notices, 24(5):93–101, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burch:1989:LBD

[1695] Carl Burch. A language-based design for portable data files. ACM SIG-PLAN Notices, 24(5):102–111, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donaho:1989:CEP

[1696] Jane Donaho. A comparison of extended Pascal and Ada. ACM SIG-PLAN Notices, 24(5):112–120, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marcus:1989:JOO

[1697] Bob Marcus. Joyce: an object-oriented decision tree builder. ACM SIG-PLAN Notices, 24(5):121–123, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:1989:CAU

[1698] R. Fischer. A case against using procedure calls for input/output. ACM SIG-PLAN Notices, 24(5):124–127, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bron:1989:BWC

[1699] C. Bron and E. J. Dijkstra. A better way to combine efficient string length encoding and zero-termination. ACM SIGPLAN Notices, 24(6):11–19, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wichmann:1989:SPI

[1700] B. A. Wichmann. Scientific processing in ISO-Pascal: a proposal to get the benefits of mixed precision floatingpoint. ACM SIGPLAN Notices, 24 (6):20–22, June 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sankar:1989:NDA

[1701] Sriram Sankar. A note on the detection of an Ada compiler bug while debugging an Anna program. ACM SIGPLAN Notices, 24(6):23–31, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Misra:1989:MSD

[1702] A. K. Misra and B. D. Chaudhary.

Modified structured decision table and its complexity. ACM SIGPLAN Notices, 24(6):32–34, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1989:PPC

[1703] Keith McC. Clarke. Procedure parameters can imitate sequence concatenation. ACM SIGPLAN Notices, 24 (6):35–36, June 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coggins:1989:MCL

[1704] James M. Coggins and Gregory Bollella. Managing C++ libraries. ACM SIGPLAN Notices, 24(6):37–48, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kanze:1989:HAT

[1705] James Kanze. Handling ambiguous tokens in LR-parsers. ACM SIG-PLAN Notices, 24(6):49–54, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mauney:1989:EPN

[1706] Jon Mauney. Extended Pascal is no problem. ACM SIGPLAN Notices, 24 (6):55–58, June 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Iannello:1989:DDC

[1707] Giulio Iannello, Antonino Mazzeo, and Giorgio Ventre. Definition of the DISC concurrent language. ACM SIG-PLAN Notices, 24(6):59–68, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davies:1989:DPF

[1708] G. L. Davies. Developments in Pascal-FC. ACM SIGPLAN Notices, 24 (6):69–76, June 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joslin:1989:EPN

[1709] David A. Joslin. Extended Pascal — numerical features. ACM SIG-PLAN Notices, 24(6):77–80, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1989:EES

[1710] Brian Meek and K. K. Siu. The effectiveness of error seeding. ACM SIG-PLAN Notices, 24(6):81–89, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ophel:1989:IMR

[1711] John Ophel. An improved mixture rule for pattern matching. ACM SIG-PLAN Notices, 24(6):91–96, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapman:1989:RAG

[1712] Nigel P. Chapman. Regular attribute grammars and finite state machines. ACM SIGPLAN Notices, 24(6):97–105, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Venkatesh:1989:FCE

[1713] G. A. Venkatesh. A framework for construction and evaluation of highlevel specifications for program analysis techniques. ACM SIGPLAN Notices, 24(7):1–12, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p1-venkatesh/

.

Cartwright:1989:SPD

[1714] Robert Cartwright and Matthias Felleisen. The semantics of program dependence. ACM SIGPLAN Notices, 24(7):13–27, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p13-cartwright/

.

Horwitz:1989:DAP

[1715] Susan Horwitz, Phil Pfeiffer, and Thomas Reps. Dependence analysis for pointer variables. ACM SIGPLAN Notices, 24(7):28–40, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p28-horwitz/

.

Balasundaram:1989:TSD

[1716] Vasanth Balasundaram and Ken Kennedy. A technique for summarizing data access and its use in parallelism enhancing transformations. ACM SIGPLAN Notices, 24(7):41–53, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p41-balasundaram/

.

Cytron:1989:AGD

[1717] Ron Cytron, Michael Hind, and Wilson Hsieh. Automatic generation of DAG parallelism. ACM SIG-PLAN Notices, 24(7):54–68, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p54-cytron/

.

Rogers:1989:PDT

[1718] Anne Rogers and Keshav Pingali.

Process decomposition through locality of reference. ACM SIG-PLAN Notices, 24(7):69–80, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p69-rogers/

.

Kranz:1989:MHP

[1719] David A. Kranz, Robert H. Halstead, Jr., and Eric Mohr. Mul-T: a high-performance parallel Lisp. ACM SIGPLAN Notices, 24(7):81–90, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p81-kranz/

.

Gross:1989:PCP

[1720] Thomas Gross, Angelika Zobel, and Markus Zolg. Parallel compilation for a parallel machine. ACM SIG-PLAN Notices, 24(7):91–100, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p91-gross/

.

Horwat:1989:ECP

[1721] Waldemar Horwat, Andrew A. Chien, and William J. Dally. Experience with CST: programming and implementation. ACM SIGPLAN Notices, 24(7):101–109, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p101-horwat/

.

Koopman:1989:FLC

[1722] Philip J. Koopman, Jr. and Peter Lee. A fresh look at combinator graph reduction (or, having a TIGRE by the tail). ACM SIG-PLAN Notices, 24(7):110–119, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p110-koopman/

.

Farrow:1989:VCB

[1723] Rodney Farrow and Alec G. Stanculescu. A VHDL compiler based on attribute grammar methodology. ACM SIGPLAN Notices, 24(7):120–130, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p120-farrow/

.

Vogt:1989:HOA

[1724] H. H. Vogt, S. D. Swierstra, and M. F. Kuiper. Higher order attribute grammars. ACM SIGPLAN Notices, 24(7):131–145, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p131-vogt/

.

Chambers:1989:COC

[1725] Craig Chambers and David Ungar. Customization: optimizing compiler technology for SELF, a dynamically-typed object-oriented programming language. ACM SIG-PLAN Notices, 24(7):146–160, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p146-chambers/

.

Cormack:1989:LSP

[1726] Gordon V. Cormack. An LR substring parser for noncorrecting syntax error recovery. ACM SIG-PLAN Notices, 24(7):161–169, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p161-cormack/

.

Salomon:1989:SNP

[1727] Daniel J. Salomon and Gordon V. Cormack. Scannerless NSLR(1) parsing of programming languages. ACM SIG-PLAN Notices, 24(7):170–178, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p170-salomon/

.

Heering:1989:IGP

[1728] J. Heering, P. Klint, and J. Rekers. Incremental generation of parsers. ACM SIGPLAN Notices, 24(7):179–191, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p179-heering/

.

Boehm:1989:TIP

[1729] Hans-J. Boehm. Type inference in the presence of type abstraction. ACM SIGPLAN Notices, 24(7):192–206, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p192-boehm/

.

OToole:1989:TRF

[1730] James William O’Toole, Jr. and David K. Gifford. Type reconstruction with first-class polymorphic values. ACM SIGPLAN Notices, 24(7):207–217, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p207-o_toole/

.

Jouvelot:1989:RAC

[1731] Pierre Jouvelot and David K. Gifford. Reasoning about continuations with control effects. ACM SIG-PLAN Notices, 24(7):218–226, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p218-jouvelot/

.

Emmelmann:1989:BGE

[1732] Helmut Emmelmann, Friedrich-Wilhelm Schröer, and Rudolf Landwehr. BEG: a generator for efficient back ends. ACM SIGPLAN Notices, 24(7):227–237, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p227-emmelmann/

.

Fraser:1989:LWC

[1733] Christopher W. Fraser. A language for writing code generators. ACM SIGPLAN Notices, 24(7):238–245, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p238-fraser/

.

Hwu:1989:IFE

[1734] Wen-Mei W. Hwu and Pohua P.

Chang. Inline function expansion for compiling C programs. ACM SIG-PLAN Notices, 24(7):246–257, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p246-chang/

.

Bernstein:1989:SCM

[1735] David Bernstein, Dina Q. Goldin, Martin C. Golumbic, Hugo Krawczyk, Yishay Mansour, Itai Nahshon, and Ron Y. Pinter. Spill code minimization techniques for optimizing compilers. ACM SIGPLAN Notices, 24(7):258–263, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p258-bernstein/

.

Gupta:1989:RAC

[1736] Rajiv Gupta, Mary Lou Soffa, and Tim Steele. Register allocation via clique separators. ACM SIG-PLAN Notices, 24(7):264–274, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p264-gupta/

.

Briggs:1989:CHR

[1737] Preston Briggs, Keith D. Cooper, Ken Kennedy, and Linda Torczon. Coloring heuristics for register allocation. ACM SIGPLAN Notices, 24(7):275–284, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p275-briggs/

.

Schonberg:1989:FDA

[1738] Edith Schonberg. On-the-fly detection of access anomalies. ACM SIG-PLAN Notices, 24(7):285–297, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p285-schonberg/

.

Sarkar:1989:DAP

[1739] Vivek Sarkar. Determining average program execution times and their variance. ACM SIGPLAN Notices, 24(7):298–312, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p298-sarkar/

.

Goldberg:1989:GRC

[1740] Benjamin Goldberg. Generational reference counting: a reducedcommunication distributed storage reclamation scheme. ACM SIG-PLAN Notices, 24(7):313–321, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p313-goldberg/

.

Atkinson:1989:ECP

[1741] Russ Atkinson, Alan Demers, Carl Hauser, Christian Jacobi, Peter Kessler, and Mark Weiser. Experiences creating a portable Cedar. ACM SIG-PLAN Notices, 24(7):322–329, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p322-atkinson/

.

Wilson:1989:DMP

[1742] Paul R. Wilson and Thomas G. Moher. Demonic memory for process histories. ACM SIGPLAN Notices, 24(7):330–343, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p330-wilson/

.

Chi:1989:UMR

[1743] Chi-Hung Chi and Hank Dietz. Unified management of registers and cache using liveness and cache bypass. ACM SIGPLAN Notices, 24(7):344–355, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/73141/p344-chi/

.

Sagri:1989:ROC

[1744] Mohommedyusuf M. Sagri. Rated and operating complexity of program — an extension to McCabe’s theory of complexity measure. ACM SIGPLAN Notices, 24(8):8–12, August 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andersen:1989:HEJ

[1745] B. Andersen. Hypercube experiments with Joyce. ACM SIGPLAN Notices, 24(8):13–22, August 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zuse:1989:SMU

[1746] H. Zuse and P. Bollmann. Software metrics: using measurement theory to describe the properties and scales of static software complexity metrics. ACM SIGPLAN Notices, 24(8):23–33, August 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baldwin:1989:CUP

[1747] W. A. Baldwin. A case for using procedure calls for I/O. ACM SIG-PLAN Notices, 24(8):34–36, August 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milanese:1989:AOS

[1748] Vitaliano Milanese. Aspects of OBERON suitable for graphics applications. ACM SIGPLAN Notices, 24 (8):37–45, August 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bothe:1989:PHS

[1749] K. Bothe, B. Hohberg, Ch. Horn, and O. Wikarski. A portable high-speed PASCAL to C translator. ACM SIG-PLAN Notices, 24(9):60–65, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:1989:SDA

[1750] Defu Zhang, Peigen Yang, and Zhongxiu Sun. SN-2: a data acquisition and processing system. ACM SIGPLAN Notices, 24(9):66–71, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hirvisalo:1989:XDR

[1751] Vesa Hirvisalo, Jari Arkko, Juha Kuusela, Esko Nuutila, and Markku Tamminen. XE design rationale: Clu revisited. ACM SIGPLAN Notices, 24(9):

72–79, September 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huntbach:1989:IGC

[1752] Matthew Huntbach. Implementing a graph-colouring algorithm in Parlog. ACM SIGPLAN Notices, 24(9): 80–85, September 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klaassen:1989:TPL

[1753] Arno J. Klaassen and Anton M. van Wezenbeek. A three-processor Lisp machine architecture based on statistical analysis of Common Lisp programs. ACM SIGPLAN Notices, 24(9): 86–91, September 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:1989:TPD

[1754] Ming-Yuan Y. Zhu. Transformational programming: the derivation of a Prolog interpretation algorithm. ACM SIGPLAN Notices, 24(9):92– 111, September 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiener:1989:PEM

[1755] Richard S. Wiener and Lewis J. Pinson. A practical example of multiple inheritance in C++. ACM SIGPLAN Notices, 24(9):112–115, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winter:1989:PIP

[1756] Kirt A. Winter and Curtis R. Cook. A prototype intelligent prettyprinter for Pascal. ACM SIGPLAN Notices, 24(9):116–125, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jung:1989:IND

[1757] Chul-Doo Jung and Ernest Sibert. Indirect naming in distributed programming languages. ACM SIGPLAN Notices, 24(9):126–132, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rapin:1989:PON

[1758] Ch. Rapin. Procedural objects in Newton. ACM SIGPLAN Notices, 24(9): 133–141, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frost:1989:APC

[1759] R. A. Frost. Application programming as the construction of language interpreters. ACM SIGPLAN Notices, 24(9):142–151, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsdell:1989:AFP

[1760] John D. Ramsdell. The Alonzo functional programming language. ACM SIGPLAN Notices, 24(9):152–157, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beck:1989:LTO

[1761] K. Beck and W. Cunningham. A laboratory for teaching object oriented thinking. ACM SIGPLAN Notices, 24 (10):1–6, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosson:1989:PSM

[1762] M. B. Rosson and E. Gold. Problemsolution mapping in object-oriented design. ACM SIGPLAN Notices, 24(10): 7–10, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberherr:1989:CTO

[1763] K. J. Lieberherr and A. J. Riel. Contributions to teaching object-oriented design and programming. ACM SIG-PLAN Notices, 24(10):11–22, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1989:DOG

[1764] P. R. Wilson and T. G. Moher. Design of the opportunistic garbage collector. ACM SIGPLAN Notices, 24(10): 23–35, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schelvis:1989:IDT

[1765] M. Schelvis. Incremental distribution of timestamp packets: a new approach to distributed garbage collection. ACM SIGPLAN Notices, 24(10):37–48, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:1989:EIS

[1766] C. Chambers, D. Ungar, and E. Lee. An efficient implementation of SELF a dynamically-typed object-oriented language based on prototypes. ACM SIG-PLAN Notices, 24(10):49–70, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirfs-Brock:1989:OOD

[1767] R. Wirfs-Brock and B. Wilkerson. Object-oriented design: a responsibility-driven approach. ACM SIGPLAN Notices, 24(10):71–75, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dietrich:1989:SLO

[1768] W. C. Dietrich, Jr., L. R. Nackman, and F. Gracer. Saving legacy with objects. ACM SIGPLAN Notices, 24(10): 77–83, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1989:GNB

[1769] W. H. Harrison, P. F. Sweeney, and J. J. Shilling. Good news, bad news: experience building software development environment using the objectoriented paradigm. ACM SIGPLAN Notices, 24(10):85–94, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanDenBos:1989:PPO

[1770] J. Van Den Bos and C. Laffra. PRO-COL: a parallel object language with protocols. ACM SIGPLAN Notices, 24 (10):95–102, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tomlinson:1989:ISE

[1771] C. Tomlinson and V. Singh. Inheritance and synchronization with enabled-sets. ACM SIGPLAN Notices, 24(10):103–112, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marques:1989:EOS

[1772] J. A. Marques and P. Guedes. Extending the operating system to support an object-oriented environment. ACM SIGPLAN Notices, 24(10):113– 122, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenton:1989:POO

[1773] J. Fenton and K. Beck. Playground: an object-oriented simulation system with agent rules for children of all ages. ACM SIGPLAN Notices, 24(10): 123–137, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:1989:NAF

[1774] S. S. Adams and A. K. Nabi. Neural agents — a frame of mind. ACM SIG-PLAN Notices, 24(10):139–149, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Block:1989:EFL

[1775] F. P. Block and N. C. Chan. An extended frame language. ACM SIG-PLAN Notices, 24(10):151–157, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walther:1989:SSP

[1776] S. S. Walther and R. L. Peskin. Strategies for scientific prototyping in Smalltalk. ACM SIGPLAN Notices, 24(10): 159–167, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jerrell:1989:FMA

[1777] M. E. Jerrell. Function minimization and automatic differentiation using C++. ACM SIGPLAN Notices, 24(10): 169–173, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McDonald:1989:OOP

[1778] J. A. McDonald. Object-oriented programming for linear algebra. ACM SIGPLAN Notices, 24(10):175–184, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrel:1989:VAA

[1779] P. J. Ferrel and R. F. Meyer. Vamp:

the Aldus application framework. ACM SIGPLAN Notices, 24(10):185–188, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shah:1989:DOR

[1780] A. V. Shah, J. H. Hamel, R. A. Borsari, and J. E. Rumbaugh. DSM: an object-relationship modeling language. ACM SIGPLAN Notices, 24(10):191– 202, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thuraisingham:1989:MSO

[1781] M. B. Thuraisingham. Mandatory security in object-oriented database systems. ACM SIGPLAN Notices, 24(10): 203–210, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dixon:1989:FMD

[1782] R. Dixon, T. McKee, M. Vaughan, and P. Schweizer. A fast method dispatcher for compiled languages with multiple inheritance. ACM SIGPLAN Notices, 24(10):211–214, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dussud:1989:TIC

[1783] Patrick H. Dussud. TICLOS: an implementation of CLOS for the Explorer family. ACM SIGPLAN Notices, 24(10):215–219, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paepcke:1989:PCR

[1784] A. Paepcke. PCLOS: a critical review. ACM SIGPLAN Notices, 24(10): 221–253, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1989:UMM

[1785] H. Muller, J. Rose, J. Kempf, and T. Stansbury. The use of multimethods and method combination in a CLOS based window interface. ACM SIG-PLAN Notices, 24(10):239–253, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barry:1989:PRT

[1786] B. M. Barry. Prototyping a realtime embedded system in Smalltalk. ACM SIGPLAN Notices, 24(10):255– 265, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russo:1989:VMB

[1787] Vincent F. Russo and Roy H. Campbell. Virtual memory and backing storage management in multiprocessor operating systems using object-oriented design techniques. ACM SIGPLAN Notices, 24(10):267–278, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conner:1989:OAM

[1788] R. C. H. Conner, A. Dearle, R. Morrison, and A. L. Brown. An object addressing mechanism for statically typed languages with multiple inheritance. ACM SIGPLAN Notices, 24(10):279– 285, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Page:1989:OOM

[1789] T. W. Page, Jr., S. Berson, W. Cheng, and R. R. Muntz. An object-oriented modelling environment. ACM SIG-PLAN Notices, 24(10):287–296, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenblatt:1989:OTT

[1790] W. R. Rosenblatt, J. C. Wileden, and Alexander L. Wolf. OROS: toward a type model for software development environments. ACM SIGPLAN Notices, 24(10):297–304, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graube:1989:MC

[1791] N. Graube. Metaclass compatibility. ACM SIGPLAN Notices, 24(10):305– 315, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferber:1989:CRC

[1792] J. Ferber. Computational reflection in class based object-oriented languages. ACM SIGPLAN Notices, 24(10):317– 326, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foote:1989:RFS

[1793] B. Foote and R. E. Johnson. Reflective facilities in Smalltalk-80. ACM SIG-PLAN Notices, 24(10):327–335, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reenskaug:1989:ELS

[1794] Trygve Reenskaug and Anne Lise Skaar. An environment for literate Smalltalk programming. ACM SIG-PLAN Notices, 24(10):337–345, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1989:EDM

[1795] Y.-P. Shan. An event-driven modelview-controller framework for Smalltalk. ACM SIGPLAN Notices, 24(10):

347–352, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shiling:1989:TSV

[1796] J. J. Shiling and P. F. Sweeney. Three steps to views: extending the objectoriented paradigm. ACM SIGPLAN Notices, 24(10):353–361, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Almarode:1989:RBD

[1797] J. Almarode. Rule-based delegation for prototypes. ACM SIGPLAN Notices, 24(10):363–370, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minsky:1989:CDE

[1798] N. H. Minsky and E. Rozenshtein. Controllable delegation: an exercise in lawgoverned systems. ACM SIGPLAN Notices, 24(10):371–380, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moloney:1989:CTU

[1799] J. Moloney, A. Borning, and B. Freeman-Benson. Constraint technology for user-interface construction in Thing-Lab II. ACM SIGPLAN Notices, 24 (10):381–388, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freeman-Benson:1989:MMC

[1800] B. N. Freeman-Benson. A module mechanism for constraints in Smalltalk. ACM SIGPLAN Notices, 24(10): 389–396, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madsen:1989:VCP

[1801] O. L. Madsen and B. Møller-Pedersen. Virtual classes: a powerful mechanism in object-oriented programming. ACM SIGPLAN Notices, 24(10):397– 406, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pedersen:1989:EOI

[1802] C. H. Pedersen. Extending ordinary inheritance schemes to include generalization. ACM SIGPLAN Notices, 24(10):407–417, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briot:1989:PEM

[1803] J.-P. Briot and P. Cointe. Programming with explicit metaclasses in Smalltalk-80. ACM SIGPLAN Notices, 24(10):419–431, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1989:DSI

[1804] W. Cook and J. Palsberg. A denotational semantics of inheritance and its correctness. ACM SIGPLAN Notices, 24(10):433–443, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ohori:1989:STI

[1805] A. Ohori and P. Buneman. Static type inference for parametric classes. ACM SIGPLAN Notices, 24(10):445– 450, October 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Canning:1989:IST

[1806] P. S. Canning, W. R. Cook, W. L.

Hill, and W. G. Olthoff. Interfaces for strongly-typed object-oriented programming. ACM SIGPLAN Notices, 24(10):457–467, October 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1989:CRD

[1807] Anonymous. Critical research directions in programming languages. ACM SIGPLAN Notices, 24(11):10– 24, November 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boone:1989:BRW

[1808] John Boone. Book review: Writing Readable Ada — A Case Study Approach. ACM SIGPLAN Notices, 24 (11):27–28, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heering:1989:SDF

[1809] J. Heering, P. R. H. Hendriks, P. Klint, and J. Rekers. The syntax definition formalism SDF — reference manual. ACM SIGPLAN Notices, 24(11): 43–75, November 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herring:1989:RNG

[1810] C. Herring and J. I. Palmore. Random number generators are chaotic. ACM SIGPLAN Notices, 24(11):76– 79, November 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salomon:1989:CPS

[1811] D. J. Salomon and G. V. Cormack. Corrections to the paper: Scannerless NSLR(1) Parsing of Programming Languages. ACM SIGPLAN Notices, 24(11):80–83, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jouvelot:1989:RPM

[1812] P. Jouvelot and B. Dehbonei. Recursive Pattern Matching on concrete data types. ACM SIGPLAN Notices, 24(11):84–93, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pezaris:1989:SRC

[1813] S. D. Pezaris. Shift-reduce conflicts in LR parsers. ACM SIGPLAN Notices, 24(11):94–95, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pokkunuri:1989:OOP

[1814] B. P. Pokkunuri. Object Oriented Programming. ACM SIGPLAN Notices, 24(11):96–101, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeKerf:1989:NPO

[1815] J. L. F. De Kerf. A note on the power operators “Loops are Harmful”. ACM SIGPLAN Notices, 24(11):102– 108, November 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grogono:1989:PTC

[1816] P. Grogono and A. Bennett. Polymorphism and type checking in objectoriented languages. ACM SIGPLAN Notices, 24(11):109–115, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Albizuri-Romero:1989:PIG

[1817] M. B. Albizuri-Romero. Programming interactively with a graphical abstract programming language. ACM SIGPLAN Notices, 24(11):116– 124, November 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanenbaum:1989:DVF

[1818] A. S. Tanenbaum, M. F. Kaashoek, K. G. Langendoen, and C. J. H. Jacobs. The design of very fast portable compilers. ACM SIGPLAN Notices, 24 (11):125–131, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanRoy:1989:UEP

[1819] P. Van Roy. A useful extension to Prolog’s definite clause grammar notation. ACM SIGPLAN Notices, 24(11):132– 134, November 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonsson:1989:PA

[1820] D. Jonsson. Pancode assessed. ACM SIGPLAN Notices, 24(12):17–20, December 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nemecek:1989:AFD

[1821] S. Nemecek. An anomaly found during the implementation of a 2dimensional dataflow editor for constructing dataflow graphs in NAKL: not another keypunch language. ACM SIGPLAN Notices, 24(12):21–23, December 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levien:1989:INP

[1822] R. L. Levien. Io: a new programming notation. ACM SIGPLAN Notices, 24(12):24–31, December 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shopiro:1989:EMI

[1823] J. E. Shopiro. An example of multiple inheritance in C++: a model of the iostream library. ACM SIGPLAN Notices, 24(12):32–36, December 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trono:1995:CTS

[1824] John A. Trono. A comparison of three strategies for computing letter oriented, minimal perfect hashing functions. ACM SIGPLAN Notices, 30 (4):29–35, April 1, 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper introduces a simple modification that produces dramatic speedups to the algorithms of Cichelli [2] and Cook and Oldehoeft [458].

Ahmad:1989:SHI

[1825] B. Ahmad and A. Adiga. Systematic high-level interrupts in forth. In ACM [1971], pages 1–7. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p1-ahmad/

.

Davis:1989:FIM

[1826] R. H. Davis. Forth interface to MS-DOS^{T M} interrupts. In ACM [1971], pages 9–12. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p9-davis/

.

Davis:1989:BFM

[1827] R. H. Davis. A 16-bit forth model for a 32-bit addressable host. In ACM [1971], pages 13–15. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p13-davis/.

Feierbach:1989:CCT

[1828] G. Feierbach. Creating conditional transform words. In ACM [1971], pages 17–19. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p17-feierbach/.

Forsley:1989:FU

[1829] L. Forsley. Forth in the USSR. In ACM [1971], pages 21–22. ISBN 0-89791-316-7. LCCN QA76.54 .S54

                                        1989. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73312/p21-forsley/.

Forsley:1989:OOT

[1830] L. Forsley. Object oriented thinking in forth. In ACM [1971], pages 23–24. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/73312/p23-forsley/

.

Frenger:1989:NNC

[1831] P. Frenger. A neural net controller. In ACM [1971], pages 25–27. ISBN 0-89791-316-7. LCCN QA76.54 .S54

                                        1989. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73312/p25-frenger/.

Glass:1989:DRM

[1832] H. Glass, M. Mellen, and T. Hand. The design of a real-time multitasking operating system kernel for the Harris RTX 2000. In ACM [1971], pages 29–33. ISBN 0-89791-316-7. LCCN QA76.54 .S54

                                        1989. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73312/p29-glass/.

Hand:1989:AIC

[1833] T. Hand. Architecture impact on compiler construction. In ACM [1971], pages 35–39. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

                                        //www.acm.org:80/pubs/citations/

proceedings/plan/73312/p35-hand/

.

Hand:1989:MRS

[1834] T. Hand. Metrics for real-time systems.

In ACM [1971], pages 41–44. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p41-hand/.

Hoselton:1989:LTF

[1835] R. F. Hoselton. Load time forth capabilities. In ACM [1971], pages 45–47. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p45-hoselton/

.

Jun:1989:ADS

[1836] M. S. Jun. Asynchronous designs for systolic arrays. In ACM [1971], pages 49–58. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p49-jun/.

Lee:1989:MS

[1837] Y. M. Lee and E. Conjura. Multitasking/multiuser systems. In ACM [1971], pages 59–62. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p59-lee/.

Leverenz:1989:HSF

[1838] H. Leverenz. A high speed foreground mulitasker for forth. In ACM [1971], pages 63–69. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p63-leverenz/.

Lisle:1989:NPC

[1839] L. G. Lisle. A new process controller: a study in modularity. In ACM [1971], pages 71–79. ISBN 0-89791-316-7. LCCN QA76.54 .S54

                                        1989. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73312/p71-lisle/.

Mikiten:1989:DFB

[1840] B. C. Mikiten and S. Mikiten. Development of a forth based remote weather station. In ACM [1971], pages 81–83. ISBN 0-89791-316-7. LCCN QA76.54 .S54

                                        1989. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73312/p81-mikiten/.

Morgenstern:1989:AQU

[1841] L. Morgenstern. Anonymous ‘Things’ used as locals. In ACM [1971], pages 85–88. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/73312/p85-morgenstern/

.

Morgenstern:1989:SNE

[1842] L. Morgenstern. Safety nets for error trapping. In ACM [1971], pages 89–91. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/73312/p89-morgenstern/

.

Neubert:1989:RTS

[1843] K.-D. Neubert and M. Ulbrich. Real time signal processing with the RTX 2000. In ACM [1971], pages 93–

97. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p93-neubert/

.

Orr:1989:FRA

[1844] J. L. Orr. Forth real-time automata for experiment control. In ACM [1971], pages 99–102. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73312/p99-orr/.

Stamps:1989:MIR

[1845] V. O. Stamps. Mammography information and reporting system in forth. In ACM [1971], pages 103–107. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p103-stamps/.

Vannorman:1989:CMR

[1846] R. Vannorman. Cooperative multitasking on the RTX 2000. In ACM [1971], pages 109–112. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p109-vannorman/.

Weinstein:1989:VMS

[1847] D. Weinstein. Virtual memory on systems without hardware support. In ACM [1971], pages 113–122. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p113-weinstein/.

Wong:1989:DRD

[1848] K. C. Wong and J. Cornacchio. Directory replication in distributed systems.

In ACM [1971], pages 123–127. ISBN 0-89791-316-7. LCCN QA76.54 .S54

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73312/p123-wong/.

Monna:1987:SVC

[1849] Z. Monna and A. Pnueli. Specification and verification of concurrent programs by A ∀. In ACM [1963], pages 1–2. ISBN ???? LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/41625/p1-monna/

.

Herlihy:1987:ACO

[1850] M. P. Herlihy and J. M. Wing. Axioms for concurrent objects. In ACM [1963], pages 13–26. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p13-herlihy/.

Widom:1987:CIT

[1851] J. Widom, D. Gries, and F. B. Schneider. Completeness and incompleteness of trace-based network proof systems. In ACM [1963], pages 27–38. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p27-widom/.

Jouvelot:1987:SPP

[1852] P. Jouvelot. Semantic parallelization: a practical exercise in abstract interpretation. In ACM [1963], pages 39–48. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p39-jouvelot/.

Saraswat:1987:CLP

[1853] V. A. Saraswat. The concurrent logic programming language CP: Definition and operational semantics. In ACM [1963], pages 49–62. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p49-saraswat/.

Allen:1987:ADS

[1854] r. Allen, D. Callahan, and K. Kennedy. Automatic decomposition of scientific programs for parallel execution. In ACM [1963], pages

63–76.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p63-allen/.

Kohlbecker:1987:MDS

[1855] E. E. Kohlbecker and M. Wand.

Macro-by-example: Deriving syntactic transformations from their specifications. In ACM [1963], pages

77–84.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p77-kohlbecker/.

Cai:1987:BPA

[1856] J. Cai and R. Paige. Binding performance at language design time. In ACM [1963], pages 85–97. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/41625/p85-cai/.

Gelernter:1987:EFC

[1857] D. Gelernter, S. Jagannathan, and T. London. Environments as first class objects. In ACM [1963], pages

98–110.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p98-gelernter/.

Jaffar:1987:CLP

[1858] J. Jaffar and J.-L. Lassez. Constraint logic programming. In ACM [1963], pages 111–119. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p111-jaffar/.

Nielson:1987:SAD

[1859] F. Nielson. Strictness analysis and denotational abstract interpretation. In ACM [1963], pages

120–131.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p120-nielson/.

Hall:1987:CSS

[1860] C. V. Hall and D. S. Wise. Compiling strictness into streams. In ACM [1963], pages 132–143. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p132-hall/.

Kuo:1987:SA

[1861] T.-M. Kuo and P. Mishra. On strictness and its analysis. In ACM [1963], pages 144–155. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p144-kuo/.

Moshier:1987:LPS

[1862] M. D. Moshier and W. C. Rounds. A logic for partially specified data structures. In ACM [1963], pages

156–167.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p156-moshier/.

Chase:1987:IBT

[1863] D. R. Chase. An improvement to bottom-up tree pattern matching. In ACM [1963], pages 168–

177.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p168-chase/.

Caspi:1987:LDL

[1864] P. Caspi, D. Pilaud, N. Halbwachs, and J. A. Plaice. LUSTRE: a declarative language for real-time programming. In ACM [1963], pages

178–188.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p178-caspi/.

Apt:1987:AFD

[1865] K. R. Apt, N. Francez, and S. Katz.

Appraising fairness in distributed languages. In ACM [1963], pages 189–198. ISBN ???? LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/41625/p189-apt/

.

Stark:1987:CTS

[1866] E. W. Stark. Concurrent transition system semantics of process networks. In ACM [1963], pages

199–210.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p199-stark/.

Oles:1987:SCP

[1867] F. J. Oles. Semantics for concurrency without powerdomains. In ACM [1963], pages 211–222. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p211-oles/.

Huizing:1987:FAR

[1868] C. Huizing, R. Gerth, and W. P. deRoever. Full abstraction of a realtime denotational semantics for an OCCAM-like language. In ACM [1963], pages 223–236. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p223-huizing/.

Breazu–Tannen:1987:CVC

[1869] V. Breazu-Tannen and A. R. Meyer. Computable values can be classical. In ACM [1963], pages 238–

245.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p238-breazu-tannen/.

Leivant:1987:SFF

[1870] D. Leivant and T. Fernando. Skinny and fleshy failures of relative completeness. In ACM [1963], pages

246–252.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p246-leivant/.

Meyer:1987:ETP

[1871] A. R. Meyer, J. C. Mitchell, E. Moggi, and R. Statman. Empty types in polymorphic lambda calculus. In ACM [1963], pages 253–262. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/41625/p253-meyer/

.

Bernstein:1987:SAL

[1872] D. Bernstein, J. M. Jaffe, and M. Rodeh. Scheduling arithmetic and load operations in parallel with no spilling. In ACM [1963], pages

263–273.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p263-bernstein/.

Neirynck:1987:CAS

[1873] A. Neirynck and A. J. Demers.

Computation of aliases and support sets. In ACM [1963], pages

274–283.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p274-neirynck/.

Lee:1987:RCG

[1874] P. Lee and U. Pleban. A realistic compiler generator based on high-level semantics: another progress report. In ACM [1963], pages 284–295. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/41625/p284-lee/

.

Cheriton:1987:EMR

[1875] D. R. Cheriton and M. E. Wolf.

Extensions for multi-module records in conventional programming languages. In ACM [1963], pages

296–306.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
41625/p296-cheriton/.

Wadler:1987:VWP

[1876] P. Wadler. Views: a way for pattern matching to cohabit with data abstraction. In ACM [1963], pages

307–313.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p307-wadler/.

Felleisen:1987:CAH

[1877] M. Felleisen and D. P. Friedman. A calculus for assignments in higherorder languages. In ACM [1963], page 314. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        41625/p314-felleisen/.

Alpern:1988:DEV

[1878] B. Alpern, M. N. Wegman, and F. K. Zadeck. Detecting equality of variables in programs. In ACM [1966], pages 1–11. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73560/p1-alpern/.

Rosen:1988:GVN

[1879] B. K. Rosen, M. N. Wegman, and F. K. Zadeck. Global value numbers and redundant computations. In ACM [1966], pages 12–27. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73560/p12-rosen/.

Mitchell:1988:EM

[1880] J. C. Mitchell and R. Harper. The essence of ML. In ACM [1966], pages 28–46. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        73560/p28-mitchell/.

Lucassen:1988:PES

[1881] J. M. Lucassen and D. K. Gifford.

Polymorphic effect systems. In ACM [1966], pages 47–57. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p47-lucassen/.

Kfoury:1988:PEM

[1882] A. J. Kfoury, J. Tiuryn, and P. Urzyczyn. A proper extension of ML with an effective type-assignment. In ACM [1966], pages 58–69. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p58-kfoury/.

Cardelli:1988:SSN

[1883] L. Cardelli. Structural subtyping and the notion of power type. In ACM [1966], pages 70–79. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p70-cardelli/.

Kamin:1988:ISD

[1884] S. Kamin. Inheritance in smalltalk-80: a denotational definition. In ACM [1966], pages 80–87. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p80-kamin/.

Stansifer:1988:TIS

[1885] R. Stansifer. Type inference with subtypes. In ACM [1966], pages

88–97.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p88-stansifer/.

Nielson:1988:ABT

[1886] F. Nielson and R. H. Nielson. Automatic binding time analysis for a typed λ-calculus. In ACM [1966], pages 98–106. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p98-nielson/.

Young:1988:CIE

[1887] J. Young and P. Hudak. A collecting interpretation of expressions. In ACM [1966], pages

107–118.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p107-young/.

Wadler:1988:SAA

[1888] P. Wadler. Strictness analysis aids time analysis. In ACM [1966], pages

119–132.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p119-wadler/.

Horwitz:1988:INV

[1889] S. Horwitz, J. Prins, and T. Reps. Integrating non-intering versions of programs. In ACM [1966], pages

133–145.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p133-horwitz/.

Horwitz:1988:APD

[1890] S. Horwitz, J. Prins, and T. Reps. On the adequacy of program dependence graphs for representing programs. In ACM [1966], pages

146–157.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p146-horwitz/.

Johnson:1988:SPC

[1891] G. F. Johnson and D. Duggan. Stores and partial continuations as first-class objects in a language and its environment. In ACM [1966], pages

158–168.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p158-johnson/.

Williams:1988:SSC

[1892] J. H. Williams and E. L. Wimmers. Sacrificing simplicity for convenience: Where do you draw the line? In ACM [1966], pages

169–179.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p169-williams/.

Felleisen:1988:TPF

[1893] M. Felleisen. The theory and practice of first-class prompts. In ACM [1966], pages 180–190. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p180-felleisen/.

Meyer:1988:TFA

[1894] A. R. Meyer and K. Sieber. Towards fully abstract semantics for local variables. In ACM [1966], pages

191–203.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p191-meyer/.

Montenyohl:1988:CFA

[1895] M. Montenyohl and M. Wand. Correct flow analysis in continuation semantics. In ACM [1966], pages

204–218.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p204-montenyohl/.

Garland:1988:IMR

[1896] S. J. Garland and J. V. Guttag. Inductive methods for reasoning about abstract data types. In ACM [1966], pages 219–228. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p219-garland/.

Bloom:1988:BCB

[1897] B. Bloom, S. Istrail, and A. R. Meyer. Bisimulation can’t be traced. In ACM [1966], pages

229–239.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p229-bloom/.

Bouge:1988:CAS

[1898] L. Bougé and N. Francez. A compositional approach to superimposition. In ACM [1966], pages

240–249.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p240-bouge/.

Vardi:1988:TFC

[1899] M. Y. Vardi. A temporal fixpoint calculus. In ACM [1966], pages

250–259.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p250-vardi/.

Debray:1988:EDA

[1900] S. K. Debray. Efficient dataflow analysis of logic programs. In ACM [1966], pages 260–273. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p260-debray/.

Carroll:1988:IDF

[1901] M. D. Carroll and B. G. Ryder. Incremental data flow analysis via dominator and attribute update. In ACM [1966], pages

274–284.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p274-carroll/.

Ruggieri:1988:LAD

[1902] C. Ruggieri and T. P. Murtagh.

Lifetime analysis of dynamically allocated objects. In ACM [1966], pages 285–293. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p285-ruggieri/.

Pelegri–Llopart:1988:OCG

[1903] E. Pelegrí-Llopart and S. L. Graham. Optimal code generation for expression trees: an application BURS theory. In ACM [1966], pages

294–308.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p294-pelegri-llopart/.

Lam:1988:COA

[1904] M. Lam. Compiler optimizations for asynchronous systolic array programs. In ACM [1966], pages 309–318. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/73560/p309-lam/

.

Irigoin:1988:SP

[1905] F. Irigoin and R. Triolet. Supernode partitioning. In ACM [1966], pages 319–329. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
73560/p319-irigoin/.

Baxter:1989:PDG

[1906] W. Baxter and H. R. Bauer, III. The program dependence graph and vectorization. In ACM [1972], pages 1–11. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p1-baxter/.

Parsons-Selke:1989:RSP

[1907] R. Parsons-Selke. A rewriting semantics for program dependence graphs. In ACM [1972], pages

12–24.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p12-parsons-selke/.

Cytron:1989:EMC

[1908] R. Cytron, J. Ferrante, B. K. Rosen, M. N. Wegman, and F. K. Zadeck. An efficient method of computing static single assignment form. In ACM [1972], pages 25–35. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p25-cytron/.

Sagiv:1989:RCA

[1909] S. Sagiv, O. Edelstein, N. Francez, and M. Rodeh. Resolving circularity in attribute grammars with applications to data flow analysis (preliminary version). In ACM [1972], pages 36–48. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p36-sagiv/.

Cooper:1989:FIA

[1910] K. D. Cooper and K. Kennedy. Fast interprocedural alias analysis. In ACM [1972], pages 49–59. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p49-cooper/.

Wadler:1989:HMA

[1911] P. Wadler and S. Blott. How to make ad-hoc polymorphism less ad hoc. In ACM [1972], pages

60–76.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p60-wadler/.

Rey:1989:TCR

[1912] D. Réy. Type checking records and variants in a natural extension of ML. In ACM [1972], pages 77–88. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/75277/p77-rey/.

Paulin-Mohring:1989:EPP

[1913] C. Paulin-Mohring. Extracting ω’s programs from proofs in the calculus of constructions. In ACM [1972], pages 89–104. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p89-paulin-mohring/.

Kanellakis:1989:PUM

[1914] P. C. Kanellakis and J. C. Mitchell. Polymorphic unification and ML typing. In ACM [1972], pages

105–115.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p105-kanellakis/.

Yelick:1989:MTS

[1915] K. A. Yelick and J. L. Zachary. Moded type systems for logic programming. In ACM [1972], pages

116–124.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p116-yelick/.

Hickey:1989:CCA

[1916] T. J. Hickey. CLP and constraint abstraction. In ACM [1972], pages

125–133.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p125-hickey/.

Gaifman:1989:FAC

[1917] H. Gaifman and E. Shapiro. Fully abstract compositional semantics for logic programs. In ACM [1972], pages 134–142. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p134-gaifman/.

Thomsen:1989:CHO

[1918] B. Thomsen. A calculus of higher order communicating systems. In ACM [1972], pages 143–154. ISBN

????  LCCN ????  URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/75277/p143-thomsen/

.

Jonsson:1989:FAT

[1919] B. Jonsson. A fully abstract trace model for dataflow networks. In ACM [1972], pages 155–165. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/75277/p155-jonsson/

.

Emerson:1989:ETR

[1920] E. A. Emerson, T. Sadler, Srinivasan, and J. Efficient temporal reasoning (extended abstract). In ACM [1972], pages 166–178. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p166-emerson/.

Pnueli:1989:SRM

[1921] A. Pnueli and R. Rosner. On the synthesis of a reactive module. In ACM [1972], pages 179–

190.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p179-pnueli/.

Attie:1989:SCS

[1922] P. C. Attie and E. A. Emerson.

Synthesis of concurrent systems with many similar sequential processes (extended abstract). In ACM [1972], pages 191–201. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p191-attie/.

Cardelli:1989:MTS

[1923] L. Cardelli, J. Donahue, M. Jordan, B. Kalsow, and G. Nelson. The Modula-3 type system. In ACM [1972], pages 202–212. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p202-cardelli/.

Abadi:1989:DTS

[1924] M. Abadi, L. Cardelli, B. Pierce, and G. Plotkin. Dynamic typing in a statically-typed language. In ACM [1972], pages 213–227. ISBN

????  LCCN ????  URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/75277/p213-abadi/

.

Meseguar:1989:RMP

[1925] J. Meseguar. Relating models of polymorphism. In ACM [1972], pages

228–241.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p228-meseguar/.

Monteleone:1989:GCT

[1926] G. Monteleone. Generalized conjunctive types. In ACM [1972], pages

242–249.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p242-monteleone/.

Dershowitz:1989:RRR

[1927] N. Dershowitz and S. Kaplan. Rewrite, rewrite, rewrite, rewrite, rewrite, . . .. In ACM [1972], pages 250–

259.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        75277/p250-dershowitz/.

Parker:1989:POP

[1928] D. S. Parker. Partial order programming (extended abstract). In ACM [1972], pages 260–266. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/75277/p260-parker/

.

Baudinet:1989:TLP

[1929] M. Baudinet. Temporal logic programming is complete and expressive. In ACM [1972], pages 267–

280.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p267-baudinet/.

Kelsey:1989:RCP

[1930] R. Kelsey and P. Hudak. Realistic compilation by program transformation (detailed summary). In ACM [1972], pages 281–292. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p281-kelsey/.

Ael:1989:CCS

[1931] A. W. Ael and T. Jim. Continuationpassing, closure-passing style. In ACM [1972], pages 293–302. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/75277/p293-appel/

.

Gopinath:1989:CEF

[1932] K. Gopinath and J. L. Hennessy.

Copy elimination in functional languages. In ACM [1972], pages

303–314.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p303-gopinath/.

Pugh:1989:ICF

[1933] W. Pugh and T. Teitelbaum. Incremental computation via function caching. In ACM [1972], pages

315–328.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p315-pugh/.

Larsen:1989:BTP

[1934] K. G. Larsen and A. Skou. Bisimulation through probabilistic testing (preliminary report). In ACM [1972], pages 344–352. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
75277/p344-larsen/.

Stevenson:1981:PSBb

[1935] David Stevenson. A proposed standard for binary floating-point arithmetic: draft 8.0 of IEEE Task P754. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. 36 pp. See [1937, 1938].

Waser:1982:IAD

[1936] Shlomo Waser and Michael J. Flynn. Introduction to Arithmetic for Digital Systems Designers. Holt, Reinhart, and Winston, New York, NY, USA, 1982. ISBN 0-03-060571-7. xvii + 308 pp. LCCN TK7895 A65 W37 1982. Master copy output on Alphatype CRS high-resolution phototypesetter. This book went to press while the IEEE 754 Floating-Point Standard was still in development; consequently, some of the material on that system was invalidated by the final Standard (1985) [1937].

IEEE:1985:AIS

[1937] ANSI/IEEE 754-1985, Standard for Binary Floating-Point Arithmetic.

IEEE, New York, NY, USA, August 12, 1985. ISBN 1-55937-653-8. 20 pp. US$35.00. URL http://

standards.ieee.org/reading/ieee/
std/busarch/754-1985.pdf. Revised

1990. A preliminary draft was published in the January 1980 issue of IEEE Computer, together with several companion articles [10, 12, 7, 9, 11, 8, 1935]. The final version was republished in [1940, 1939]. See also [1936]. Also standardized as IEC 60559 (1989-01) Binary floating-point arithmetic for microprocessor systems.

IEEE:1985:ASI

[1938] IEEE Task P754. ANSI/IEEE 754-1985, Standard for Binary Floating-Point Arithmetic. IEEE, New York, August 12 1985. A preliminary draft was published in the January 1980 issue of IEEE Computer, together with several companion articles [10, 12, 7, 9, 11, 8, 1935]. Available from the IEEE Service Center, Piscataway, NJ, USA.

IEEE:1985:ISBb

[1939] IEEE Computer Society Standards Committee.Working group of the Microprocessor Standards Subcommittee and American National Standards Institute. IEEE standard for binary floating-point arithmetic. ANSI/IEEE Std 754-1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. 18 pp. See [1937].

IEEE:1987:ISB

[1940] IEEE. IEEE standard for binary floating-point arithmetic. ACM SIG-PLAN Notices, 22(2):9–25, February 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shriver:1987:RDO

[1941] Bruce Shriver and Peter Wegner, editors. Research directions in objectoriented programming. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-19264-0. 585 pp. LCCN QA76.76.D47 R47 1987.

ACM:1980:PAS

[1942] ACM, editor. Proceedings of the ACM-SIGPLAN Symposium on the Ada Programming Language. Boston, MA, USA, 9–11 December, 1980, volume 15(11) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, November 1980. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Sammet:1980:HCE

[1943] J. E. Sammet and J. A. N. Lee, editors. HOPL conference excerpts [ACM SIG-PLAN History of Programming Languages Conference (Los Angeles, June 1–3, 1978)]. ACM Press, New York, NY, USA, 1980. ACM Member: Audio US$17, Video (U-Matic) US$105; Non-member: Audio US$20, Video (U-Matic) US$125. 2 cassettes, 54 and 49 min.

Abrahams:1981:TMA

[1944] P. Abrahams, editor. Text manipulation: Proceedings of the ACM SIG-PLAN/SIGOA symposium (Portland, OR, June 8–10, 1981). ACM Press, New York, NY, USA, 1981. ISBN 0-89791-043-5, 0-89791-050-8. LCCN QA76.7 .S54 v.16:6. Published as ACM SIGPLAN Notices, v. 16, no. 6, and ACM SIGOA newsletter, vol. 2, no. 1/2, spring/summer 1981.

ACM:1981:PAS

[1945] ACM, editor. Proceedings of the ACM SIGPLAN SIGOA Symposium on Text Manipulation. Portland, OR, USA, 8– 10 June, 1981, volume 16(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1981:HPL

[1946] R. L. Wexelblat, editor. History of programming languages. Proceedings of the ACM SIGPLAN conference (Los Angeles, Calif., June 1–3, 1978). Academic Press Inc., New York, USA, June 1981. ISBN 0-12-745040-8. LCCN QA76.7 .H56 1978. US$45.

ACM:1982:PSS

[1947] ACM, editor. Proceedings of the SIG-PLAN ’82 Symposium on Compiler Construction. Boston, MA, USA, 23– 25 June, 1982, volume 17(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1982:LP

[1948] K. L. (Keith L.) Clark and S.-Å. (Sten-Åke) Tarnlund, editors. Logic programming, volume 16 of A.P.I.C. studies in data processing. Academic Press Inc., New York, USA, 1982. ISBN 0-12-175520-7. LCCN QA76.6 .L587 1982.

Levy:1982:ACP

[1949] Beth Helene Levy, editor. ACM ’81. Conference proceedings (Los Angeles, CA, Nov. 9–11, 1981), ACM, New York, 1982. ACM Press, New York, NY, USA, 1982. ISBN 0-89791-049-4. LCCN TK 7885 A84p 1981, QA76 .A8 1981.

ACM:1983:ASS

[1950] ACM, editor. ACM SIGSOFT/ SIGPLAN Software Engineering Symposium on High-Level Debugging. Pacific Grove, CA, USA, 20–23 March, 1983, volume 18(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, August 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1983:PSS

[1951] ACM, editor. Proceedings of the SIG-PLAN ’83 Symposium on Programming Language Issues in Software Systems. San Francisco, CA, USA, 27–29 June, 1983, volume 18(6) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, June 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1983:PAS

[1952] Mark Scott Johnson, editor. Proceedings of the ACM SIGSOFT/SIGPLAN Software Engineering Symposium on High-Level Debugging: Pacific Grove, California, March 20–23, 1983, ACM Software Engineering Notes, vol. 8, no. 4, August, 1983; ACM SIGPLAN Notices, vol. 18, no. 8, August, 1983. ACM Press, New York, NY, USA, 1983. ISBN 0-89791-111-3. LCCN ???? US$13.50 members; US$18 nonmembers. Published in joint issue of Software engineering notes, v. 8, no. 4, August, ACM order no. 593830.

Wexelblat:1983:PLI

[1953] R. L. Wexelblat, editor. Programming language issues in software systems: Proceedings of the [ACM] SIG-PLAN ’83 symposium (San Francisco, CA, June 27–29, 1983). ACM Press, New York, NY, USA, 1983. ISBN 0-89791-108-3. LCCN QA76.7 .S54 v.18:6. members US$14; others US$19. ACM SIGPLAN Notices, vol. 18, no. 6, June 1983.

ACM:1984:PAS

[1954] ACM, editor. Proceedings of the ACM SIGSOFT/SIGPLAN Software Engineering Symposium on Practical Software Development Environments. Pittsburgh, PA, USA, 23–25 April, 1984, volume 19(5) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1984:PSS

[1955] ACM, editor. Proceedings of the SIGPLAN 84 Symposium on Compiler Construction. Montréal, Qué., Canada, 17–22 June, 1984, volume 19(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henderson:1984:PAS

[1956] Peter Henderson, editor. Proceedings of the ACM SIGSOFT/SIGPLAN Software Engineering Symposium on Practical Software Development Environments: Pittsburgh, Pennsylvania, April 23–25, 1984. ACM Press, New York, NY, USA, 1984. ISBN 0-89791-131-8. LCCN QA76.6 .A14 1984. US$13.00 (members); US$17.00 non-members. Published in joint issue of Software engineering notes, v. 9, no. 3, May 1984.

VanDeusen:1984:CCP

[1957] M. Van Deusen, editor. Compiler construction: Proceedings of the ACM SIGPLAN ’84 symposium (Montreal, Canada, June 17–22, 1984), volume 19(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1984. ISBN 0-89791-139-3. LCCN QA 76.7 A1095 v.19 no.6. US$19 members; US$25 non-members.

Deutsch:1985:PAS

[1958] L. Peter Deutsch, editor. Proceedings of the ACM SIGPLAN 85 Symposium on Language Issues in Programming Environments: papers presented at the symposium in Seattle, Washington, 25–28 June, 1985, volume 20(7) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 vol. 20, no. 7. US$16 members; US$21 non-members. SIGPLAN Notices, Vol. 20, No. 7, July 1985.

ACM:1986:OOO

[1959] ACM, editor. OOPSLA ’86. Object-Orientated Programming Systems, Languages and Applications. Conference Proceedings. Portland, OR, USA, 29 September–2 October 1986, volume 21(11) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1986:OOP

[1960] ACM, editor. Object-Orientated Programming Workshop. Yorktown Heights, NY, USA, 9–13 June, 1986, volume 21(10) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1986:PSS

[1961] ACM, editor. Proceedings of the SIG-PLAN ’86 Symposium on Compiler Construction. Palo Alto, CA, USA, 25–27 Junek 1986, volume 21(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1986:CC

[1962] Richard L. Wexelblat, editor. Proceedings of the SIGPLAN ’86 Symposium on Compiler Construction, Palo Alto, Calif., June 25–27, 1986, volume 21(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 1986. ISBN 0-89791-197-0 (paperback). ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA 76.7 A1095 v.21 no.7 1986. US$23. ACM order number: 548860.

ACM:1987:PFA

[1963] ACM, editor. POPL ’87. Fourteenth Annual ACM SIGACT-SIGPLAN Symposium on Principles of programming languages, January 21– 23, 1987, Munich, W. Germany. ACM Press, New York, NY, USA,

1987.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/contents/proceedings/plan/
41625/index.html.

Henderson:1987:PSD

[1964] Peter Henderson, editor. Practical Software Development Environments: Proceedings of the ACM SIGSOFT/ SIGPLAN Software Engineering Symposium (Palo Alto, California, Dec. 9– 11, 1986), volume 22(1) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, January 1987. ISBN 0-89791-212-8 (paperback). LCCN QA 76.7 A1095 v.22 no.1 1987. US$19.00.

Wexelblat:1987:IIT

[1965] Richard L. Wexelblat, editor. Proceedings of the SIGPLAN ’87 Symposium on Interpreters and Interpretive Techniques, St. Paul, Minnesota, June 24–26, 1987. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-235-7. LCCN QA76.7 .S54 v.22:7. US$23.00. SIGPLAN Notices, v. 22, no. 7 (July 1987).

ACM:1988:PPC

[1966] ACM, editor. POPL ’88. Proceedings of the conference on Principles of programming languages, January 13–15, 1988, San Diego, CA. ACM Press, New York, NY, USA,

1988.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/contents/proceedings/plan/
73560/index.html.

ACM:1988:PSC

[1967] ACM, editor. Proceedings of the SIG-PLAN ’88 Conference on Programming Language Design and Implementation, Atlanta, Georgia, June 22–24, 1988, volume 23(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 1988. ISBN 0-89791-269-1 (paperback). LCCN QA 76.7 A1095 v.23 no.7 1988.

Henderson:1988:PSD

[1968] Peter Henderson, editor. Proceedings of the ACM SIGSOFT/SIGPLAN Software Engineering Symposium on Practical Software Development Environments: Boston, Massachusetts, November 28–30, 1988, volume 24(2) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, February 1988. ISBN 0-89791-290-X. LCCN QA76.7 .S54 v.24:2. US$21.00. ACM order number: 594880.

Miller:1988:PAS

[1969] Barton Paul Miller and Thomas LeBlanc, editors. Proceedings of the ACM SIGPLAN and SIGOPS Workshop on Parallel and Distributed Debugging, May 5–6, 1988, University of Wisconsin, Madison, Wisconsin 53706, volume 24(1) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, January 1988. ISBN 0-89791-296-9 (paperback). LCCN QA76.7 .S54 v.24:1.

Wexelblat:1988:PPE

[1970] Richard L. Wexelblat, editor. Parallel programming: experience with applications, languages and systems: Proceedings of the ACM/SIGPLAN PPEALS 1988 (New Haven, CT, July 19–21, 1988). ACM Press, New York, NY, USA, 1988. ISBN 0-89791-276-4. LCCN QA76.7 .S54 v.23:9. SIGPLAN Notices, Vol. 23, No. 9 (Sept. 1988).

ACM:1989:PFA

[1971] ACM, editor. Proceedings of the first annual workshop for the ACM special interest group on Forth: SIGForth ’89 [Four Seasons Hotel, Austin, TX, Feb. 17–19, 1989]. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-316-7. LCCN QA76.54 .S54 1989.

ACM:1989:PPS

[1972] ACM, editor. POPL ’89. Proceedings of the sixteenth annual ACM symposium on Principles of programming languages, January 11–13, 1989, Austin, TX. ACM Press, New York, NY, USA, 1989. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/contents/proceedings/plan/
                                        75277/index.html.

ACM:1989:SCP

[1973] ACM, editor. SIGPLAN ’89 Conference on Programming Language Design and Implementation, volume 24(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1989:SIP

[1974] ACM, editor. SPLOS-III proceedings / Third International Conference on Architectural Support for Programming Languages and Operating Systems, Boston, Massachusetts, April 3–6, 1989. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-300-0. LCCN QA76.9.A73 C64 v.17:2.

US$24.00. Also published as Computer Architecture News, v. 17, no. 2 (Apr. 1989), Operating Systems Review, v. 23, special issue (Apr. 1989), and SIGPLAN Notices, v. 24, special issue (May 1989).

Agha:1989:PAS

[1975] Gul Agha, Peter Wegner, and Akinori Yonezawa, editors. Proceedings of the ACM SIGPLAN Workshop on Object-Based Concurrent Programming, volume 24(4) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bosworth:1989:OPS

[1976] George Bosworth, editor. Objectoriented programming: systems, languages and applications OOPSLA ’89 conference proceedings (New Orleans, LA, Oct. 1–6, 1989 ). ACM Press, New York, NY, USA, 1989. ISBN 0-89791-333-7. LCCN QA76.7.O2 1989. ACM SIGPLAN Notices, vol. 24, no. 10, 1989.

Gul:1989:OCP

[1977] Agha Gul, Peter Wegner, and Akinori Yonezawa, editors. Proceedings of the ACM SIGPLAN Workshop on Object-Based Concurrent Programming, San Diego, California, September 26–27, 1988, volume 24(4) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, April 1989. ISBN 0-89791-304-3 (paperback). LCCN QA76.7 .S54 v.24:4. US$18.

Wexelbalt:1989:PDD

[1978] Richard L. Wexelbalt, editor. Parallel and Distributed Debugging: Proceedings of the ACM SIGPLAN and SIGOPS Workshop (Madison, WI, May 5–6, 1989). ACM Press, New York, NY, USA, 1989. ISBN 0-89791-296-9. LCCN QA 76.7 A1095 v.24 no.1 1989. US$24. SIGPLAN Notices, vol. 24, no. 1, Jan. 1989.

Wexelblat:1989:PLD

[1979] R. L. Wexelblat, editor. Proceedings of the SIGPLAN ’89 Conference on Programming Language Design and Implementation, Portland, Oregon, June 21– 23, 1989. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-306-X (paperback). LCCN QA 76.7 A1095 v. 24 no. 7 1989. ACM SIGPLAN Notices, vol. 24 no. 7, July 1989.